A Complete Bibliography of Publications in the Journal of Informetrics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

26 November 2024 Version 1.23

Title word cross-reference

16 [CG10]. 17 [SMP12]. 2 [Cam15, DGDG13, YDW^{+}23]. 20 [CA11]. 37 [BMHD11]. 404 [KK13]. 5 [Cam15]. _{2} [RYY10]. _{m} [Sch08]. A [Bur07b]. C [YZF13]. F^{3} [BGS19]. g [AK15, ER19b, Gag09, GR09, Rou08, Sch10a, Sch10b, Sch13e, Tol08, WTR^{+}22, Woe08a, vEW08, vEW09]. H [ZTG11, ZYCK12, ACHVH09, AOdFC12, BBMP12, BI10, BBL15, BD07a, BD10b, BMD10, BMHD11, BM12, Bor13a, Bor14b, BGT22, BN21, Bur07a, Bur07b, Bur13a, Cam14, zChHY13, CGD^{+}21, CMMT13, CB07, Die15, Egg08a, Egg08b, Egg10b, Egg10a, ER19a, ER19b, ER21, Fas24, FM11b, Gag09, GM12, GP12, GS10, GR09, Kos13a, KHC11a, KHC11b, KHC13, LR08, LF12, LRWY13, LY14, LJHH23, Mal16, MR13, NT12, NO10, PSZZ12, Pra13a, Pra13b, Rou08, RYY10, RY12b, RGZSC13, Sch08, SMP12, Sch13c, Sch13d, Sch15, Smo21b, Van08, WWY21, WTR^{+}22, WLH11, YWZ^{+}23, YR08, ZYZ18, ZRY11, ZY12, vEW08, vEW09]. h^{(}2) [Rou08]. hg [MT10]. k [CPN24, YWZ^{+}23, ZLL^{+}18]. L [LY14]. L^{2}[0, T] [ER19b]. N [Khr09, BO19]. p [WTR^{+}22]. π [Vin13b]. ψ [Lat21, WTR^{+}22]. Q [dABJGMC21, RLG13, RLG14]. q^{2} [CAHVH10]. R [Gag11, LCY19]. R^{2} [Rou08]. S [KK21, Pra13c]. w [Per17b]. Z [PS13, Lun07].

-authors [BO19]. -based [FM11b]. -bubble [Pra13b, RGZSC13]. -core [Cam14, zChHY13, KHC11a]. -cores [Fas24, LRWY13]. -Degree [ZRY11, ZY12]. -gram [Khr09]. -Index [ACHVH09, AK15, AOdFC12, BBMP12, BBL15, BN21, Bur07a, Bur07b, Bur13a, CMMT13, CB07, Egg08a, GM12, GS10, GR09, KHC11b, KHC13, Lat21, LR08, LF12, LY14, LJHH23, Mal16, MR13, MT10, NT12, NO10, PSZZ12, Per17b, Rou08, RYY10, RY12b, Sch08, Sch10b, Sch13c, Sch13e, Sch13d, Sch15, Tol08, Van08, Vin13b, WTR^{+}22, Woe08a, WLH11, YWZ^{+}23, YR08, ZYZ18, Bor14b, CAHVH10, BGS19, Bur07b, PS13, Pra13c, Rou08, YZF13]. -indices [ER19b, Gag09, vEW08, vEW09]. -like [CGD^{+}21]. -means [CPN24, ZLL^{+}18]. -measures [RLG14, RLG13]. -page [KK13]. -related [ZTG11]. -score [Lun07]. -Sequence [LY14]. -step [YDW^{+}23, YWZ^{+}23]. -subnet [WWY21]. -tail [zChHY13, KHC11a]. -tails [LRWY13]. -type [BI10, Egg08b, Egg10b, Egg10a, ER19a, ER21, SMP12, Smo21b]. -year [Cam15, DGDG13].

10 [KE17, Ron13, YAY^{+}17]. 101316 [BGCGS23]. 11 [BCGM17a, FM17b]. 12 [Lin18a, SV24]. 12/1 [SV24]. 13 [BTZY24, HLB^{+}21]. 13/4 [BTZY24]. 14 [MdMAGBLI21b]. 14/Article [WZHH23]. 145 [AORC12a]. 17 [Lin24, ZSF^{+}24]. 17/1 [Lin24]. 18 [DGRCDG24a]. 18/3 [DGRCDG24a]. 19 [KYhK23, LZH^{+}22, RGZSC23, SLLM23, UKL23]. 19/101395 [ZSF^{+}24].

2.0 [TS12]. 2010 [BdMA11]. 2011 [BdMAL12]. 2014 [FM17a]. 2022 [Ano22a, Ano22f, Ano22g]. 2023 [Ano23a, Ano23f, Ano23g, Ano23h]. 2024 [Ano24a, Ano24f, Ano24g, Ano24h]. 21st [BI08]. 2nd [Kos10a].

357 [BCGM17b].

5 [AORC12a, AORC12b]. 5G [TWC^{+}23]. 5SQual [MGLF09].

6 [Ron13]. 63 [AORC12b].

7 [RLG14]. 787 [BCGM17a]. 799 [BW22].

9 [RGRE16].

abilities [CLW24]. ability [WTR^{+}22]. above [dNL15]. above-journal [dNL15]. Abramo [GTD16, Zit16]. abrupt [BSdA^{+}21]. absolute [Vin12]. absorptive [CMN08]. abstracts [LPM16, SD22, SCZ23, BMS^{+}09]. academia [HG14, NLSM23, YPL^{+}22]. Academic [BPI22, Ort14, RL18, ADR16, ADS16, ALR^{+}23, CH23, CtRlY21, Cso21, DFHCT21, DFCGB15, DFFBR16, DGV18b, FC10, FM11b, FCZ24, Hag15, HZW19, HWL22, HHC14, wH16, KT18, KBP23, LBML23, LYYY09, LHWS18, Luo21, Ma12, MMOMLC18, MD19a, MKM^{+}22, Ort15, OZZW24, PYSM24, RA15, RDF^{+}18, SCBB11, SZZ22, SCML23, Szi21, TV21, VZMFAB16, WG10a, WHL17, WZ21b, WLLL18, WOLD22, WKCC23, XWW^{+}24, YVW^{+}13, YLCC22, YF23, ZXS21, ZW21, ZZL^{+}21, ZXH^{+}23, ZWL^{+}24, ZZ21b, ZY17, KTA18, KT18, The17b, The18c]. academic-network [ZXS21]. academics [Yur17a]. Academy [XlYcG18]. accelerator [WCG21]. acceptance [SLNC13]. accepted [BMS^{+}09]. Access [SZZ23, CPM^{+}07, ED15, Fra09, LB16, LLW21, MMCvLLC18, MdMAGBLI21b, MdMAGBLI21a, SLB13, SZ10, Zha24, vV19]. Accidentality [MGS22]. according [Cam14, Cam15, OA10, OLRF11, WHL17]. account [HSSM22, WZ21a]. accounting [ADR13, ADG16, ADD21a, ADDWD24]. accounts [PS13, Sch08]. accreditation [Hag14b]. accuracies [CJW14]. Accuracy [Mil13, The18c, ACD11a, BMD10, TF17a, WW16, XLSW24, vZ13]. achievements [ZJMX23]. acknowledgement [HKXS23]. acknowledgements [PHMSML17]. ACM [PXS24]. across [ASS22, AN18, DLGC13, HZW19, KBCR10, KP24, MO24, Ort15, Per10b, RC12a, RL22, WZ14, Wu15]. Actions [TDCM24]. activities [HDC12, KS16]. Activity [WHW^{+}14, Pet21, RY12a, WG10a, YZ18, SK14]. actor [Let16]. actors [BSW17, Lat21]. ad [ZY21]. adaptive [CSH18, KPB09, LF23]. add [BD10b]. added [Wan24]. addendum [RLG14]. Adding [LDS^{+}12]. address [BO12, LHT18]. addressing [WRGL23]. adjusted [BW13, GLY21, KSS16, NMNJM13]. adopters [Zha24]. adoption [LBP24]. advance [CS22, KRM14]. advanced [BL13]. advancement [MJ13]. advantage [ADD22, BM13b, HG14, LPM16, LWWZ22, Sch13e, TS14, WZ18]. Advantages [CB07, LH22, SHL24]. advisee [RFMC17]. advisor [RFMC17]. advisor-advisee [RFMC17]. affect [AD17a, AD17c, HTZS23, KMJ^{+}19, Web17, WOLD22]. affecting [CH23, HKXS23, SLL^{+}12, TT22, Van13]. affects [LTH18]. affiliation [GKR^{+}24, RL18, XZKS22]. affiliations [KCH24]. Africa [GW17]. after [BLW14]. again [SG12]. against [CLR^{+}17, Kos12a]. age [CH23, LBML23, NLC17, RA15, ZLYC21]. age-based [ZLYC21]. ageing [GG19, TLV^{+}21]. agent [SG12]. agent-based [SG12]. ages [Sta21]. agglomeration [TLZM24]. aggregated [BdMA11, LBZ16, NF12, dNL15]. Aggregating [GM12]. aggregation [Csa21, LG11]. aging [HGJZ17, WFC14]. agreements [RJDD08]. AI [GKR^{+}24, Ser10]. aid [DG21]. AIDS [XLG21]. al [KE17, YAY^{+}17]. al. [But17, FM17b, Glä17, MCD16]. algorithm [CXMR07, JS22, RY23, WZ21b, WXZ24]. algorithmic [CCL18]. Algorithmically [HSMB18, RCW15, SA18]. algorithms [DVG16, RS10b]. alignment [ETGS21]. alike [SVKM21]. all-author [WJW^{+}19, WDLX23, ZS08]. Alliance [RW10]. allocating [LF12]. allocation [KK15, XZS21, XWZY21]. allows [GP12]. along [BSdA^{+}21]. alphabet [LT13]. alphabetical [Wal12]. Alphabetization [LT13]. alphanumeric [BvR11]. also [ACD14]. alternative [Bor13a, GL11, JZ21, JY21, Kon14, LRCRC13, Pra13a, SW12, The17c]. alternatives [NO07, RC14]. alters [Hag15]. Altmetric [DGRCDG24a, DGRCDG24b, Ort15, Ort18, Ort21b, SKM22, Bor14c]. Altmetric.com [TN18, YZYW21]. Altmetrics [LH24, ZDJ^{+}16, AAK^{+}21, BGLS22, Bor14a, Bor14c, BH18, BHA19, CWT^{+}24, YXX^{+}17]. altruism [KB14]. always [CySlY18, HR16, WvEW13]. American [AYL14, HZW19]. among [ADM17, AF15, LWB19b, LF21, LAL09, SYP^{+}16, SLB13, WTR^{+}22, YDCL13]. analyses [GWMP16, LHTW15, Ort14, ZS08]. Analysing [Fra07, GHvdB^{+}14]. Analysis [BPI22, Cam14, CH23, CK22, IST10, PSG^{+}21, SJGCGB^{+}24, AAH11, Abb16, ACD12e, ACD12d, ADM13a, ADD19a, ADD21a, ADD21b, ADD21d, Ada18, ASG12, AYL14, AKML18, AC17, BGLS22, BBL19, BRA19, BHCH17, BBS21, BMD07, BP11, BMHD11, BM12, BM13a, BLM13, BL17, BH18, BHM21, BH22, BvECW18, BRN18, BN21, BNbH17, BvR11, But11, CMN08, CX16, CXZ^{+}22a, CD23, CML24, CHD^{+}24, CY22, Cho21, Cho23, CRS14, Cop19, CT16, Cso18, DLM^{+}17, DN17, DFCGB15, Din11b, DLGC13, DLGT19, DMCLH22, EAG22, ED15, Eom08, FGW^{+}22, Fra10a, FM11a, FMM16a, FJW^{+}15, FHH17, GG09, GAW21, GKR^{+}24, Glä07, GÁCC17, HB23b, HLD^{+}23, HTZS23, HPNM22, HYC15, JWL17, JSD14, JS14, JHK^{+}16, JY21, KM22, Ke21a, KYC12, KJS16, KPS22, KB17, KZH^{+}24, KTA18, Kua24, LG11, LL15, LBSA13]. analysis [LLW21, LH24, LR08, LJM13, LHWS18, LH21, LCL^{+}22, LGS^{+}24, LBP24, LDSD18, Ma12, Mag12, Mal16, Mal10, MMCvLLC18, Mar15, MMY17, MF21, MBD17, OG13, Ort21a, PWY21, PYCH19, PXS24, PMAT13, Pis22, PD24, PT09, QCVQdMA10, RS16, RNB18a, RNB24, RS10b, RA15, San13b, SMP12, SG07, SLK^{+}23, SYP^{+}16, SY24, SHD15, SZ10, SLNC13, TMK16, The18c, The19a, TWH08, TD10, TLV^{+}21, Tsa14, VZAMMFAB15, VTPF15, Vas09, VCT22, Wal12, Wal16a, Wal16c, Wal16d, WW16, WCG21, WML^{+}21, Wil16, WB16a, WB16b, Woe08a, WZ14, WW15, WLLL18, XLLZ14, XZDS21, XWK^{+}21, YG14, YF24, YVW^{+}13, YW15, YHWZ16, YHS18, YWW^{+}23, YSS23, YSPW09, YS21, YY23, Yur17a, ZH18, ZLJ^{+}10, ZSH^{+}16, ZWM17, ZLW21, ZWDS21a, ZWDS21b, ZMZ24, ZKJ^{+}24, SK21, TSM09]. analytic [CCL16]. analytical [BGGB13]. analytics [SAB^{+}16, ZWM^{+}21]. analyze [BTZY21, BTZY24, CGD^{+}21, RFMC17]. Analyzing [LBD^{+}19, WYC23, FM10b, vEW14]. Angewandte [BMS^{+}09, BD09, BD10a]. Annual [BH22, PHANPP13]. anomalous [Pet21]. ANVUR [BCGM17a, AD17b]. Any [FMM14, Cla16, MBD17, vdBHS17a]. anyway [GTD16]. API [TS12]. apples [ACD13b]. applicability [Jar07b, LH24]. Application [CtRlY21, CY22, ETGS21, OG13, San14, ACD11b, BBMP12, BSW17, BBL19, BSdMAM16, BGCG^{+}24, CCC^{+}18, CLHHVH11, DLM^{+}19, GPNA13, HG14, HR18, Jar07a, JFZZ23, MF21, MdGNS22, RSRP19, RYY10, San11b, KdBOK15b]. applications [AORC11a, AORC12a, BD07b, BLV10, Cho21, Cho23, Egg07a, KCH21b, LRE17, MJB11, RH13, WKCC23, vdBM22]. applied [BMD08b, LBMM14, Tol08]. Applying [HS11]. Approach [BGS21, AHRCV17, AF17b, BL21, BMD08a, BW11, BLWSE11, BLW13, Bor19, BM11b, BM14, BM16, Bur13b, CCL16, CMCMM22, CLC11, CtRlY21, CML24, CKKY21, CLHHVH11, CCL18, DBS15, DFHCT21, FLB22, FC17, GG19, GFGALG21, GPGBMA10, HM21a, HB23a, HLD^{+}23, HR18, HLHC22, JLAL23, Khr09, Kos10a, Kos12a, KR24, LL15, MLCL21, MD19b, OLRF11, PD21, PT09, PCCN21, QZY^{+}24, QM23, RGRE15, RGRE16, Reh21, RNB24, Ron12, Ron13, Sch14c, Sch12, SW12, SLK^{+}23, Ser10, SAP18, The16f, Tol13a, TD10, TdR21, VPAE24, Vîi17, WvEN10, WRGL23, WXZ24, XMW^{+}24, YPYJ23, ZSZ24, ZCM^{+}24, ZF22, HCS19]. Approaches [WRB^{+}11, ADZ18, AC09, BMD10, BL12, BLW13, CT16, HDB22, HZL23, TLM23, WvE13]. appropriate [WvE15]. approximate [dSSOH07]. approximating [vZ13]. approximation [Ron18]. Arabic [Khr09, VGRB16]. Arbitrage [Smo21a]. arbitrariness [Sch13c, Sch13e]. Archetypal [SW13]. area [AMKA21, Ley12]. areas [BH22, TBTB19]. arguments [Gau17]. arising [SKM22]. arithmetic [The16f]. armed [YXXZ24]. arrays [LRE17]. Article [CKZ^{+}23, JJ24, WZHH23, BD10a, LH24, NMNJM13, Ono16, PPFKB19, The19a, TLV^{+}21, VG10a, WNH^{+}23, ZWBA21, Wal14]. article-level [LH24, PPFKB19]. Articles [GAB11, ACD14, AD15c, AH22, BGGB13, BPU^{+}14, BvECW18, BSdA^{+}21, CGXW22, CKZ^{+}23, CH14, CPM^{+}07, DLM^{+}17, GLY21, HGJZ17, HWL22, HCL13, KTA18, LQLJ14, LQLJ15, LLW21, LHDH21, RS10a, RY12b, San11b, San15, SL15, TW14a, TF15a, TS16, The16b, The18b, The18c, VG10a, Vin13a, Wal17a, WS21, WZ21b, WZCC24, WWXL24, YP22, ZWG^{+}19, ZXH^{+}23, dlP11]. Artificial [SD11, HSE22, QLS21, RV24, TLD^{+}21, VRK23, WOLD22, ZYCK12]. arXiv [Sil23]. Asian [NLSM23, WG10b]. ASIS&T [RW10]. aspect [KZH^{+}24]. aspects [Mut16]. assembly [CW18, LHC14, WH15]. assertion [YSS23]. assess [AD21, BHA19, DG21, ETGS21, KTA18, KT18, KT19]. Assessing [ACD11a, AD14, Mar16, PYWH15, RWL12, WKCC23, XlYcG18, ZZ21a, ZMZ24, ADR13, ADG16, CtRlY21, Meh19, WLLL18]. assessment [ACD11b, ACD12c, AD17b, ADD19b, AAB^{+}21, BCGM17a, BCGM17b, BRN18, CCC^{+}18, DLM^{+}19, FC11, FM17a, FM17b, Gag11, Gag13, Hag15, HB16, HM21b, Kos12a, KCH24, RNB18a, Vin13a, WLH19, XZS21, ZL11, ZZ21c]. assessments [AW14, BD07b, Bor13b, BL13, CMCMM22, RNB24, Sch13b]. assignee [KHC11b]. Assignees [KLC21]. Assigning [ZWBA21]. assignment [FN10, MLDP21, PCCN21]. assignments [BT21]. assistive [SCMG16]. associate [ZHS^{+}22]. associated [MdMALIV11]. association [UG24, WJ23]. association-based [UG24]. Associations [BSA21, PC10]. Assortative [Pen15]. astronomical [HKA^{+}09]. Asymmetry [OMFF23]. attachment [AHL12, IPS21, Pen15, WFC14]. attack [Jal11]. Attention [PPG^{+}15, BD10a, CBT21, GBDG23, KW21, KK12, LR08, XWW^{+}24, Yur17b, KK12]. Attitudes [MJ13]. attractors [GP11]. attribute [Let16]. attributes [BD07b]. attribution [GKR^{+}24]. Audience [Zit10]. augmented [SRP15]. August [BW22, Ano22a, Ano23a, Ano24a]. Australia [Van12b]. Australian [But11, But17, Van11, Van12a, vdBHS17b]. Author [DGV18a, Ma12, MMOMLC18, NCJ15, YW15, ALW11, AOdFC12, BO12, BK08, BSA21, BNbH17, CH14, DCG^{+}19, DVG16, DGV19a, Eom08, FGW^{+}22, FSZB15, HM21a, HLB^{+}19, HLB^{+}21, HDY12, JSD14, JXYS21, KJS16, Kos12d, KK07, LF12, LHT18, LF23, LLH^{+}21, Mil13, Per17a, RRKH17, Reh21, RL18, Sch10a, SKK15, UK16, WJW^{+}19, WDLX23, Wil16, XZS21, YG14, YHWZ16, YWZ^{+}23, ZSS21, ZS08]. author- [YG14]. author-entity [JXYS21]. Author-level [MMOMLC18, BSA21, Wil16]. author-selected [UK16]. authored [LF12, PRRC15, Sch08, Sch10b]. AuthorRank [ABS21]. authors [ADR13, ACD14, AD15d, ADZ18, ADD24, Bor19, BO19, BM10, BM16, BADFT11, CMCMM21, CSdFCA17, DT13, DWW21, Hag14b, HCS19, Kos12d, LWWL14, LZR15, LLH^{+}21, RV24, RS10a, Ron18, San11a, San11b, San13b, San14, Smo21a, SDL16, TS16, VW16, JY21]. Authorship [ADD19a, AvdB24, AAH11, BAGADF15, DN17, IPS21, KMJ^{+}19, LT13, Ort14, PHMSML17, SJGCGB^{+}24, Wal12, ZKJ^{+}24, dMAGBLIM18]. authorships [AF17b]. autism [ZSWD17]. automate [SK21]. automated [PCCN21]. Automatic [MGLF09, ZXH^{+}23, BHH^{+}07, The18c, UG24]. Automating [GWMP16]. autopoiesis [LAL09]. autoregressive [Cop19]. Availability [GMFGVZ16]. Average [HRC12, ALB24, RNB18b]. Average-based [HRC12]. averaged [DGV19a, Wan24]. Averages [Egg12a, LG11, Ant18, BM11a, HSB16, vZ13]. avoidance [BM11a]. awakening [CTL22]. award [FT17, Van12a]. award-winning [FT17, Van12a]. awarded [BLV10]. awards [DVG16]. aware [Fia12, HLRD24, JS22, ZWZZ21]. axiom [Woe08b]. axiomatic [BM11b, BM14, BM16, HM21a, HCS19, Woe08a]. Axiomatics [Que11]. axiomatisation [Rou08]. axiomatization [Kon14]. axiomatizations [AK15]. Axiomatizing [Mir13].

back [WWML14]. backbone [TSRFAM19]. background [WOLD22]. backgrounds [TDCM24]. backward [HRC11a]. bafflegab [AH22]. balance [ADC19, LJB^{+}22]. balanced [MMZ16]. bare [BH16c]. Barycenter [VE14]. barycenters [RGRE15, RGRE16]. base [LBZ16]. based [AS17, ADG21, AA19, AW14, AF17a, AAB^{+}21, AvdB24, AKML18, ALX^{+}21, BZLX23, BSHK21, BRA19, BLWSE11, BLW13, BLW14, BSdMAM16, BW17a, BL17, BHA19, BTZY21, BHM21, BW22, BGT22, BTZY24, BH24, BM11b, BADFT11, BRN18, CAHVH10, CCC^{+}18, CLC11, CXZ^{+}22b, CFK21, CCL18, Egg10a, EGR18, FGW^{+}22, FLB22, FSZB15, FT17, FM10a, FM11b, FM14, GG19, GLW^{+}21, Glä07, Glä10, GFGALG21, GBDG23, HHAA16, HSB16, HSMB18, HLB^{+}19, HLB^{+}21, HB23b, HRC12, HY21, HWL22, HTZS23, HW14, HTLC21, HYS21, HLHC22, JWL17, JKHS21, JJ24, JSD14, JS14, JY21, JS22, KS16, KPB09, KJS16, KKK21, KP13, KKK17, KR21, LW17, LBZ16, LRC13, LWWL14, LR10, LHC13, LYYY09, LF12, LF21, LWCH22, Ma12, MLCL21, MLDP21, Mil13, MT10, NCJ15, Ort21b, Paj15, PWY21, PY18, PD21, PHANPP13]. based [PRCRVQ^{+}09, PSI^{+}17, PCCN21, QCVQdMA10, RGRE15, RGRE16, RY23, Ron13, RC14, SKLR16, SK21, SW14, SLK^{+}23, Ser10, SG12, SDL16, SWHC15, TS14, TWY24, UG24, VPAE24, VCG14, Vîi17, Wal17b, WYX^{+}12, WDLX23, WRGL23, WLY^{+}24, Woe14, XWZL21, XZS21, XlYcG21, YZ15, YHS18, lYFyS18, YWW^{+}23, YSPW09, YX24, ZLJ^{+}10, ZWG^{+}19, ZSZ24, ZZ12, ZLYC21, ZKJ^{+}24, vdBHS17b, Ron12]. baseline [CA11]. baselines [Pis22]. basic [LR07, LY14, ZRY11]. basics [YAL21]. basis [SW12, XMW^{+}24]. battle [CBT21]. Bayesian [BSdMAM16, Mal16, MBD17, MD18, PHANPP13]. be [ACD15, BLW13, BW17a, BHM21, BW22, BGT22, Cla16, CCSM21, Gag13, KP24, OOB10, The19b, Vin13b, Wan24]. beauties [HY21, LSZY14, LMF24, YBL^{+}22]. beauty [FLQ23]. Become [YSJ21, ES15]. becoming [MdMAGBLI21b, MdMAGBLI21a, Pet18]. before [SHL24]. begets [Tol23]. beginning [BI08]. behavior [ADD19b, ADG21, HHB^{+}23, LCL^{+}22, LLH^{+}21, MKCM07, RBV07, Sch14a, WG10b, vEW09]. behaviors [ADM13a]. behaviour [AvdB24, LY24, MMOMLC18, YL22]. behaviours [DTG12]. behind [HHB^{+}23]. Beijing [ZL11]. benchmarking [HHAL17]. Benedetto [FM17b]. benefiting [CXZ^{+}22b]. benefits [AD18, BFS11, Bor14a]. Benford [TV21]. BERT [HLRD24, XZDS21]. besides [BSMD12]. Besselaar [But17, Glä17, Lin18c, Mar17]. Best [The16d, BL12, BM15, SB13]. better [Bor13a, BRN19, CMN08, MdMALIV11, WvEW13, YSJ21, ZMBP24, ZZ18]. between [ACD12a, AD15d, AAD21b, ADD21b, AL08, ALB24, AW14, AZS^{+}14, Ant22, BSK10, BAGADF15, BMHD11, BADFT11, BSA21, Bur13b, CCH12, Che17, CML24, CZG16, DGDGSPSV14, DGV19b, Egg13a, Fin13, Fra10a, GPNA13, Glä17, GSV^{+}21, Hag15, HDD^{+}21, wH16, JZY^{+}18, KK21, LW09, LY19, LCL^{+}22, LJB^{+}22, MEG^{+}16, OA10, Ort15, PRWvE16, PSZZ12, PD24, RL18, RP08, RY12b, Rou16, RGRE17, SLL^{+}12, SYW18, SVTdFC11, TF15b, TSVBAM22, VPAE24, WWW10, WZ21a, WY21, YW24, YSJ21, ZH18, ZSS21, ZBK^{+}21, dMAGBLIM18]. between-field [DGDGSPSV14]. Betweenness [AHL12]. beyond [TWC^{+}23]. Bi [ZYZ18, ZWM^{+}21]. Bi-directional [ZYZ18]. Bi-layer [ZWM^{+}21]. Bias [DGV19a, BMD08a, CCC^{+}18, DGV19b, GMPRSFV23, Hag14a, Hag14b, LBZ^{+}22, MD12, Pet21, RL18, SB13, SY24, SPL24, TLV^{+}21, VMWM17, WYLW21]. biased [LQLJ14, LQLJ15, SL15]. Bibliographic [Jar07a, LWWZ22, NLK^{+}21, BM21, CHHL11, FZDW15, Fia12, Jar07b, JLYM16, LJJ21, Ma12, NF12, PSG^{+}21, PUKR21, SZR^{+}19, SZ10, Yun22, ZH18, ZY16]. Bibliographically [KCH21a]. Bibliometric [BM11b, CH14, Gag11, Per17a, Pet18, Ron18, WKCC23, Ada18, BMS^{+}09, BD10b, BO12, BM13b, BLM13, BL13, BTZY21, BTZY24, BM10, BM14, BADFT11, CMCMM21, CG11, CHD^{+}24, CB07, CT16, Fra10a, FC11, FM11a, FMM14, FJW^{+}15, Glä07, GWMP16, HB23a, HTLC21, JZY^{+}18, Kos13b, LP17, Lin18c, LHTW15, Mal10, MN15, MD18, MD19b, NLC17, Ort15, PYCH18, PD21, PXS24, PRWvE16, Pis22, RGMJC17, RGSM^{+}19, SMP12, Sch14a, TWH08, VZMFAB19, VCG14, Vîi16, WvEN10, WB16a, WB16b, YVW^{+}13, ZWG^{+}19, ZXS21]. Bibliometrics [GSV^{+}21, Ley17a, TSOMDV^{+}24, BW13, BM13a, BL14, BM15, BM18, BHM21, BGT22, CX16, Cla16, KRM14, Sch14c, SAP18, MMOMLC18]. bibliometrics-based [BGT22]. bibliometrix [AC17]. Biblioranking [ST19]. BibSonomy [ZDJ^{+}16]. bid [RBV07]. Big [GKR^{+}24, PL13]. bigger [BP11]. bilingual [LQLJ14]. binary [CM14]. Bing [TS12]. biographical [RS16]. biology [Gol14, HR18, McC10]. Biomedical [MCKM08, Ke18, LYZ^{+}17, LTC22, LTL24, STP17, SHL24, ZMBP24]. biomedicine [Ke21a]. bioRxiv [PYSM24]. biotechnology [HR18, RJDD08]. bioventure [GHvdB^{+}14]. bipartite [CB19, JXYS21]. Biplot [DFCGB15]. bird’s [YDCL13, Tan13]. bird’s-eye [YDCL13]. birth [LHDH21]. block [HKK18]. blockchain [YS21, YP21]. blog [WBL12]. blogosphere [CLC11]. Blogs [LAT07, Ort21a]. BMLR [BSdMAM16]. Board [Ano15d, Ano16a, Ano16b, Ano16d, Ano17a, Ano17b, Ano17c, Ano17d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15b, Ano15c, Ano16c, Ano18a, Ano18b, Ano18c, Ano18d, Ano19a, Ano19b, Ano19c, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano22b, Ano22c, Ano22d, Ano22e, Ano23b, Ano23c, Ano23d, Ano23e, Ano24b, Ano24c, Ano24d, Ano24e]. boards [BB11]. body [HCL13]. Book [Pet18, vL18, Chi16, Moe18, TSM09, TSRSRG^{+}13, VE14, WLH19, WHH21, WZHH23, ZZ21a, ZZ21c, TSRSRG^{+}13]. bookmarking [HS11, PB16]. books [KT18]. boost [CGXW22]. boosting [TDCM24]. bootstrap [CJW14]. bootstrapped [PYWH15]. border [HDC12]. Bornmann [Mut16, Nan16, Pet17, Pra13a, Rou19, Sch16, SM18]. bottom [LL15]. bottom-up [LL15]. bound [Hür15]. Boundary [HZW^{+}23, ZZ21b]. Boundary-spanning [HZW^{+}23]. BP [RZLC21]. BQ [MD18]. Brazil [BPI22]. Brazilian [DFFBR16, PSI^{+}17, PBIR24]. breadth [SHZ23, ZZ21a]. Breaking [ZZ21b]. breaks [HPNM22]. breakthrough [LMF24]. breakthroughs [MBZS23]. BRICS [WW15]. Bridge [MF21]. Bridging [BSW17, MF21]. broad [RCC18]. broadening [BM13a]. broader [Bor14a]. bubble [Pra13b, RGZSC13]. bundles [ER22, ER23, Egg24a]. Burgers [JSSK14]. burstiness [YSS23]. business [CH23, KP21, WYH17, Zha24]. Butler [Glä17, Mar17, vdBHS17b]. byline [Hag13, Hag14b].

C [GSV^{+}21]. calculate [BW13]. calculated [HDB22]. calculating [BLW13]. calculation [LB21, PHANPP13]. calendar [Mag13]. Calibration [Kos12a]. Can [BW17a, BW22, BGT22, KBCR10, KTA18, KT18, KT19, MMOMHLC17, OOB10, SK21, The17a, TBTB19]. cancer [JHK^{+}16]. cannot [Gag13]. Canonical [DFCGB15]. Capability [Cha22, The17a]. capacity [CMN08, Mêg14, MD18, VRK23, ZWDS21a, ZWDS21b]. capital [MBZS23, SCBB11, Tu21]. capture [GHvdB^{+}14]. Capturing [ZMBP24, KHC18]. carbon [MF21]. career [LHDH21, Lin18a, Lin18b, LRH^{+}21, MKM^{+}22, NLSM23, TDCM24, WLH11, YR08, vdBS15]. career-boosting [TDCM24]. careers [ALR^{+}23, CTGM15, KMS23a, vdBS15]. CAS9 [WY21]. cascades [FZZ23]. Case [TMK16, TWC^{+}23, WHW^{+}14, YY23, AD15b, AL08, AvdB24, BD09, BHA19, Bue21, Bur13a, CMN08, CySlY18, CTL22, CA11, Cso21, EI10, FZDW15, Fra09, Gol14, GAB11, GHvdB^{+}14, HRC12, HYS21, LBK^{+}21, LZR15, Lin18a, Lin18b, MLDP21, Mar16, MSL^{+}14, MÉVQR21, OLRF11, Ort15, OMFF23, PYSM24, PRRC15, PBIR24, QLS21, RC12a, RL22, Sch10a, Sch13c, Sch13d, Sch15, SSAGB11, SFF23, Vin12, WYX^{+}12, WJW^{+}19, WML^{+}21, WDLX23, WWXL24, XlYcG18, XLG21, YHS18, YWW^{+}23, YP21, vdBHS17b]. case-based [WYX^{+}12]. Cassidy [vL18]. Castellano [Sir12]. Catalog [TSM09]. categories [HSMB18, KP24, MMOMTLC18]. categorisation [XWZL21]. categorization [Gau17, HDB22, SCMG16, UG24]. category [SVTdFC11, ZLJ^{+}10]. causa [Kos12c]. causality [HDD^{+}21]. causation [ZSWD17]. causes [AS17]. Caveats [OL10, Sch13b]. CDS [Vin13b]. CDS-index [Vin13b]. cells [HYC15, LBK^{+}21, WHW^{+}14]. Central [Ant18]. centralities [XLSW24]. centrality [AHL12, BSW17, LR11, MMZ16, YZF13, ZYZ18]. Centre [DK22]. centric [LRH^{+}21]. centroids [KHC11a]. century [BI08, WLG09]. certain [Bor19]. certainty [YSS23]. Challenges [LCY19, CGM15]. Chambers [Woe14]. Chandos [Pet18]. change [BGCGS22, BGCGS23, Coc08, HLB^{+}19, HLB^{+}21, SFF23, YSJ21]. changes [ASS22, BSdA^{+}21, LF21]. Changing [Bue21]. channels [AW14, SKLR16]. chapters [TSRSRG^{+}13]. Characteristic [Egg10a, Glä07, Vîi17, Vîi18]. characteristics [ADD24, AH22, BGR21, BSdMAM14, CXZ^{+}22b, DHLW17, Glä07, GS10, HZZY24, KHC18, LY24, LW10, MN15, VZMFAB19, XLG21, ZH18]. characterization [CJF^{+}15, Egg13b]. Characterizations [Egg11]. Characterizing [BvECW18, HKXS23, KLC21, LT21, Sma18, SX21, ZLW21, RZ15, ZWM^{+}21]. characters [BvR11]. charging [DCM21]. Chemical [BMS^{+}09]. Chemie [BMS^{+}09, BD09, BD10a]. chemistry [BMS^{+}09, BSMD12, BH22]. Chief [Egg15, Wal15]. Chile [PVH21]. China [FH13, HYS21, Kos08, LWB14, LHTW15, LBP24, SZZ22, SY24, Tan13, Tan24, TYWG24, WG10a, YW15, YZ22, ZBK^{+}21]. China/USA [Kos08]. Chinese [LQLJ15, AYL14, CySlY18, GG09, GGSG10, GWL24, LQLJ14, Ma12, SL15, WWW10, XlYcG18, lYFyS18, ZWDS21a, ZWDS21b, ZZL^{+}21]. Chinese-language [LQLJ15, LQLJ14, SL15]. choice [ACD12b, SAP18, WvE15]. Choices [SZZ23]. circuits [HZL23]. circumvent [CZZ24]. citability [AD17a, AD17c, Web17]. CITAN [Gag11]. citance [WML^{+}21]. Citation [BMS^{+}09, BD10a, BH16a, Cam14, CH14, Cop19, DG21, EBG11, FH13, HZW^{+}23, LBZ^{+}22, MRBI17, San13a, TDKZ21, The16c, TSVBAM22, Tsa14, ACD11a, ACD12e, ACD12d, ADG21, AA19, AW14, ALW11, ASG12, AORC11a, AORC11b, AORC12a, AORC12b, AHRCV17, Als21, AOdFC12, AN18, Ant18, BILL07, BBL19, BD10b, BM11a, BP11, BSMD12, BW13, Bor13b, BLW13, BLW14, BM15, BL17, Bor19, BvR11, CMN08, CGSZS22, CM14, CPN24, CR10, CHHL11, CML24, Chi16, CY22, CRS14, CF16, CA11, CCL18, CPM^{+}07, CS22, Din11b, DLGC13, DWW21, Don18, DGDG13, DGDGSPSV14, DLH^{+}21, EI10, FT15a, FGW^{+}22, FSZB15, FM10a, GG19, GP12, Gau17, GNZ^{+}19, Glä07, GDDG24, HHAA16, HM21a, HSMB18, HKXS23, HRC12, HTZS23, HLZ24, HKK18, HRC11a, HCL13, HTX^{+}18, HDD^{+}21, HYC15, HCZ21, JKHS21]. citation [JJ24, JSD14, JZ21, Ke21b, KJS16, KKK21, KP13, KZH^{+}24, KTA18, KT18, LL14, LBB17, LBZ16, LWB19b, LSW^{+}10, LYF17, LTC22, LTL24, LH24, LR07, LHC13, LR08, LY14, LF21, LYC21, LCL^{+}22, LBP24, LIdMAM08, Lun07, Mal16, MD19a, Moe10a, Moe10b, MGS22, MD12, NLSM23, Nie17, NFH13, NJFD14, Ono16, Ort21a, Paj15, PYCH18, PSG^{+}21, PY18, PW19, PD21, Pen15, PSZZ12, Per10c, PS13, PPFKB19, RC12a, RWL12, RZLC21, RCW15, RCC18, SKLR16, San13b, Sch14a, Sch14b, Sch14c, Sch15, SD11, SYW18, SXAC22, SKY^{+}19, SSAGB11, Sir12, Sma09, STP17, Sma18, SLB13, SLW15, SZ10, TYWG24, TS14, TW14b, TF15b, The16e, The16d, The16a, TS16, The16f, The17a, The17b, TN18, The18c, Tok23, TSRSRG^{+}13, TLV^{+}21, UK16, VMWM17, Van13, Vîi16, WLG09, Wal14, Wal17a, WvE13]. citation [WvE15, Wal16a, Wal16b, Wal16c, Wal16d, WJW^{+}19, WZ21a, WXML21, WY21, WDLX23, WYC23, WLY^{+}24, Wan24, XMLM21, YW15, YL22, YWZ^{+}23, YYW14, YS21, YP21, YY23, Yun22, ZH18, ZWBA21, ZLW21, ZXS21, ZWDH24, ZS08, ZF22, ZZ12, ZSCSG23, dMAGBLIM18, dNL15, dlP11, vEW14, Fin13, SYW18]. citation-based [ADG21, AW14, HHAA16, LF21, Paj15, PD21, SKLR16, ZZ12]. citation-rank [BLW13, Sch14a, Sch14c]. Citations [CGD^{+}21, VG10a, ACD12b, ACD12c, AD15d, ADF19, ADD21a, ADD21b, ADD21c, AAK^{+}21, BZL19, BSS22, BAL18, BHM21, BM16, BvECW18, CMCMM21, CM14, Che17, CWT^{+}24, DFHCT21, FLQ23, FT17, Fin14, FC17, FMM14, Fra07, GAD19, GYZ17, HGJZ17, HWX24, HCL13, HCS19, KW21, Kos10a, KCH21b, KK13, LO10, LR11, LQLJ14, LQLJ15, LCY19, LWCH22, MMOMTLC18, MSPD16, NMNJM13, Ng17, PWY21, PD24, PML21, San11a, San14, San15, SL15, SC21, Smo16, The18b, The19b, Vin13b, WvEW13, Wan14, WS21, WFNH24, WNH^{+}23, WWXL24, ZYCK12, ZL11, OA13]. cited [ACD14, AD15c, ADD21b, BL12, BM13a, BH16a, BTZY21, BTZY24, BK14, BADFT11, Cam14, CAL15, CQAL21, Cho21, Cho23, CH14, HB23b, HW14, HR16, HTX^{+}18, HTLC21, LWB14, LYF17, MMOMHLC17, MGC18, MAS23, PD21, RNB24, Sch13c, Sma18, Sun24, TB18b, TW14a, TF15a, TMLB16, WYX^{+}12]. citedness [BL15]. CitedReferencesExplorer [TMLB16]. CiteScore [CS22, Meh19]. cities [BP11, CT16, Cso18, MEG^{+}16]. Citing [Zit10, ADD21b, BAL18, BADFT11, Cam14, CMCMM21, DTG12, GAD19, SD22, TB18a, WZ14]. Citing-side [Zit10]. CitNetExplorer [vEW14]. city [BMML21]. city-level [BMML21]. claims [FJOROMT22, RC12b]. clarifications [FGMM12]. clarifying [ZY21]. class [AS21, ADDWD24]. classes [BLM13]. Classification [BPU^{+}14, BSdA^{+}21, YYW14, CLC11, CXZ^{+}22b, DU21, FMM16a, GNBVQ^{+}14, Khr09, KZH^{+}24, LRC13, MLDP21, PRRC17, RC14, RCW15, WW16, XLSW24, ZLJ^{+}10, ZXH^{+}23]. classification-system-based [RC14]. classifications [LBZ16, SJZ^{+}19, SA18]. classifier [SD22, WYX^{+}12]. classify [OLRF11]. climate [HLB^{+}19, HLB^{+}21, SFF23]. clinical [Ke21b, LTC22, LTL24, MCKM08, SD22]. ClinicalTrials.gov [JHK^{+}16]. close [BP11]. closer [GHP21]. closure [Pen15]. cloud [WZCC24]. Clues [ZSP^{+}22]. cluster [CPN24, Wil16]. Clustering [CF16, VW16, YAL21, CB19, CGM15, CPN24, FLB22, JS22, LY18, NPS13, WvEN10, ZLJ^{+}10, ZLL^{+}18]. clustering-based [FLB22]. clusters [Sma09]. CNCIs [PSA21]. Co [KM22, LY18, MCLV23, AAH11, ADR13, BSW17, BAGADF15, CB19, CML24, DN17, IPS21, JSD14, KJS16, KMJ^{+}19, KSJL15, KK07, LW11, Ort14, PRRC15, PVH21, Sma09, VY10, WW15, XLSW24, Yun22, ZS08, ZKJ^{+}24, dMAGBLIM18]. co-author [KK07]. co-authored [PRRC15]. co-authors [ADR13]. co-authorship [AAH11, BAGADF15, DN17, IPS21, KMJ^{+}19, Ort14, ZKJ^{+}24, dMAGBLIM18]. co-citation [CML24, JSD14, KJS16, Sma09, Yun22, ZS08]. co-clustering [CB19]. co-evolution [WW15]. co-invention [PVH21]. Co-keywords [KM22]. Co-mention [LY18]. co-occurrences [VY10]. Co-patents [MCLV23]. co-publication [BSW17]. co-readership [KSJL15]. co-reference [XLSW24]. co-words [LW11]. Coauthor [ZLY21, Hag13, Per17a]. Coauthors [ZLY21, Xie21b]. coauthorship [Din11b, GFGALG21, Kar14, KK15, OMFF23, RP08, XOL16, XZKS22]. cocitation [BNbH17, Eom08]. code [Als21]. coefficient [LWB19b, YDW^{+}23]. coercive [YYW14]. cognitive [ADD21a, Jar07b, Mil15, RGRE17]. cohesion [TLZM24, ZMZ24]. cohorts [Tol13b]. coin [BRN19]. cold [XWZY21]. cold-start [XWZY21]. collaborate [ADS11, CGFÍ^{+}19]. collaborating [BSdMAM16, KdBOK15a, KdBOK15b, RV24]. Collaboration [WHW^{+}14, WDLX23, AHL12, Abb16, ADM13a, ADM13b, ADM17, ÁBB21, BMML21, BZLX23, BKK09, BRA19, BADFT11, CZZ24, CY23, Din11b, DKK^{+}17, ES15, FGW^{+}22, FC10, GPZC24, GYZ17, GW17, HDT^{+}13, HRL12, HPZ21, HNT22, JCWX24, KD15, KP21, KMJ^{+}19, KCH24, KR21, LW08, LY24, LY12, LRG13, LZH^{+}22, LH22, LJB^{+}22, LGX^{+}23, LRH^{+}21, MAS14, MSY21, OA13, PHMSML17, Per10a, PRCRVQ^{+}09, PVH21, PSA21, QZY^{+}24, SJGBMA17, Sch13a, SXAC22, SCJX22, SY24, SFCK19, Tan13, Tan24, Tu24, Vas09, VAdFC13, VCT22, WRA^{+}21, WFNH24, YZF13, ZY22, ZZL^{+}21, ZMZ24, ZKJ^{+}24, DT13]. collaboration-Collaboration [HRL12]. collaborations [AMKA21, CZG16, DK21, LWB14, MEG^{+}16, NFH13, ÖSY21, RSRP19, TS14, TSAH22, VPAE24, VRK23, YG14, ZZ18]. Collaborative [YAT^{+}22, AAB^{+}21, CRS15, HDC12, JZY^{+}18, LY12, LGS^{+}24, PTV17, Tu21]. collaborators [LH21]. Collective [LJM13]. college [CRS14, CRS15]. Combination [LLHW22, ADF19]. combinations [BTZY21, BTZY24, CY23]. combinative [KPB09]. combined [KCH21a, PT09, ZZ12]. Combining [BNbH17, CMN08, SZ10, YX24, PHMSML17, Ron18, YKKS16]. commands [BO12]. Comment [Glä17, KE17, Mut16, Web17, BW17b, BHM19, Cla16, Lin18c, Nan16, Pet17, Zit16, vdBS18, YAY^{+}17]. commentary [CBT21]. Comments [BCGM17a, BCGM17b, GL11, Mar17, Rou14a, Sch16, AD17c, Bor10, Bur08, FM17b, GTD16, WvE08, WvE09]. commercial [MCKM08, SZ10]. committee [vdB12]. Common [ADD22, JY21]. Communication [HMU13, KMS23b, TWC^{+}23, WCG21, YXX18, ZZ21b]. Communications [CWT^{+}24]. Communities [LP09, CW18, HRL12, RP08, TLM23, YP22, YDMS12]. Community [CR10, Din11a, KLI21, CB19, DKK^{+}17, GNBVQ^{+}14, Let16, LDS^{+}12, MMOMLC18, SG12, WWML14, ZZ21b]. Community-guided [KLI21]. comparable [BHA19, Hag15]. Comparative [Vin13a, ADD22, ASG12, BL13, CM14, EI10, Eom08, LY12, PYCH18, Vin14, WG10b]. compare [BDR14, BGT22, MMCR21]. compared [Egg12a]. Comparing [CHD^{+}24, Ke18, NO07, SD11, SJZ^{+}19, SSAGB11, WV24, ZS08, ZZ21b, ACD13b, BILL07, BLV10, Hag14a, HLRD24, LL15, MBIH16, Zha24]. Comparison [AAD21a, AF15, KW21, Kos08, LSW^{+}10, PSZZ12, PPFKB19, San13b, AD16a, ADZ18, AC09, BLW13, BL15, BH16c, CX16, DTG12, GSV^{+}21, GW17, HLB^{+}19, HLB^{+}21, KK15, LRC13, LF21, MMOMTLC18, MRBI17, PRWvE16, PRRC17, QCVQdMA10, RC12a, RC14, SLB13, The17b, Vîi17, WvE13, WY21, WZ14, XWZL21, YZ15, YXX^{+}17]. comparisons [DGDGSPSV14, FT15a]. competence [YZF13]. Competition [LWB14, SZZ23, ZWZZ21]. competition-aware [ZWZZ21]. competitive [ACD12a, ADS16, HG14, HPNM22]. Competitiveness [PNCG23]. competitor [The18a]. competitors [WJ23]. complement [TWY24]. complementarity [QZY^{+}24]. complementary [ZSZ24]. complementing [BM13a]. Complete [Jar07b, The16d]. Completing [Die15]. complex [Jal11, YDW^{+}23]. complexity [Don17, LBML23, LBD^{+}19]. component [Sma09, TMK16]. components [CY23]. composed [Ort21b]. composite [EGR18, MdGNS22]. composition [SXAC22, ZWL^{+}24]. comprehensive [AC17, BGCG^{+}24, DMCLH22, HHAA16, WLZJ24, ZZ21c]. Compressing [CH15]. computation [ACHVH09]. computational [CCH^{+}09, MMY17]. compute [AD15b]. Computer [SSAGB11, BFM^{+}18, FCZ24, HbHB^{+}23, LK17, Meh19, RV24, Tsa14]. computing [SZR^{+}19, VPAE24]. concentration [Chi16, GDDG24, HCC12, MdMALIV11, YLZ18]. concept [Egg24b, LY19, LYC21, SRP15, VY10]. concept-matchedness [LY19]. concepts [BGCG^{+}24]. Conceptual [LZL23, TB18a, Wal16a, RZ15, VZMFAB19]. conceptualization [Abr18]. conceptualizations [BS09, Sku09]. conceptualized [OOB10]. Conceptualizing [Gol14]. Concerns [But11, MCLV23]. conclusions [Egg08a]. ConfAssist [SCMG16]. conference [LK17, PCCN21, RBV07, RW10, SCMG16, WWXL24]. conference-journal [WWXL24]. conferences [HbHB^{+}23, Meh19]. Confidence [The17a, TF17a]. configure [MBZS23]. conflict [PCCN21, SCMG16]. confronted [Vîi16]. Congress [SDAJ17]. congruent [RGRE15, RGRE16]. Conjugate [Egg10b]. Connecting [YW17]. connection [FM14]. connections [BFM^{+}18]. Conscientiousness [Lin23, Lin24]. Consensus [NF12]. consequences [Egg08a, KG07]. conservation [BMML21]. consider [BP11, MD19b]. consideration [HTLC21, LY14]. considerations [AS17, WvEvL^{+}11]. considering [FLQ23, WFC14, XZS21, YF24, YP21, YY23]. consistency [Pra13c, XOMW21]. Consistent [BM10, XMLM21]. constituting [BM12]. constraint [HSB16]. constraints [Cho21]. constructed [RCW15, SA18]. Constructing [PRWvE16]. Construction [LBZ16]. contemporaneous [KHC18]. contemporary [GÁCC17]. Content [JSD14, KJS16, AAB^{+}21, BL15, KPB09, KZH^{+}24, MLCL21, PYCH19, TLZM24, WZ21b, YY23, ZLW21]. Content- [KJS16]. Content-based [JSD14]. Contents [XMM14, ZZ21a]. context [ALB24, But17, EI10, KPB09]. contexts [LW11, STP17, Sma18]. contextual [JJ24, Moe10b, Sch13a]. continuity [SXAC22, Yan14]. continuous [AL08, Egg10b, ER22, Sku09]. contribute [Cam15]. contributing [VTPF15]. contribution [CSdFCA17, Dan16, DVG16, HTZS23, LF12]. contributions [KHC13, RRKH17, SRZ19]. Contributory [Hag15]. Controversy [BPI22]. conventional [CZZ24]. convergence [DLGT19, KZDK23, LY24, PY18, YXXZ24]. Convergent [BD07a, BMS^{+}09, BT21, BHA19, HLHC22]. Convexity [SFCK19]. convolutional [TWC^{+}23]. cooperation [FZDW15, WWW10, ZSS21]. cooperative [IST10]. coordinates [ER21, Smo21b]. Coping [MÉVQR21]. copy [Bur08, EP07]. copying [PSG^{+}21]. Core [LBSA13, McC10, TB18a, YAT^{+}22, Bur07b, CAHVH10, Cam14, zChHY13, FCAD21, HYS21, Jar07a, KHC11a, LW08, MD12, OOB10, WWY21, Mêg18]. Core-periphery [LBSA13]. core/periphery [OOB10]. cores [Fas24, Jar07b, LRWY13]. coronavirus [LZH^{+}22]. coronavirus-related [LZH^{+}22]. corporate [CT16, ZYCK12]. corporation [CCH12]. corpus [PTA07]. correcting [TF15a]. Correction [vdBBL14, HCS19, RLG14]. correlated [CySlY18]. correlates [CWT^{+}24, KYhK23]. correlating [LSW^{+}10]. Correlation [Fin13, GPNA13, YSJ21, AAH11, AW14, BvR11, wH16, Rou16, WZ21a, ZSS21]. correlations [BMHD11, RSRP19, TF15b, TSVBAM22]. correspondence [BCGM17a]. corresponding [dMAGBLIM18]. Corrigendum [AORC12a, AORC12b, BGCGS23, BW22, BTZY24, DGRCDG24a, HLB^{+}21, Lin24, MdMAGBLI21b, RGRE16, Ron13, SV24, WZHH23, ZSF^{+}24]. cost [ALB24, MRR13]. Could [TN18]. Council [DK22, Van11]. Count [BD16, HZW^{+}23, AF15, BHM18, BHM19, CS22, HM21a, JZ21, LTC22, LHC13, MBD17, The16c, ZXS21]. counted [The19b]. counterintuitive [Sch14a]. Counting [Hag14a, Ng17, Sch10a, WvEW13, ASG12, BM16, Don24, Gau17, HCS19, KK19, LO10, LP17, LHC13, LF23, MD19b, PWY21, PRRC15, PRWvE16, SRZ19, WvE15, ZL11]. countries [AAD21b, Cha22, CG10, Hag15, HM21b, KE17, MO24, San13a, The18b, YAY^{+}16, YAY^{+}17, YZ22]. countries/territories [KE17, YAY^{+}16, YAY^{+}17]. Country [GNBVQ^{+}14, AD18, BGR21, BSdMAM14, CH23, LBSA13, LHDH21, OA13, PSA21, VCT22, WFNH24, YG14, YLCC22]. country-level [YG14]. country-scale [PSA21]. country-specific [BSdMAM14]. counts [AA19, BMS^{+}09, BSMD12, BH16c, DWW21, FSZB15, HB16, HKXS23, HDD^{+}21, RC12a, RZLC21, SYW18, TF15b, The17a, The17b, TN18, UK16, Wan24, ZH18, ZF22]. coupled [KCH21a]. coupling [BL21, CHHL11, Jar07a, Jar07b, LWWZ22, Ma12, NF12, Ort21a, PUKR21, SZR^{+}19, YHWZ16, Yun22]. covariates [BL17]. Cover [Ano14a, Ano14b, Ano14c, Ano14d, Ano15d, Ano15a, Ano15b, Ano15c, Ano16a, Ano16b, Ano16c, Ano16d, Ano17a, Ano17b, Ano17c, Ano17d, BSS22, KW21]. Coverage [LIdMAM08, RGMJC17, AN18, Egg14, Laf07, LA15, NO07, Rou14a, ZSWD17]. covers [GP12]. COVID [KYhK23, LZH^{+}22, RGZSC23, SLLM23, UKL23]. COVID-19 [KYhK23, LZH^{+}22, RGZSC23, SLLM23, UKL23]. create [The18b]. creation [CZZ24, LP09, MSL^{+}14, Tu21, Tu24]. creative [TB18b]. Creativity [TB18b, ALR^{+}23, HDC12, Sol07, YW24]. credibility [PHANPP13]. credit [DS21, Hag13, KK15, LF12, XZS21, XWZY21, ZWBA21, FN10]. credits [BO19]. CRExplorer [TMLB16]. crisis [HMU13, WBL12]. CRISPR [WY21]. CRISPR/CAS9 [WY21]. criteria [CCSM21]. criterion [BW17a, BW17b, BW22, Pet17]. Critical [BCGM17b, BM18, FM17a, FM17b, Sma09, CPM^{+}07, Wil16]. critique [And17]. Cross [KHC13, CM14, HDC12, HYC15, KCH24, LBSA13, MLCL21, SLNC13, SDL16, TLZM24, VCT22, ZL11]. cross- [ZL11]. cross-border [HDC12]. cross-country [LBSA13, VCT22]. cross-disciplinary [MLCL21, SLNC13]. cross-domain [CM14]. Cross-field [KHC13, SDL16]. cross-national [KCH24]. Crown [LO11, GL11, Lun07, Moe10a, Smo16, WvEvL^{+}11, vRvLV^{+}10]. crown-citation [Lun07]. CSIC [OLRF11, Ort15]. CSNCR [BH16a]. cubes [CM14]. cue [CSH18]. cues [CBT21]. cultural [HPZ21]. culture [KB21]. cumulative [FHH09, LY14]. Curie [TDCM24]. Current [PW19, GLW^{+}21]. curve [BSK10, BBL19, GP12, Sar08]. curves [BBGSZS24b, Egg10b, GDDG24, LBB17, SGDSP10]. CV [PSI^{+}17]. CWTS [FHH17, LO11, Moe10a, OL10]. cycle [LLGM21, LHWS18, Ort18].

DAC [JS22]. D’Angelo [Zit16, GTD16]. danger [Spa10]. Danish [BRA19, FJW^{+}15]. Data [DS21, XLSW24, dABJGMC21, BRA19, BMS^{+}09, BLWSE11, BdMA11, BO12, BLM13, BL13, Bor14c, BH15, BW17a, BL17, BHM18, BH18, BHA19, BHM19, BTZY21, BW22, BTZY24, CCL16, Cop19, DGV18b, ER22, FLB22, FHH17, GAW21, HMBI17, HB23b, HS11, JPL22, KD15, KK12, KR24, LL15, LBSA13, LR12, LB21, LSW^{+}10, MJZ^{+}24, MAS14, MBD17, MD19b, NV22, NLC17, PL13, QCVQdMA10, RS16, RGSM^{+}19, Sch16, SW12, SKY^{+}19, SKM22, TW14b, The16d, The16a, The16f, VY13, XLLZ14, XWK^{+}21, XMLM21, YKKS16, YDWC08, YSPW09, ZWM17, ZY16, dSSOH07, PW19]. data-based [FLB22]. database [ABS21, ED15, FMM16a, GP11, KP13, Ma12, Mal10, PSI^{+}17, WHL17, YDS^{+}15]. databases [BILL07, DS21, FZDW15, FMM14, SSAGB11, SZ10, VZMFAB16, YDWC08]. DataCite [RGMJC17]. dataset [HWZW22, KJKS22]. datasets [PC10, ZWBA21]. date [BLW14]. DEA [KE17, YAY^{+}17, ACD11b, CtRlY21, XlYcG18, XlYcG21, YAY^{+}16, lYFyS18]. DEA-based [XlYcG21]. dead [The16e]. debate [Mar17, PT08]. decades [KR24, PVH21, SCZ23, XLG21]. decay [ADD21a, PPG^{+}15]. decayed [KK13]. decision [Pet21, SLLM23]. decisions [BD07a, BD09, BD10a, BFM^{+}18, VCG14, vdBS15]. Decomposing [PL13, WYLL18]. decoupling [Tan24]. Deep [JFZZ23, JLH23, ZF22, AA19, CXZ^{+}22b, EAG22, FZZ23, WXZ24, ZLL^{+}18, ZSZ24]. defined [Egg21]. Defining [HY21]. Definition [HLR17, ER23, HRC11a, Sar08]. Definitions [LR08]. Degree [ZRY11, ADD21b, Egg24b, FZZ23, LY12, ZY12]. delay [BS13, GLY21]. Delayed [ZWDH24, DLH^{+}21, LL14, MSPD16]. delays [JZDZ24]. delimit [SZ10, The17a]. delineating [MÉVQR21]. demographics [WHL17]. demonstration [KMS23a]. density [BW11, GNZ^{+}19, SGGMF18]. departments [BC16, BGT22, CA11, TF17b, Tol08]. depend [Csa21]. dependence [MR13, VG10a]. Dependency [YDCL13, ZKJ^{+}24, CZZ24]. dependent [Ant18, HW14, Sch13d]. depth [ZZ21a]. derivation [Egg09, WvE09]. derive [Sch13e]. Deriving [AKML18]. Descendant [JS22]. Descendant-aware [JS22]. describe [KBCR10]. description [ER23]. descriptors [KHC11b]. detail [But11]. detect [BMD08a, FLQ23]. Detecting [EAG22, GW17, PTV17, YDS^{+}15, CLHHVH11, WHHW22, YYW14]. detection [BW11, CB19, CH15, Din11a, GNBVQ^{+}14, XHA^{+}21, YWW^{+}23, YKKS16, YX24]. determinants [ADM17, PBIR24, SHZ23]. determine [BBL19, BSMD12]. determined [KHC13]. develop [CGFÍ^{+}19]. developed [BLW13]. Developing [GAW21, KPS22, YF23, Wu13]. Development [BGCG^{+}24, WYX^{+}12, CMN08, CT16, IST10, PCPG17, Ser10, VRK23, WLY^{+}24, WG12, ZSF^{+}23, ZSF^{+}24]. developmental [CXZ^{+}22a, McC10]. developments [KHC18, LB21, dNL15]. devil [But11]. diagrams [FM10a]. differ [GAB11, SVKM21]. difference [BK14, Fra10a, YS21, ZY21]. Differences [HZW19, OA10, ADM13b, AAD21b, AN18, BGGB13, BMD07, BW13, FJW^{+}15, HSSM22, Luo21, MJB11, RL22, TBTB19, YZ18]. Different [BLV10, AD15a, ACHVH09, BMHD11, CT16, FGW^{+}22, Fin14, FN10, GM12, HDB22, LRC13, LH24, PSZZ12, San13b, San14, SXAC22, SSAGB11, TW14b, WvE13, YY23, ZLW21]. differential [GBDG23]. Differing [Chi16]. difficulties [Wal16a, Hic17]. Diffusion [TD10, ADD21c, CTL22, Gol14, HTZS23, HLZ24, KBCR10, LP09, LRG13, MSL^{+}14, NPS13, PYCH18, RLY12, SHZ23, WZ18, WBL12, WYH17, XLLZ14, YL22, YBL^{+}22, ZTG11, ZCM^{+}24, ZLYC21]. digit [Hür15]. digital [GMFGVZ16, MGLF09, YP22]. dimension [WZ18, YW17]. dimensional [AL17, BBGSZS24a]. dimensions [Ort21b, VCT22, The18a]. direct [Per10c, YW15]. directed [ZYZ18, ZY22, ZY12]. direction [YSJ21]. directional [DBS15, lYRyYbL14, ZYZ18]. directions [LEW22]. disadvantage [Ke21b]. disadvantages [Bor14a]. Disaggregated [Vîi17]. disagree [Mar17]. disagreements [Ley17b]. disambiguation [Mil13, Reh21, SKK15]. disaster [Mag12]. disciples [ALW11]. Disciplinary [WYH17, ADZ18, BDR14, Chi16, LY24, MLCL21, SLNC13, TS16]. discipline [ADS17, ADD21a, BH16c, Kos15, KTA18, LCL^{+}22, LBZ^{+}22, QLS21, SCBB11, Szi21, TF15b, WZ14]. discipline-level [ADD21a, LCL^{+}22]. disciplines [ÁBB21, AN18, FN10, HZW19, KBCR10, RL22, SGGMF18, San14, VPAE24, Wal14, WZ14, Wu15, ZLW21]. Disclosing [HHB^{+}23]. discometrics [Sch12]. disconnection [LEW22]. discontinuance [HLZ24]. discord [LEW22]. discourse [KM22]. discourses [Ant22]. Discoverers [SKY^{+}19]. discoveries [STP17]. Discovering [RY23, STP17, SHD15, XYP^{+}22]. Discovery [CFK21, BKK09, CCH^{+}09, KPB09, LDSD18, PY18]. Discrete [Sku09, Egg10b]. discretised [The16d, The16a, The16b]. discriminative [ZWDS21a, ZWDS21b]. discursive [Ley17b, LAL09]. discuss [HLB^{+}19, HLB^{+}21]. discussants [WB16b]. discussed [YXX^{+}17]. discussion [RYY10]. disjoined [Mir13]. disparity [Kos15]. dispersion [ACD12a, SHZ23]. dispute [Glä17]. disrupt [LZL23]. disrupting [Zha24]. Disruption [LB21, TWY24]. Disruptive [TLZM24, WLY^{+}24, YDW^{+}23, HLRD24, JZDZ24, LSSP24, WYC23, ZWL^{+}24]. disruptiveness [BT21]. dissatisfaction [GFGALG21]. dissemination [HWZW22]. dissertations [KT19]. distance [BADFT11, CH23, CY22, DBS15, JZY^{+}18, ÖSY21, RGRE17, VCT22]. distances [VPAE24]. distant [NFH13]. distinction [KMS21]. distinctive [ADD22]. distinguished [ACD15]. Distinguishing [SD22]. distributed [BHCH17]. Distribution [AL17, BBGSZS24b, Brz14, Cam14, Cam15, DLGC13, DS21, Egg09, Egg24b, FZZ23, HW14, KSS16, KK07, MKCM07, PSZZ12, SJGCGB^{+}24, San13b, SPT12, SYW18, VG10a, Vîi18, WvE09, YXX18]. Distributions [San14, TW14a, Wu15, ADS17, AORC11b, AORC12b, BL17, DD19, GS10, HCL13, Hür15, KP24, MD12, Per10b, RCC18, San13a, The16c, The16d, The16a, The16b, WLG09, Wu13]. Disturbance [YPL^{+}22]. diverse [CTGM15]. diversification [LWWZ22]. diversified [ADD19a]. Diversity [LR11, LWB19a, WLCL23, AJ13, ADZ18, CZZ24, HDT^{+}13, LWB19b, QM23, Rou19, Yan24a]. divided [LZH^{+}22]. DNA [YS21]. Do [Bor14a, BHA19, BTZY21, BTZY24, Che17, CGFÍ^{+}19, CPM^{+}07, FSZB15, GAB11, HNT22, LYZ^{+}17, LGS^{+}24, NFH13, The18b, Wal14, Wal17a, WXP^{+}12, WWXL24, Yur17b, ZLY21, vdBHS17a, GTD16, HWZW22, JCWX24, LLH^{+}21, MMCR21, Ort21a, RC12b, SVKM21]. doctoral [Lin23, Lin24]. Document [AC09, DT13, The18c, XMM14]. Document-document [AC09]. documented [Vîi18]. documents [Bor19, Bur08, EP07, Fra09, Jar07a, Kua24, MMOMHLC17, ZH18, ZLR16]. Does [AD17a, AD17c, ADD21a, BL15, Bor19, CP23, FGW^{+}22, GBDG23, HLB^{+}19, HLB^{+}21, HTZS23, HWX24, Ley13a, MJZ^{+}24, Tan13, TLZM24, Web17, XWW^{+}24, YLZ18, ZLL^{+}18, ZHS^{+}22, BK14, CGXW22, KMJ^{+}19, LBP24, WWML14, WOLD22, vdBS18]. dog [Bur13a]. DOIs [GMFGVZ16, HL19]. domain [BNbH17, CM14, CX16, CTDM17, CTL22, Sku09, SCZ23, VCT22, WZ21b, WXZ24, YS21, YP21, ZXH^{+}23, ZSF^{+}23, ZSF^{+}24]. domains [QCVQdMA10, RC12a, YY23]. domestic [KP13, MEG^{+}16, PSA21]. doors [Zit16]. Double [RNB18a, BRN18]. download [HDD^{+}21]. downloads [Kak16]. dressed [SB13]. driven [KPS22, LDSD18, YPYJ23]. driver [AHL12]. Drivers [DFFBR16]. drives [FHH17]. drug [JHK^{+}16, JXYS21, LT21, LDSD18, SX21]. drug-research [JHK^{+}16]. drugs [YSS23]. dual [Yan24b]. Dubious [KCH24]. duplex [ZY21]. duplicate [CGXW22, VZAMMFAB15]. duplicated [CK22]. durability [HLZ24, Ono16]. Durfee [Egg10b]. during [FH13, GGSG10, HMU13, LZH^{+}22, RGZSC23, SPL24]. dyadic [Tu21]. dye [WHW^{+}14]. dye-sensitized [WHW^{+}14]. dynamic [AF17a, zChHY13, FC17, LDS^{+}12, PTV17, XlYcG18, YDMS12, ZW21]. Dynamics [KM22, BH22, Fas24, GZRMLSC18, HGJZ17, HMU13, JSSK14, JCWX24, LB21, LJM13, LAL09, MEG^{+}16, RWL12, San15, Vas09, WLH11, Yan14, YXXZ24, dNL15].

e-Research [MS09]. each [The19b]. Earlier [ZJMX23]. Early [AAK^{+}21, LMF24, WRA^{+}21, vdBS15, ADF19, DLH^{+}21, KTA18, LT13, Lin18a, Lin18b, MN15]. Early-stage [WRA^{+}21, MN15]. easily [HL19]. East [NLSM23]. Easter [Sir12]. Eastern [Cso21]. Eck [RC14]. economic [DCM21, GGSG10, HPZ21, SW14]. Economics [TSM09, AH22, BW14, CySlY18, Fra09, OG13, Tol08, Tol23, WLLL18, WKCC23]. economies [DBS15, WG10b, WG12]. economist [ÖSY21]. economists [Bue21, LYZ^{+}17, SW12, Tol13b]. ecosystem [HSE22]. edge [LTL24]. edge-labeled [LTL24]. editing [MBZS23, WJW^{+}19, WDLX23]. Edition [BMS^{+}09, BD09, BD10a]. editor [Pet21, Egg15, Wal15]. Editor-in-Chief [Egg15, Wal15]. Editorial [Ano08d, Ano15d, Ano16a, Ano16b, Ano16d, Ano17a, Ano17b, Ano17c, Ano17d, BB11, BD09, BD10a, BFM^{+}18, SLLM23, ZSS21, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15b, Ano15c, Ano16c, Ano18a, Ano18b, Ano18c, Ano18d, Ano19a, Ano19b, Ano19c, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano22b, Ano22c, Ano22d, Ano22e, Ano23b, Ano23c, Ano23d, Ano23e, Ano24b, Ano24c, Ano24d, Ano24e]. editorship [WLLL18, XWZL21]. educated [Yur17a]. Education [BGR21, LTH18, ACD12a, LBSA13, MMCR21, SLL^{+}12, VL16, WOLD22]. educational [AZS^{+}14, TDCM24]. Effect [AS21, Don18, SYK22, SZZ23, WWW10, ADD21a, BMML21, BGCGS22, BGCGS23, BSdMAM14, Bue21, CTL22, FC10, GLY21, Hag14b, Hic17, HM21b, JXYS21, KMS23a, KD15, KYhK23, LLW21, Lia21, LH22, MSY21, MdMAGBLI21b, MdMAGBLI21a, PSA21, SB11, SB13, SZZ22, SCJX22, WFNH24, YL22, vV19, vdBM22, BW14, Tol13b, Wan14]. effective [WZCC24, YZYW21]. Effectiveness [DK22, ADR16, ZCTN23]. Effects [BD07b, OZZW24, AS17, AAH11, ADG21, AOdFC12, BSS22, BW13, BAL18, BFM^{+}18, CA11, Cop19, Fra09, HZZY24, HCZ21, Let16, Mag12, ZYCK12, vdBHS17b]. Efficiency [BH16b, BGdMAW23, DBS15, Mêg14, PMAT13, GAW21, GTD16, LS14, OG13, XlYcG18, MBD17]. efficient [QM23, RNB18b, YP22]. effort [ALB24, LA15, AL08]. Egghe [Bur08, Pet18, AK15, Que11, Rou14a, Rou14b, WvE09, Woe08a]. eggs [Sir12]. Ego [Ort14, LRH^{+}21, RZ15]. ego-centric [LRH^{+}21]. eHealth [WML^{+}21]. elements [TB18a]. elephant [WvEVW16]. elite [HZZY24, Yur17a]. Elo [LW17]. Elsevier [Pet18, Ley13b]. Elsevier/Chandos [Pet18]. elsewhere [BMS^{+}09]. embeddedness [JCWX24]. embedding [AS21, CY22, MA21, XZS21, YF24, ZLL^{+}18]. embedding-based [XZS21]. embeddings [JFZZ23, XZDS21]. emerge [LEW22]. emergence [GW17, HLRD24, JS22, Sma09]. emergencies [MJZ^{+}24]. Emerging [LJM13, XHA^{+}21, Ant22, CG11, EAG22, Gol14, JYG24, MO24, McC10, MN15, MÉVQR21, PL13, WLY^{+}24, WG12, ZWM^{+}21, DG21]. Empirical [AORC11a, AORC12a, BH22, Brz14, FMM16a, KMS23a, LY14, WLH11, And17, AS21, BGLS22, BW13, BLW13, BW17a, BL17, BH18, BHA19, BHM21, BW22, CG10, CFvL19, DGRCDG24a, DGRCDG24b, Eom08, FGW^{+}22, Fin14, FM11a, GG19, GFGALG21, HSMB18, HW14, KB21, KdBOK15a, LAT07, LG11, LH24, Pis22, RL18, SZZ22, SY24, Wal12, WvE13, WCG21, WLLL18, ZZL^{+}21]. endeavor [ZXS^{+}22]. Endogenous [SZZ23]. endorsement [Din11b]. Energetics [Pra13b]. energy [BMML21, HYS21, LBP24]. engagement [DMS18, MdGNS22]. engineering [CRS14, CRS15, LK17]. English [Cha22, LQLJ14]. enhance [BMD10]. enhances [ZJMX23]. Enhancing [TWZ^{+}24, Tu21, Tu24]. ensemble [CPN24]. enterprise [QZY^{+}24]. entire [PS13]. entities [ALX^{+}21, DGV18b, PYWH15, TWY24, WZ21b, YZ15, YF23]. entity [JXYS21, WRGL23, YF24, YX24]. Entropy [AL08, Hur24, LLGM21, QM23]. envelopment [GAW21, LL15]. environmental [Van13]. enzymes [HR18]. epidemic [KBCR10]. epistemic [HRL12]. epistemological [Glä17]. eponyms [VZMFAB19]. Equal [Kar14]. equality [Yur16]. equalizing [Hag14a, Hag14b]. equation [LYYY09]. equations [Egg24a]. Equity [TDCM24]. Equivalence [BBGSZS24a]. equivalent [CSH18, Kar14]. ERA [Van12a]. eras [NLK^{+}21]. Erratum [Lin18a]. Error [Jal11, TLD^{+}21]. errors [BD07a, BD09, FMM16a, FMM16b, MCD16]. errors/horrors [FMM16b, MCD16]. ESI [TYWG24]. ESI’s [HTX^{+}18]. essential [WYC23, ZYCK12]. estimate [DS21]. Estimating [CJW14, KSS16, WBL12, GG19, Mag12]. Estimation [SYK22, Bur08, Cho23, CJF^{+}15, EP07, IPS21]. ESTs [YX24]. ethics [OOB10]. EU [RNB18b]. Euclidean [And17]. Eugene [DKK^{+}17, Sma17]. Europe [APF09, Cso21, MCLV23]. European [AMKA21, BSW17, BHCH17, BBS21, DBS15, DK22, LTH18, LWB14, SLL^{+}12]. Europeans [RNB18b]. evaluate [DGV18b, GAW21, HS11, PRCRVQ^{+}09, WZ21b, YF23]. Evaluating [AD15a, BGR21, BRN19, Dan16, DVG16, EGR18, KS16, KP13, LK17, WLZJ24, XlYcG21, dDGFFI^{+}21]. Evaluation [HZL23, ADD19b, ADG21, BLV10, BH15, CAHVH10, CRS15, CCSM21, Cso21, DGV18a, FM11b, GNZ^{+}19, HMBI17, HPNM22, KK19, Kos11, KG07, KTR10, KHC13, LRCRC13, LH24, LYYY09, Mar16, MRR13, MGLF09, NJFD14, Per17a, RGRE15, RGRE16, RYY10, SV18, SV24, SC21, SPZ^{+}11, TW14b, The17c, TDCM24, Van11, Vîi16, Vîi17, Vin14, XMLM21, YZYW21, YSPW09, ZL11, ZZ21c]. evaluations [Bor13b, OL10, Sch14a, Sch14b]. evaluative [BM13b, BW13, BM13a, BHM21, WvEVW16]. events [Mag13, PTA07]. evergreens [ZWM17]. Everyone [vL18]. Everything [KK12]. Evidence [AD15d, Ant22, ALR^{+}23, BMML21, BSS22, BPI22, CBT21, CWT^{+}24, GWL24, KKK17, LBML23, LZH^{+}22, LBZ^{+}22, MMCvLLC18, SCZ23, SHL24, Tan13, TYWG24, TS16, XWW^{+}24, YZ22, YS21, Che17, DLH^{+}21, Glä10, LAT07, LHTW15, SD22, WKCC23, CP23, HNT22, JCWX24, SW14, WOLD22, YXX18]. evidence-based [Glä10]. Evidences [LQLJ14, NLK^{+}21]. Evolution [CFK21, AHL12, AMKA21, BDR14, CTDM17, CLHHVH11, EAG22, Egg07a, Fin14, FCZ24, Gol14, HKK18, HR18, JKHS21, JLYM16, JY21, Ke21a, KD15, QLS21, SDL16, TLM23, WHHW22, WW15]. Evolutionary [KLC21, Egg07a, LB21, McC10]. evolving [PSG^{+}21]. examination [HHAA16, JZY^{+}18, LCY19, LY19, RL18, YZ18]. examined [VY13]. Examining [AN18, JXYS21, PYCH18, XLG21, ZY17, ALW11]. Example [SD11, AL17, BP11, BW13, BH22, Per10b]. Examples [Egg08a, Sch14a]. Excellence [BSdMAM16, MSY21, BHCH17, BLWSE11, BHA19, HHAL17, KKK17, Lin18a, Lin18b, Lin18c, Szi21, ZJMX23, BdMAL12]. excellence-based [KKK17]. excellent [PXS24, SB11, Tol13a]. Except [KK12]. exchanges [LAL09]. excluding [ADD21b]. exemplified [BMD08a]. exercise [AD17b, BCGM17a, BCGM17b, DLM^{+}19, FC11, FM17a, FM17b, KK19]. exercises [ADD19b]. exhaustivity [ALW11]. exist [HL19]. Expanded [CH14, FH13]. Expected [Smo16, GW17]. Expedited [SLLM23]. experience [HJ15, YW24]. Experimental [AC09, Fin13, The16f]. Expert [SKLR16, AW14, HHAA16, LSW^{+}10, SK21, SD11, WXML21, WS21]. Expert-based [SKLR16, AW14, SK21]. expert-selected [WXML21]. expertise [RGRE15, RGRE16]. experts [BM15, BT21]. explain [TBTB19]. explains [Vîi18]. explanatory [CCH^{+}09, DGRCDG24a, DGRCDG24b]. Exploitation [CD23, GWL24]. Exploration [GWL24, BNbH17, CD23, CY22, RV24]. explorative [HLB^{+}19, HLB^{+}21, LHT18]. Exploratory [MBZS23, SMP12, CTDM17, HB23b, KTA18, TSM09, YDMS12, ZZZ^{+}23]. explore [DFCGB15, HR18]. Exploring [BMML21, CQAL21, CTL22, CT16, DHLW17, JKHS21, KJKS22, LSSP24, LWCH22, McC10, SKK15, THQL24, WXP^{+}12, WPZ^{+}13, WJW^{+}19, WLH19, WXZ24, XZKS22, YW24, YSS23, YLCC22, ZYCK12, ZY12, HZW19, ZY22, ÁBB21]. exponential [HLD^{+}23, Laf07, LA15]. exported [ZDJ^{+}16]. Exposing [RS10b]. exposure [CGXW22, LLW21, SB11]. expressed [RA15]. Extension [YAT^{+}22, GP12]. extensions [LJHH23]. Extent [BD07a, BD09, Mil15, TS16]. external [BD07b]. Extracting [FLB22, WWY21]. extraction [KJKS22, WRGL23, ZLL^{+}18]. extreme [Vîi16]. eye [YDCL13]. eyes [TSRFAM19].

F1000 [BL13, BH15]. F1000Prime [BTZY24, Bor14c, BTZY21, WS21]. faces [RGSM^{+}19]. facets [MKM^{+}22, XZS21]. facilitate [DHLW17]. factor [ACD12b, AD15d, ADF19, BW17a, BW17b, Bor19, BW22, BRN19, Brz14, Bur13b, CJW14, DGDG13, DGDGSPSV14, Egg09, Egg13a, GLY21, HY08, HHC14, JLAL23, MKCM07, MD12, PHANPP13, Pet17, RXL15, San13a, SPT12, SMP12, SW14, WvE08, WvE09, WWW10, Fin13, TSVBAM22, UKL23, Zit10]. Factors [BM11b, SLL^{+}12, TT22, Van13, BSMD12, BL15, Cam15, CH23, DT13, DGRCDG24a, DGRCDG24b, Fra10b, HKXS23, HPZ21, JZY^{+}18, MD12, Rou16, SSAGB11, TF15a, VTPF15, XWK^{+}21, ZXS21, Ant18, SYW18]. faculties [AvdB24]. faculty [WS21, ZY17, ZZN19]. fading [SPL24]. failure [Moe18, Wou18]. fair [Gag13, Sir12]. fairness [RC12a, Sir12]. fame [WHLT18]. familiar [SB11]. familiarity [JZDZ24]. Family [Kos13b, BH18, Hür15, LBK^{+}21, Sch10c]. Family-tree [Kos13b]. far [WWML14]. Farewell [Zit16, AD16b, AD16c, GTD16]. fastText [JLAL23]. feather [Tan13]. feature [AOdFC12]. Features [CFK21, ED15, ADD22, BGGB13, HY21, HTLC21, KPB09, SD22, Sun24, VG10a]. February [Ano23f, Ano24f]. fee [DCM21]. fee-charging [DCM21]. feeds [LAT07]. fellows [TDCM24, PXS24]. fellowship [BD07b]. Female [TT22, KdBOK15b]. females [The18b]. Ferrers [Egg10b, FM10a]. fiber [MF21]. Field [Ron12, Ron13, RCW15, WvE15, ACD11b, ADM13a, ADS17, AD21, ADD21b, BMML21, BMS^{+}09, BM11a, BM18, BHM21, CMN08, CH23, CTDM17, CtRlY21, CLHHVH11, CA11, DFCGB15, DGDGSPSV14, FZDW15, FM14, GAB11, GBDG23, HDB22, HRC12, HZL23, Jar07b, KSS16, KHC13, Kua24, LO10, LRCRC13, MBZS23, McC10, MCKM08, MD19b, ÖSY21, OL10, PD21, PPFKB19, RYY10, RCC18, SD11, SFF23, SDL16, TS16, TF17a, The17c, TdR21, Wan24, WNH^{+}23, Wu13, WLLL18, WOLD22, XZS21, YDMS12, KLC21, PPFKB19]. field-adjusted [KSS16]. field-categorization [HDB22]. field-independent [PPFKB19, Wu13]. field-normalization [BM11a]. Field-normalized [RCW15, WvE15, BM18, BHM21, HDB22, XZS21]. field-standardized [ACD11b]. Field-Weighted [PPFKB19]. fields [AD14, ASS22, ACHVH09, BAGADF15, BH22, Fin14, Gol14, HGJZ17, JKHS21, Let16, MF21, MÉVQR21, Per10b, RCC18, San11b, SYP^{+}16, SROdFC13, SZ10, SDL16, YDCL13]. fight [Glä17]. film [LBK^{+}21]. final [VCG14]. financial [ADDWD24]. Finding [CXMR07, PT08, Sir12, ZLR16, CSH18, LWWL14]. Findings [PVH21]. fine [HHAL17, KZH^{+}24]. fine-grained [HHAL17, KZH^{+}24]. Finnish [SK21]. firm [SLK^{+}23, WJ23]. firm-specific [SLK^{+}23, WJ23]. First [SB13, dABJGMC21, Eom08, FC11, Hür15, LT13, ZS08, EBG11]. first-author [ZS08]. First-Citation-Speed-Index [EBG11]. fitness [CLR^{+}17]. Five [Egg12b, AC09, AN18, BLW13, The18b]. flights [HNT22]. flipping [BRN19]. flock [Tan13]. flow [DHLW17, JYG24, LBZ^{+}22, MLCL21, PY18]. flows [ADC19]. fluidness [WH15]. focus [MF21]. focused [ACHVH09, BSdMAM14, BGdMAW23]. foothold [Cha22]. FOP [TWZ^{+}24]. force [CJF^{+}15]. Forecasting [JSSK14]. foreign [YZ22, Yur17a]. foreign-educated [Yur17a]. forest [Van12a]. forestry [PMAT13, Van08]. form [LRG13]. formal [RY12b]. formation [Abb16, HLD^{+}23, LW08, LJB^{+}22, NF12, SVKM21]. formed [YZ18]. formula [BBL15]. formulae [The17c]. formulas [PHANPP13]. Forward [JZ21, GBMA12, HRC11a, JKHS21]. foster [CP23, HNT22, TLZM24]. found [KP24, KK13]. Foundations [TSOMDV^{+}24]. Four [HTX^{+}18, BFM^{+}18, LG11, MF21, XZS21, YF23]. four-entities [YF23]. fractal [ADS17]. Fractional [LO10, ZL11, BM16, Don24, HCS19, LP17, MD19b, PWY21, PRRC15, PRWvE16, RC12a, SRZ19]. fractionalised [ASG12]. fractionally [Sch10a]. framework [AAB^{+}21, BHA19, BGT22, CPN24, KPB09, LRG13, RZ15, SCMG16, WZCC24, YXXZ24]. Framing [SL12]. Franca [YXX18]. France [CP23]. Franceschini [BCGM17a, MCD16, AD17b]. free [Moe18]. frequencies [CM14, KK07, Sch14b]. frequency [ER22, Khr09]. frequently [LWB14]. freshness [LJB^{+}22, YW24]. Front [Ano14a, Ano14b, Ano14c, Ano14d, Ano15d, Ano15a, Ano15b, Ano15c, Ano16a, Ano16b, Ano16c, Ano16d, Ano17a, Ano17b, Ano17c, Ano17d, BM12, Jar07a, WV24]. front-page [WV24]. frontier [MBD17, OG13, YWW^{+}23]. frontiers [KE17, MBD17, YAY^{+}16, YAY^{+}17]. FSS [AD16a]. fuel [HYC15]. fuels [ZSP^{+}22]. fulfilling [LY19]. Full [LP17, PYWH15, PYCH19, PRWvE16, WZ21b, Zha24]. full-text [PYWH15, PYCH19, WZ21b]. function [ACD11a, Laf07, San14]. Functional [Cho21, Cho23, Egg13a, TWZ^{+}24, ZWM17]. functionality [WG12]. functions [ER19b, IPS21, NV22, San13b, dSSOH07, dlP11]. Fund [GWL24]. fundamental [ST19]. fundamentals [CZG16]. funded [DK22, ETGS21, KP21, LY19, MBD17, Yan24b]. Funding [SV18, SV24, AS17, ÁBB21, AvdB24, CP23, DLGT19, GWL24, GHS21, HM21b, HSSM22, KKK17, LW09, Lia21, MLDP21, OZZW24, VS13, Wal17b, Wu15, WYLL18, YVW^{+}13, YLZ18, vdBHS17b]. Funnel [ADG15, ADG16]. Further [AK15, BM11a, FGMM12, LZH^{+}22, GBMA12, Mut16]. fusion [BBB^{+}22, JFZZ23]. future [GLW^{+}21, GHS21, RNB18b, YW17, ZW21, ZXH^{+}23, ZJMX23]. Fuzzy [CLHHVH11]. FWCI [PPFKB19].

G [KE17, YAY^{+}17]. G.-l [KE17, YAY^{+}17]. Game [BGS21, Mêg18, ZCM^{+}24]. gap [Hag15, JS14, KK21]. gaps [LZH^{+}22]. Garfield [Sma17]. Gatekeepers [BD07b]. Gazing [SGKM16]. gems [CXMR07]. Gender [ADM13b, AAD21b, BMD07, FJW^{+}15, HSE22, Kos15, KR21, LHDH21, MJB11, Nie17, TBTB19, TBM^{+}19, WYLW21, ACD15, AN18, AF17b, BMD08a, GÁCC17, GBDG23, HSSM22, LZH^{+}22, NLSM23, RA15, SJGCGB^{+}24, SXAC22, SCJX22, ZHS^{+}22, ZWL^{+}24, CCSM21]. Gender-based [KR21]. gender-neutral [ZHS^{+}22]. Gender-specific [HSE22]. gendered [KYhK23]. genders [HZW19, ZZL^{+}21]. gene [MBZS23, WJW^{+}19, WDLX23]. Genealogical [CFK21, RFMC17]. genealogy [FCZ24, KBP23, RDF^{+}18]. General [Egg07a, RZ15, Sar08, SGDSP10, ZWM^{+}21]. general-purpose [ZWM^{+}21]. generalised [BBL19, BBGSZS24b, GS10]. Generalization [Yun22]. generalizations [vEW09]. Generalized [WFC14, BMD08a, Cho21, Egg11, ER19b, ER21, Gag09, GM12, LBB17, LGS^{+}24, Smo21b, LH21]. Generalizing [Ng17, vEW08]. generally [KP13]. generated [HSMB18]. Generating [PTA07, LR12, SGDSP10]. Generation [JLYM16, AAB^{+}21, BM15, Kos10a, MO24, RY23, WRGL23]. generations [HRC11a]. generative [Pen15]. Geographic [LW10, ADD21a, HZW19]. Geographical [MAS14, AJ13, ADD21c, CT16]. Geography [PNCG23, BW11, BLWSE11]. Geometric [TF15a, AL17, LF23, PS13, TS16, The16f, XOL16]. geometrical [CMCMM22]. geopolitical [HPZ21]. geotemporal [WHL17]. German [AvdB24, Bue21]. Germany [BAL18]. ghost [Smo21a]. giant [Sma09]. giants [ADX^{+}17, ZZN19]. Gini [BBGSZS24b, LWB19b]. Gini-stable [BBGSZS24b]. GIS [TWH08]. glass [BH16b]. Global [ER23, TWH08, WFNH24, YXX^{+}17, CT16, GHP21, HHAL17, HRL12, LR12, LHTW15, PCPG17, SJGCGB^{+}24, SL12, WPZ^{+}13, ZZ21b, RW10]. Globalisation [WTvE11]. Globalised [DGV19a]. Globalization [HDC12, TSAH22]. glossary [BGCG^{+}24]. gold [ED15]. golden [NLK^{+}21]. good [VCG14]. Goodreads [WLH19]. Google [AN18, BMS^{+}09, CXMR07, HMBI17, KT19, MMOMHLC17, MMCvLLC18, MMOMTLC18, MBIH16, MT10, OA13]. Government [KP21, DFFBR16, LL15, Mêg14, Mêg18, ZSP^{+}22]. government-restructuring [DFFBR16]. government-subsidized [ZSP^{+}22]. GP [KPB09]. GP-adaptive [KPB09]. grade [KE17, YAY^{+}16, YAY^{+}17]. Graduates [FJW^{+}15]. graduating [HZW19]. grained [HHAL17, KZH^{+}24]. gram [Khr09]. Grand [LW17]. Granger [HDD^{+}21]. Granger-causality [HDD^{+}21]. Grant [BGS21, BMD07, BMD08b, CP23, DK21, HPNM22, MJB11, TDCM24, vdBS15, vdBM22]. grantee [KK21]. grants [DK22, LY19, PBIR24, TA22, YWW^{+}23, vdBS15]. Granularity [SA18]. Graph [QCVQdMA10, CB19, HLD^{+}23, MA21, TdR21, TWC^{+}23, XOL16, ZY16]. Graph-based [QCVQdMA10]. graphene [NLK^{+}21, YHS18]. Graphs [CFK21, Egg10b, GNZ^{+}19, Jar07b, RDF^{+}18, ZY16]. Great [SVKM21, WHLT18]. greater [CPM^{+}07]. greedy [PCCN21]. group [LW08, Moe10a, RL18, WYLL18]. groups [FM11b, PSZZ12, RGRE15, RGRE16]. growing [GHP21]. Growth [Per10a, RS10a, WG10b, CGD^{+}21, HGJZ17, LHTW15, LW10, PXS24, San11a, San11b, San15, SLB13, VTPF15, VL16, WLH11]. guarantor [SJGBMA17]. Guide [Pet18]. guided [KLI21]. Guns [Pet18].

H. [DKK^{+}17]. Habibzadeh [WvE08]. Haenszel [Smo19]. halo [Lia21]. Halt [Tan24]. hand [DCG^{+}19]. handbooks [MSL^{+}14]. Handelsblatt [Bue21]. handling [XWW^{+}24]. hapaxes [FCAD21]. happens [But17]. harder [GMPRSFV23]. Harmonic [Hag13]. hashtags [HLB^{+}19, HLB^{+}21]. Haunschild [SM18]. Hawkes [JWL17]. headings [The19a, SDAJ17]. health [FJW^{+}15, MJZ^{+}24, YWW^{+}23]. heartbeat [LSZY14]. Helix [Mêg14, Mêg18]. help [DT13, KT18, KT19, ZLL^{+}18]. Herrero [AHRCV17]. Heterogeneity [ADDWD24, HLR17, Hur24]. Heterogeneous [YF24, AS21, BZLX23, JLYM16, WWY21, XlYcG21, XMW^{+}24, KKK21]. heuristics [BGT22]. Hidden [BFM^{+}18, RS16, Tok23, Sir12]. Hierarchical [DKK^{+}17, QLS21, UG24]. hierarchy [Hag13, Hag14b]. High [AORC11a, AORC12a, HKXS23, AORC11b, AORC12b, GÁCC17, HZL23, TBM^{+}19, WFNH24]. High- [AORC11a, AORC12a]. high-impact [AORC11b, AORC12b, GÁCC17]. high-value [HZL23]. Higher [LTH18, ACD12a, ÁBB21, LBSA13, MMCR21, SLL^{+}12, The18b, VL16, Wal14, BGR21]. higher-impact [Wal14]. highest [DT13]. highly [ACD14, AD15c, BL12, CQAL21, Cho21, Cho23, HTX^{+}18, HTLC21, MMOMHLC17, PD21, Ron12, Ron13, Sch13c, Sma18, SZ10, TF15a, WYX^{+}12]. highly-cited [AD15c, Cho21, HTLC21, MMOMHLC17]. highly-cited-publications [Sch13c]. hiring [ZY17, ZZN19]. Hirsch [Bur07a, Bur07b, CAHVH10, Egg13b, Egg24a, FCAD21, FM10b, GS10, Kon14, Kos09, Kos10a, Kos10b, Mir13, Que09, Que11, Rou07, Sch10c, SMP12, Sch13d, Sch15, SG07, Vîi16, Vin13b]. Hirsch-index [Egg13b, Vin13b]. Hirsch-type [Egg24a, GS10, Kos09, Kos10a, Kos10b, SG07, Vîi16]. HistCite [BM12, Gar09]. historical [LB21]. histories [CF16]. history [Gar09]. HIV [RYY10, XLG21]. hoc [ZY21]. holes [HYS21]. home [WFNH24]. homophily [KR21]. hooked [The16d, The16a, The16b]. horrors [FMM16b, MCD16]. Horse [DCM21]. hot [BW11]. HTTP [KK13]. human [CKKY21, RS16]. Humanitie [WZHH23]. humanities [DG21, TSM09, TDCM24, VW16, YZYW21, TSRFAM19, VE14, WHH21]. hurt [Ley13a]. Hybrid [LB16, PRCRVQ^{+}09, ZSH^{+}16]. hypothesis [MJB11].

I3 [Ley12, RY12b]. iable [AAB^{+}21]. I’d [KP24]. idea [Kos11]. ideal [Bor10, BM11a]. idealist [Sma17]. ideas [GKR^{+}24]. identifiable [HL19]. Identification [HB23a, HTLC21, MMZ16, SA18, BSdMAM16, BGCG^{+}24, HZL23, Jar07b, LMF24, MN15]. identifiers [GMFGVZ16]. identify [CXZ^{+}22a, JYG24, LLGM21, MMOMHLC17, MA21, WZ21b]. Identifying [AAH11, CHHL11, FCZ24, HYS21, LBK^{+}21, Per10c, STP17, Tol13a, WJ23, YZ15, YBL^{+}22, ZWZZ21, Cho21, Cho23, FCAD21, Hic17, HY21, YDW^{+}23, ZSZ24, ZCTN23]. identities [YXXW19]. identity [HWX24]. IDR [WRB^{+}11]. IEEE [FM14]. IFC [Ano07d, Ano08d]. ignoring [KCH21b]. II [BD07a, BD09, KdBOK15b]. illegal [HL19]. illuminating [CHHL11]. image [JPL22]. imaginarium [Sch16]. imbalance [AS21, AF17b, NLSM23, WRGL23]. iMetrics [LBMM14]. Impact [BGGB13, BM11b, DGDG13, Egg14, Fin13, HPZ21, KTR10, PPFKB19, Rou14a, TSVBAM22, UKL23, ZWL^{+}24, dlP11, AJ13, ACD11a, ACD13a, AD15d, Abr18, ADF19, ADD21d, AAK^{+}21, AORC11a, AORC11b, AORC12a, AORC12b, AHRCV17, ÁBB21, Ant22, BZLX23, BBGSZS24a, BRA19, BP11, BW13, Bor13b, BLW13, Bor14a, BLW14, Bor14c, BM15, BH16a, BH16c, BW17a, BL17, BW17b, BHA19, Bor19, BW22, BRN19, Brz14, Bur13b, BvR11, CAHVH10, Cam15, CCL16, CM14, CZZ24, CG10, CJW14, CKZ^{+}23, CA11, CH15, CP23, CPM^{+}07, CZFD21, DG21, DN17, DT13, Don18, DGDGSPSV14, DGRCDG24a, DGRCDG24b, DS21, DVG16, DGV19b, Egg09, Egg13a, Egg21, ER22, ER23, EI10, FT15a, FT15b, FT17, FC10, Gag11, Gag13, GLW^{+}21, GTD16, GAB11, GÁCC17, GYZ17, GLY21, HY08]. impact [HB16, HHC14, HHB^{+}23, Hur24, JWL17, JZ21, JLH23, JZDZ24, KRM13, KFM^{+}15, Ke18, KP21, KTA18, KT18, KCH21b, LBK^{+}21, LBB17, Ley12, LY18, LCY19, LSSP24, LCL^{+}22, LJB^{+}22, LGS^{+}24, LIdMAM08, LBD^{+}19, LHC14, MAS14, MKCM07, Moe10a, Moe10b, MKM^{+}22, MD12, Nie17, NFH13, Ort14, Ort18, Ort21b, OZZW24, PYWH15, PYSM24, PHANPP13, PBIR24, Pet17, PS13, PD24, RWB17, RWL12, RXL15, Rou16, RCW15, San13a, SPT12, SW14, SD11, SB11, SXAC22, SSAGB11, SFF23, SLW15, TYWG24, TF15a, TS16, The18b, THQL24, Tsa14, UK16, VG10b, Vin12, Wal14, Wal17a, WvE08, WvE09, WvE13, WvEvLV13, WvE15, Wal16b, WWW10, WLH19, WZ21a, WYC23, WFNH24, WLZJ24, Woe08b, WLLL18, XZS21, XZKS22, XWW^{+}24, XOMW21, YZYW21, Yan24b, ZXS21, ZHS^{+}22, ZWDH24, ZZ12, ZZ21c, Zit10, dMAGBLIM18, Ant18, SYW18]. impactful [XLG21]. Impacts [ZZ21c, FGW^{+}22, KT19, LYF17, WKCC23]. Implementation [ABS21]. Implications [Sku09, DLGC13, DCG^{+}19, KHC18, PW19, Van12a]. importance [ADR13, Lin18a, Lin18b, Sch13a]. important [ACD12b, ES15, PYCH19]. importing [BO12]. improve [BLW14, MJZ^{+}24, Sch14b]. improved [ALX^{+}21, YHS18]. improvements [BL12]. improves [XLSW24]. Improving [CLR^{+}17, SRP15, WZ21a, BNbH17, PMAT13]. In-code [Als21]. in-group [RL18]. in-press [KTA18]. in-text [BvECW18, PWY21, WV24]. Inbreeding [BPI22]. InCites [BL13]. Including [BK14, ADD21b, BLW13]. inclusion [TMK16]. inconsistency [BN21, MAS23]. increase [RGZSC23, Vin13b]. increases [LBML23]. Increasing [HYC15, Vin13b]. increment [Rou14b]. incremental [SD22]. increments [Egg13b]. Independent [FH13, AD16b, AD16c, GG09, GGSG10, GTD16, Kos09, PPFKB19, Wal16c, Wu13, XlYcG21, Zit16]. Index [ACHVH09, BH24, CH14, FH13, NMNJM13, AK15, AOdFC12, And17, BBMP12, BBL15, BGS19, BD07a, BD10b, BMD10, BMHD11, BM12, Bor13a, BGT22, BADFT11, BN21, Bur07a, Bur07b, Bur13a, CMMT13, CB07, Don17, Egg08a, Egg13b, Egg14, FCAD21, FGMM12, FGMM13, GM12, GLW^{+}21, GP12, GS10, GR09, HKXS23, Kak16, KP13, Kon14, Kos09, Kos10b, Kos12b, KHC11b, KHC13, Lat21, LR08, LF12, LY14, LJHH23, Mal16, MR13, MRBI17, Mir13, MT10, MdGNS22, NT12, Ng17, NO10, Ono16, PSZZ12, Per17b, PS13, Pra13a, Pra13c, Que09, Que11, RFMC17, Rou07, Rou08, RYY10, RY12b, RY12a, Rou14a, Sch08, Sch10a, Sch10b, Sch10c, SMP12, Sch13c, Sch13e, Sch13d, Sch15, Tol08, Van08, Vin13b, Vin14, WTR^{+}22, Woe08a, WLH11, Wu13, WYLL18, WLLL18, YZF13, YWZ^{+}23, YR08, ZYZ18]. index [ZYCK12, BMS^{+}09, Bor14b, CAHVH10, DG21, EBG11, PW19, SYW18, SK14, TSRSRG^{+}13, WHW^{+}14]. Indexed [TDKZ21, MdMAGBLI21b, MdMAGBLI21a]. indexes [CGD^{+}21, SZ10]. India [TBM^{+}19]. indicate [Che17]. indications [BMD08a]. Indicator [LO11, ACD12a, AD15a, ADS16, BSW17, BD10b, BH16a, BH18, BGT22, BGdMAW23, BH24, GPGBMA10, GBMA12, HDB22, HHC14, JZ21, Ley12, Ley13b, Min14, Moe10a, SK21, Sch13c, Sch14a, Smo16, Sol07, TS16, The17c, TWY24, VG10b, Vin12, WvEvL^{+}11, WvEvLV13, XZS21, ZZ12, BdMAL12, TSVBAM22, KTR10]. Indicator-JCI [TSVBAM22]. Indicators [LR11, LB21, AD16b, AD16c, AAK^{+}21, AZS^{+}14, ASG12, APF09, BL13, BM18, BHM18, BHM19, BT21, BTZY21, BTZY24, CCC^{+}18, CMCMM21, CB07, Cso21, DLM^{+}17, EGR18, EI10, FT15b, FM11b, Gau17, Glä10, GTD16, HSB16, HRC12, HTX^{+}18, HZL23, HCZ21, JYG24, LW09, LH24, Lin18c, LF21, Mêg14, Mêg18, Ort15, Ort21b, PRCRVQ^{+}09, RC12a, RY12a, RCW15, RC16, SMP12, SW12, Sir12, SM18, SKM22, SY24, The16e, TF17a, Tok23, VCG14, Vîi16, WvE13, WvE15, Wal16b, Wal16c, WZ21a, Wil16, Wu13, Zit16]. indices [ABS21, BI10, BBGSZS24a, BM14, CGSZS22, Egg08a, Egg08b, Egg10b, Egg10a, Egg11, ER19a, ER19b, ER21, Gag09, GPNA13, HRC11b, KS16, Kos13b, LB21, SG07, Smo21b, Woe08b, Woe14, vEW08, vEW09]. indirect [FLQ23, JZ21]. Individual [ACD13b, RCC18, ADR13, BGT22, BPU^{+}14, CKZ^{+}23, Cso21, DFCGB15, DWW21, FM11a, Lin18a, Lin18b, Lin18c, Mar16, PRCRVQ^{+}09, RZLC21, San11a, San13b, San13a, San14, TW14a, TF15a, TDCM24]. individuals [HB23a]. industrial [Han07]. industry [FZDW15, Mêg14, Mêg18]. inefficiency [lYFyS18]. inequality [BBGSZS24a, DWW21, Hag15, PNCG23, TBM^{+}19, TDCM24, WYLL18, ZY17]. infection [RYY10]. infer [Nan16]. inference [BBMP12, HDD^{+}21, Sch16, SKM22, Wal16a, Wal16d]. Infinite [ER19a]. inflated [BHM18, BH18, BHM19, The16b]. Inflation [PPPF18, HGJZ17, HCZ21]. inflationary [Hag14a]. Influence [HLZ24, Ort14, SFF23, YS21, BGR21, Egg08b, Fra10b, GNZ^{+}19, GKR^{+}24, HR16, JCWX24, KBP23, LLH12, PCPG17, PML21, Rou07, SC21, TF15b, XZKS22, YF23, ZW21, ZZZ^{+}23, Wal14]. influencers [MdGNS22]. influences [LHC13, SYP^{+}16]. influential [HR16, HNT22, LWWL14, XWK^{+}21]. informal [YXX18, ZKJ^{+}24]. informatics [YWW^{+}23]. Information [Ada18, AYL14, GAB11, Sma17, ALB24, BB11, BI10, BO12, BL12, CGM15, CTDM17, CTL22, Egg07a, FZZ23, HMBI17, LP09, LSW^{+}10, LZR15, LBML23, LHT18, LBP24, MMY17, MRR13, OOB10, PYCH19, PS13, PB16, WWY21, WBL12, XWZL21, XMW^{+}24, YDMS12, YW15, YHWZ16, YWW^{+}23, YAL21, ZY21, ZS08, ZY17, ZZN19, XWZL21]. Informetr [Lin18a, BCGM17a]. informetric [AF15, BBGSZS24a, ER23, FGMM13, LA15, NV22, YDWC08]. Informetrics [BI08, BCGM17b, BGCGS23, BW22, BTZY24, DGRCDG24a, HLB^{+}21, Lin24, MdMAGBLI21b, SV24, Spa10, WZHH23, ZSF^{+}24, AL17, BD16, Bur08, EP07, Egg10b, Egg24a, LW10, Moe18, AORC12a, AORC12b, Egg12b, FM17b, KE17, RGRE16, Ron13, YAY^{+}17, Egg07b, LBMM14]. infrastructure [BFS11]. initials [Mil13]. initials-based [Mil13]. Initiative [MSY21]. innovation [AMKA21, BMML21, CG10, Don17, DLGT19, GG09, GGSG10, HCC12, HHC14, KHC11a, LSSP24, Mêg18, MCKM08, RY23, SL12, TT22, WYC23, XWY^{+}21, Yan24b, ZSP^{+}22]. innovations [TLD^{+}21]. innovator [Sma17]. input [BGdMAW23, LR10, SYP^{+}16]. input-output [SYP^{+}16]. insight [ZMBP24]. Insights [WW15, BdMA11]. institute [WYLL18]. institutes [KHC13, LL15, OLRF11, XlYcG18, ZZL^{+}21]. institution [ACD12d, YG14, YWZ^{+}23]. institution- [YG14]. Institutional [OA13, BW13, Bor13b, BGdMAW23, RYY10, YZ18]. Institutions [BdMAL12, LTH18, ADG15, AD15c, AYL14, BBS21, BSdMAM14, BSdMAM16, BGdMAW23, CJF^{+}15, HZW19, Per10b, SLL^{+}12, Szi21, VG10b, VL16, WLLL18, YLZ18, ZZ18]. instruments [HYS21]. integer [CtRlY21]. integrated [HZL23, KJKS22, Ley12, WZCC24, KTR10]. Integrating [Kua24, GWMP16, ZZ21c]. integration [WML^{+}21]. Integrative [XYP^{+}22]. intellectual [MBZS23, RRKH17, SCBB11, YHWZ16, YL22]. Intelligence [SD11, AKML18, HSE22, JPL22, QLS21, RV24, TLD^{+}21, VRK23, WOLD22]. intensities [GW17]. intensity [ACD12a]. inter [SCJX22]. inter-gender [SCJX22]. interaction [BK08, CML24, HbHB^{+}23, JSSK14, TLZM24, TWC^{+}23]. interactions [DMS18, OMFF23, WY21]. Interactive [LR12, HHB^{+}23]. interdependence [WH15]. Interdisciplinarity [LWB19b, CQAL21, HLZ24, KPS22, LR11, LWB19a, LT21, QM23, SROdFC13, SFF23, ZZ18]. interdisciplinary [ADZ18, CAL15, EI10, FLB22, Gol14, LHC14, McC10, SZ10, WRB^{+}11, WML^{+}21, YP21, ZWDH24, ZL11]. Interest [JY21, PCCN21, SAFR22]. interesting [BdMA11]. interests [AF17a, RGRE15, RGRE16]. interface [MCKM08, ZY16]. Interfield [Yur16]. interlinking [LBSA13]. internal [CY23, Fas24]. International [BMS^{+}09, BD09, BD10a, GPZC24, HZZY24, LW08, WFNH24, YZ22, ADS11, BAL18, Cha22, CZG16, DTG12, DK21, HPZ21, HNT22, JZY^{+}18, Kos10b, KCH24, MEG^{+}16, MAS14, MdMAGBLI21b, MdMAGBLI21a, PSA21, RSRP19, SY24, Tu24, VCT22, WWW10, ZZL^{+}21, dMAGBLIM18, WHW^{+}14]. internationality [MMCR21]. internationalization [ADD21b, VE14]. Internet [TMK16]. interplay [CZG16, DGV19b]. Interpolated [RXL15]. Interpretable [ZSCSG23, Sun24, TLM23]. Interpretation [SKM22, LGX^{+}23]. Interpreting [CMCMM22, PSA21]. Interrelations [SYP^{+}16]. intervals [TF17a, The17a]. Introducing [Cso21, ER22, MLCL21, MMY17, TMLB16, YHWZ16]. introduction [BH16a, BGdMAW23, HDD^{+}21]. intuitive [MRBI17]. invention [PVH21]. inventions [KZDK23]. Inventor [CKKY21, TT22, WLCL23, BK08]. inventor-author [BK08]. inventors [TT22, WG10a]. investigate [vZ13]. Investigating [CZG16, DMS18, KCH21b, LTL24, LHWS18, SZZ22, SVTdFC11, Wu13]. investigation [ADS11, ADS17, BD10a, BP11, CG11, CFvL19, Eom08, HTX^{+}18, KB21, LGX^{+}23, RLY12, YXXW19, ZBK^{+}21]. Investigations [Woe14]. investment [ZSP^{+}22]. investments [CZG16]. invisible [DCG^{+}19]. Ireland [Tol08]. IS-LS [XWZL21, SCZ23]. ISI [ZLJ^{+}10]. ISSI [RW10]. Issue [BGCGS23, BW22, RLG14]. issues [Cla16, WB16a, WB16b, ZCTN23]. Italian [BCGM17a, BCGM17b, FM17b, ADS11, AD17b, AAD21a, DLM^{+}19, FC11, FM17a, Mar15, Mar16]. Italy [ADM13a, AD15b, AD15d, AAD21b]. item [Laf07]. items [BK14, GAD19]. iterative [KCK21]. iudex [Kos12c].

J [HLB^{+}21, Lin18a, MdMAGBLI21b, SV24]. J. [AORC12a, AORC12b, BCGM17a, Ron13]. Japan [Mag12]. Japanese [IST10]. JASIST [LBMM14]. Jazz [Sch12]. JCI [TSVBAM22]. JCR [Cam15, PHANPP13, TSVBAM22]. JIF [BI10, Fra07]. Jin [Bur07b]. Joint [IPS21, Cho23]. Jointly [HM21a]. Journal [BCGM17b, BGCGS23, BTZY24, Csa21, DGDGSPSV14, DGRCDG24a, Fra10b, Fra07, GNBVQ^{+}14, HY08, Lin24, SZZ23, SLNC13, WZHH23, WLLL18, ZSF^{+}24, AD15d, ADF19, BW17a, BW17b, BW22, BRN19, But11, Cam14, Cam15, CB19, Cha22, CLW24, CJW14, Chi16, DTG12, DT13, DCG^{+}19, DGDG13, Fas24, FMM14, GPGBMA10, GBMA12, GLY21, HHAA16, HM21a, HS11, KP24, LS14, LR12, LBZ16, LR10, Lin18a, Lin18b, LF21, Mal16, MKCM07, McC10, MRBI17, MGS22, MD12, NLK^{+}21, OL10, Paj15, PD21, Pen15, Pet17, PMAT13, PML21, RL18, RXL15, Rou16, SK21, SW14, Ser10, SD11, SB11, SB13, SJZ^{+}19, SHL24, TF15a, The18c, TSVBAM22, Van11, Van12a, Van12b, Vin13a, Wal17a, WvE08, WvEvLV13, WWW10, WW16, WZ14, WWXL24, XWZL21, XZS21, XOMW21, YSPW09, ZG12, ZWDS21a, ZWDS21b]. journal [Zha24, Zit10, dNL15, Cam14, Egg12b, Fin13, FM17b, KE17, LW17, RGRE16, TSVBAM22, UKL23, YAY^{+}17, Egg07b, LBMM14, TDKZ21]. journal-journal [dNL15]. Journals [TDKZ21, Yur16, AH22, BBB^{+}22, BI10, BGS19, BGCGS22, BGCGS23, BS13, Bor19, BM10, BM11b, BFM^{+}18, Cam14, Cam15, CK22, CySlY18, CtRlY21, CS22, DFHCT21, DCM21, Dem18, ED15, FM10b, FM14, GÁCC17, HM21b, HW14, wH16, KMS21, KP13, KK12, LBB17, LR11, LWB19b, LQLJ14, LYYY09, LHWS18, LIdMAM08, MMCR21, Moe10b, MdMAGBLI21b, MdMAGBLI21a, MT10, OG13, PSI^{+}17, PMAT13, RS10a, San11b, San13a, SG07, SROdFC13, SLB13, SPZ^{+}11, SAP18, The16c, The17b, Tsa14, Van08, Vin14, Wal14, Wal17a, WWW10, WHL17, YYW14, ZG12, ZSS21]. judgements [BM15]. judgment [HHAA16]. judgments [LSW^{+}10]. junior [dABJGMC21, BW17a, BW17b, BW22, GWL24, Pet17]. just [The16e].

KCI [KP13]. kernel [MYZ^{+}22, ZLL^{+}18]. key [Kua24, Ron18, TWC^{+}23]. key-node [Kua24]. keyword [BSHK21, CHD^{+}24, HTLC21, LLH^{+}21, YHWZ16, YBL^{+}22]. keywords [CX16, HLB^{+}19, HLB^{+}21, KM22, LLH^{+}21, UK16]. kilometres [WTvE11]. king [GL11]. Know [vL18]. Knowledge [ADD21a, CFK21, KZDK23, LHDH21, XLLZ14, ADC19, ADD21c, BMML21, BL21, BNbH17, CCC^{+}18, CZZ24, Che17, CML24, CTL22, CY22, DHLW17, DLGT19, DLH^{+}21, FCZ24, GLW^{+}21, GYZ17, GHvdB^{+}14, Han07, HWL22, HTZS23, HR18, JYG24, JCWX24, Kua24, LP09, LJM13, LRG13, LAL09, LBZ^{+}22, MLCL21, MS09, MSL^{+}14, PPPF18, PY18, SCBB11, Sku09, SHD15, TWZ^{+}24, TWY24, TdR21, Tu21, Tu24, WZ18, WML^{+}21, WYC23, WYH17, YF24, YL22, YBL^{+}22, Yan24a, YLZ18, YSS23, YPYJ23, ZMBP24]. Kolmogorov [KMS21]. Korea [KP13]. Kosmulski [Egg11]. Kosmulski-indices [Egg11].

l [KE17, YAY^{+}17]. labeled [LTL24]. labeling [XLSW24]. laboratory [MCKM08]. landmark [CCL18]. landscape [TWC^{+}23]. Langmuir [San14]. Langmuir-type [San14]. language [LQLJ14, LQLJ15, MdMAGBLI21b, MdMAGBLI21a, SL15, SCML23, WZ21b, WXZ24, YXX18]. Large [WW16, BW17a, BL17, BW22, BK14, BvECW18, DGRCDG24a, DGRCDG24b, KJKS22, KR21, LYZ^{+}17, LCL^{+}22, LBP24, MMCvLLC18, PSI^{+}17, RA15, The16c, VMWM17, ZH18, ZLR16]. Large-scale [WW16, BW17a, BL17, BW22, BK14, BvECW18, DGRCDG24a, DGRCDG24b, KJKS22, KR21, LCL^{+}22, LBP24, MMCvLLC18, RA15, ZH18, ZLR16]. Larivière [vL18]. last [LT13]. Latent [BMD08b, BMD08a, TD10]. laureates [CD23]. law [AL17, CCC^{+}18, Per10b, The16d, The16a, The16b, TV21, WRGL23, YR08]. law-based [WRGL23]. Laws [BGLS22]. layer [ZWM^{+}21, ZY21]. layered [LRG13, RLY12]. layers [LWB14]. LDA [XZDS21]. lead [FGW^{+}22]. Leadership [SC21, HLD^{+}23, MAS14]. leading [Egg24b, WFNH24, ZCTN23]. Learning [LJJ21, AA19, CZZ24, CXZ^{+}22b, ETGS21, EAG22, HZL23, JFZZ23, JLH23, Khr09, KZH^{+}24, LMF24, MA21, PYWH15, Reh21, SLK^{+}23, Sun24, TLM23, TA22, VRK23, WXZ24, WKCC23, Xie21b, XHA^{+}21, ZLL^{+}18, ZSZ24, ZF22]. Least [AL08, ALB24]. leave [ADD19b]. LED [CHHL11]. Leiden [FHH17, OL10, PRRC15]. Leimkuhler [BSK10, GDDG24, Sar08, SGDSP10]. length [ACD11a]. Leo [Pet18, Rou14a]. less [RNB18b]. level [ACD11a, ADS11, ADM13a, ADR13, ADS17, ADD21a, AD21, ADD21b, BMML21, BSdMAM14, BPU^{+}14, BSA21, CY22, CB07, Csa21, DGV19a, Jar07b, KE17, LW09, LO10, LH24, Lin18a, Lin18b, Lin18c, LCL^{+}22, LBZ^{+}22, MMOMLC18, PRRC17, PPFKB19, SJZ^{+}19, SA18, TLV^{+}21, Wan24, Wil16, YG14, YZ18, YAY^{+}16, YAY^{+}17, ZZ21a, dNL15]. levels [LG11, YWZ^{+}23]. Leveraging [YP22]. Leydesdorff [Bor10, Rou19, vRvLV^{+}10]. libraries [Als21, MGLF09, YP22]. Library [XWZL21, BB11, BI10, CTL22, LSW^{+}10, LZR15, LBML23, PYCH19, PB16, ZY17, AYL14, GAB11, SDAJ17, TSM09]. life [ADR13, BSK10, HB23a, LLGM21, LHWS18, MO24, Ort18]. lifecycle [LL14]. Lifetime [SHZ23]. Lifting [Lun07]. like [AD16b, AD16c, CGD^{+}21, GTD16, KP24]. Limit [Ant18, Gag09]. limitations [CB07, RGMJC17]. limiting [vEW09]. limits [BLM13]. Lindahl [vdBS18]. line [LBZ16, TDCM24]. linear [AF15, Cho21, Per17a]. linearity [LCL^{+}22]. Lingua [YXX18]. Linguistic [AH22, LBD^{+}19, PD24]. link [Abb16, BSHK21, KLI21, OMFF23, PD24, RL18, TB18b, YDS^{+}15]. linkage [BL21, Che17, HYC15, Ke21a, SWHC15, ZMBP24]. linkages [DLGT19, DLH^{+}21]. linked [XWY^{+}21, XYP^{+}22]. links [ADDWD24, CHHL11, GW17, KHC18, LTL24, Per10c, SLL^{+}12]. List [Ano10e, Ano11e, Ano12e, Ano13e, Ano14e, Ano15e, ADZ18]. listed [Cam14]. lists [Kos12d]. Literature [CFK21, HMBI17, ADD21c, Chi16, CPM^{+}07, HKA^{+}09, KCK21, LYC21, LZH^{+}22, LW10, McC10, NPS13, NO07, PD21, PB16, TB18a, VZMFAB19, WRB^{+}11, Wal16b, XLLZ14, XMM14, XMW^{+}24, YAL21, ZTG11]. Literature-based [CFK21]. literatures [McC10]. live [WHLT18]. lme4 [LCY19]. local [JLAL23, RWL12, WXML21, YXX^{+}17]. located [HCL13]. location [JJ24]. location-based [JJ24]. locations [HCL13, YXXW19]. log [Per10b]. log-normal [Per10b]. logistic [BSdMAM16]. Lognormal [SYW18, AF15, The16d, The16a, The16b, Vîi18]. Long [RSRP19, WHLT18, ADF19, Glä07, LGX^{+}23, SAFR22, SLW15, WZ21a, ZKJ^{+}24]. Long-term [RSRP19, ADF19, Glä07, LGX^{+}23, SAFR22, SLW15, WZ21a, ZKJ^{+}24]. longer [TN18]. longitudinal [Abb16, LB16, Ort18, PD21, Pen15, PB16, SLB13, Tan13]. Look [WLCL23, Mag13, WWML14]. Lorenz [BBL19, BBGSZS24b, Egg10b, LBB17, SGDSP10]. lost [Bur08, EP07]. Lotka [KK07]. Lotkaian [Egg11]. Low [MRR13, AORC11a, AORC11b, AORC12a, AORC12b, KP13]. low- [AORC11b, AORC12b]. Low-cost [MRR13]. low-impact [AORC11a, AORC12a]. LS [XWZL21, SCZ23].

machine [ETGS21, HZL23, KMS21, Khr09, LMF24, MA21, Reh21, SLK^{+}23, Sun24, TLM23, TA22, VRK23, WKCC23, XHA^{+}21]. Macro [DLH^{+}21, LW09]. Macro-level [LW09]. made [VCG14]. magnitude [TF15b]. Main [YY23, CXZ^{+}22a, HLHC22, Kua24, XLLZ14, YF24, YDS^{+}15, YS21, YP21]. mainstream [LLH12]. Maisano [BCGM17a, AD17b]. major [PTA07, SGGMF18]. make [BK14, LH21]. makes [HPNM22, HbHB^{+}23]. malaria [RS10a]. male [KdBOK15b]. man [KR21]. man-woman [KR21]. management [CH23, CtRlY21, KSJL15, Nie17, SCBB11]. Mantel [Smo19]. Manuscript [Pet21, BGCGS22, BGCGS23, SHL24]. manuscripts [CSdFCA17, Sch08, Sch10b]. many [MKM^{+}22, RGSM^{+}19, SW12, The16b]. map [BSdMAM14, BK14]. Mapping [BLWSE11, CLR^{+}17, GHP21, KB14, RBV07, SDAJ17, SDL16, TSRSRG^{+}13, TSRFAM19, AC09, AC17, BL12, CMN08, LW11, MCKM08, PYCH18, WvEN10, YX24]. mappings [BNbH17]. maps [BW11, HKK18, LR12, LBZ16, NPS13, SRP15]. marginal [BW13]. Marie [TDCM24]. marketing [MT10]. Markov [BMD08b, DFHCT21]. Markov-switching [DFHCT21]. Marx [BHM18]. Master [LW17]. matched [LLW21]. matched-pair [LLW21]. matchedness [LY19]. matching [BHM21, FGW^{+}22, FJW^{+}15, LGS^{+}24, dSSOH07]. material [RJDD08]. materials [Mag12]. Mathematical [Egg09, Egg12a, Egg24a, RLG13, RLG14, BN21, Egg13b, San13b, KdBOK15a]. mathematicians [LYZ^{+}17, Lin18a, Lin18b]. mathematics [WCG21]. matrices [LR07]. matrix [LR10, ZKJ^{+}24]. matrix-based [ZKJ^{+}24]. MatrixSim [WHHW22]. Matter [ZLY21, BL15, Tan13, VCT22]. matters [GPZC24, SJGBMA17, vdBS18]. Matthew [BW14, KMS23a, Lia21, SG12, Tol13b, Wan14]. maturity [DGDG13]. maximum [DGDG13, AL08]. may [BBL19, Ano22f, Ano23g, Ano24g]. McCall [Ley12]. MDNRS [BH16c]. mean [BSK10, BBL19, BH16c, The16f, Vîi17]. mean-based [Vîi17]. meaning [RH13]. meaningful [BL14]. means [CPN24, Egg13b, HSB16, ZLL^{+}18]. measure [ADR16, BSW17, BDR14, BD10a, BTZY21, BTZY24, GTD16, HHC14, MD18, RGSM^{+}19, VY10, YZF13, ZYZ18, ZSH^{+}16, ZRY11, dlVV18]. measured [ADD21c]. Measurement [ZWDS21a, ZWDS21b, ACD11a, Abr18, AORC11b, AORC12b, ALX^{+}21, BMD10, HLR17, HCZ21, LWB19a, LLHW22, MBD17, OG13, Rou19, VPAE24, WLY^{+}24]. measurements [SHZ23]. measures [AAH11, AORC11a, AORC12a, BILL07, zChHY13, Egg21, GM12, LY12, LLH12, Moe10a, Per10b, RLG13, RLG14, SRP15, SLNC13, VY13, YDW^{+}23]. Measuring [BGCGS22, BK08, CKZ^{+}23, CY22, Coc08, DLGT19, GFGALG21, Han07, HZW19, JLAL23, LBB17, LS14, LL15, LR10, Mal16, MD19a, MSPD16, Moe10b, MdGNS22, RGRE17, Sch13a, SRZ19, SWHC15, WKY^{+}22, Yan14, lYFyS18, YWS23, ZXS21, ZY21, dSSOH07, ADZ18, BL21, Bor10, BM11a, Bor14c, BT21, GBMA12, HWL22, LCY19, LBZ^{+}22, QM23, WRB^{+}11, WZCC24, ZZZ^{+}23, ZZ12, BGCGS23, vL18]. mechanism [Fra07, HLD^{+}23, HHB^{+}23, San11b]. mechanisms [GHvdB^{+}14, Pen15, WKY^{+}22]. Media [ZSWD17, CGXW22, FLB22, HY21, LLW21, MdGNS22, Ort21a, VGRB16, WLZJ24]. media-based [HY21]. median [Vîi17]. median-based [Vîi17]. medians [Pis22]. mediating [ÁBB21, TYWG24]. medical [AvdB24, VS13, dDGFFI^{+}21]. medicine [MO24]. Megajournal [Pet21]. members [Ort15]. membership [vdB12]. memes [MLCL21]. Memetic [GKR^{+}24]. memory [PPPF18]. Mendeley [BH15, BH16c, DD19, FT15b, HB16, KT19, PB16, The17b, The18b, ZH18]. mention [LY18, Ort21a, WXZ24]. mentioned [HLSH17, LWCH22]. mentions [YLCC22]. mentorship [SZZ22, ZKJ^{+}24]. merging [Egg08b]. MeSH [ZMBP24]. meso [HDT^{+}13]. Message [Egg15, Wal15]. meta [AH22, BGGB13, BMD07, BLV10, BMHD11, WWY21]. meta-analysis [BMD07, BMHD11]. meta-analytical [BGGB13]. meta-evaluation [BLV10]. meta-path [WWY21]. Metaknowledge [YSS23, MMY17]. method [Bor10, BM11a, BM15, CCC^{+}18, CSH18, CXZ^{+}22b, CXZ^{+}22a, DS21, GL11, HWL22, KPS22, KCK21, LR12, LF23, MCKM08, SGDSP10, SZR^{+}19, Sma18, SPZ^{+}11, WvE15, WHHW22, Wan24, XlYcG21, YHS18, YYW14, ZSH^{+}16, ZLL^{+}18]. Methodi [GSV^{+}21]. Methodological [MCLV23, BGCG^{+}24, MJB11, Siv16]. methodology [AD15b, ADR16, AD21, MBIH16, ZWM^{+}21]. Methods [BM15, GHvdB^{+}14, ASG12, CX16, CHD^{+}24, FT15a, Gau17, HPNM22, KK19, Kos12a, KdBOK15a, KdBOK15b, LHC13, LHWS18, LLHW22, Mil13, MÉVQR21, PRRC15, SD11, SAP18, TW14b, TBTB19, YZ15]. Metric [Pet18, GPGBMA10, MRBI17, RFMC17, TSVBAM22, WLZJ24, AAB^{+}21]. Metric-based [AAB^{+}21]. Metric-Wise [Pet18]. Metrics [WKY^{+}22, BSA21, CRS14, DGV19b, GNZ^{+}19, HHAA16, KRM13, KG07, MMOMLC18, Ort18, PD21, PPFKB19, RDF^{+}18, RW10, TSVBAM22, TLV^{+}21, XMLM21, ZZZ^{+}23]. MHq [BHM18, BHM19]. micro [CB07, DLH^{+}21, HSMB18, KR24, WBL12]. micro-data [KR24]. microarrays [MCKM08]. Microsoft [KTA18, KT18, Ort14, The17b, The18c]. migrant [CH23]. migration [CH23]. milestone [BT21, MMZ16, WXML21, ZCM^{+}24]. Miller [Woe14]. minds [SVKM21, WHLT18]. minimal [Egg13b]. minimum [ER19b, HSB16]. Mining [CFK21, HDY12, HDT^{+}13, CKKY21, QCVQdMA10, WJ23, ZZ21a]. misinformation [Ada18]. mismanagement [Pet21]. Missing [KHC18, LHT18, CHHL11, Rou07]. misunderstood [But17]. misuse [Wan24]. Mitigating [TLV^{+}21]. mixed [HPNM22]. mixed-methods [HPNM22]. mixing [Pen15]. mixture [GDDG24]. MNCS [GTD16, LO11, AD16a, AD16b, AD16c]. MNRS [BH16c]. mobile [TWC^{+}23]. mobility [GHP21, GPZC24, HZZY24, HZW19, LH22, MKM^{+}22, RGSM^{+}19, YZ22, ZBK^{+}21]. modality [MdMAGBLI21b, MdMAGBLI21a]. mode [DN17, FGW^{+}22]. model [AOdFC12, BBGSZS24a, BHCH17, Bur07a, CtRlY21, Cop19, ER22, FGMM13, HLD^{+}23, HKK18, HTLC21, JSSK14, JS14, JY21, KdBOK15a, LWWL14, MJB11, MD18, OOB10, SG12, TF17b, VCG14, WZCC24, XOL16, Xie21b, Xie21a, XlYcG18, XYP^{+}22, YZ15, lYFyS18, YR08, YF23, ZSF^{+}23, ZSF^{+}24, ZLYC21]. model-based [JS14, YZ15]. modeled [NF12]. Modeling [GDDG24, SPT12, VGRB16, WLG09, BMD08a, BMD08b, Bor19, Brz14, CTGM15, LYYY09, MD19b, WZCC24, YKKS16, YX24]. Modelling [AF15, Mêg18, NV22, BP22, The16d, The16f, YDWC08]. models [AF15, BS09, BSdMAM14, BD16, CGSZS22, Cho21, GAW21, Glä10, KE17, KBCR10, LDS^{+}12, MLDP21, OG13, WBL12, XHA^{+}21, YAY^{+}16, YAY^{+}17, ZSCSG23]. moderating [ADD24]. moderation [Sil23]. Modesty [Kos12b]. Modesty-index [Kos12b]. modification [Sch08]. modifications [WvEvLV13]. modified [Don24, Sch10a, SRZ19]. modify [Sch10b]. Modifying [LF12]. MONO [DU21]. monodisciplinary [The16c]. monograph [TS14]. monograph-based [TS14]. Monolingual [XZDS21]. Monotonicity [Que09]. month [Don18]. most [ACD14, LWB14, MGC18, PD21, RNB24]. most-frequently [LWB14]. motif [BZLX23]. motif-based [BZLX23]. motivate [GMPRSFV23]. motivations [WXZ24]. mountain [Szi21]. move [CH23, LH21]. movement [RGSM^{+}19]. Mover [LH22]. Moving [CMCMM21, GHP21]. multi [BSdMAM14, Bur08, EP07, KE17, KZH^{+}24, Kos12d, KTA18, LJJ21, LWWL14, LF12, Moe18, RRKH17, RS10b, Sch08, Sch10a, Sch10b, YF24, YAY^{+}16, YAY^{+}17, ZZ21a]. multi-author [Kos12d, RRKH17, Sch10a]. multi-authored [LF12, Sch08, Sch10b]. multi-copy [Bur08, EP07]. multi-discipline [KTA18]. multi-entity [YF24]. multi-level [BSdMAM14, KE17, YAY^{+}16, YAY^{+}17, ZZ21a]. multi-paradigmatic [Moe18]. multi-parameter [LWWL14]. multi-relational [RS10b]. multi-resolution [LJJ21]. multi-task [KZH^{+}24]. Multidimensional [JYG24, HZL23, KR24]. multidisciplinary [BBS21, HDT^{+}13, LS14, The17b, The19a, ZZ18]. multilayer [KCK21, LTC22, WDLX23]. multilevel [BBS21, BMHD11, BSdMAM16, Bor19, LY24, MD19b]. multilingual [XZDS21]. multiple [ASS22, BNbH17, CPN24, CXZ^{+}22a, CH15, XHA^{+}21, ZZ21c]. multiplex [KLI21, ZY21]. Multiplicative [PRRC15, HRC12]. Multiplicity [ZSWD17]. multiply [HLSH17]. Multivariate [OLRF11, ADD21d, ZKJ^{+}24]. Munich [BP11]. museum [FMM16b, MCD16]. MVPs [KHC13]. my [Sch14b]. myopia [PPPF18]. mystery [Cla16]. myth [HRL12].

name [BP11, FN10, Mil13, SKK15, ZHS^{+}22]. names [BK08, LT13, LF12]. nanobiopharmaceuticals [CG11]. Nanoscience [MÉVQR21]. nanotechnology [MÉVQR21, WG10a, BSW17]. Narin [PML21]. Narrative [TSOMDV^{+}24]. National [DK22, FT15b, Mar15, Mar16, TS16, WLCL23, ACD11b, AD14, ASG12, ADDWD24, BAL18, CZG16, FT15a, HM21b, KK19, KCH24, LBSA13, SV18, SV24, dMAGBLIM18]. national-scale [ACD11b]. nationally [MdMAGBLI21b, MdMAGBLI21a]. nations [AD21, ADD22, AZS^{+}14, HG14, PCPG17, TF17b]. natural [BAL18, CA11, HB23a, HSSM22, Wal14, WZ21b, WXZ24]. natural-sciences [BAL18]. Nature [XWW^{+}24, ADS17, LJHH23, BSS22, CWT^{+}24]. navigation [YP22]. necessary [BP11]. need [RRKH17, TSVBAM22]. Needs [vL18, CLR^{+}17]. negative [BAL18, THQL24]. neighborhood [Cop19]. neighbourhood [SJGBMA17]. neither [GL11]. Nemo [Kos12c]. nested [WYLL18]. Netherlands [BAL18]. Network [BFM^{+}18, CW18, Din11b, HZW^{+}23, KYC12, Let16, AAH11, AA19, AS21, AF17b, BL21, BB11, BBB^{+}22, BSW17, CMN08, CR10, CLC11, CY22, CRS14, DKK^{+}17, FLB22, HMU13, HLR17, JJ24, JFZZ23, JCWX24, JS22, KBP23, LWWL14, LY18, LTC22, LY24, LRE17, LJHH23, MEG^{+}16, MMZ16, MMY17, Ort14, PPPF18, PY18, Per10a, Per17a, RP08, RS10b, RZ15, RZLC21, Sch12, SAB^{+}16, TMK16, TD10, TLV^{+}21, TWC^{+}23, VMWM17, WWY21, WLY^{+}24, XZS21, XMW^{+}24, YZF13, YF24, YHS18, lYFyS18, YPYJ23, Yun22, ZY22, ZYCK12, ZWM^{+}21, ZXS21, ZW21, ZCM^{+}24, ZF22, dlP11]. network-based [JS22, LWWL14, YHS18]. network-driven [YPYJ23]. networking [YZ18]. Networks [Egg24b, AAH11, AHL12, Abb16, AOdFC12, BMML21, BZLX23, BM21, BSHK21, BSW17, BKK09, BAGADF15, BSdMAM16, CY23, DN17, Din11b, DCG^{+}19, ES15, Egg07a, FZZ23, Fia12, GLW^{+}21, GYZ17, HLB^{+}19, HLB^{+}21, HDT^{+}13, HWL22, IST10, IPS21, Jal11, JLYM16, JXYS21, KLI21, KD15, KKK21, KMJ^{+}19, KdBOK15b, LJJ21, LP17, LWWZ22, LTL24, LJM13, MD19a, NJFD14, Ort14, Ort21a, OMFF23, PTV17, PSG^{+}21, PL13, Pen15, PRWvE16, PVH21, RS10b, RWL12, Sch13a, SVTdFC11, SFCK19, TV21, VZMFAB19, VAdFC13, WvEN10, WJW^{+}19, WXML21, WDLX23, XOL16, XLSW24, YZ18, YDW^{+}23, YWZ^{+}23, YDS^{+}15, YS21, YF23, ZWBA21, ZYZ18, ZWG^{+}19, ZY21, ZMZ24, ZRY11, ZY12, ZY17, vEW14]. networks-based [KKK21]. neural [AA19, CLC11, FZZ23, LTC22, RZLC21, YX24, ZYCK12, ZXS21]. neuroscience [GÁCC17]. neutral [ZHS^{+}22]. never [HW14]. never-cited [HW14]. newly [BLW13]. news [Ort21a, YLCC22]. nexus [YW24]. NIH [LY19, YVW^{+}13]. NIH-funded [LY19]. No [TS14, KP24]. Nobel [CD23, LHDH21, San15, Tol23]. node [Kua24, LJHH23, YZF13, YXXZ24, Yun22, ZYZ18]. nodes [YDW^{+}23]. Non [BvR11, LCL^{+}22, BL14, BK14, Cha22, Cho21, HWZW22, IPS21, KW21, RSRP19, Sma18, SWHC15, XMW^{+}24, Zha24]. non-adopters [Zha24]. Non-alphanumeric [BvR11]. non-cover [KW21]. non-English-speaking [Cha22]. non-linear [Cho21]. Non-linearity [LCL^{+}22]. non-meaningful [BL14]. non-method [Sma18]. non-parametric [IPS21]. non-patent [SWHC15, XMW^{+}24]. non-scholars [HWZW22]. non-science-based [SWHC15]. non-source [BK14]. non-stationary [RSRP19]. Nondeterministic [BC16]. nonlinear [Wan24, ZYCK12]. Nonparametric [Cho23]. nonstop [HNT22]. nor [GL11]. normal [MJ13, Per10b]. normalisation [DGV19b]. normalised [TS16, TF17a, The17a, The17c]. Normalization [BH16c, BH18, HB16, LO10, Ron13, BM11a, CA11, FM14, GL11, HRC12, LRC13, LRCRC13, MD19b, RC14, Wan24, Zit10, Ron12]. normalizations [OL10]. Normalized [NMNJM13, BLW13, BM15, BH16a, BH16c, BM18, Bor19, BHM21, DGDGSPSV14, HDB22, RCW15, WvE15, Wu13, XZS21]. normalizing [HTX^{+}18, WvE13]. Norway [AAD21b]. Norwegian [AAD21a, AW14, RA15]. Note [Per17b, Ano17e]. Novel [YDW^{+}23, AD21, BL21, CtRlY21, DU21, FM10b, QZY^{+}24, RGMJC17, SAP18, TWY24]. novelty [BTZY21, BTZY24, JLAL23, LLHW22, WZCC24, ZXS21, dDGFFI^{+}21]. November [Ano22g, Ano23h, Ano24h]. nuclear [HYS21]. nucleation [San11b]. nucleolus [Mêg18]. null [MJB11]. number [ADR13, AD15d, AD15c, Bor19, CAHVH10, EP07, LF23, Smo16, Vin13b, Xie21b]. numbers [BGdMAW23, HL19].

obesity [CLR^{+}17]. Object [YDWC08, GMFGVZ16]. Object-relational [YDWC08]. obscure [KCH24]. observation [ACD12e, ACD12d, Glä07, ZSF^{+}23, ZSF^{+}24]. observations [vdBHS17a]. occurrence [BvR11]. occurrences [VY10]. Odds [SM18, Smo19]. oeuvre [ALW11, Moe10a]. off [ADD24]. offset [WTR^{+}22]. offset-ability [WTR^{+}22]. often [SVKM21, WXP^{+}12]. old [Sch10c]. omitted [FMM14]. oncological [LIdMAM08]. one [CS22, Pet21, CWT^{+}24]. ones [ACD14]. oneself [Sir12]. ongoing [PT08, Tan24]. Online [CCL18, KTR10, CBT21, FJOROMT22, KSJL15, MMOMLC18, MS09, OA13, SAFR22]. only [Vin13b]. Ontario [GAW21]. Open [CWT^{+}24, SZZ23, Als21, CPM^{+}07, ED15, Fra09, HWX24, LB16, LLW21, MJZ^{+}24, MMCvLLC18, SLB13, SZ10, SL12, SPL24, Zha24, Zit16, vV19]. open-access [SZ10]. Openness [Tu24]. operationalised [Vas09]. operations [CtRlY21]. Opinion [YKKS16, Glä10]. opinions [WS21]. opportunistic [ADD19b]. Opportunities [BLM13, SLK^{+}23, Smo21a, WJ23]. opportunity [AZS^{+}14, PY18, RY23]. Opthof [Bor10, vRvLV^{+}10]. optimal [CGD^{+}21]. optimization [BHH^{+}07]. optimize [MLDP21]. Optimizing [GNBVQ^{+}14]. options [The16d, ZZ12]. oranges [ACD13b]. ORCID [GHP21]. order [ADR13, GPNA13, Kos12d, MKCM07, SB13, ZY22]. order-effect [SB13]. orders [HDY12, LRE17]. Ordinatio [GSV^{+}21]. organization [Dan16, DKK^{+}17, VY13]. organizational [BHH^{+}07, HMU13]. Organizations [WLCL23, KS16, VY10]. orientation [BAL18, Luo21]. orientations [TBTB19]. oriented [MdMAGBLI21b, MdMAGBLI21a]. Origin [VZMFAB19, OA10]. originality [HWL22]. originating [The19b]. other [CB07, Egg08a, HLB^{+}19, HLB^{+}21, Jar07a]. our [BL12]. outcome [JFZZ23, Sch14b]. outcomes [AZS^{+}14, CMCMM22, JZY^{+}18]. outgrow [HRC11b]. outlier [JLAL23]. outperform [FSZB15]. output [ADZ18, ADD19a, BGCL07, BMD10, CM14, Cso18, ETGS21, FM10a, FM11a, FM11b, GHS21, Hag14a, Hag15, Hic17, Kos11, Kos12a, LW09, Per10b, PSI^{+}17, RC16, SJGCGB^{+}24, San11a, SYP^{+}16, YLZ18, ZWL^{+}24, vdBHS17b]. output-based [vdBHS17b]. outputs [LYF17, OLRF11, XlYcG21]. overall [ACD13a, HSMB18, Lat21, RGZSC23]. Overcitation [MGC18]. Overcoming [JZDZ24]. overestimate [Wal14]. overlap [SVKM21]. overlays [LR12]. overlooked [KCH21b]. overrepresentation [MGC18]. overtime [WXP^{+}12]. Overton [DGRCDG24a, DGRCDG24b]. overview [Bor14a, FM10a, PTA07, TB18a]. Oxford [vL18].

P100 [BLW13, Sch14c, Sch14a]. pace [TLD^{+}21]. package [Gag11, LCY19]. packages [LY18]. page [KK13, WV24]. PageRank [CXMR07, Fia12, FSZB15, FT17, MD19a, NJFD14, NCJ15]. PageRank-based [FSZB15, FT17]. Pages [BW22, Pet18]. pair [JXYS21, KK07, LLW21, MRBI17]. pair-wise [MRBI17]. pairs [ASS22]. pancreatic [JHK^{+}16]. pandemic [KYhK23, LZH^{+}22, SLLM23]. pandemics [ZZ21b]. panel [AZS^{+}14, YSPW09]. panels [RGRE15, RGRE16]. Paper [DLH^{+}21, HZW^{+}23, BZL19, BCGM17b, Bor10, BH16c, DHLW17, DVG16, DGV19b, FGW^{+}22, FM17b, GM12, Glä10, GTD16, HKXS23, JLH23, KKK21, LPM16, LHC13, PCCN21, RRKH17, SJZ^{+}19, Wan24, YWZ^{+}23, YX24, ZF22]. Paper-patent [DLH^{+}21]. paper-side [BH16c]. PaperRank [ABS21]. Papers [WG12, BZLX23, BMS^{+}09, BM12, BL12, BH15, BAL18, BT21, CAHVH10, CAL15, CQAL21, Cho21, Cho23, DMS18, FLQ23, FC10, Fra09, GNZ^{+}19, GAB11, Han07, HSMB18, HB23b, HLZ24, HW14, HTX^{+}18, HTLC21, HYC15, KMS23b, KW21, Kos11, Kos12d, LL14, LTC22, LH24, LZR15, LF12, LRWY13, MMZ16, MSPD16, MGC18, MAS23, Ort21a, PYWH15, PD21, PYSM24, RZLC21, San13b, San14, Sch14b, Sma18, SCZ23, SPZ^{+}11, Sun24, TB18b, WYX^{+}12, WXML21, WY21, WLZJ24, XWW^{+}24, XLSW24, YLCC22, ZG12, ZW21, ZCM^{+}24, dDGFFI^{+}21]. para [BL15]. para-textual [BL15]. paradigm [LY24, Moe18, Wou18]. paradigmatic [Moe18]. parallel [XlYcG21]. parameter [LWWL14]. parameterisation [Sta21]. parametric [IPS21, SGDSP10]. Pareto [BBGSZS24b]. parsimonious [Hag13]. Part [KdBOK15a, KdBOK15b]. Partial [LRE17, Zha24]. participate [HWZW22]. Partition [Ron13, Ron12].

Partition-based [Ron13, Ron12]. partitions [Egg10b]. partner [Sch13a]. partners [QZY^{+}24]. past [GHS21, SCZ23, YW17]. Patent [CCL18, HCZ21, KLC21, TWC^{+}23, TSAH22, ALX^{+}21, CCH12, CHHL11, Che17, CY22, DK21, DLH^{+}21, FJOROMT22, HTZS23, HZL23, HDC12, JPL22, JFZZ23, KZDK23, KHC11b, KCH21b, LBK^{+}21, NLK^{+}21, SHZ23, SWHC15, TMK16, WV24, WLY^{+}24, WW15, XMW^{+}24, ZYCK12, ZSH^{+}16]. patenting [KS16, WG10a, WYLW21]. patents [AKML18, Che17, CXZ^{+}22b, CCL18, Han07, HZL23, HYC15, IST10, JKHS21, JCWX24, KCH21a, LRWY13, MCLV23, MF21, MCKM08, PVH21, SWHC15, TWZ^{+}24, WY21, WG10b, WG12, ZSP^{+}22, ZSZ24]. path [CZZ24, CGD^{+}21, CXZ^{+}22a, HR18, JZ21, KJKS22, Kua24, SHD15, WWY21, WLY^{+}24, XLLZ14, YF24, YS21, YP21, YY23]. paths [HLHC22, WHHW22, XWY^{+}21, YDS^{+}15]. pattern [CD23, KCH21a, Vîi18]. Patterns [CSdFCA17, TDKZ21, ADS17, APF09, ALR^{+}23, CQAL21, Chi16, DMCLH22, HSE22, HDY12, HKA^{+}09, HRL12, HKK18, JKHS21, KSJL15, LJJ21, LWB19b, LGX^{+}23, MSY21, MGS22, PYCH18, PHMSML17, PXS24, PUKR21, TSRSRG^{+}13, Van12a, VW16]. Paul [Moe18]. Paving [Zit16]. pay [Yur17b]. PCS [CCL18]. Peer [BGS21, Cop19, NO10, BS13, BD07a, BMD07, BMD08a, BMD08b, BL13, BGCG^{+}24, CWT^{+}24, DCM21, DCG^{+}19, HWX24, MJB11, SG12, SPL24, SHL24, Sun24, ZXS^{+}22, ZCTN23]. peer-review [DCM21]. peer-reviewed [BS13]. peers [VCG14]. people [BH15, Luo21, TBTB19]. perceived [RWB17, SB11]. Percentage [Vin14, HW14]. percentile [BLM13, BLW13, BRN18, PD21]. percentile-based [BLW13, BRN18]. percentiles [BLM13, BLW14, Pis22, Sch14b, The16f]. perception [GFGALG21, LWCH22]. perceptron [LTC22]. Performance [CZG16, ZLY21, AS17, AAH11, ACD12a, ADR13, ACD13b, ACD13a, AD15a, ADG15, ACD15, AD16a, ADG16, AAD21a, AAD21b, AvdB24, BBMP12, BAGADF15, Bor10, BM11a, BSdMAM14, BH16b, CCC^{+}18, CCH12, CG11, CRS14, Coc08, CA11, Cso21, Dan16, DFCGB15, DFFBR16, DGV19a, DGV19b, EGR18, FJW^{+}15, FHH17, GG09, GTD16, GHS21, KP21, Kos08, KHC11a, KHC11b, KR24, LS14, LL15, Let16, LY14, LHC14, MdMALIV11, MBD17, MD18, OL10, RL22, SV18, SV24, Sch14b, SCJX22, SC21, TT22, VY13, Wal17b, Wil16, WOLD22, XMLM21, YSJ21, ZYCK12, ZBK^{+}21, vdBS15, Fin13]. performance-based [AS17, AvdB24, Wal17b]. performers [ADS11]. performing [BL12]. periods [BSA21]. peripheral [Let16]. periphery [LBSA13, OOB10]. perish [LHWS18]. Perry [Ng17]. persistence [ZJMX23]. persistent [McC10]. personal [ADD24]. personalized [NCJ15]. Perspective [JYG24, BL21, BM13a, CY23, FC11, FCZ24, HDT^{+}13, HLZ24, HHB^{+}23, HbHB^{+}23, KMS21, KZH^{+}24, LS14, LT21, LTL24, LY24, LGX^{+}23, LRH^{+}21, MLCL21, SJGCGB^{+}24, SX21, SHZ23, Sun24, WJ23, YBL^{+}22, ZLW21, ZMZ24, Zha24]. perspectives [CMCMM21]. Perverse [vdBHS17b]. Peters [BW17b]. pharmaceutical [DLGT19, KS16]. phenomenon [HR16, HCZ21, SCBB11]. physical [CR10, CF16]. physicists [LYZ^{+}17].

Physics [CD23, KMS21, BT21, KM22, ST19, WHLT18, WWXL24]. piecewise [Xie21a]. pilot [HW14, PC10]. Pinski [PML21]. pitfalls [HTX^{+}18]. Plan [KK21]. platform [CBT21]. platforms [MMOMLC18]. plausible [The16a]. player [ES15]. plentiful [KK12]. PLOS [Pet21, CWT^{+}24]. plots [ADG15, ADG16]. plotting [BGCL07]. PlumX [Ort18]. Plus [YAT^{+}22]. PNAS [XWW^{+}24]. Poincaré [SYK22]. point [Bor14a, EGR18, EI10]. pointwise [Egg21]. Poisson [MD18, MD19b, Xie21a]. Poland [KKK17, KR24]. Polar [ER21, Smo21b]. polarity [YKKS16]. policies [HM21b, LBP24, VL16]. policy [DGRCDG24a, DGRCDG24b, Hic17, HYS21, LBP24, RNB18b, VS13, vdBHS17a]. Polish [DK22, Kos15]. political [FCAD21]. popularity [FZZ23, Fra10a, HTLC21, XWK^{+}21, YP22, Yan14, ZLYC21]. population [JSSK14, Sch16]. portfolio [KB17, SLK^{+}23, ZSH^{+}16]. portfolios [RGRE17]. Position [Vin14, BAGADF15, CT16, RA15]. Positioning [KHC11a]. Positions [KLC21, Cam14, KMJ^{+}19]. Positive [wH16, Rou16, Ley17b]. positively [CySlY18, HWX24]. possible [Vin13b]. Post [WS21, HM21b, LWCH22]. Post-publication [WS21]. post-retraction [LWCH22]. post-socialist [HM21b]. Posted [ZDJ^{+}16]. posts [YXXW19]. Potential [GLW^{+}21, JPL22, ACD12a, BMD08a, DGDGSPSV14, HB23a, HZL23, JLH23, LSSP24, MJZ^{+}24, TB18b]. Power [BGLS22, CCC^{+}18, CBT21, Mêg14, Sch13d, The16d, The16a, The16b, YR08]. pp [BCGM17b, RLG14]. practical [BW13, CLHHVH11, HHC14, The16e, The17c]. practice [NLSM23]. practices [Als21, FN10, KK21, PW19, WNH^{+}23]. pragmatic [LBZ16]. pragmatism [Siv16]. Praise [Ley17a]. Prathap [RYY10]. pre [HZL23, KD15, RW10]. pre-conference [RW10]. pre-processing [KD15]. pre-screening [HZL23]. precise [FT15a]. precision [The16f]. Predatory [Dem18]. predict [BD10a, Sun24, TN18, ZF22]. predictability [OMFF23]. Predicting [ADF19, AA19, AAK^{+}21, BZL19, dABJGMC21, BSHK21, CS22, HLRD24, LTC22, Lin18a, Lin18b, Lin18c, NLC17, RZLC21, SLW15, TLM23, Xie21b, Xie21a, YG14, ZW21, ZLYC21, DLH^{+}21, JFZZ23, TA22, VCG14, WYX^{+}12, YXXZ24, ZKJ^{+}24]. Prediction [CFK21, BSHK21, BLW14, FZZ23, FT17, JS22, KLI21, KKK21, KBP23, PY18, WDLX23]. predictions [BW13]. predictive [Sch13d, vdBS15]. predictor [GHS21]. predicts [Lin23, Lin24, SCML23]. predominance [VZMFAB19]. predominant [HCC12]. preference [Ser10, YP21, YLCC22]. preferences [ZZ21b]. preferential [AHL12, IPS21, Pen15, WFC14]. preferred [BLW13]. preliminary [BNbH17, LLH^{+}21, RLY12]. Preposition [AKML18]. Preposition-based [AKML18]. preprint [ASS22, PYSM24]. preprint-publication [ASS22]. Preprints [WCG21, XOMW21]. presence [SB13]. present [GMPRSFV23]. Press [vL18, BHM19, KTA18]. prestige [CKZ^{+}23, Fra10a, GPGBMA10, GBMA12]. PrestigeRank [SPZ^{+}11]. prestigious [AH22]. prevalence [Ort21b, ZZ18]. previous [BL12]. primary [Hag14b]. princes [FLQ23]. principle [ALB24, AL08]. prior [HTZS23, Kua24]. priorities [CLR^{+}17, RS16, VS13]. prioritize [NMNJM13]. private [ADD21d, BRA19]. prize [ZWZZ21, BH24, LHDH21, San15]. prize-winning [ZWZZ21, LHDH21]. prizes [BH24]. probabilistic [SPT12]. probe [zChHY13]. Probing [WY21]. problem [Bor13b, ER19b, WvEVW16, XWZY21]. Problems [CGM15, Min14, LO11, MD12]. procedure [UG24]. procedures [LRC13, LRCRC13, Mar16, RC14]. Proceeding [ZG12]. proceedings [GAB11]. process [DCM21, JWL17, LA15, SPL24, SHL24, TB18a, VL16, WYLW21, YBL^{+}22]. process-based [JWL17]. processes [Egg07a]. processing [BO12, KD15, WZ21b, WXZ24]. produce [DT13, SAB^{+}16]. producer [CGM15]. production [Egg07a, GÁCC17, KSS16, OG13, Rou16, TF17b, TWH08]. productive [ACD14]. Productivity [GTD16, ACD12e, ACD12d, AD15b, ADS17, ADM17, BBS21, GPZC24, HHB^{+}23, KYC12, KSS16, KYhK23, Lat21, Lin23, Lin24, LZH^{+}22, LH22, PBIR24, PS13, RCC14, RC16, TA22, TDCM24, WvEVW16, Xie21a, XZKS22, YXXW19]. professors [AAD21a, HZW19]. profile [CH14, CKKY21, MMOMLC18, PS13]. profiles [BDR14, BSdA^{+}21, OA13]. Profiling [RV24]. profit [Bor19]. program [DFFBR16, FHH09, TMLB16]. progress [BGR21, Hur24, Ley13a, Mag12]. Progressive [San11b]. project [SZZ22]. projects [CGFÍ^{+}19, LWWZ22, MLDP21, MCKM08, MBD17]. Proknow [GSV^{+}21]. Proknow-C [GSV^{+}21]. prominence [KB17]. promote [YLZ18]. promotes [ZJMX23]. promotion [CCSM21]. prompt [HLRD24]. prompt-aware [HLRD24]. Proot [Bur08]. Propensity [LH21, BHM21, FGW^{+}22, LGS^{+}24]. proper [vEW09]. Properties [Ono16, Bor19, DT13, Gag09, Pis22, PD24, RLG13, RLG14]. prophet [YX24]. Proposal [DLM^{+}17, HSB16, Ort21b, ACD13b, BM13a, BH24, EBG11, HY08, Rou19]. Proposals [BGS21, BLV10, HPNM22]. proposed [LY19, WvE08]. prospective [CKKY21]. Providing [Cam15]. Proximal [WZ18]. proximity [ADD21a, ADD21c, BP11, KJS16, SRP15, YF24]. proximity-based [KJS16]. proxy [CGSZS22]. pseudo [Cop19]. pseudo-spatial [Cop19]. pseudoscience [Spa10]. psychiatric [SX21]. psychology [BGGB13]. Public [BRA19, PC10, Coc08, DMS18, HLB^{+}19, HLB^{+}21, MJZ^{+}24, VL16, ZSP^{+}22]. Public-private [BRA19]. Publication [CH15, EI10, GLY21, KK19, LBMM14, RA15, SYK22, Van12a, AD15d, ADF19, ADD24, AW14, ALW11, ASS22, AvdB24, dABJGMC21, BSW17, BRA19, Bor13b, BLW14, Bue21, CX16, Cso18, CZFD21, Don18, FM10a, Gau17, Hag14a, HM21b, KK21, LR07, LR10, Lin18a, Lin18b, Mal16, Moe10a, MdMAGBLI21b, MdMAGBLI21a, NZZ^{+}16, PRRC17, Ron12, Ron13, RGRE17, SKLR16, SK21, San11a, Sch15, SA18, Smo21a, The16e, TMLB16, UG24, VW16, WS21, WW15, Xie21a, YVW^{+}13, Yur17b]. publication-citation [LR07]. publication-level [PRRC17, SA18]. publication-reference [LR10]. Publications [RGZSC23, AD17a, AD17c, ADD21b, BSMD12, BO19, BH22, BvR11, CK22, CTL22, DWW21, HDY12, HR16, HHC14, JLAL23, KHC13, LWB14, LY19, Lin18a, Lin18b, LY14, LHTW15, LGX^{+}23, LGS^{+}24, MMCvLLC18, MSY21, MCKM08, NLK^{+}21, PYCH19, PRRC15, RWB17, RNB24, Rou07, Sch10a, Sch13c, SA18, SLW15, SHL24, TB18a, TYWG24, TF17b, Vin13b, WvEW13, Web17, WG10b, XOMW21, YZ15, ZH18, ZG12, ZSS21, ZLYC21]. publicly [SHD15]. publish [CGFÍ^{+}19, KKK21, KP24]. published [BMS^{+}09, BdMA11, Bor19, Fra09, KCH21b, MMCR21, PSI^{+}17, San13a, Sta21, TSM09, XZS21]. Publisher [Ano17e]. publishers [FMM14, WWW10, WHH21, WZHH23]. publishes [Dem18]. Publishing [Pet18, AD15d, BS13, Bor19, Cha22, DTG12, Luo21, MJ13, San14, TBM^{+}19, TV21, VE14, Wal12, YPL^{+}22, ZXS^{+}22]. PubMed [ALR^{+}23, JHK^{+}16, SKK15, VTPF15]. PubMed-scale [SKK15]. purpose [ZWM^{+}21]. pushing [Zit16]. PWI [BH24].

Q1 [KP24]. Qualification [Mar15, Mar16]. Qualitative [Egg08a, CAHVH10]. Quality [CySlY18, KFM^{+}15, SZZ23, ADD24, AW14, AAB^{+}21, BSMD12, BL15, DMS18, FC10, GM12, HHAA16, wH16, LGS^{+}24, LK17, Meh19, Mir13, ÖSY21, PMAT13, RWL12, SB11, TF15b, XLLZ14, dSSOH07, vV19]. quantification [Hag13]. quantify [PRCRVQ^{+}09, RRKH17]. Quantifying [BZLX23, CZFD21, LY12, MLCL21, Mil15, NLSM23, RWB17, SAFR22, SROdFC13, VMWM17, YPYJ23, CCL16, CLHHVH11, WvEVW16]. quantile [BGCL07, GG19, NV22]. Quantitative [CAHVH10, LRCRC13, Mar15, BDR14, Egg08a, FZDW15, Kos12a, LSSP24, OG13, RGRE15, RGRE16, WNH^{+}23]. Quantity [Mir13, vdBS18, ADD24, CySlY18, wH16, KFM^{+}15]. quartile [KP24]. query [ZY16]. question [ADG16]. questionable [YPL^{+}22]. questions [LLHW22]. quicksand [Glä17]. quotient [MD18]. quotients [Smo19].

R [vL18, AC17, LYF17, LY18]. R&D [KP21, LL15, QZY^{+}24, RSRP19, XlYcG18]. R-tool [AC17]. Radicchi [Sir12]. Raf [Pet18]. random [HLD^{+}23, Hür15]. Rank [ER22, RNB24, Vin14, BBL19, BLM13, BLW13, BSdMAM14, BRN18, DFCGB15, LF12, MKCM07, PS13, RNB18a, Sch14a, Sch14c, Vin13a, YF23, GNBVQ^{+}14]. rank-citation [PS13]. Rank-frequency [ER22]. rank-order [MKCM07]. ranked [Tsa14]. Ranking [AD15c, ASG12, BM16, GZRMLSC18, KHC11b, LSW^{+}10, MT10, SW12, SAP18, Szi21, Van08, ZWG^{+}19, dlVV18, BC16, Csa21, DVG16, DGV18a, DGV19a, FCAD21, GP12, JJ24, KPB09, LW17, MAS14, NCJ15, SKLR16, Ser10, SD11, SB13, VMWM17, Van11, Van12a, Van12b, XMM14, XMLM21, ZWZZ21, ZY17, Bue21, FHH17, HCS19, PRRC15]. Rankings [BI10, BdMAL12, ACD12e, ACD12d, AD15a, ADG16, BM10, BM11b, BM14, Bue21, But11, CLW24, DGV18b, FSZB15, LHC13, Paj15, RXL15, SK21, SW14, Tsa14, BGR21]. ranks [AD16a, HZW19]. Rao [LWB19b]. rare [BK08]. Rasch [MD18]. rate [FC17, KP13, Lin18a, Lin18b, RA15]. rated [ED15]. rates [HSSM22, SLB13, SLNC13, Van13]. rating [LW17]. ratings [Wal17a]. Ratio [PPFKB19, ADS16, BM11a, NF08]. rational [GR09, Sol07, Tol08]. rationalism [BM18]. Ratios [LRWY13, Egg12a, LG11, SM18, Smo19, vZ13]. RCR [PPFKB19]. readability [Ant22, SHL24]. reader [BH16c, HB16]. reader- [BH16c]. readers [FT15b, The18b]. readership [DD19, KSJL15, ZH18]. Real [GR09, WLH11]. reality [RC12b]. realized [GW17]. really [Sch14c]. reason [SYW18]. reasons [BD07a]. rebels [MA21]. reboot [Tan24]. rebuttal [HbHB^{+}23]. recalibrated [Cso21]. receive [WWXL24]. received [Fin14]. receives [BD10a]. recency [LHDH21]. reciprocity [WRA^{+}21, ZKJ^{+}24]. recitation [ALW11]. Recognition [YAT^{+}22, BAL18, JZDZ24, Kos10b, LL14, MYZ^{+}22, MSPD16, ZXH^{+}23, ZJMX23]. recognized [DLH^{+}21]. recombination [WYC23]. recommend [YP22]. Recommendation [AF17a, XMW^{+}24]. recommendations [WS21]. Recommended [WS21]. Recommending [QZY^{+}24, SYK22, YG14]. reconsidered [BW14]. records [Ron12, Ron13, VZAMMFAB15]. Recovery [KK13]. recruitment [ADR16]. Reddit [SX21]. reduce [CCC^{+}18]. reduction [HCC12]. REF [BHA19]. referee [DCG^{+}19]. referees [RBV07]. Reference [SGGMF18, ADZ18, BM13a, BL14, EGR18, Kos12a, KSJL15, LR10, PSG^{+}21, TMLB16, XLSW24, CH15, NF08]. Referenced [LBMM14]. references [BH16a, BTZY21, BTZY24, Cam15, DLGC13, HLRD24, HLSH17, LAL09, Ron18, Sta21, SWHC15, TB18b, TMLB16, WV24, ZLW21]. referencing [ASS22, LCL^{+}22]. refined [SZR^{+}19]. refinement [Rou14b]. reflected [Luo21, TB18b]. reflecting [BGdMAW23]. Reflections [RY12a]. reflects [BM15]. regarding [AD17b, BCGM17a, ZZ21b]. regionalization [GHP21]. regions [BW11, BL12]. Regression [TW14b, AAH11, AF15, BD16, BSdMAM16, CRS14, GG19, MD19b, Per17a, The16d, YKKS16]. regular [ZG12]. Regularity [FM11a, HW14]. regulations [AvdB24]. reinforced [YF24, YF23]. reinforcements [MF21]. rejected [BMS^{+}09, BLV10]. Rejoinder [AD16c, LQLJ15, BW17b, FM17b, Lin18c]. rel [AAB^{+}21]. relate [HWX24, XWW^{+}24]. related [ÁBB21, BH22, CCC^{+}18, KB21, LZH^{+}22, RY12a, TMK16, Wal14, ZTG11]. relatedness [SRP15, VY10]. relation [ADD21b, AL08, BSK10, BBGSZS24b, Bur13b, CXZ^{+}22b, CB07, Egg13a, HHAA16, JXYS21, Rou08, RY12b, SYW18]. relational [Mal10, RS10b, YDWC08]. Relations [WTR^{+}22, ALX^{+}21, HSMB18, HLD^{+}23, LW09, LBZ16, YDCL13, dNL15]. Relationship [ALB24, AZS^{+}14, Ort15, AD15d, ADM17, ÁBB21, Ant22, BAGADF15, FMM14, HDD^{+}21, RP08, ZBK^{+}21]. Relationships [VY13, BNbH17, CCH12, JZY^{+}18, Mêg14, Mêg18, RFMC17, SVTdFC11, SLNC13, ZH18, dMAGBLIM18]. Relative [KLC21, CA11, HCZ21, LWB19b, LF21, RRKH17, SYP^{+}16, Vin12, PPFKB19]. relevance [GAB11, HLHC22, KCH21a]. relevance-based [HLHC22]. relevant [CHHL11, LBMM14, Sch13d]. reliability [Glä10, WZ21a, dDGFFI^{+}21]. Remaining [LO11]. remark [Gag09]. remarkable [AL17, Vîi18]. Remarks [Don24, BCGM17b, FM17a, FM17b]. Reny [Ng17]. reparameterisations [ZSCSG23]. repeat [LJB^{+}22]. RePEc [SW12]. replacement [TLZM24]. replicability [BP22]. Reply [YAY^{+}17, vRvLV^{+}10, Moe18, vdBHS17a]. Report [BdMA11, BdMAL12, HKA^{+}09]. reporting [BMHD11, WNH^{+}23]. reports [BGCG^{+}24, HPNM22, Cam14, TDKZ21]. repositories [XWK^{+}21]. represent [Wal17a]. representation [CHD^{+}24, DKK^{+}17, JLH23, PS13, VE14, dlVV18]. representations [LJJ21]. representative [NZZ^{+}16, ZLR16]. represented [QCVQdMA10]. reproduce [AOdFC12]. reputation [JZDZ24, MD19a]. requirements [HM21b, KK21]. Research [AJ13, BRN18, DN17, GÁCC17, GPZC24, GHS21, HLD^{+}23, KB17, LBML23, LYYY09, MS09, NLSM23, NPS13, PW19, RC16, SZZ23, Van11, XZS21, Yan14, YZ18, YWW^{+}23, ZLY21, vL18, AS17, AHL12, ACD11b, ACD12a, ACD12c, ACD12d, ADM13b, ADR13, ACD13b, ACD13a, AD14, AD15a, ADG15, AD15c, ACD15, ADG16, ADS17, ADM17, AD17b, ADZ18, AD18, ADD19a, ADD19b, AAD21a, ADG21, AAD21b, ADD21d, ALW11, AZS^{+}14, ASG12, Als21, ÁBB21, AH22, AYL14, AMKA21, BSHK21, BGCL07, BCGM17a, BBS21, BDR14, BAGADF15, BMD10, BLV10, Bor10, BM12, Bor14a, Bor14b, BSdMAM14, BH16b, BHA19, BH22, BGdMAW23, BPU^{+}14, BRN19, Bue21, CCC^{+}18, CMCMM22, CLR^{+}17, CM14, CG11, CtRlY21, CD23, CHD^{+}24, CRS15, CZG16, CLHHVH11, Coc08, CFvL19, CP23, CJF^{+}15, CT16, Cso21, Dan16, DLM^{+}19]. research [DT13, DGRCDG24a, DGRCDG24b, DK22, ETGS21, FT15a, FT15b, FLB22, FC11, FM11a, FM11b, FM17a, FM17b, FHH09, FHH17, FH13, GLW^{+}21, GNZ^{+}19, GKR^{+}24, GAB11, HHAL17, HJ15, HKXS23, HGJZ17, HM21b, HPZ21, HPNM22, HNT22, HCC12, IST10, JSSK14, Jar07a, JHK^{+}16, JZY^{+}18, JZDZ24, KS16, KMS23a, Ke18, Ke21b, KYC12, KK19, KSS16, KTR10, KHC11a, KHC13, KCH24, KKK17, KR21, KR24, KYhK23, LS14, LL15, LJJ21, LW09, LYF17, LY19, LT21, LMF24, LTL24, LY12, Lia21, Lin18a, Lin18b, Lin18c, Lin23, Lin24, LHTW15, LZH^{+}22, LLHW22, MYZ^{+}22, MBZS23, MAS14, MLDP21, McC10, MMY17, Moe10a, MdMALIV11, MCKM08, MBD17, Nie17, OL10, OLRF11, Ort14, Ort21a, OZZW24, PYCH19, PL13, PD21, Per10b, PRCRVQ^{+}09, PBIR24, Per10c, PC10, RGRE15, RGRE16, RS10a, RNB18a, RNB18b, RNB24]. research [RL22, SV18, SV24, Sch13b, SAFR22, SCJX22, SD22, SA18, Sta21, SZ10, SL12, TS12, The17a, TF17b, The17c, The18b, TBTB19, TBM^{+}19, TWH08, TA22, TD10, VPAE24, VS13, Vas09, VG10b, VCT22, Vîi17, VRK23, WRB^{+}11, Wal17b, WG10a, WHHW22, Woe14, Wu15, WYH17, WYLL18, XZKS22, XlYcG18, XLG21, XHA^{+}21, YDMS12, lYFyS18, YZYW21, Yan24b, YLZ18, YZ22, YSJ21, YP21, YX24, ZZL^{+}21, ZHS^{+}22, ZWDH24, ZBK^{+}21, ZL11, ZKJ^{+}24, ZZ18, dMAGBLIM18, vV19, vdBHS17b, SVKM21]. research-focused [BSdMAM14, BGdMAW23]. research-front [Jar07a]. Researcher [KBP23, ACD12e, BH16b, CLW24, CRS14, GFGALG21, KMJ^{+}19, MD18, VPAE24, WOLD22, XZKS22]. ResearcherID [BW22, BW17a]. researchers [ADS11, ACD13a, dABJGMC21, BW17a, BW17b, BW22, BGT22, BRN19, CTGM15, CH23, CGFÍ^{+}19, FT17, GP12, GBDG23, HLB^{+}19, HLB^{+}21, Let16, LSW^{+}10, LLH12, Mar16, NLC17, Pet17, PUKR21, RJDD08, Tol13a, Xie21b, Xie21a, Yur16, Yur17b, ZW21, Pet18]. ResearchGate [LTH18, YZ18]. residual [BSK10]. ResIP [KBP23]. resolution [LJJ21, SCMG16]. resource [CKKY21, FJOROMT22]. resourcefulness [PUKR21]. resources [ADDWD24, CGM15]. respect [CA11, Mal16]. Response [AD17c, But17, WB16b, AF15, BHM18, BHM19, MJZ^{+}24, MCD16]. responses [SX21]. resting [The16e]. Restricted [SZZ23]. Restricting [Sch15]. restriction [GPNA13, HR18]. restructuring [DFFBR16]. results [AORC11b, AORC12b, Egg12a, PRCRVQ^{+}09, TBTB19, THQL24]. Rethinking [KK15, SCJX22]. retiring [Egg15]. Retracted [ZWDS21b, ZWDS21a]. retraction [LWCH22]. retractions [ZCTN23]. retrieval [CTDM17, CCL18, MRR13, YDMS12]. Return [NF08, YAL21]. returns [lYRyYbL14]. reuse [PW19]. Revalidation [LH24]. reveal [LWWZ22]. revealed [Ser10, SYP^{+}16]. Revealing [ADD22, GLW^{+}21, PHMSML17]. reveals [HKK18, JLH23, VS13, XMLM21]. Reversing [Hag14b]. Review [BGS21, HMBI17, Pet18, vL18, ACHVH09, BI08, BD07a, BMD07, BMD08a, BMD08b, BGCG^{+}24, CR10, CWT^{+}24, CPM^{+}07, DCM21, DCG^{+}19, KCK21, MGC18, Moe18, MRR13, NO10, SG12, SPL24, SHL24, Sun24, WRB^{+}11, Wal16b, ZXS^{+}22, ZCTN23, CF16]. reviewed [BS13]. reviewer [BD07b, GMPRSFV23, GHS21]. Reviewers [Ano10e, Ano11e, Ano12e, Ano14e, Ano15e, Ano13e, BGS19, BGCGS22, BGCGS23, HWX24, SCML23]. reviews [BGGB13, GSV^{+}21, MJB11, PD21, WLH19]. revised [Ley13b]. Revision [PYSM24]. revisited [Egg13a, GKR^{+}24, Kos12d, vdBHS17b]. Revisiting [ACD12c, Abr18, MAS23]. revolution [RS16]. rewarded [CCSM21]. RG [LTH18]. rhetorical [The19a]. rhythm [LR10]. rising [PTV17]. risk [SM18, YWS23]. risks [RNB18b]. Rivals [vRvLV^{+}10]. roadmap [TWC^{+}23]. Robust [Sta21, SW14, Zit10]. Robustness [BP22, Glä10, Mal16]. role [ADD21c, ADD24, ÁBB21, BSW17, DK21, HCC12, MSL^{+}14, OMFF23, SJGBMA17, Sma18, Tu21, WG10a, vdBHS17a]. Ronald [Pet18]. room [ADD19b, WvEVW16]. Rooted [GNZ^{+}19]. Rousseau [Pet18]. Royal [BGCGS23, BGCGS22]. RSS [LAT07, PTA07]. rule [MLDP21, SLK^{+}23, WJ23]. rule-based [MLDP21, SLK^{+}23]. rural [CFvL19]. Russia [SY24]. Russian [CK22, MSY21].

Saint [SG12]. Same [AD15a, Dan16, GAD19]. sample [BGCGS22, BGCGS23]. samples [BM13b]. Sampling [WB16a, WB16b, Cla16, KRM14, Mut16]. Sandström [Lin18c]. SAO [YHS18]. scalability [BP22]. scalable [CSH18, AAB^{+}21]. scale [ACD11b, Ant18, BW17a, BL17, BW22, BK14, BvECW18, DBS15, DGRCDG24a, DGRCDG24b, DGV18a, GG09, GGSG10, KJKS22, KR21, LCL^{+}22, LBP24, MMCvLLC18, PSA21, RA15, SKK15, WW16, lYRyYbL14, ZH18, ZLR16]. scale-independent [GG09, GGSG10]. scales [CH15, Egg10a, Glä07, Tok23, Vîi17, Vîi18]. scaling [ACD12b, ACD12c]. schemas [BHH^{+}07]. scheme [DU21, ZWG^{+}19]. schemes [ADG21, KK15]. Scholar [OA13, AJ13, ALR^{+}23, CB19, FC10, LGS^{+}24, MKM^{+}22, XZS21, AN18, BMS^{+}09, HMBI17, KT19, MMOMHLC17, MMCvLLC18, MMOMTLC18, MBIH16, MT10]. Scholar-based [MT10]. scholar/journal [CB19]. Scholarly [YSPW09, ADD21d, AF17a, BBB^{+}22, BZL19, BGS19, BS13, BNbH17, CGXW22, CBT21, DCG^{+}19, JLH23, KRM13, KYC12, KT19, LLW21, SKLR16, WCG21, WHH21, WZHH23, WLLL18, YP22]. scholars [AAH11, DFCGB15, FC17, HWZW22, PXS24, SCJX22, dlVV18]. scholarship [MMCR21]. scholarships [PBIR24]. schools [ZZN19]. Science [APF09, DK22, SSAGB11, XWZL21, XWW^{+}24, AC09, AC17, BL21, BB11, BI10, BFS11, BS09, BD07b, BW11, BLWSE11, BL12, BSdMAM16, BK08, BPU^{+}14, BK14, BFM^{+}18, CtRlY21, CTL22, CA11, Dan16, DG21, DHLW17, DLGT19, FCZ24, Gar09, HRL12, HR18, HYC15, HbHB^{+}23, Ke21a, KB21, KZDK23, Kos15, Kos08, Lat21, Let16, LW08, LSW^{+}10, LT21, LTL24, LZR15, LBML23, LEW22, LHTW15, LZL23, LK17, MJZ^{+}24, Mag12, MO24, MJ13, MMY17, Meh19, MSL^{+}14, Mil15, MAS23, MN15, NLK^{+}21, NF12, NO07, PYWH15, PPPF18, PYCH19, PPG^{+}15, PNCG23, PRRC17, PB16, RV24, RY23, RCW15, SDAJ17, SJZ^{+}19, SAB^{+}16, Sku09, SWHC15, TB18b, The19a, Tsa14, Van13, Wal14, WTvE11, WV24, WG12, WW15, WKY^{+}22, WKCC23, XWY^{+}21, XMLM21, XYP^{+}22, YDCL13, YW15, YHWZ16, Yan24a, YW17, YWS23, YXX^{+}17, ZWM17, ZS08, ZMBP24]. science [ZY17, vdBHS17a, AYL14, AN18, BMS^{+}09, BL17, CKZ^{+}23, FMM16a, GP11, GAB11, GMFGVZ16, LR12, LB21, LQLJ14, LHT18, LIdMAM08, LBZ^{+}22, MMOMTLC18, NO07, PRRC17, Reh21, The18a, WW16, ZH18, CH14, FH13]. science-based [SWHC15]. science-technology [BL21, BK08, ZMBP24]. science/computer [LK17]. Sciences [VE14, WHH21, WZHH23, ADR13, BAL18, Fra10a, FJW^{+}15, HB23a, HSSM22, Ley17b, LSSP24, SW14, TSM09, TDCM24, VW16, YZYW21, XlYcG18]. Scientific [ALR^{+}23, BKK09, CB19, Din11b, FMM14, Gag13, HR16, LY18, LRH^{+}21, Mar15, Mar16, PD24, RL22, WH15, AD18, AD21, ADD21c, ADD22, AAK^{+}21, ACHVH09, AN18, Ant22, AF17b, BZLX23, BSHK21, BRA19, BLV10, BSdMAM14, BH15, BH24, BvECW18, BSdA^{+}21, BvR11, CMN08, CCL16, CTGM15, CH23, CXMR07, CCH^{+}09, CTDM17, CKZ^{+}23, CY23, CML24, CW18, Coc08, CSdFCA17, DLM^{+}17, DFCGB15, DMS18, DGRCDG24a, DGRCDG24b, DKK^{+}17, ES15, Fin14, FM10b, FM11b, FM14, FJW^{+}15, GHP21, GPGBMA10, GPZC24, GBMA12, GWMP16, GW17, GHS21, HSE22, HMBI17, HDY12, HDT^{+}13, HGJZ17, HPZ21, HLZ24, HCL13, HNT22, HCC12, IPS21, JLAL23, JLH23, KMS23a, KMS23b, KFM^{+}15, Ke18, KJKS22, KK12, Kos11, Kos12a, LWWZ22, LLH12, LRWY13, LGX^{+}23, LBD^{+}19, LAL09, LLHW22, MBZS23, MEG^{+}16, MMCvLLC18, Moe10b, OA13, OG13, OZZW24]. scientific [PCPG17, Per10a, Per10b, PRCRVQ^{+}09, PSI^{+}17, QLS21, QCVQdMA10, RC12a, RWB17, Ron18, RCC14, RCC18, SKLR16, SJGBMA17, SJGCGB^{+}24, San11b, San13a, San14, Sch13a, SYP^{+}16, SXAC22, SKY^{+}19, SROdFC13, SAB^{+}16, SC21, SRZ19, SVKM21, Sol07, SFF23, SCZ23, SG12, SFCK19, SDL16, SHL24, Tan13, Tan24, TWH08, TLM23, THQL24, TdR21, VG10a, VCT22, WRB^{+}11, Wal12, WWW10, WG10a, WWML14, WHLT18, WRA^{+}21, WZCC24, WLZJ24, Woe08b, Wu15, XWW^{+}24, XLSW24, XWZY21, XMM14, XMW^{+}24, YDCL13, YZ15, YLZ18, YZ22, YXX18, YXXW19, YSJ21, YX24, YAL21, ZH18, ZWBA21, ZY22, ZWG^{+}19, ZXS^{+}22, ZLYC21, ZJMX23, dDGFFI^{+}21, dlP11]. Scientist [LY24, ZZL^{+}21, AD15c, BBMP12, FM10a, HHB^{+}23, LBML23, LH22, WXP^{+}12, WPZ^{+}13, ZHS^{+}22, ZBK^{+}21]. scientists [ADM13a, AD15b, ACD15, ADS16, BAGADF15, Dan16, DFCGB15, FM11a, GWL24, HZZY24, LYZ^{+}17, LH21, Mar17, RGSM^{+}19, SW13, SZZ22, SGKM16, TN18, Van12a, WXP^{+}12, WHLT18, XZKS22, ZWZZ21]. scientograms [QCVQdMA10]. scientometric [Abr18, Cso18, GPNA13, Mag13]. scientometricians [Vin12]. Scientometrics [Gar09, VS13, Bor14b, FHH09, GL11, Glä10, Ley13a, Vîi18, WvEVW16, LBMM14]. SciKGraph [TdR21]. SCImago [BdMAL12, GNBVQ^{+}14]. SCiMet [AAB^{+}21]. scope [TYWG24]. Scopus [MCD16, MdMAGBLI21b, ABS21, BMS^{+}09, BLWSE11, ED15, FMM16a, FMM16b, GMFGVZ16, HB23b, HL19, Ley13b, LQLJ14, LIdMAM08, MMOMTLC18, Meh19, MBIH16, MdMAGBLI21a, The17b, The18a, The18b, VZAMMFAB15, WW16]. score [BH16a, BH16c, BHM21, FGW^{+}22, LTH18, LGS^{+}24, Lun07, RY12b, LH21]. Scores [HDB22, AD16a, BM15, Egg10a, Glä07, GHS21, TF15b, TN18, Vîi17, Vîi18, Wal14, YZYW21]. Scoring [BGCL07, PUKR21]. scouting [CKKY21]. Screening [PD21, HZL23]. Search [TS12, ZWM17, BM18, JZ21, Kua24, Ort14]. searches [The18c]. Searching [ZY16]. Seats [BB11]. secondary [Hag14b]. section [The19b, Wal16c, Wal16d, Wal17b]. sector [ADD21d]. securing [Cha22, DK21]. segregation [TSAH22]. select [LLH^{+}21]. selected [UK16, WXML21, WG10b]. Selecting [CX16]. Selection [vdB12, BW17a, BW17b, BW22, HTX^{+}18, HHB^{+}23, LLH^{+}21, Pet17]. Self [SC21, TDKZ21, WH15, ADG21, ADD21a, ADD21b, Fra07, PML21, Vîi16, YYW14, vdB12]. Self-assembly [WH15]. Self-Citation [TDKZ21, ADG21, Vîi16, YYW14]. Self-citations [SC21, ADD21a, ADD21b, Fra07, PML21]. self-service [vdB12]. Semantic [SHD15, AKML18, ALX^{+}21, CXZ^{+}22b, CXZ^{+}22a, CHD^{+}24, HWL22, HTZS23, LW11, PL13, TWZ^{+}24, TD10]. semantic-based [HTZS23]. semantically [CSH18]. semantics [YP22]. semi [BHH^{+}07]. semi-automatic [BHH^{+}07]. seminar [RW10]. SemPathFinder [SHD15]. senior [Hag14b, VW16]. seniority [Kos09]. seniority-independent [Kos09]. senses [KM22]. sensitive [XMM14]. sensitivity [ACD12e, ACD12d, CGD^{+}21, Mal16, XWK^{+}21]. sensitized [WHW^{+}14]. sentence [RY23]. sentences [KJS16, SD22, ZXH^{+}23]. Sentiment [PT09, CLC11, KZH^{+}24]. separately [The19b]. Sequence [LY14]. sequences [ER19a, RXL15, WLH11, YR08]. series [HRC11b, LR08, RSRP19, RH13, Vas09, ZKJ^{+}24]. Service [vdB12, OA13]. Services [GWMP16]. sessions [OA10]. set [Kos12a, RY12b]. Sets [CLHHVH11, BL14, LRWY13]. setting [Egg10b]. several [ÁBB21, BT21]. Sex [Luo21]. shape [KHC11a, KHC11b, LBP24]. shapes [AF17b]. Shapley [DLM^{+}19, Kar14, Mêg18]. share [TF17b]. shared [HLRD24, XWK^{+}21]. sharing [Kar14, LRG13, PW19, PC10]. shift [YSS23]. short [Mag12, RSRP19, WZ21a]. short- [WZ21a]. short-term [Mag12, WZ21a]. shortly [BLW14]. Should [ACD15, BHM21, CZZ24, The19b, BLW13, Cla16, CCSM21, Csa21, KKK21, Wan24]. shoulders [ADX^{+}17, ZZN19]. show [GBDG23]. shrinkage [YKKS16]. side [BH16c, JXYS21, Ley17b, Zit10]. SIF [RXL15]. signal [EAG22]. Signatures [VRK23, MAS14]. significance [BW13, Sch13b]. significant [Hür15]. signs [DLH^{+}21]. similar [CGFÍ^{+}19, GP12, HDB22, PXS24, KCK21]. Similarities [BGGB13, Che17]. Similarity [BBB^{+}22, NMNJM13, AC09, ALX^{+}21, QZY^{+}24, SRP15, TWZ^{+}24, VPAE24, WZ14, ZSH^{+}16, ZY21, dSSOH07]. simple [BSK10, Don17, Egg08a, FSZB15, LPM16, Mil13]. Simulation [BGT22, The16f, vZ13]. single [CH14, RS10b]. single-author [CH14]. single-relational [RS10b]. Sino [HNT22]. SIR [BdMA11]. Sirtes [RC12b]. site [YZ18]. sites [Ort15]. six [HW14, Ort18, WZ14]. Size [ZLY21, AD16b, AD16c, BBL19, Bur07b, CCC^{+}18, CW18, GTD16, JZDZ24, Wal16c, Zit16]. size-independent [AD16b, AD16c, GTD16, Wal16c, Zit16]. size-related [CCC^{+}18]. sizes [PSZZ12]. SJR [AW14, GPGBMA10]. SJR2 [GBMA12]. Skewed [MD12]. skewing [LT13]. Skewness [BL17, ADS17, RCC14]. Skłodowska [TDCM24]. Skłodowska-Curie [TDCM24]. skyline [SGKM16]. sleeping [FLQ23, HY21, LSZY14, LMF24, YBL^{+}22]. Slovenia [Per10a, Per10b, RL22]. slow [MAS23]. slow-cited [MAS23]. small [BP11, CG10, Egg24b, Jal11, KP13, Let16]. small-worldness [Jal11]. SME [KP21]. SMIAltmetric [WLZJ24]. Smirnov [KMS21]. Smolinsky [BHM18, BHM19]. SNIP [AW14, Ley13b, Min14, WvEvLV13]. Social [HY21, WZHH23, AAH11, CGXW22, CBT21, CRS14, DG21, FLB22, Fra10a, HS11, HbHB^{+}23, KK12, Let16, Ley17b, LLW21, LSSP24, LHTW15, Mar17, MdGNS22, NO07, Ort15, OMFF23, PL13, SAP18, TD10, TSM09, TDCM24, Tu21, VW16, Wal14, WLZJ24, WKCC23, XWW^{+}24, YZ18, YZYW21, ZZ21b, ZDJ^{+}16, VE14, WHH21]. socialist [HM21b]. Socializing [CMMT13]. socialomics [SX21]. Societal [DGRCDG24a, DGRCDG24b, Bor14c, BHA19, CLR^{+}17, LSSP24, OZZW24].

Societies [FM14]. Society [BGCGS23, GBDG23, BGCGS22]. socioacademic [RP08]. sociology [MS09]. soft [HLRD24]. Software [MMY17, Als21, Gar09, LCY19, PYWH15, PYCH18, PYCH19, PW19, vEW14]. solar [LBK^{+}21, WHW^{+}14]. solitude [BHCH17]. Solution [ER19b]. Solutions [MD12]. Solving [XWZY21]. Some [AS17, BILL07, Bor10, BdMA11, Bur08, DLGC13, GTD16, Mut16, WvE08, WvE09, WvEvL^{+}11, WvEvLV13, DGRCDG24a, DGRCDG24b, Pis22, RH13, San15, SMP12, BL12]. Source [DG21, BK14, HMBI17, JPL22, Laf07, LA15, RGMJC17, RC14]. source-effort [LA15]. source-item [Laf07]. sources [BMD08a, LSW^{+}10, LRWY13, Ron18, VY13, ZZ21c]. Spanish [DMCLH22]. spanning [HZW^{+}23, WYC23]. Sparkles [HZW^{+}23]. sparse [SKM22]. Spatial [FHH09, Cop19, Cso18]. speaking [Cha22]. Special [Wal16c, Wal16d, Wal17b, LR08]. specialisms [CFvL19]. Specialization [ÖSY21, DBS15]. specialized [ADD19a, Ron12, Ron13]. Specialty [dNL15, FLB22, Ron18]. specific [AF17b, BSdMAM14, HSE22, HDB22, HGJZ17, SLK^{+}23, WJ23, ZXH^{+}23]. specification [Ley13a]. specificity [CW18]. spectra [LSZY14]. Spectroscopy [LBMM14, CCL18, TMLB16, CH15]. spectrum [FM10b]. Speed [EBG11, TLD^{+}21, BD10b, BD10a]. spillovers [ADD21a]. split [Yun22]. splitting [HSMB18]. spot [WNH^{+}23]. spurs [YZ22]. square [RC12b]. squares [Egg10b]. SSCI [LHTW15]. SSRN [Kak16]. Stabilisation [Vas09]. stability [CA11, Paj15, ZMZ24]. stable [BBGSZS24b, AAB^{+}21]. staff [ADS16, BGdMAW23, RA15]. stage [LHDH21, MN15, WRA^{+}21, WYLL18, XlYcG18, lYFyS18]. stages [LRH^{+}21]. standard [BGGB13, BM18, LYF17]. standardization [ACHVH09, SK14, TMLB16]. standardized [ACD11b]. standardizing [ACD12b]. Standing [ADX^{+}17, ZZN19, AD21]. Stanley [DKK^{+}17]. star [SGKM16]. stars [BHCH17, PTV17]. start [XWZY21]. Stata [BO12]. States [LWB14, SY24, Yur17a]. stationary [RSRP19]. statistic [KMS21]. Statistical [BBMP12, dMAGBLIM18, BGCL07, BSdMAM14, BGCG^{+}24, LJHH23, MD19b, Sch13b, Sch16, TF17b, Wal16a, Wal16d]. statistics [KRM14, Khr09, LAT07, SKM22, Tok23]. status [LWCH22, SCML23, SPL24, ZSF^{+}23, ZSF^{+}24]. steep [Szi21]. step [ER19b, GBMA12, Woe14, YDW^{+}23, YWZ^{+}23]. step-based [Woe14]. Steps [LWB19a, BM11a]. still [HL19]. Stirling [LWB19b]. Stochastic [HKK18, Bur07a, Bur13b, Glä10, MBD17, NT12, OG13]. stock [JCWX24]. stocks [Han07]. Strange [GP11]. strategic [Coc08]. strategies [RV24]. strategy [HZL23]. strength [ADS16, SHZ23, TLZM24, Tu24, WWY21, ZYCK12]. strengths [AD14, RGMJC17, SZR^{+}19]. strikes [SG12]. strong [KB21]. stronger [GHS21]. strongly [LYZ^{+}17, MF21]. Structural [KLC21, BGGB13, DLM^{+}17, HYS21, LYYY09, QLS21, RP08, YAL21, ZMZ24, Ort21a]. Structure [YAT^{+}22, AL17, CR10, KMS23b, LY18, Per10a, PVH21, SG12, SL12, The19a, TdR21, WWY21, WXML21, YHWZ16, YL22, YY23, ZF22, ZMBP24, LYF17]. Structured [FM11b]. structures [FLB22, LBSA13]. student [Lin23, Lin24, YZ22]. students [YZ22]. studies [AS17, AF15, BGGB13, BMHD11, BM13b, BHA19, MSL^{+}14, NO10, PD21, Rou14b, TSM09, WNH^{+}23, YY23, vdBHS17a]. Study [KP24, ASS22, AvdB24, AS21, BGCGS22, BGCGS23, BMD08a, BD09, BL13, Bor14c, BW17a, BTZY21, BW22, BTZY24, CM14, CG10, CTDM17, CySlY18, CTL22, CA11, Cso21, DFHCT21, DGRCDG24a, DGRCDG24b, EI10, FZDW15, Fin13, Fin14, Fra09, GAB11, HSMB18, HCL13, HW14, HYS21, KR21, LB16, LSZY14, LSSP24, LZR15, LY12, LRG13, LY14, LHT18, LLH^{+}21, MBZS23, MLDP21, MÉVQR21, Ort18, PYCH18, Pen15, PYSM24, PSI^{+}17, PC10, PB16, QLS21, Sch10a, Sch13c, Sch13d, Sch15, SZZ22, SY24, TMK16, Tan13, TSM09, WG10a, WHW^{+}14, WJW^{+}19, WHH21, WML^{+}21, WDLX23, WZHH23, WG10b, WLH11, WWXL24, XlYcG18, YDMS12, lYRyYbL14, YHS18, YWW^{+}23, ZZL^{+}21, ZZZ^{+}23, vZ13, vdBS15]. studying [YZ22]. style [vdBM22]. styles [HSSM22, SCZ23]. stylised [GHvdB^{+}14]. sua [Kos12c]. Sub [FM14, HRC12, RXL15, WYLL18]. Sub-field [FM14, HRC12]. sub-group [WYLL18]. sub-impact [RXL15]. subfields [Wal14]. subgraph [HDT^{+}13]. Subject [SDAJ17, ZLJ^{+}10, GPNA13, Glä07, HSMB18, KP24, MMOMTLC18, UG24, XWZL21]. subjective [Wal17a]. subjects [ADD19b, PT08, TW14a]. submission [SHL24]. submissions [BGCGS22, BGCGS23]. subnet [WWY21]. subsequent [BSA21]. subset [LBMM14, ZLR16]. subsidies [Yur17b]. subsidized [ZSP^{+}22]. Substantial [HL19]. substantive [MJB11]. substantive-methodological [MJB11]. subunits [XlYcG21]. suburbs [BP11]. Success [SYW18, Egg14, GWL24, KKK21, MRBI17, Rou14a, SZZ22, vdBM22, FGMM12, FGMM13]. success-index [Egg14, Rou14a, FGMM12, FGMM13]. Successful [Kos11, HbHB^{+}23]. successfully [BSdMAM16]. successive [Tol08]. such [Bor19]. Suggestions [PMAT13]. Sugimoto [vL18]. Suitability [HMBI17, The18c]. sum [LRE17, PHMSML17, QM23]. summary [Bor14b]. summation [MSPD16]. supercomputing [BFS11]. superstar [SCBB11]. supervised [Reh21]. supervision [DMCLH22]. support [BHH^{+}07, FJOROMT22, MJB11]. supported [GWL24, LY19]. supporters [HKXS23]. supposed [GL11]. suppressing [VMWM17]. surname [AD17a, AD17c, Web17]. survey [SD11, WPZ^{+}13]. surveys [SB13, SAB^{+}16]. survival [LHWS18]. Survive [LHWS18]. sustainability [WG12]. sustainable [WRA^{+}21]. sweet [WNH^{+}23]. switching [DFHCT21]. symbiosis [LGX^{+}23]. symmetry [Woe08b]. synchronization [YPYJ23]. synergy [MJB11, Mêg18]. Synthetic [PRCRVQ^{+}09]. System [BGR21, dDGFFI^{+}21, ADS11, GG09, GGSG10, HHAL17, KSJL15, KKK17, LW17, Mêg18, PCCN21, RLY12, RC14, Sil23, ZDJ^{+}16]. Systematic [KCK21, GSV^{+}21, MMOMTLC18, PD21, SG07, VZAMMFAB15, WvE13]. systems [AS17, ACD12a, ASG12, AAB^{+}21, AvdB24, ADDWD24, BDR14, CZG16, Egg11, EI10, Gol14, LBSA13, LRC13, MRR13, PRRC17, RCW15, SV18, SV24, Wal17b, WW16, dMAGBLIM18].

table [BB11]. tables [ZZ21a]. tagging [ZDJ^{+}16]. tail [Bur13a, zChHY13, GP12, GS10, KHC11a]. tails [LRWY13]. taking [WZ21a]. talent [TLD^{+}21]. talented [HB23a]. task [CJF^{+}15, Dan16, KZH^{+}24]. task-force [CJF^{+}15]. tat [AD17b, BCGM17a]. Taverna [GWMP16]. Taxonomy [BHH^{+}07]. Team [LJB^{+}22, YAT^{+}22, ZLY21, JZDZ24, MYZ^{+}22, RV24, YW24]. Teams [WLCL23, CY23, LHC14, MBZS23, SVKM21, WH15, ZWL^{+}24]. teamwork [CZFD21]. Tech [GKR^{+}24]. Technical [AORC11b, AORC12b, FM14, HSSM22, LRWY13, LZL23, QZY^{+}24]. technique [JS14, TMK16]. techniques [AA19, ETGS21, MRR13, VPAE24]. Technological [HHC14, PY18, RNB18b, CML24, HCC12, JYG24, Ke18, KHC18, LBK^{+}21, LLGM21, PCPG17, QZY^{+}24, SHZ23, ZYCK12]. technologies [EAG22, HLRD24, JYG24, TMK16, TWC^{+}23, WLY^{+}24, ZWM^{+}21]. technology [AKML18, Ant22, BL21, BK08, CHHL11, DHLW17, DLGT19, FLQ23, HTZS23, HR18, HCC12, HYC15, Hur24, JWL17, JKHS21, JPL22, KS16, Ke21a, KZDK23, KG07, Kos08, Lat21, LLGM21, MSL^{+}14, NLK^{+}21, PY18, RY23, SLK^{+}23, SHZ23, TWC^{+}23, WY21, WG12, WW15, WJ23, XWY^{+}21, XMLM21, XYP^{+}22, YHS18, YXXZ24, ZMBP24]. temper [HCZ21]. Temporal [BM21, CFK21, BH22, JS14, KYC12, KCH21a, TS16, Yan24a]. tendencies [LGS^{+}24]. tension [WYC23]. tenure [CCSM21]. TeraGrid [BFS11]. term [ADF19, DU21, Glä07, LGX^{+}23, Mag12, RSRP19, SAFR22, SLW15, TN18, WZ21a, ZKJ^{+}24]. terms [BL12, BAL18, Sch14b]. territorial [AD15b]. territories [CFvL19, KE17, YAY^{+}16, YAY^{+}17]. test [BHA19, DGV18b, Jar07b, KP13, Ma12, Sir12]. Testing [DD19, RC12a, KdBOK15a]. tests [Sch13b]. text [BvECW18, Che17, DU21, JFZZ23, Khr09, PWY21, PYWH15, PYCH19, SAB^{+}16, TWZ^{+}24, WZ21b, WRGL23, WV24, YX24]. texts [DLGC13, FCAD21, LAL09]. Textual [Sun24, BL15, SD22]. TF [DU21]. TF-MONO [DU21]. Theil [WYLL18]. Their [KCH21a, KLC21, ADR13, ADM17, AOdFC12, BBGSZS24b, BDR14, BAGADF15, BSMD12, Bor19, BvR11, Cam14, Che17, CRS14, CFvL19, CA11, Dan16, EAG22, ER19a, Egg24b, FM14, HKK18, HR18, Hur24, KHC13, KHC18, LTH18, LB21, LY19, LF21, LGS^{+}24, OA10, OLRF11, RY12b, RH13, SYP^{+}16, SCML23, TB18b, Tsa14, WW15, XWW^{+}24, XOMW21, ZBK^{+}21]. them [Dem18]. themes [McC10, Per10c]. Theorem [Ant18]. theorems [Smo21b]. Theoretic [BGS21]. theoretical [And17, SY24, Vîi16, WvEvL^{+}11]. theory [Bur08, CCH^{+}09, EP07, Egg07a, Egg21, LRE17, Mêg18, Per17b, SAP18, ZCM^{+}24, CLHHVH11]. therapy [RYY10]. There [Cla16, GL11, KP24, MBD17, The16b, TSVBAM22, YXX18]. theses [DMCLH22]. thin [LBK^{+}21]. thing [TBTB19]. things [Luo21, TMK16]. Things-related [TMK16]. think [SVKM21]. Three [AOdFC12, The17c, AL17, BBGSZS24a, BAGADF15, CX16, HHAA16, KR24, PYCH18, PVH21, SCZ23, WYLL18, XLG21]. three-dimensional [AL17, BBGSZS24a]. Three-feature [AOdFC12]. three-stage [WYLL18]. thresholds [Pis22, Sma09]. tie [Pen15, Tu24]. tie-generative [Pen15]. tier [SCMG16]. Time [Fia12, HRC11b, JS14, ACD12e, ACD12d, ADD21a, Ada18, BDR14, BL15, BSA21, DGDG13, Fas24, Fin14, HKK18, HW14, LR08, LF21, MR13, MMZ16, RSRP19, RH13, Sch13d, Sch15, TT22, TF15b, Vas09, VAdFC13, WZ18, XZS21, XWW^{+}24, XMM14, YW17, ZKJ^{+}24]. Time-aware [Fia12]. time-balanced [MMZ16]. time-dependent [HW14, Sch13d]. time-trend [Ada18]. time-varying [VAdFC13]. timed [Sch15]. timelines [PTA07]. TimeRank [FC17]. times [BM13a, DGDG13, RGZSC23]. timetable [WXP^{+}12, WPZ^{+}13]. Timing [KHC18]. tit [AD17b, BCGM17a]. title [Ron18]. titles [BvR11, TSM09]. to/from [ZBK^{+}21]. together [Tan13]. tolerance [Jal11]. too [The16b]. tool [AC17, CGD^{+}21, FM10b, QCVQdMA10, TD10, TSM09, vEW14]. tools [FM11a, KK12, PYCH18]. Top [KR24, ADS11, ACD13a, ADS16, AZS^{+}14, BBMP12, CAL15, CH14, ED15, GP12, LWB14, Lin18a, Lin18b, MMCR21, SCMG16, Sun24, Tsa14, WOLD22, XZKS22]. top-1 [LWB14]. top-10 [LWB14]. top-cited [CAL15, CH14, Sun24]. top-rated [ED15]. top-scientist [BBMP12]. top-tier [SCMG16]. Topic [XWY^{+}21, BP22, Bue21, CTDM17, DGDGSPSV14, HTLC21, HHB^{+}23, JLYM16, JS14, JY21, JS22, KPS22, KJKS22, KB17, LDS^{+}12, SAB^{+}16, SVKM21, WZCC24, XZDS21, XZS21, XZKS22, YKKS16, YF23, ZLL^{+}18]. topic-driven [KPS22]. topic-integrated [KJKS22]. Topic-linked [XWY^{+}21]. topic-model-based [HTLC21]. topical [AYL14, Din11a]. Topics [YDMS12, GLW^{+}21, HLB^{+}19, HLB^{+}21, KBCR10, KP24, LYZ^{+}17, LZR15, MO24, MN15, SA18, TBTB19, WHHW22, XLG21, XHA^{+}21, XYP^{+}22, Yan14]. Topological [Din11a, RDF^{+}18, BKK09]. Topology [LDSD18, AOdFC12]. Topology-driven [LDSD18]. Total [LLH12, MSPD16, YR08]. trace [LYZ^{+}17]. Tracing [YDS^{+}15, YP21]. Tracking [WHLT18, PT08]. trade [ADD24]. trade-off [ADD24]. trading [YDCL13]. traditional [SMP12]. traditions [HM21b]. traffic [VY13]. Trajectories [WW15, CPN24, CXZ^{+}22a, KJKS22, LBK^{+}21, LH24]. Trajectory [JHK^{+}16, KLC21, LBK^{+}21]. Transdisciplinary [KZH^{+}24]. transfer [RJDD08]. transformation [Ley12]. Transformations [LR07, Egg08a]. transformative [SD22]. transition [GGSG10, HR18]. transitions [BKK09]. transitivity [IPS21]. translation [DLGT19]. translational [LT21, LTL24]. transmission [Mêg14]. Transparency [WNH^{+}23, VZMFAB16]. Treatment [SYK22]. Tree [ZMBP24, Kos13b]. trees [JLYM16, MLDP21]. trend [Ada18, HCC12, LDSD18, YHS18]. Trends [MO24, RL22, BSHK21, Bue21, ETGS21, JSSK14, JHK^{+}16, KYC12, LJM13, MCLV23, SGGMF18, VZMFAB16]. triad [FM10a]. triadic [Pen15]. triads [dNL15]. trial [TLD^{+}21]. trial-and-error [TLD^{+}21]. Triangular [LYC21, AMKA21]. tribes [CFvL19]. tribute [Sma17]. triggering [MBZS23]. tripartite [WJW^{+}19, WDLX23]. triple [Mag12, Mêg14, Mêg18]. Trojan [DCM21]. truncated [KSS16]. Tsinghua [ZL11]. turnover [ADR16]. tutorial [GWMP16]. tweet [WHH21, WZHH23]. tweeters [YXXW19]. tweeting [CGXW22]. tweets [HB23b, YXX18, YXXW19]. twenty [AZS^{+}14]. Twitter [WZHH23, DMS18, HWZW22, MA21, MdGNS22, WHH21, WLZJ24, YKKS16, YXX^{+}17, ZZZ^{+}23]. Two [BMD10, NO10, RH13, Smo21b, ADZ18, CT16, DN17, EI10, Let16, MD12, PPFKB19, SSAGB11, XlYcG18, lYFyS18]. two-mode [DN17]. two-stage [XlYcG18, lYFyS18]. type [BI10, BD07a, BD09, Bur08, CH23, EP07, Egg08b, Egg10b, Egg10a, ER19a, ER21, Egg24a, GS10, JSSK14, Kos09, Kos10a, Kos10b, PSA21, San14, SMP12, SG07, Smo21b, Vîi16]. types [FJOROMT22, LH24]. typical [Wal17a].

U.S. [HCC12]. UK [BAL18, BHA19]. Ukraine [SY24]. Ulrich [WHL17]. un-published [Fra09]. Unbiased [XMLM21]. uncertainty [ADG15, ADG16, CSH18, Ley13a, Sma18, YSS23, ZSWD17]. Uncited [KMS23b, LZR15, LRWY13, The16b]. uncitedness [Bur13b, Egg13a]. Unconnected [TMK16]. uncovered [BSS22]. Uncovering [DMCLH22, HHAL17, LHTW15, ZCM^{+}24]. uncovers [WXML21]. under-cited [HR16]. underlying [The17a]. understand [Vas09]. Understanding [CTGM15, CTDM17, HLSH17, LHC14, QLS21, SCZ23, WML^{+}21, ZXS^{+}22, ZSF^{+}23, ZSF^{+}24, CMN08, KD15, WRB^{+}11, vEW09]. undirected [HLR17]. Uneven [MSPD16]. unexpected [RGZSC23]. unfolding [MCLV23]. Ungendered [HSSM22]. unified [WvEN10]. uniform [Hür15]. uniformity [MAS23]. Uniformly [TBM^{+}19]. unintended [KG07]. Union [LWB14]. United [SY24, LWB14, Yur17a]. units [Coc08, HWL22]. unity [Vin13b]. universal [LAT07, Wu13]. Universality [KRM13, DD19]. Universities [GAW21, ACD11b, ACD12a, ADS16, AAD21a, ADDWD24, BHCH17, BBS21, BSdMAM14, Bor19, BGdMAW23, DBS15, PRRC15, SZZ22, Wu15, YZ18, lYFyS18, ZZL^{+}21]. University [MSY21, vL18, ADS11, ACD13a, AD15a, AD16a, ADG16, BC16, BGT22, CCC^{+}18, DFFBR16, EGR18, FZDW15, FHH17, Hic17, LHC13, Mêg14, Mêg18, MdMALIV11, WYLL18, ZL11]. university-industry [FZDW15]. unknown [SHD15]. unlikely [HSSM22]. unmeasurable [WKY^{+}22]. Unpacking [Wan14, WYC23]. unproductive [ACD13a]. Unraveling [HGJZ17, PD24]. Unravelling [DFCGB15]. Unsupervised [CJF^{+}15]. Untangling [MS09]. unused [Mêg14]. unusual [BTZY21, BTZY24]. Unveiling [Yan24b]. Unweighted [Wal17a]. upon [HM21b]. upper [Hür15]. urban [CFvL19]. URL [KK13]. USA [Kos08]. usability [MN15]. usage [HS11, HKA^{+}09, PYCH18, YDS^{+}15]. Use [BW17b, DFCGB15, HKA^{+}09, MLDP21, TV21, AL17, BM13b, BLM13, BL14, BH15, BH18, BGT22, HM21b, KRM14, MMOMHLC17, Pet17, SCML23, TN18, Vin14, Wal12, Wal16a, WHH21, WZHH23]. used [BW17a, BW22, BGT22, PD21]. useful [BD10b]. user [AF17a, ZZZ^{+}23]. users [SX21]. Using [KE17, KTR10, LLGM21, MA21, Meh19, RGSM^{+}19, SAB^{+}16, Vas09, WZ21b, YAY^{+}16, YAY^{+}17, ASG12, dABJGMC21, BGCL07, BD07a, BMS^{+}09, BW11, BP11, BW13, BL13, Bor14c, BM16, BK08, CHHL11, Cop19, DGRCDG24a, DGRCDG24b, DVG16, DGV18b, EAG22, ER19b, ER23, EGR18, FC17, FHH17, GHP21, GWMP16, HSMB18, HTLC21, HZL23, HCS19, JSSK14, JXYS21, Khr09, KJS16, KJKS22, KHC11a, KHC13, KBP23, LBSA13, LB21, LTC22, LMF24, LHWS18, MT10, MdGNS22, NV22, NLC17, NPS13, OOB10, Ort21a, PUKR21, RCW15, Sch13b, SW12, SD22, SDAJ17, STP17, Sma18, SAP18, TLV^{+}21, TWC^{+}23, Van08, WWY21, WZ14, WJ23, XZDS21, XlYcG18, YAL21, Yun22, ZWBA21, ZYCK12, ZXS21, ZY16, vZ13, BSdMAM14]. USPTO [ZSP^{+}22]. Utilizing [ZF22, Kua24]. UTPL [MLDP21].

Validating [CGSZS22]. validation [BHH^{+}07, BD07a, BL13, BTZY21, BTZY24]. Validity [Bor14c, SY24, BMS^{+}09, BD10a, BHA19, BT21, vdBS15]. value [BO19, DLM^{+}19, GAD19, HJ15, HZL23, Kar14, Mêg18, Moe18, SHZ23, The16e]. values [BLW13, Mal16]. valuing [HM21a, BGS19, GAD19]. Vanclay [But11]. variability [SLNC13]. variable [BMD08a]. variables [AF15, GPNA13]. variant [Zit10]. variants [ACHVH09, BP11, BMHD11, GR09, NJFD14, Sch10c, SMP12, The16b]. variations [TS16]. Variety [ZLY21, LWB19b]. varying [ACD11a, CW18, VAdFC13]. Venues [SYK22, AF17a, ZW21]. versus [AD17b, BCGM17a, KFM^{+}15, Ley12, PRRC15, SKLR16, Yur16, ZG12]. very [KP13]. via [BZLX23, CB19, FZZ23, HLRD24, MD19a, RZLC21, SYK22, TA22, YSS23, ZZ21a]. view [GGSG10, YDCL13]. viewing [BSdA^{+}21]. Villar [AHRCV17]. Vincent [vL18]. visibility [AOdFC12, HM21a, LSSP24, OZZW24, RJDD08]. visible [HR16, LTH18]. visited [ZDJ^{+}16]. visits [OA10]. Visual [AYL14, BS09, ZY16]. Visualization [FZDW15, KSJL15, MAS14, BHH^{+}07, BW11, Sku09, VGRB16]. Visualizing [YW15, YHWZ16, ADG15, CLHHVH11, Gar09, vEW14]. vital [YDW^{+}23]. vocabulary [YZ15]. vocabulary- [YZ15]. vol [RLG14]. volatility [GZRMLSC18]. Volume [BGCGS23, BW22, WZHH23]. VQR [BCGM17a, FM17b, AD17b, DLM^{+}19, FM17a]. vs [ADZ18, ADD19a, AAD21b, AW14, Din11a, DGV19a, GAB11, LG11, PD21].

wagging [Bur13a]. Wagner [Rou19]. Waltman [RC14]. was [Van12b]. water [HSMB18]. way [BHA19, CGFÍ^{+}19, Zit16]. ways [BLV10]. weak [EAG22]. weaknesses [AD14]. web [KPB09, KK12, MS09, OA10, SLL^{+}12, AN18, BSdMAM16, BL17, CKZ^{+}23, FMM16a, GP11, GMFGVZ16, GWMP16, KTR10, LR12, LB21, LQLJ14, LHT18, LIdMAM08, LBZ^{+}22, MMOMTLC18, NO07, PRRC17, Reh21, The18a, VY13, WW16, ZH18]. Web-based [BSdMAM16]. Web-of-Science [LR12, LB21]. Webometric [TS12]. Webometrics [VY10]. Webpages [VY10]. Weibo [YXX^{+}17]. weight [JJ24]. weighted [DKK^{+}17, HY08, MSPD16, Per10c, Wal17a, WvE08, YZF13, ZYZ18, ZRY11, ZY12, PPFKB19]. weighting [DU21, ZWG^{+}19]. weights [Kar14, PML21]. Welcome [Egg07b, Siv16]. were [BMS^{+}09]. Where [HCL13, KKK21, BFS11]. Which [BL12, BLW13, BM15, BH15, DT13, FJOROMT22, HB23b, LTH18, NZZ^{+}16, VCT22, PS13, TS16]. Who [AD18, ALW11, Dem18, KdBOK15a, KdBOK15b, LW17, LZL23, YXXW19]. whole [ASG12, Wal17a]. whom [KdBOK15a, KdBOK15b]. whose [KP13, LF12]. Widespread [PD21]. Wikipedia [RS16, SVTdFC11, TSRFAM19]. Will [LH21, LZL23]. Williams [Mut16, Nan16, Pet17, Sch16]. window [ACD11a, Sch15]. windows [DGDG13, AZS^{+}14]. Winner [BH24]. winners [San15]. winning [FT17, LHDH21, Van12a, ZWZZ21]. Wise [Pet18, MRBI17]. Within [PML21, ACD12a, AAD21b, BGT22, Mêg18, PT08, SVTdFC11, Szi21, WLCL23]. Within-journal [PML21]. without [ADD21a, Hag14a]. Woeginger [Rou08]. woman [KR21]. Word [LAT07, VY10, ZLL^{+}18, ZSF^{+}23, ZSF^{+}24]. Words [FCAD21, CSH18, LW11, Ron18]. work [GMPRSFV23, LZL23, NZZ^{+}16, WXP^{+}12, ZXH^{+}23, vdBS18]. workflows [GWMP16]. working [Fra09, WXP^{+}12, WPZ^{+}13]. world [ADDWD24, CG10, HCC12, MEG^{+}16, RNB18b, YW15, BdMA11, BdMAL12]. world-class [ADDWD24]. worldness [Jal11]. worlds [Egg24b]. worldwide [BSdMAM14, BGdMAW23, Cso18]. worth [GAD19]. Would [Vin13b]. Wouters [Moe18]. writing [HSSM22, SCZ23, vdBM22]. wrong [Van12b]. Wu [Egg11]. Wu- [Egg11].

X [WLZJ24].

Yadollahie [WvE08]. Yang [KE17]. Year [CH15, Cam15, CS22, DGDG13, Hic17, Kos15, TMLB16]. yearly [LY14]. Years [LBMM14, dABJGMC21, Bor13b, BLW14, Egg12b, Per10b, TW14a]. Young [GWL24, HB23a, SZZ22].

zero [BHM18, BH18, BHM19, KSS16, LRE17, The16b]. zero-inflated [BHM18, BH18, BHM19]. zero-sum [LRE17]. zero-truncated [KSS16]. Zipf [Per10b, WRGL23]. Zipfian [ALB24].

References

Abrishami:2019:PCC

[AA19] Ali Abrishami and Sadegh Aliakbary. Predicting citation counts based on deep neural network learning techniques. Journal of Informetrics, 13(2):485–499, May 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718303821.

Allahbakhsh:2021:STA

[AAB^{+}21] Mohammad Allahbakhsh, Haleh Amintoosi, Behshid Behkamal, Amin Beheshti, and Elisa Bertino. SCiMet: Stable, sCalable and reliable Metric-based framework for quality assessment in collaborative content generation systems. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720306441.

Abramo:2021:CRP

[AAD21a] Giovanni Abramo, Dag W. Aksnes, and Ciriaco Andrea D’Angelo. Comparison of research performance of Italian and Norwegian professors and universities. Journal of Informetrics, 14(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157719301981.

Abramo:2021:GDR

[AAD21b] Giovanni Abramo, Dag W. Aksnes, and Ciriaco Andrea D’Angelo. Gender differences in research performance within and between countries: Italy vs Norway. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157721000158.

Abbasi:2011:IEC

[AAH11] Alireza Abbasi, Jörn Altmann, and Liaquat Hossain. Identifying the effects of co-authorship networks on the performance of scholars: a correlation and regression analysis of performance measures and social network analysis measures.

Journal of Informetrics, 5(4):594–607, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000630.

Akella:2021:EIS

[AAK^{+}21] Akhil Pandey Akella, Hamed Alhoori, Pavan Ravikanth Kondamudi, Cole Freeman, and Haiming Zhou. Early indicators of scientific impact: Predicting citations with altmetrics. Journal of Informetrics, 15(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720306453.

Abbasi:2016:LAL

[Abb16] Alireza Abbasi. A longitudinal analysis of link formation on collaboration networks. Journal of Informetrics, 10(3):685– 692, August 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157716300815.

Alvarez-Bornstein:2021:FRH

[ÁBB21] Belén Álvarez-Bornstein and María Bordons. Is funding related to higher research impact? Exploring its relationship and the mediating role of collaboration in several disciplines. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157720301309.

Abramo:2018:RSC

[Abr18] Giovanni Abramo. Revisiting the scientometric conceptualization of impact and its measurement. Journal of Informetrics, 12(3):590–597, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157718301925.

Amodio:2021:IPA

[ABS21] Pierluigi Amodio, Luigi Brugnano, and Filippo Scarselli. Implementation of the PaperRank and AuthorRank indices in the Scopus database. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000778.

Ahlgren:2009:DDS

[AC09] Per Ahlgren and Cristian Colliander. Document-document similarity approaches and science mapping: Experimental comparison of five approaches. Journal of Informetrics, 3(1): 49–63, January 2009. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157708000680.

Aria:2017:BRT

[AC17] Massimo Aria and Corrado Cuccurullo. bibliometrix: an R-tool for comprehensive science mapping analysis. Journal of Informetrics, 11(4):959–975, November 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717300500.

Abramo:2011:AVL

[ACD11a] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. Assessing the varying level of impact measurement accuracy as a function of the citation window length. Journal of Informetrics, 5(4):659–667, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771100068X.

Abramo:2011:FSA

[ACD11b] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. A field-standardized application of DEA to national-scale research assessment of universities. Journal of Informetrics, 5(4):618–628, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000654.

Abramo:2012:DRP

[ACD12a] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. The dispersion of research performance within and between universities as a potential indicator of the competitive intensity in higher education systems. Journal of Informetrics, 6(2):155–168, April 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711001064.

Abramo:2012:HIC

[ACD12b] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. How important is choice of the scaling factor in standardizing citations? Journal of Informetrics, 6(4):645– 654, October 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712000533.

Abramo:2012:RSC

[ACD12c] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. Revisiting the scaling of citations for research assessment. Journal of Informetrics, 6(4):470–479, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000272.

Abramo:2012:SARb

[ACD12d] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. A sensitivity analysis of research institutions’ productivity rankings to the time of citation observation. Journal of Informetrics, 6(2):298–306, April 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711001040.

Abramo:2012:SARa

[ACD12e] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. A sensitivity analysis of researchers’ productivity rankings to the time of citation observation. Journal of Informetrics, 6(2):192–201, April 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711001155.

Abramo:2013:IUT

[ACD13a] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. The impact of unproductive and top researchers on overall university research performance. Journal of Informetrics, 7(1):166–175, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000892.

Abramo:2013:IRP

[ACD13b] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. Individual research performance: a proposal for comparing apples to oranges. Journal of Informetrics, 7(2): 528–539, April 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000151.

Abramo:2014:AHC

[ACD14] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. Are the authors of highly cited articles also the most productive ones? Journal of Informetrics, 8(1):89– 97, January 2014. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000886.

Abramo:2015:SRP

[ACD15] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. Should the research performance of scientists be distinguished by gender? Journal of Informetrics, 9(1):25– 38, January 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157714000960.

Alonso:2009:IRF

[ACHVH09] S. Alonso, F. J. Cabrerizo, E. Herrera-Viedma, and F. Herrera. h-index: a review focused in its variants, computation and standardization for different scientific fields. Journal of Informetrics, 3(4):273–289, October 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000339.

Abramo:2014:ANS

[AD14] Giovanni Abramo and Ciriaco Andrea D’Angelo. Assessing national strengths and weaknesses in research fields. Journal of Informetrics, 8(3):766–775, July 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000625.

Abramo:2015:EUR

[AD15a] Giovanni Abramo and Ciriaco Andrea D’Angelo. Evaluating university research: Same performance indicator, different rankings. Journal of Informetrics, 9(3):514–525, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715000462.

Abramo:2015:MCT

[AD15b] Giovanni Abramo and Ciriaco Andrea D’Angelo. A methodology to compute the territorial productivity of scientists: the case of Italy. Journal of Informetrics, 9(4):675–685, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715200363.

Abramo:2015:RRI

[AD15c] Giovanni Abramo and Ciriaco Andrea D’Angelo. Ranking research institutions by the number of highly-cited articles per scientist. Journal of Informetrics, 9(4):915–923, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715300109.

Abramo:2015:RBN

[AD15d] Giovanni Abramo and Ciriaco Andrea D’Angelo. The relationship between the number of authors of a publication, its citations and the impact factor of the publishing journal: Evidence from Italy. Journal of Informetrics, 9(4):746–761, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715200557.

Abramo:2016:CUP

[AD16a] Giovanni Abramo and Ciriaco Andrea D’Angelo. A comparison of university performance scores and ranks by MNCS and FSS. Journal of Informetrics, 10(4):889–901, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716301730.

Abramo:2016:FMLa

[AD16b] Giovanni Abramo and Ciriaco Andrea D’Angelo. A farewell to the MNCS and like size-independent indicators. Journal of Informetrics, 10(2):646–651, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S1751157716300955. See comments [GTD16, Zit16] and rejoinder [AD16c].

Abramo:2016:FMLb

[AD16c] Giovanni Abramo and Ciriaco Andrea D’Angelo. A farewell to the MNCS and like size-independent indicators: Rejoinder. Journal of Informetrics, 10(2):679–683, May 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S1751157716300992. See [AD16b, GTD16, Zit16].

Abramo:2017:DYS

[AD17a] Giovanni Abramo and Ciriaco Andrea D’Angelo. Does your surname affect the citability of your publications? Journal of Informetrics, 11(1):121–127, February 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S1751157716302073. See comment [Web17] and response [AD17c].

Abramo:2017:TTF

[AD17b] Giovanni Abramo and Ciriaco Andrea D’Angelo. On tit for tat: Franceschini and Maisano versus ANVUR regarding the Italian research assessment exercise VQR 2011– 2014. Journal of Informetrics, 11(3):783–787, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157717302183. See [FM17a] and com-

ment [BCGM17a].

Abramo:2017:RCS

[AD17c] Giovanni Abramo and Ciriaco Andrea D’Angelo. Response to comments on: “Does your surname affect the citability of your publications?”. Journal of Informetrics, 11(3):855–858, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157717302705. See [Web17].

Abramo:2018:WBC

[AD18] Giovanni Abramo and Ciriaco Andrea D’Angelo. Who benefits from a country’s scientific research? Journal of Informetrics, 12(1):249–258, February 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717302420.

Abramo:2021:NMA

[AD21] Giovanni Abramo and Ciriaco Andrea D’Angelo. A novel methodology to assess the scientific standing of nations at field level. Journal of Informetrics, 14(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719302743.

Adams:2018:IMB

[Ada18] Jonathan Adams. Information and misinformation in bibliometric time-trend analysis. Journal of Informetrics, 12(4):1063–1071, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157718300877.

Abramo:2019:BKF

[ADC19] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Massimiliano Carloni. The balance of knowledge flows. Journal of Informetrics, 13(1):1–9, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302736.

Abramo:2019:AAS

[ADD19a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Authorship analysis of specialized vs diversified research output. Journal of Informetrics, 13(2):564–573, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718303699.

Abramo:2019:WRA

[ADD19b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. When research assessment exercises leave room for opportunistic behavior by the subjects under evaluation. Journal of Informetrics, 13(3):830–840, August 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719300896.

Abramo:2021:KSD

[ADD21a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Knowledge spillovers: Does the geographic proximity effect decay over time? A discipline-level analysis, accounting for cognitive proximity, with and without selfcitations. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115772030095X.

Abramo:2021:RBD

[ADD21b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. On the relation between the degree of internationalization of cited and citing publications: a field level analysis, including and excluding self-citations. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720303965.

Abramo:2021:RGP

[ADD21c] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. The role of geographical proximity in knowledge diffusion, measured by citations to scientific literature. Journal of Informetrics, 14(1):??, February 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719303347.

Abramo:2021:SIP

[ADD21d] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. The scholarly impact of private sector research: a multivariate analysis. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721000626.

Abramo:2022:RSC

[ADD22] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Revealing the scientific comparative advantage of nations: Common and distinctive features. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721001152.

Abramo:2024:MRP

[ADD24] Giovanni Abramo, CiriacoAndrea D’Angelo, and Flavia Di Costa. The moderating role of personal characteristics of authors in the publications’ quality for quantity trade-off. Journal of Informetrics, 18(1):??, February 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000087.

Avenali:2024:HNA

[ADDWD24] Alessandro Avenali, Cinzia Daraio, Simone Di Leo, and Joanna Wolszczak-Derlacz. Heterogeneity of national accounting systems, world-class universities and financial resources: What are the links? Journal of Informetrics, 18 (2):??, May 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000154.

Abramo:2019:PPL

[ADF19] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Giovanni Felici. Predicting publication long-term impact through a combination of early citations and journal impact factor. Journal of Informetrics, 13(1):32–49, February 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718300208.

Abramo:2015:FPV

[ADG15] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Leonardo Grilli. Funnel plots for visualizing uncertainty in the research performance of institutions. Journal of Informetrics, 9(4):954–961, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157715300353.

Abramo:2016:RFP

[ADG16] Giovanni Abramo, Andrea Ciriaco D’Angelo, and Leonardo Grilli. From rankings to funnel plots: the question of accounting for uncertainty when assessing university research performance. Journal of Informetrics, 10(3):854–862, August 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716300736.

Abramo:2021:ECB

[ADG21] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Leonardo Grilli. The effects of citation-based research evaluation schemes on self-citation behavior. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157721000754.

Abramo:2013:CBS

[ADM13a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Gianluca Murgia. The collaboration behaviors of scientists in Italy: a field level analysis. Journal of Informetrics, 7(2):442–454, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157713000114.

Abramo:2013:GDR

[ADM13b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Gianluca Murgia. Gender differences in research collaboration. Journal of Informetrics, 7(4):811–822, October 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000552.

Abramo:2017:RAR

[ADM17] Giovanni Abramo, Andrea Ciriaco D’Angelo, and Gianluca Murgia. The relationship among research productivity, research collaboration, and their determinants. Journal of Informetrics, 11(4):1016–1030, November 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717301475.

Abramo:2013:IAN

[ADR13] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Francesco Rosati. The importance of accounting for the number of co-authors and their order when assessing research performance at the individual level in the life sciences. Journal of Informetrics, 7(1):198–208, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000934.

Abramo:2016:MME

[ADR16] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Francesco Rosati. A methodology to measure the effectiveness of academic recruitment and turnover. Journal of Informetrics, 10(1):31–42, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157715300924.

Abramo:2011:RCM

[ADS11] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Marco Solazzi. Are researchers that collaborate more at the international level top performers? An investigation on the Italian university system. Journal of Informetrics, 5(1):204–213, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000982.

Abramo:2016:RTS

[ADS16] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Anastasiia Soldatenkova. The ratio of top scientists to the academic staff as an indicator of the competitive strength of universities. Journal of Informetrics, 10(2):596–605, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115771630027X.

Abramo:2017:ISP

[ADS17] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Anastasiia Soldatenkova. An investigation on the skewness patterns and fractal nature of research productivity distributions at field and discipline level. Journal of Informetrics, 11(1):324–335, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716303728.

Amjad:2017:SSG

[ADX^{+}17] Tehmina Amjad, Ying Ding, Jian Xu, Chenwei Zhang, Ali Daud, Jie Tang, and Min Song. Standing on the shoulders of giants. Journal of Informetrics, 11(1):307–323, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716302942.

Abramo:2018:CTA

[ADZ18] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Lin Zhang. A comparison of two approaches for measuring interdisciplinary research output: the disciplinary diversity of authors vs the disciplinary diversity of the reference list. Journal of Informetrics, 12(4):1182–1193, November 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718301639.

Ajiferuke:2015:MCR

[AF15] Isola Ajiferuke and Felix Famoye. Modelling count response variables in informetric studies: Comparison among count, linear, and lognormal regression models. Journal of Informetrics, 9(3):499–513, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000498.

Alhoori:2017:RSV

[AF17a] Hamed Alhoori and Richard Furuta. Recommendation of scholarly venues based on dynamic user interests. Journal of Informetrics, 11(2):553–563, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303406.

Araujo:2017:SSG

[AF17b] Tanya Araújo and Elsa Fontainha. The specific shapes of gender imbalance in scientific authorships: a network approach. Journal of Informetrics, 11(1):88–102, February 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302140.

Amon:2022:IAB

[AH22] Julian Amon and Kurt Hornik. Is it all bafflegab? — linguistic and meta characteristics of research articles in prestigious economics journals. Journal of Informetrics, 16(2): ??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722000360.

Abbasi:2012:BCD

[AHL12] Alireza Abbasi, Liaquat Hossain, and Loet Leydesdorff. Betweenness centrality as a driver of preferential attachment in the evolution of research collaboration networks. Journal of Informetrics, 6(3):403–412, July 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771200003X.

Albarran:2017:HVA

[AHRCV17] Pedro Albarrán, Carmen Herrero, Javier Ruiz-Castillo, and Antonio Villar. The Herrero–Villar approach to citation impact. Journal of Informetrics, 11(2):625–640, May 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303716.

Abbasi:2013:RIS

[AJ13] Alireza Abbasi and Ali Jaafari. Research impact and scholars’ geographical diversity. Journal of Informetrics, 7(3): 683–692, July 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000400.

Adachi:2015:FAE

[AK15] Tsuyoshi Adachi and Takumi Kongo. Further axiomatizations of Egghe’s g-index. Journal of Informetrics, 9(4):839– 844, October 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715200223.

An:2018:DTI

[AKML18] Jaehyeong An, Kyuwoong Kim, Letizia Mortara, and Sungjoo Lee. Deriving technology intelligence from patents: Preposition-based semantic analysis. Journal of Informetrics, 12(1):217–236, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S1751157716303777.

Agouzal:2008:RBM

[AL08] Abdelatif Agouzal and Thierry Lafouge. On the relation between the Maximum Entropy Principle and the Principle of Least Effort: the continuous case. Journal of Informetrics, 2(1):75–88, January 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157707000727.

Agouzal:2017:RET

[AL17] Abdellatif Agouzal and Thierry Lafouge. A remarkable example in three-dimensional informetrics. The geometric law: Distribution of use or distribution of structure? Journal of Informetrics, 11(4):1003–1015, November 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717300238.

Agouzal:2024:RBP

[ALB24] Abdellatif Agouzal, Thierry Lafouge, and Marc Bertin. Relationship between the principle of least effort and the average cost of information in a Zipfian context. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723001037.

Ao:2023:SCP

[ALR^{+}23] Weiyi Ao, Dongqing Lyu, Xuanmin Ruan, Jiang Li, and Ying Cheng. Scientific creativity patterns in scholars’ academic careers: Evidence from PubMed. Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000883.

Alsudais:2021:CCP

[Als21] Abdulkareem Alsudais. In-code citation practices in open research software libraries. Journal of Informetrics, 15(2): ??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721000109.

Ajiferuke:2011:WRD

[ALW11] Isola Ajiferuke, Kun Lu, and Dietmar Wolfram. Who are the research disciples of an author? Examining publication recitation and oeuvre citation exhaustivity. Journal of Informetrics, 5(2):292–302, April 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000022.

An:2021:IPS

[ALX^{+}21] Xin An, Jinghong Li, Shuo Xu, Liang Chen, and Wei Sun. An improved patent similarity measurement based on entities and semantic relations. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721000067.

Angelou:2021:ETR

[AMKA21] K. Angelou, M. Maragakis, K. Kosmidis, and P. Argyrakis. The evolution of triangular research and innovation collaborations in the European area. Journal of Informetrics, 15 (3):??, August 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000638.

Andersen:2018:GSW

[AN18] Jens Peter Andersen and Mathias Wullum Nielsen. Google Scholar and Web of Science: Examining gender differences in citation coverage across five scientific disciplines. Journal of Informetrics, 12(3):950–959, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718300282.

Andersen:2017:ETC

[And17] Jens Peter Andersen. An empirical and theoretical critique of the Euclidean index. Journal of Informetrics, 11(2):455–465, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716303613.

Anonymous:2007:EBa

[Ano07a] Anonymous. Editorial Board. Journal of Informetrics, 1 (2):co2, April 2007. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157707000399.

Anonymous:2007:EBb

[Ano07b] Anonymous. Editorial Board. Journal of Informetrics, 1 (3):co2, July 2007. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157707000491.

Anonymous:2007:EBc

[Ano07c] Anonymous. Editorial Board. Journal of Informetrics, 1 (4):co2, ???? 2007. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157707000648.

Anonymous:2007:IEB

[Ano07d] Anonymous. IFC: Editorial Board. Journal of Informetrics, 1 (1):co2, January 2007. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157706000149.

Anonymous:2008:EBa

[Ano08a] Anonymous. Editorial Board. Journal of Informetrics, 2(1): co2, January 2008. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157708000047.

Anonymous:2008:EBb

[Ano08b] Anonymous. Editorial Board. Journal of Informetrics, 2 (2):co2, April 2008. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157708000175.

Anonymous:2008:EBc

[Ano08c] Anonymous. Editorial Board. Journal of Informetrics, 2(4): co2, October 2008. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157708000576.

Anonymous:2008:IEB

[Ano08d] Anonymous. IFC Editorial Board. Journal of Informetrics, 2(3):co2, July 2008. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157708000370.

Anonymous:2009:EBa

[Ano09a] Anonymous. Editorial Board. Journal of Informetrics, 3(1): co2, January 2009. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157709000030.

Anonymous:2009:EBb

[Ano09b] Anonymous. Editorial Board. Journal of Informetrics, 3 (2):co2, April 2009. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157709000182.

Anonymous:2009:EBc

[Ano09c] Anonymous. Editorial Board. Journal of Informetrics, 3 (3):co2, July 2009. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157709000455.

Anonymous:2009:EBd

[Ano09d] Anonymous. Editorial Board. Journal of Informetrics, 3(4): co2, October 2009. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157709000649.

Anonymous:2010:EBa

[Ano10a] Anonymous. Editorial Board. Journal of Informetrics, 4(1): co2, January 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115770900090X.

Anonymous:2010:EBb

[Ano10b] Anonymous. Editorial Board. Journal of Informetrics, 4 (2):co2, April 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000155.

Anonymous:2010:EBc

[Ano10c] Anonymous. Editorial Board. Journal of Informetrics, 4 (3):co2, July 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000489.

Anonymous:2010:EBd

[Ano10d] Anonymous. Editorial Board. Journal of Informetrics, 4(4): co2, October 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000763.

Anonymous:2010:LR

[Ano10e] Anonymous. List of reviewers 2009. Journal of Informetrics, 4(2):137, April 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000209.

Anonymous:2011:EBa

[Ano11a] Anonymous. Editorial Board. Journal of Informetrics, 5(1): co2, January 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710001069.

Anonymous:2011:EBb

[Ano11b] Anonymous. Editorial Board. Journal of Informetrics, 5 (2):co2, April 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711000204.

Anonymous:2011:EBc

[Ano11c] Anonymous. Editorial Board. Journal of Informetrics, 5 (3):co2, July 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771100054X.

Anonymous:2011:EBd

[Ano11d] Anonymous. Editorial Board. Journal of Informetrics, 5(4): co2, October 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711000903.

Anonymous:2011:LR

[Ano11e] Anonymous. List of reviewers 2010. Journal of Informetrics, 5(2):231, April 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711000162.

Anonymous:2012:EBa

[Ano12a] Anonymous. Editorial Board. Journal of Informetrics, 6(1): co2, January 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711001106.

Anonymous:2012:EBb

[Ano12b] Anonymous. Editorial Board. Journal of Informetrics, 6 (2):co2, April 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712000156.

Anonymous:2012:EBc

[Ano12c] Anonymous. Editorial Board. Journal of Informetrics, 6 (3):co2, July 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712000351.

Anonymous:2012:EBd

[Ano12d] Anonymous. Editorial Board. Journal of Informetrics, 6(4): co2, October 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712000685.

Anonymous:2012:LR

[Ano12e] Anonymous. List of reviewers 2011. Journal of Informetrics, 6(2):iii, April 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771200020X.

Anonymous:2013:EBa

[Ano13a] Anonymous. Editorial Board. Journal of Informetrics, 7(1): co2, January 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712001083.

Anonymous:2013:EBb

[Ano13b] Anonymous. Editorial Board. Journal of Informetrics, 7 (2):co2, April 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000308.

Anonymous:2013:EBc

[Ano13c] Anonymous. Editorial Board. Journal of Informetrics, 7 (3):co2, July 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000588.

Anonymous:2013:EBd

[Ano13d] Anonymous. Editorial Board. Journal of Informetrics, 7(4): co2, October 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000990.

Anonymous:2013:LR

[Ano13e] Anonymous. List of reviewers. Journal of Informetrics, 7 (2):iii, April 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000369.

Anonymous:2014:IFCa

[Ano14a] Anonymous. Inside front cover: Editorial Board. Journal of Informetrics, 8(1):co2, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000200.

Anonymous:2014:IFCb

[Ano14b] Anonymous. Inside front cover: Editorial Board. Journal of Informetrics, 8(2):co2, April 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714000327.

Anonymous:2014:IFCc

[Ano14c] Anonymous. Inside front cover: Editorial Board. Journal of Informetrics, 8(3):ifc, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714000716.

Anonymous:2014:IFCd

[Ano14d] Anonymous. Inside front cover: Editorial Board. Journal of Informetrics, 8(4):ifc, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771400100X.

Anonymous:2014:LR

[Ano14e] Anonymous. List of reviewers. Journal of Informetrics, 8 (2):iii, April 2014. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157714000376.

Anonymous:2015:IFCa

[Ano15a] Anonymous. Inside front cover: Editorial Board. Journal of Informetrics, 9(1):ifc, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000115.

Anonymous:2015:IFCb

[Ano15b] Anonymous. Inside front cover: Editorial Board. Journal of Informetrics, 9(2):ifc, April 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157715000346.

Anonymous:2015:IFCc

[Ano15c] Anonymous. Inside front cover: Editorial Board. Journal of Informetrics, 9(3):ifc, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157715000589.

Anonymous:2015:IFC

[Ano15d] Anonymous. Inside front cover: Editorial board. Journal of Informetrics, 9(4):ifc, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771500067X.

Anonymous:2015:LR

[Ano15e] Anonymous. List of reviewers. Journal of Informetrics, 9 (2):iii, April 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715000395.

Anonymous:2016:IFCa

[Ano16a] Anonymous. Inside front cover: Editorial board. Journal of Informetrics, 10(1):ifc, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716000055.

Anonymous:2016:IFCb

[Ano16b] Anonymous. Inside front cover: Editorial board. Journal of Informetrics, 10(2):ifc, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S175115771630133X.

Anonymous:2016:IFCc

[Ano16c] Anonymous. Inside front cover: Editorial Board. Journal of Informetrics, 10(3):ifc, August 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302164.

Anonymous:2016:IFCd

[Ano16d] Anonymous. Inside front cover: Editorial board. Journal of Informetrics, 10(4):ifc, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303224.

Anonymous:2017:IFCa

[Ano17a] Anonymous. Inside front cover: Editorial board. Journal of Informetrics, 11(1):ifc, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717300573.

Anonymous:2017:IFCb

[Ano17b] Anonymous. Inside front cover: Editorial board. Journal of Informetrics, 11(2):ifc, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157717301852.

Anonymous:2017:IFCc

[Ano17c] Anonymous. Inside front cover: Editorial board. Journal of Informetrics, 11(3):ifc, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717303036.

Anonymous:2017:IFCd

[Ano17d] Anonymous. Inside front cover: Editorial board. Journal of Informetrics, 11(4):ifc, November 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S1751157717304133.

Anonymous:2017:PN

[Ano17e] Anonymous. Publisher’s note. Journal of Informetrics, 11(4): 976, November 2017. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157717301220.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board. Journal of Informetrics, 12 (1):ii, February 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157718300592.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board. Journal of Informetrics, 12 (2):ii, May 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157718302001.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board. Journal of Informetrics, 12 (3):ii, August 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157718303353.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Journal of Informetrics, 12(4): ii, November 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157718304590.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board. Journal of Informetrics, 13 (1):ii, February 2019. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157719300549.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. Journal of Informetrics, 13 (2):ii, May 2019. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157719301749.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. Journal of Informetrics, 13 (3):ii, August 2019. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157719303189.

Anonymous:2021:EBa

[Ano21a] Anonymous. Editorial Board. Journal of Informetrics, 13(4): ??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157719304705.

Anonymous:2021:EBb

[Ano21b] Anonymous. Editorial Board. Journal of Informetrics, 14(1): ??, February 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157720301097.

Anonymous:2021:EBc

[Ano21c] Anonymous. Editorial Board. Journal of Informetrics, 14 (2):??, May 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157720303084.

Anonymous:2021:EBd

[Ano21d] Anonymous. Editorial Board. Journal of Informetrics, 14 (3):??, August 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157720305198.

Anonymous:2021:EBe

[Ano21e] Anonymous. Editorial Board. Journal of Informetrics, 14(4): ??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157720306349.

Anonymous:2021:EBf

[Ano21f] Anonymous. Editorial Board. Journal of Informetrics, 15(1): ??, February 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000201.

Anonymous:2021:EBg

[Ano21g] Anonymous. Editorial Board. Journal of Informetrics, 15 (2):??, May 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000535.

Anonymous:2021:EBh

[Ano21h] Anonymous. Editorial Board. Journal of Informetrics, 15 (3):??, August 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000821.

Anonymous:2021:EBi

[Ano21i] Anonymous. Editorial Board. Journal of Informetrics, 15(4): ??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721001024.

Anonymous:2022:A

[Ano22a] Anonymous. August 2022. Journal of Informetrics, 16(3): ??, August 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

Anonymous:2022:EBd

[Ano22b] Anonymous. Editorial Board. Journal of Informetrics, 16(1): ??, February 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722000232.

Anonymous:2022:EBe

[Ano22c] Anonymous. Editorial Board. Journal of Informetrics, 16 (2):??, May 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722000530.

Anonymous:2022:EBf

[Ano22d] Anonymous. Editorial Board. Journal of Informetrics, 16 (3):??, August 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722000797.

Anonymous:2022:EBg

[Ano22e] Anonymous. Editorial Board. Journal of Informetrics, 16(4): ??, November 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722001158.

Anonymous:2022:M

[Ano22f] Anonymous. May 2022. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

Anonymous:2022:N

[Ano22g] Anonymous. November 2022. Journal of Informetrics, 16(4): ??, November 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

Anonymous:2023:A

[Ano23a] Anonymous. August 2023. Journal of Informetrics, 17(3): ??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

Anonymous:2023:EBa

[Ano23b] Anonymous. Editorial Board. Journal of Informetrics, 17(1): ??, February 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000123.

Anonymous:2023:EBb

[Ano23c] Anonymous. Editorial Board. Journal of Informetrics, 17 (2):??, May 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000421.

Anonymous:2023:EBc

[Ano23d] Anonymous. Editorial Board. Journal of Informetrics, 17 (3):??, August 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000652.

Anonymous:2023:EBd

[Ano23e] Anonymous. Editorial Board. Journal of Informetrics, 17(4): ??, November 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000974.

Anonymous:2023:F

[Ano23f] Anonymous. February 2023. Journal of Informetrics, 17(1): ??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

Anonymous:2023:M

[Ano23g] Anonymous. May 2023. Journal of Informetrics, 17(2):??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

Anonymous:2023:N

[Ano23h] Anonymous. November 2023. Journal of Informetrics, 17(4): ??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

Anonymous:2024:A

[Ano24a] Anonymous. August 2024. Journal of Informetrics, 18(3): ??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

Anonymous:2024:EBa

[Ano24b] Anonymous. Editorial Board. Journal of Informetrics, 18(1): ??, February 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000245.

Anonymous:2024:EBb

[Ano24c] Anonymous. Editorial Board. Journal of Informetrics, 18 (2):??, May 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115772400049X.

Anonymous:2024:EBc

[Ano24d] Anonymous. Editorial Board. Journal of Informetrics, 18 (3):??, August 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000907.

Anonymous:2024:EBd

[Ano24e] Anonymous. Editorial Board. Journal of Informetrics, 18(4): ??, November 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724001342.

Anonymous:2024:F

[Ano24f] Anonymous. February 2024. Journal of Informetrics, 18(1): ??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

Anonymous:2024:M

[Ano24g] Anonymous. May 2024. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

Anonymous:2024:N

[Ano24h] Anonymous. November 2024. Journal of Informetrics, 18(4): ??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

Antonoyiannakis:2018:IFC

[Ant18] Manolis Antonoyiannakis. Impact Factors and the Central Limit Theorem: Why citation averages are scale dependent. Journal of Informetrics, 12(4):1072–1088, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718301238.

Ante:2022:RBR

[Ant22] Lennart Ante. The relationship between readability and scientific impact: Evidence from emerging technology discourses. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722000049.

Amancio:2012:TFM

[AOdFC12] Diego Raphael Amancio, Osvaldo Novais Oliveira, Jr., and Luciano da Fontoura Costa. Three-feature model to reproduce the topology of citation networks and the effects from authors’ visibility on their h-index. Journal of Informetrics, 6(3):427–434, July 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000211.

Albarran:2011:HLI

[AORC11a] Pedro Albarrán, Ignacio Ortuño, and Javier Ruiz-Castillo. High- and low-impact citation measures: Empirical applications. Journal of Informetrics, 5(1):122–145, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115771000088X. See corrigen-

dum [AORC12a].

Albarran:2011:MLH

[AORC11b] Pedro Albarrán, Ignacio Ortuño, and Javier Ruiz-Castillo. The measurement of low- and high-impact in citation distributions: Technical results. Journal of Informetrics, 5(1): 48–63, January 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000829. See corri-

gendum [AORC12b].

Albarran:2012:CHL

[AORC12a] Pedro Albarrán, Ignacio Ortuño, and Javier Ruiz-Castillo. Corrigendum to “High- and low-impact citation measures: Empirical applications” J. Informetrics 5 (2011) 122–145. Journal of Informetrics, 6(4):469, October 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000296. See [AORC11a].

Albarran:2012:CML

[AORC12b] Pedro Albarrán, Ignacio Ortuño, and Javier Ruiz-Castillo. Corrigendum to “The measurement of low- and high-impact in citation distributions: Technical results” J. Informetrics 5 (2011) 48–63. Journal of Informetrics, 6(4):515, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157712000077. See [AORC11b].

Almeida:2009:SIS

[APF09] J. A. S. Almeida, A. A. C. C. Pais, and S. J. Formosinho. Science indicators and science patterns in Europe. Journal of Informetrics, 3(2):134–142, April 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115770900008X.

Aagaard:2017:SCA

[AS17] Kaare Aagaard and Jesper W. Schneider. Some considerations about causes and effects in studies of performancebased research funding systems. Journal of Informetrics, 11(3):923–926, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1751157717301967.

Anil:2021:ECI

[AS21] Akash Anil and Sanasam Ranbir Singh. Effect of class imbalance in heterogeneous network embedding: an empirical study. Journal of Informetrics, 14(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719301051.

Aksnes:2012:RNR

[ASG12] Dag W. Aksnes, Jesper W. Schneider, and Magnus Gunnarsson. Ranking national research systems by citation indicators. A comparative analysis using whole and fractionalised counting methods. Journal of Informetrics, 6(1):36–43, January 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711000745.

Akbaritabar:2022:SRC

[ASS22] Aliakbar Akbaritabar, Dimity Stephen, and Flaminio Squazzoni. A study of referencing changes in preprint-publication pairs across multiple fields. Journal of Informetrics, 16(2): ??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722000104.

Aman:2024:ARP

[AvdB24] Valeria Aman and Peter van den Besselaar. Authorship regulations in performance-based funding systems and publication behaviour — a case study of German medical faculties. Journal of Informetrics, 18(2):??, May 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000130.

Ahlgren:2014:CBC

[AW14] Per Ahlgren and Ludo Waltman. The correlation between citation-based and expert-based assessments of publication channels: SNIP and SJR vs. Norwegian quality assessments. Journal of Informetrics, 8(4):985–996, October 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000911.

An:2014:VTA

[AYL14] Lu An, Chuanming Yu, and Gang Li. Visual topical analysis of Chinese and American Library and Information Science research institutions. Journal of Informetrics, 8(1):217–233, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157713001119.

Akhmat:2014:RBE

[AZS^{+}14] Ghulam Akhmat, Khalid Zaman, Tan Shukui, Yasir Javed, and Muhammad Mushtaq Khan. Relationship between educational indicators and research outcomes in a panel of top twenty nations: Windows of opportunity. Journal of Informetrics, 8(2):349–361, April 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771400008X.

Bras-Amoros:2011:BIB

[BADFT11] Maria Bras-Amorós, Josep Domingo-Ferrer, and Vicenç Torra. A bibliometric index based on the collaboration distance between cited and citing authors. Journal of Informetrics, 5(2):248–264, April 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000970.

Bordons:2015:RBR

[BAGADF15] María Bordons, Javier Aparicio, Borja González-Albo, and Adrián A. Díaz-Faes. The relationship between the research performance of scientists and their position in co-authorship networks in three fields. Journal of Informetrics, 9(1):135– 144, January 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157714001138.

Bornmann:2018:NEC

[BAL18] Lutz Bornmann, Jonathan Adams, and Loet Leydesdorff. The negative effects of citing with a national orientation in terms of recognition: National and international citations in natural-sciences papers from Germany, The Netherlands, and the UK. Journal of Informetrics, 12(3):931–949, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718300336.

Baccini:2011:STN

[BB11] Alberto Baccini and Lucio Barabesi. Seats at the table: the network of the editorial boards in information and library science. Journal of Informetrics, 5(3):382–391, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157711000137.

Baccini:2022:SNF

[BBB^{+}22] Federica Baccini, Lucio Barabesi, Alberto Baccini, Mahdi Khelfaoui, and Yves Gingras. Similarity network fusion for scholarly journals. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721000973.

Bertoli-Barsotti:2024:EII

[BBGSZS24a] Lucio Bertoli-Barsotti, Marek Gagolewski, Grzegorz Siudem, and Barbara Zogała-Siudem. Equivalence of inequality indices in the three-dimensional model of informetric impact. Journal of Informetrics, 18(4):??, November 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000798.

Bertoli-Barsotti:2024:GSL

[BBGSZS24b] Lucio Bertoli-Barsotti, Marek Gagolewski, Grzegorz Siudem, and Barbara Zogała-Siudem. Gini-stable Lorenz curves and their relation to the generalised Pareto distribution. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000129.

Bertoli-Barsotti:2015:FI

[BBL15] Lucio Bertoli-Barsotti and Tommaso Lando. On a formula for the h-index. Journal of Informetrics, 9(4):762–776, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715300572.

Bertoli-Barsotti:2019:HMR

[BBL19] Lucio Bertoli-Barsotti and Tommaso Lando. How mean rank and mean size may determine the generalised Lorenz curve: With application to citation analysis. Journal of Informetrics, 13(1):387–396, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771830470X.

Baccini:2012:SII

[BBMP12] A. Baccini, L. Barabesi, M. Marcheselli, and L. Pratelli. Statistical inference on the h-index with an application to topscientist performance. Journal of Informetrics, 6(4):721–728, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000600.

Bonaccorsi:2021:RPU

[BBS21] Andrea Bonaccorsi, Paola Belingheri, and Luca Secondi. The research productivity of universities. A multilevel and multidisciplinary analysis on European institutions. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720306465.

Bonaccorsi:2016:NRU

[BC16] Andrea Bonaccorsi and Tindaro Cicero. Nondeterministic ranking of university departments. Journal of Informetrics, 10(1):224–237, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157716000031.

Benedetto:2017:CCS

[BCGM17a] Sergio Benedetto, Daniele Checchi, Andrea Graziosi, and Marco Malgarini. Comments on the correspondence “On tit for tat: Franceschini and Maisano versus ANVUR regarding the Italian research assessment exercise VQR 2011–2014”, J. Informetr., 11 (2017), 783–787. Journal of Informetrics, 11(3):838–840, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717302481. See [FM17a, AD17b].

Benedetto:2017:CPS

[BCGM17b] Sergio Benedetto, Daniele Checchi, Andrea Graziosi, and Marco Malgarini. Comments on the paper “Critical remarks on the Italian assessment exercise”, Journal of Informetrics, 11 (2017), pp. 337–357. Journal of Informetrics, 11(2):622– 624, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157717300949. See [FM17a]

and rejoinder [FM17b].

Bornmann:2007:CVP

[BD07a] Lutz Bornmann and Hans-Dieter Daniel. Convergent validation of peer review decisions using the h index: Extent of and reasons for type I and type II errors. Journal of Informetrics, 1(3):204–213, July 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157707000302.

Bornmann:2007:GSE

[BD07b] Lutz Bornmann and Hans-Dieter Daniel. Gatekeepers of science — effects of external reviewers’ attributes on the assessments of fellowship applications. Journal of Informetrics, 1(1):83–91, January 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157706000113.

Bornmann:2009:ETT

[BD09] Lutz Bornmann and Hans-Dieter Daniel. Extent of type I and type II errors in editorial decisions: a case study on Angewandte Chemie International Edition. Journal of Informetrics, 3(4):348–352, October 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000406.

Bornmann:2010:CSM

[BD10a] Lutz Bornmann and Hans-Dieter Daniel. Citation speed as a measure to predict the attention an article receives: an investigation of the validity of editorial decisions at Angewandte Chemie International Edition. Journal of Informetrics, 4(1): 83–88, January 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157709000728.

Bornmann:2010:CSI

[BD10b] Lutz Bornmann and Hans-Dieter Daniel. The citation speed index: a useful bibliometric indicator to add to the h index. Journal of Informetrics, 4(3):444–446, July 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000295.

Bornmann:2016:CRM

[BD16] Lutz Bornmann and Hans-Dieter Daniel. Count regression models in informetrics. Journal of Informetrics, 10(1):29– 30, February 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715301644.

Bornmann:2011:SII

[BdMA11] Lutz Bornmann and Felix de Moya Anegon. Some interesting insights from aggregated data published in the World Report SIR 2010. Journal of Informetrics, 5(3):486–488, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711000368.

Bornmann:2012:NEI

[BdMAL12] Lutz Bornmann, Felix de Moya Anegón, and Loet Leydesdorff. The new Excellence Indicator in the World Report of the SCImago Institutions Rankings 2011. Journal of Informetrics, 6(2):333–335, April 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711001052.

Bongioanni:2014:QMC

[BDR14] Irene Bongioanni, Cinzia Daraio, and Giancarlo Ruocco. A quantitative measure to compare the disciplinary profiles of research systems and their evolution over time. Journal of Informetrics, 8(3):710–727, July 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000595.

Bravo:2018:HCN

[BFM^{+}18] Giangiacomo Bravo, Mike Farjam, Francisco Grimaldo Moreno, Aliaksandr Birukou, and Flaminio Squazzoni. Hidden connections: Network effects on editorial decisions in four computer science journals. Journal of Informetrics, 12(1):101–112, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S1751157717303176.

Bollen:2011:HWT

[BFS11] Johan Bollen, Geoffrey Fox, and Prashant Raj Singhal. How and where the TeraGrid supercomputing infrastructure benefits science. Journal of Informetrics, 5(1):114–121, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157710000878.

Buljan:2024:DAC

[BGCG^{+}24] Ivan Buljan, Daniel Garcia-Costa, Francisco Grimaldo, Richard A. Klein, Marjan Bakker, and Ana Marušić. Development and application of a comprehensive glossary for the identification of statistical and methodological concepts in peer review reports. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157724000683.

Bianchi:2022:MER

[BGCGS22] Federico Bianchi, Daniel García-Costa, Francisco Grimaldo, and Flaminio Squazzoni. Measuring the effect of reviewers on manuscript change: a study on a sample of submissions to Royal Society journals (2006–2017). Journal of Informetrics, 16(3):??, August 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S1751157722000682. See corrigendum [BGCGS23].

Bianchi:2023:CME

[BGCGS23] Federico Bianchi, Daniel Garcia-Costa, Francisco Grimaldo, and Flaminio Squazzoni. Corrigendum to ‘Measuring the effect of reviewers on manuscript change: a study on a sample of submissions to Royal Society journals (2006–2017)’ Journal of Informetrics, Volume 16, Issue 3, 2022, 101316. Journal of Informetrics, 17(1):??, February 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000876. See [BGCGS22].

Beirlant:2007:SRO

[BGCL07] Jan Beirlant, Wolfgang Glänzel, An Carbonez, and Herlinde Leemans. Scoring research output using statistical quantile plotting. Journal of Informetrics, 1(3):185–192, July 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157707000454.

Bornmann:2023:EUR

[BGdMAW23] Lutz Bornmann, Sabine Gralka, Félix de Moya Anegón, and Klaus Wohlrabe. Efficiency of universities and researchfocused institutions worldwide: the introduction of a new input indicator reflecting institutional staff numbers. Journal of Informetrics, 17(2):??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157723000251.

Barrios:2013:ISF

[BGGB13] Maite Barrios, Georgina Guilera, and Juana Gómez-Benito. Impact and structural features of meta-analytical studies, standard articles and reviews in psychology: Similarities and differences. Journal of Informetrics, 7(2):478–486, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115771300014X.

Banshal:2022:PLA

[BGLS22] Sumit Kumar Banshal, Solanki Gupta, Hiran H. Lathabai, and Vivek Kumar Singh. Power laws in altmetrics: an empirical analysis. Journal of Informetrics, 16(3):??, August 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115772200061X.

Benito:2021:EIC

[BGR21] M. Benito, P. Gil, and R. Romera. Evaluating the influence of country characteristics on the Higher Education System Rankings’ progress. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719301282.

Bianchi:2019:IVR

[BGS19] Federico Bianchi, Francisco Grimaldo, and Flaminio Squazzoni. The F^{3}-index. Valuing reviewers for scholarly journals. Journal of Informetrics, 13(1):78–86, February 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718301275.

Bayindir:2021:GTA

[BGS21] Esra Eren Bayindir, Mehmet Yigit Gurdal, and Ismail Saglam. A game theoretic approach to peer review of grant proposals. Journal of Informetrics, 13(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719300355.

Bornmann:2022:SIU

[BGT22] Lutz Bornmann, Christian Ganser, and Alexander Tekles. Simulation of the h index use at university departments within the bibliometrics-based heuristics framework: Can the indicator be used to compare individual researchers? Journal of Informetrics, 16(1):??, February 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721001085.

Bornmann:2015:WPU

[BH15] Lutz Bornmann and Robin Haunschild. Which people use which scientific papers? An evaluation of data from F1000 and Mendeley. Journal of Informetrics, 9(3):477–487, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715000450.

Bornmann:2016:CSN

[BH16a] Lutz Bornmann and Robin Haunschild. Citation score normalized by cited references (CSNCR): the introduction of a new citation impact indicator. Journal of Informetrics, 10(3):875–887, August 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157716301535.

Bornmann:2016:ERP

[BH16b] Lutz Bornmann and Robin Haunschild. Efficiency of research performance and the glass researcher. Journal of Informetrics, 10(2):652–654, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716300967.

Bornmann:2016:NMR

[BH16c] Lutz Bornmann and Robin Haunschild. Normalization of Mendeley reader impact on the reader- and paper-side: a comparison of the mean discipline normalized reader score (MDNRS) with the mean normalized reader score (MNRS) and bare reader counts. Journal of Informetrics, 10(3):776– 788, August 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715302042.

Bornmann:2018:NZI

[BH18] Lutz Bornmann and Robin Haunschild. Normalization of zero-inflated data: an empirical analysis of a new indicator family and its use with altmetrics data. Journal of Informetrics, 12(3):998–1011, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717303978.

Bornmann:2022:EAR

[BH22] Lutz Bornmann and Robin Haunschild. Empirical analysis of recent temporal dynamics of research fields: Annual publications in chemistry and related areas as an example. Journal of Informetrics, 16(2):??, May 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000050.

Bornmann:2024:PWI

[BH24] Lutz Bornmann and Robin Haunschild. The Prize Winner Index (PWI): a proposal for an indicator based on scientific prizes. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724000737.

Bornmann:2019:DAA

[BHA19] Lutz Bornmann, Robin Haunschild, and Jonathan Adams. Do altmetrics assess societal impact in a comparable way to case studies? An empirical test of the convergent validity of altmetrics based on data from the UK research excellence framework (REF). Journal of Informetrics, 13(1):325–340, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718302700.

Bonaccorsi:2017:SSA

[BHCH17] Andrea Bonaccorsi, Peter Haddawy, Tindaro Cicero, and Saeed-Ul Hassan. The solitude of stars. An analysis of the distributed excellence model of European universities.

Journal of Informetrics, 11(2):435–454, May 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303091.

Borner:2007:TVS

[BHH^{+}07] Katy Börner, Elisha Hardy, Bruce Herr, Todd Holloway, and W. Bradford Paley. Taxonomy visualization in support of the semi-automatic validation and optimization of organizational schemas. Journal of Informetrics, 1(3):214–225, July 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157707000375.

Bornmann:2018:MIZ

[BHM18] Lutz Bornmann, Robin Haunschild, and Rüdiger Mutz. MHq indicators for zero-inflated count data — a response to Smolinsky and Marx (2018). Journal of Informetrics, 12(3):1012–1014, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157718303109. See [SM18, BHM19].

Bornmann:2019:MIZ

[BHM19] Lutz Bornmann, Robin Haunschild, and Rüdiger Mutz. MHq indicators for zero-inflated count data — a response to the comment by Smolinsky (in press). Journal of Informetrics, 13(1):464–465, February 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771930029X. See [BHM18].

Bornmann:2021:SCF

[BHM21] Lutz Bornmann, Robin Haunschild, and Rüdiger Mutz. Should citations be field-normalized in evaluative bibliometrics? An empirical analysis based on propensity score matching. Journal of Informetrics, 14(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720303564.

Bar-Ilan:2008:IBC

[BI08] Judit Bar-Ilan. Informetrics at the beginning of the 21st century — a review. Journal of Informetrics, 2(1):1–52, January 2008. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157707000740.

Bar-Ilan:2010:RIL

[BI10] Judit Bar-Ilan. Rankings of information and library science journals by JIF and by h-type indices. Journal of Informetrics, 4(2):141–147, April 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000844.

Bar-Ilan:2007:SMC

[BILL07] Judit Bar-Ilan, Mark Levene, and Ayelet Lin. Some measures for comparing citation databases. Journal of Informetrics, 1(1):26–34, January 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157706000058.

Boyack:2008:MST

[BK08] Kevin W. Boyack and Richard Klavans. Measuring sciencetechnology interaction using rare inventor-author names. Journal of Informetrics, 2(3):173–182, July 2008. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157708000151.

Boyack:2014:ICN

[BK14] Kevin W. Boyack and Richard Klavans. Including cited nonsource items in a large-scale map of science: What difference does it make? Journal of Informetrics, 8(3):569–580, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157714000406.

Bettencourt:2009:SDT

[BKK09] Luís M. A. Bettencourt, David I. Kaiser, and Jasleen Kaur. Scientific discovery and topological transitions in collaboration networks. Journal of Informetrics, 3(3):210–221, July 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157709000285.

Bornmann:2012:WBP

[BL12] Lutz Bornmann and Loet Leydesdorff. Which are the best performing regions in information science in terms of highly cited papers? Some improvements of our previous mapping approaches. Journal of Informetrics, 6(2):336–345, April 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157711001015.

Bornmann:2013:VAB

[BL13] Lutz Bornmann and Loet Leydesdorff. The validation of (advanced) bibliometric indicators through peer assessments: a comparative study using data from InCites and F1000. Journal of Informetrics, 7(2):286–291, April 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771200106X.

Bornmann:2014:MNM

[BL14] Lutz Bornmann and Loet Leydesdorff. On the meaningful and non-meaningful use of reference sets in bibliometrics. Journal of Informetrics, 8(1):273–275, January 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713001156.

Bornmann:2015:DQC

[BL15] Lutz Bornmann and Loet Leydesdorff. Does quality and content matter for citedness? A comparison with para-textual factors and over time. Journal of Informetrics, 9(3):419–429, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715000425.

Bornmann:2017:SCI

[BL17] Lutz Bornmann and Loet Leydesdorff. Skewness of citation impact data and covariates of citation distributions: a large-scale empirical analysis based on Web of Science data.

Journal of Informetrics, 11(1):164–175, February 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303108.

Ba:2021:NAM

[BL21] Zhichao Ba and Zhentao Liang. A novel approach to measuring science-technology linkage: From the perspective of knowledge network coupling. Journal of Informetrics, 15(3): ??, August 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000389.

Bornmann:2013:UPP

[BLM13] Lutz Bornmann, Loet Leydesdorff, and Rüdiger Mutz. The use of percentiles and percentile rank classes in the analysis of bibliometric data: Opportunities and limits. Journal of Informetrics, 7(1):158–165, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000831.

Bornmann:2010:MES

[BLV10] Lutz Bornmann, Loet Leydesdorff, and Peter Van den Besselaar. A meta-evaluation of scientific research proposals: Different ways of comparing rejected to awarded applications. Journal of Informetrics, 4(3):211–220, July 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000789.

Bornmann:2013:WPB

[BLW13] Lutz Bornmann, Loet Leydesdorff, and Jian Wang. Which percentile-based approach should be preferred for calculating normalized citation impact values? An empirical comparison of five approaches including a newly developed citation-rank approach (p100). Journal of Informetrics, 7(4):933–944, October 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157713000758.

Bornmann:2014:HIP

[BLW14] Lutz Bornmann, Loet Leydesdorff, and Jian Wang. How to improve the prediction based on citation impact percentiles for years shortly after the publication date? Journal of Informetrics, 8(1):175–180, January 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713001065.

Bornmann:2011:MEG

[BLWSE11] Lutz Bornmann, Loet Leydesdorff, Christiane Walch-Solimena, and Christoph Ettl. Mapping excellence in the geography of science: an approach based on Scopus data. Journal of Informetrics, 5(4):537–546, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000617.

Bouyssou:2010:CBR

[BM10] D. Bouyssou and T. Marchant. Consistent bibliometric rankings of authors and of journals. Journal of Informetrics, 4(3): 365–378, July 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000258.

Bornmann:2011:FST

[BM11a] Lutz Bornmann and Rüdiger Mutz. Further steps towards an ideal method of measuring citation performance: the avoidance of citation (ratio) averages in field-normalization. Journal of Informetrics, 5(1):228–230, January 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000969.

Bouyssou:2011:BRJ

[BM11b] Denis Bouyssou and Thierry Marchant. Bibliometric rankings of journals based on impact factors: an axiomatic approach. Journal of Informetrics, 5(1):75–86, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157710000842.

Bornmann:2012:HAP

[BM12] Lutz Bornmann and Werner Marx. HistCite analysis of papers constituting the h index research front. Journal of Informetrics, 6(2):285–288, April 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711001003.

Bornmann:2013:PBP

[BM13a] Lutz Bornmann and Werner Marx. The proposal of a broadening of perspective in evaluative bibliometrics by complementing the times cited with a cited reference analysis. Journal of Informetrics, 7(1):84–88, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000788.

Bornmann:2013:AUS

[BM13b] Lutz Bornmann and Rüdiger Mutz. The advantage of the use of samples in evaluative bibliometric studies. Journal of Informetrics, 7(1):89–90, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000624.

Bouyssou:2014:AAB

[BM14] Denis Bouyssou and Thierry Marchant. An axiomatic approach to bibliometric rankings and indices. Journal of Informetrics, 8(3):449–477, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000297.

Bornmann:2015:MGN

[BM15] Lutz Bornmann and Werner Marx. Methods for the generation of normalized citation impact scores in bibliometrics: Which method best reflects the judgements of experts? Journal of Informetrics, 9(2):408–418, April 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715000073.

Bouyssou:2016:RAU

[BM16] Denis Bouyssou and Thierry Marchant. Ranking authors using fractional counting of citations: an axiomatic approach. Journal of Informetrics, 10(1):183–199, February 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S1751157715301796. See correction [HCS19].

Bornmann:2018:CRS

[BM18] Lutz Bornmann and Werner Marx. Critical rationalism and the search for standard (field-normalized) indicators in bibliometrics. Journal of Informetrics, 12(3):598–604, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718301937.

Batagelj:2021:TBN

[BM21] Vladimir Batagelj and Daria Maltseva. Temporal bibliographic networks. Journal of Informetrics, 14(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719301439.

Bornmann:2007:GDG

[BMD07] Lutz Bornmann, Rüdiger Mutz, and Hans-Dieter Daniel. Gender differences in grant peer review: a meta-analysis. Journal of Informetrics, 1(3):226–238, July 2007. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157707000363.

Bornmann:2008:HDI

[BMD08a] Lutz Bornmann, Rüdiger Mutz, and Hans-Dieter Daniel. How to detect indications of potential sources of bias in peer review: a generalized latent variable modeling approach exemplified by a gender study. Journal of Informetrics, 2(4):280–287, October 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157708000485.

Bornmann:2008:LMM

[BMD08b] Lutz Bornmann, Rüdiger Mutz, and Hans-Dieter Daniel. Latent Markov modeling applied to grant peer review. Journal of Informetrics, 2(3):217–228, July 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157708000278.

Bornmann:2010:IRO

[BMD10] Lutz Bornmann, Rüdiger Mutz, and Hans-Dieter Daniel. The h index research output measurement: Two approaches to enhance its accuracy. Journal of Informetrics, 4(3):407–414, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000271.

Bornmann:2011:MMA

[BMHD11] Lutz Bornmann, Rüdiger Mutz, Sven E. Hug, and Hans-Dieter Daniel. A multilevel meta-analysis of studies reporting correlations between the h index and 37 different h index variants. Journal of Informetrics, 5(3):346–359, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157711000071.

Ba:2021:EEC

[BMML21] Zhichao Ba, Jin Mao, Yaxue Ma, and Zhentao Liang. Exploring the effect of city-level collaboration and knowledge networks on innovation: Evidence from energy conservation field. Journal of Informetrics, 15(3):??, August 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000699.

Bornmann:2009:CVB

[BMS^{+}09] Lutz Bornmann, Werner Marx, Hermann Schier, Erhard Rahm, Andreas Thor, and Hans-Dieter Daniel. Convergent validity of bibliometric Google Scholar data in the field of chemistry — citation counts for papers that were accepted by Angewandte Chemie International Edition or rejected but published elsewhere, using Google Scholar, Science Citation Index, Scopus, and Chemical Abstracts.

Journal of Informetrics, 3(1):27–35, January 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157708000667.

Brito:2021:IHI

[BN21] Ricardo Brito and Alonso Rodríguez Navarro. The inconsistency of h-index: a mathematical analysis. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720306234.

Bu:2017:CMS

[BNbH17] Yi Bu, Shaokang Ni, and Win bin Huang. Combining multiple scholarly relationships with author cocitation analysis: a preliminary exploration on improving knowledge domain mappings. Journal of Informetrics, 11(3):810–822, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716303674.

Bornmann:2012:SCI

[BO12] Lutz Bornmann and Adam Ozimek. Stata commands for importing bibliometric data and processing author address information. Journal of Informetrics, 6(4):505–512, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000302.

Bornmann:2019:VCA

[BO19] Lutz Bornmann and António Osório. The value and credits of n-authors publications. Journal of Informetrics, 13(2): 540–554, May 2019. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157718301056.

Bornmann:2010:TIM

[Bor10] Lutz Bornmann. Towards an ideal method of measuring research performance: Some comments to the Opthof and Leydesdorff (2010) paper. Journal of Informetrics, 4(3):441–443, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000374.

Bornmann:2013:BAI

[Bor13a] Lutz Bornmann. A better alternative to the h index. Journal of Informetrics, 7(1):100, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771200079X. See comment [Pra13a].

Bornmann:2013:PCI

[Bor13b] Lutz Bornmann. The problem of citation impact assessments for recent publication years in institutional evaluations. Journal of Informetrics, 7(3):722–729, July 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771300045X.

Bornmann:2014:DAP

[Bor14a] Lutz Bornmann. Do altmetrics point to the broader impact of research? An overview of benefits and disadvantages of altmetrics. Journal of Informetrics, 8(4):895–903, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157714000868.

Bornmann:2014:IRS

[Bor14b] Lutz Bornmann. h-Index research in scientometrics: a summary. Journal of Informetrics, 8(3):749–750, July 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000649.

Bornmann:2014:VAD

[Bor14c] Lutz Bornmann. Validity of altmetrics data for measuring societal impact: a study using data from Altmetric and F1000Prime. Journal of Informetrics, 8(4):935–950, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157714000881.

Bornmann:2019:DNC

[Bor19] Lutz Bornmann. Does the normalized citation impact of universities profit from certain properties of their published documents — such as the number of authors and the impact factor of the publishing journals? A multilevel modeling approach. Journal of Informetrics, 13(1):170–184, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718302657.

Bornmann:2011:INC

[BP11] Lutz Bornmann and Andrew Plume. Is it necessary to consider suburbs (or small cities in the close proximity) and name variants in a citation impact analysis for bigger cities? An investigation using Munich as an example. Journal of Informetrics, 5(4):695–697, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000460.

Ballester:2022:RRS

[BP22] Omar Ballester and Orion Penner. Robustness, replicability and scalability in topic modelling. Journal of Informetrics, 16 (1):??, February 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115772100095X.

Borenstein:2022:AIC

[BPI22] Denis Borenstein, Marcelo S. Perlin, and Takeyoshi Imasato. The academic inbreeding controversy: Analysis and evidence from Brazil. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722000396.

Boyack:2014:CIA

[BPU^{+}14] Kevin W. Boyack, Michael Patek, Lyle H. Ungar, Patrick Yoon, and Richard Klavans. Classification of individual articles from all of science by research level. Journal of Informetrics, 8(1):1–12, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000825.

Bloch:2019:PPC

[BRA19] Carter Bloch, Thomas K. Ryan, and Jens Peter Andersen. Public-private collaboration and scientific impact: an analysis based on Danish publication data for 1995–2013. Journal of Informetrics, 13(2):593–604, May 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718304188.

Brito:2018:RAP

[BRN18] Ricardo Brito and Alonso Rodríguez-Navarro. Research assessment by percentile-based double rank analysis. Journal of Informetrics, 12(1):315–329, February 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717304297.

Brito:2019:ERR

[BRN19] Ricardo Brito and Alonso Rodríguez-Navarro. Evaluating research and researchers by the journal impact factor: Is it better than coin flipping? Journal of Informetrics, 13(1):314– 324, February 2019. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157718304097.

Brzezinski:2014:EMI

[Brz14] Michal Brzezinski. Empirical modeling of the impact factor distribution. Journal of Informetrics, 8(2):362–368, April 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714000169.

Borner:2009:VCM

[BS09] Katy Börner and Andrea Scharnhorst. Visual conceptualizations and models of science. Journal of Informetrics, 3(3): 161–172, July 2009. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157709000315.

Bjork:2013:PDS

[BS13] Bo-Christer Björk and David Solomon. The publishing delay in scholarly peer-reviewed journals. Journal of Informetrics, 7(4):914–923, October 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000734.

Brito:2021:ABA

[BSA21] Ana C. M. Brito, Filipi N. Silva, and Diego R. Amancio. Associations between author-level metrics in subsequent time periods. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157721000894.

Brito:2021:CAC

[BSdA^{+}21] Ana C. M. Brito, Filipi N. Silva, Henrique F. de Arruda, Cesar H. Comin, and Diego R. Amancio Luciano da F.Costa. Classification of abrupt changes along viewing profiles of scientific articles. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157721000298.

Bornmann:2014:WEC

[BSdMAM14] Lutz Bornmann, Moritz Stefaner, Felix de Moya Anegón, and Rüdiger Mutz. What is the effect of country-specific characteristics on the research performance of scientific institutions? Using multi-level statistical models to rank and map universities and research-focused institutions worldwide. Journal of Informetrics, 8(3):581–593, July 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000479.

Bornmann:2016:ENS

[BSdMAM16] Lutz Bornmann, Moritz Stefaner, Felix de Moya Anegón, and Rüdiger Mutz. Excellence networks in science: a Webbased application based on Bayesian multilevel logistic regression (BMLR) for the identification of institutions collaborating successfully. Journal of Informetrics, 10(1):312–327, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715301152.

Behrouzi:2021:PSR

[BSHK21] Saman Behrouzi, Zahra Shafaeipour Sarmoor, Khosrow Hajsadeghi, and Kaveh Kavousi. Predicting scientific research trends based on link prediction in keyword networks. Journal of Informetrics, 14(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720300456.

Balakrishnan:2010:SRB

[BSK10] N. Balakrishnan, José María Sarabia, and Nikolai Kolev. A simple relation between the Leimkuhler curve and the mean residual life. Journal of Informetrics, 4(4):602–607, October 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000635.

Bornmann:2012:WFD

[BSMD12] Lutz Bornmann, Hermann Schier, Werner Marx, and Hans-Dieter Daniel. What factors determine citation counts of publications in chemistry besides their quality? Journal of Informetrics, 6(1):11–18, January 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000769.

Battiston:2022:CEC

[BSS22] Pietro Battiston, Pier Luigi Sacco, and Luca Stanca. Cover effects on citations uncovered: Evidence from Nature. Journal of Informetrics, 16(2):??, May 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000451.

Berge:2017:BCI

[BSW17] Laurent Bergé, Thomas Scherngell, and Iris Wanzenböck. Bridging centrality as an indicator to measure the ‘bridging role’ of actors in networks: an application to the European Nanotechnology co-publication network. Journal of Informetrics, 11(4):1031–1042, November 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717300081.

Bornmann:2021:CVS

[BT21] Lutz Bornmann and Alexander Tekles. Convergent validity of several indicators measuring disruptiveness with milestone assignments to physics papers by experts. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000304.

Bornmann:2021:DWM

[BTZY21] Lutz Bornmann, Alexander Tekles, Helena H. Zhang, and Fred Y. Ye. Do we measure novelty when we analyze unusual combinations of cited references? A validation study of bibliometric novelty indicators based on F1000Prime data. Journal of Informetrics, 13(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S1751157718304371. See corrigendum [BTZY24].

Bornmann:2024:CSD

[BTZY24] Lutz Bornmann, Alexander Tekles, Helena H. Zhang, and Fred Y. Ye. Corrigendum to “Do we measure novelty when we analyze unusual combinations of cited references? A validation study of bibliometric novelty indicators based on F1000Prime data” [Journal of Informetrics 13/4 (2019) 100979]. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724000579. See [BTZY21].

Buehling:2021:CRT

[Bue21] Kilian Buehling. Changing research topic trends as an effect of publication rankings — the case of German economists and the Handelsblatt Ranking. Journal of Informetrics, 15 (3):??, August 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000705.

Burrell:2007:HIS

[Bur07a] Quentin L. Burrell. Hirsch’s h-index: a stochastic model. Journal of Informetrics, 1(1):16–25, January 2007. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157706000046.

Burrell:2007:ISH

[Bur07b] Quentin L. Burrell. On the h-index, the size of the Hirsch core and Jin’s A-index. Journal of Informetrics, 1(2):170– 177, April 2007. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157707000314.

Burrell:2008:SCE

[Bur08] Quentin L. Burrell. Some comments on “The estimation of lost multi-copy documents: a new type of informetrics theory” by Egghe and Proot. Journal of Informetrics, 2(1):101– 105, January 2008. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157707000570.

Burrell:2013:ICT

[Bur13a] Quentin L. Burrell. The h-index: a case of the tail wagging the dog? Journal of Informetrics, 7(4):774–783, October 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157713000515.

Burrell:2013:SAR

[Bur13b] Quentin L. Burrell. A stochastic approach to the relation between the impact factor and the uncitedness factor. Journal of Informetrics, 7(3):676–682, July 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000217.

Butler:2011:DDC

[But11] Linda Butler. The devil is in the detail: Concerns about Vanclay’s analysis of Australian journal rankings. Journal of Informetrics, 5(4):693–694, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000381.

Butler:2017:RVB

[But17] Linda Butler. Response to van den Besselaar et al.: What happens when the Australian context is misunderstood. Journal of Informetrics, 11(3):919–922, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717301955.

Boyack:2018:CTC

[BvECW18] Kevin W. Boyack, Nees Jan van Eck, Giovanni Colavizza, and Ludo Waltman. Characterizing in-text citations in scientific articles: a large-scale analysis. Journal of Informetrics, 12(1):59–73, February 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717303516.

Buter:2011:NAC

[BvR11] R. K. Buter and A. F. J. van Raan. Non-alphanumeric characters in titles of scientific publications: an analysis of their occurrence and correlation with citation impact. Journal of Informetrics, 5(4):608–617, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000642.

Bornmann:2011:DHR

[BW11] Lutz Bornmann and Ludo Waltman. The detection of “hot regions” in the geography of science — a visualization approach by using density maps. Journal of Informetrics, 5(4):547–553, October 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157711000435.

Bornmann:2013:HCP

[BW13] Lutz Bornmann and Richard Williams. How to calculate the practical significance of citation impact differences? An empirical example from evaluative institutional bibliometrics using adjusted predictions and marginal effects. Journal of Informetrics, 7(2):562–574, April 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000205.

Birkmaier:2014:MEE

[BW14] Daniel Birkmaier and Klaus Wohlrabe. The Matthew effect in economics reconsidered. Journal of Informetrics, 8(4):880– 889, October 2014. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157714000819.

Bornmann:2017:CJI

[BW17a] Lutz Bornmann and Richard Williams. Can the journal impact factor be used as a criterion for the selection of junior researchers? A large-scale empirical study based on ResearcherID data. Journal of Informetrics, 11(3):788–799, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157717300378. See corrigen-

dum [BW22].

Bornmann:2017:UJI

[BW17b] Lutz Bornmann and Richard Williams. Use of the journal impact factor as a criterion for the selection of junior researchers: a rejoinder on a comment by Peters (2017). Journal of Informetrics, 11(4):945–947, November 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717302948. See [Pet17].

Bornmann:2022:CSC

[BW22] Lutz Bornmann and Richard Williams. Corrigendum to “Can the journal impact factor be used as a criterion for the selection of junior researchers? A large-scale empirical study based on ResearcherID data” Journal of Informetrics Volume 11, Issue 3, August 2017, Pages 788–799. Journal of Informetrics, 16(1):??, February 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721001073. See [BW17a].

Bai:2019:PCS

[BZL19] Xiaomei Bai, Fuli Zhang, and Ivan Lee. Predicting the citations of scholarly paper. Journal of Informetrics, 13(1):407– 418, February 2019. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157718301767.

Bai:2023:QIS

[BZLX23] Xiaomei Bai, Fuli Zhang, Jiaying Liu, and Feng Xia. Quantifying the impact of scientific collaboration and papers via motif-based heterogeneous networks. Journal of Informetrics, 17(2):??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157723000226.

Colliander:2011:ETS

[CA11] Cristian Colliander and Per Ahlgren. The effects and their stability of field normalization baseline on relative performance with respect to citation impact: a case study of 20 natural science departments. Journal of Informetrics, 5(1):101– 113, January 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000866.

Cabrerizo:2010:IQQ

[CAHVH10] F. J. Cabrerizo, S. Alonso, E. Herrera-Viedma, and F. Herrera. q^{2}-Index: Quantitative and qualitative evaluation based on the number and impact of papers in the Hirsch core. Journal of Informetrics, 4(1):23–28, January 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115770900056X.

Chen:2015:TCP

[CAL15] Shiji Chen, Clément Arsenault, and Vincent Larivière. Are top-cited papers more interdisciplinary? Journal of Informetrics, 9(4):1034–1046, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715300201.

Campanario:2014:ADC

[Cam14] Juan Miguel Campanario. Analysis of the distribution of cited journals according to their positions in the hcore of citing journal listed in Journal Citation Reports.

Journal of Informetrics, 8(3):534–545, July 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000467.

Campanario:2015:PID

[Cam15] Juan Miguel Campanario. Providing impact: the distribution of JCR journals according to references they contribute to the 2-year and 5-year journal impact factors. Journal of Informetrics, 9(2):398–407, April 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000061.

Costas:2007:IAL

[CB07] Rodrigo Costas and María Bordons. The h-index: Advantages, limitations and its relation with other bibliometric indicators at the micro level. Journal of Informetrics, 1(3): 193–203, July 2007. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157707000338.

Carusi:2019:SCD

[CB19] Chiara Carusi and Giuseppe Bianchi. Scientific community detection via bipartite scholar/journal graph co-clustering. Journal of Informetrics, 13(1):354–386, February 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718304073.

Chan:2021:PSC

[CBT21] Ho Fai Chan, Sohel Md Bodiuzzman, and Benno Torgler. The power of social cues in the battle for attention: Evidence from an online platform for scholarly commentary. Journal of Informetrics, 14(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719303438.

Calabrese:2018:PLB

[CCC^{+}18] Armando Calabrese, Guendalina Capece, Roberta Costa, Francesca Di Pillo, and Stefania Giuffrida. A ‘power law’ based method to reduce size-related bias in indicators of knowledge performance: an application to university research assessment. Journal of Informetrics, 12(4):1263–1281, November 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157717303401.

Chen:2009:TEC

[CCH^{+}09] Chaomei Chen, Yue Chen, Mark Horowitz, Haiyan Hou, Zeyuan Liu, and Donald Pellegrino. Towards an explanatory and computational theory of scientific discovery. Journal of Informetrics, 3(3):191–209, July 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000236.

Chang:2012:RBP

[CCH12] Ke-Chiun Chang, Dar-Zen Chen, and Mu-Hsuan Huang. The relationships between the patent performance and corporation performance. Journal of Informetrics, 6(1):131–139, January 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711000782.

Cao:2016:DAA

[CCL16] Xuanyu Cao, Yan Chen, and K. J. Ray Liu. A data analytic approach to quantifying scientific impact. Journal of Informetrics, 10(2):471–484, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715301346.

Comins:2018:PCS

[CCL18] Jordan A. Comins, Stephanie A. Carmack, and Loet Leydesdorff. Patent citation spectroscopy (PCS): Online retrieval of landmark patents based on an algorithmic approach. Journal of Informetrics, 12(4):1223–1231, November 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717304492.

Cruz-Castro:2021:WSR

[CCSM21] Laura Cruz-Castro and Luis Sanz-Menendez. What should be rewarded? Gender and evaluation criteria for tenure and promotion. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157721000675.

Chen:2023:EEA

[CD23] Yifan Chen and Jingda Ding. Exploitation and exploration: an analysis of the research pattern of Nobel laureates in physics. Journal of Informetrics, 17(3):??, August 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000536.

Colavizza:2016:CCH

[CF16] Giovanni Colavizza and Massimo Franceschet. Clustering citation histories in the Physical Review. Journal of Informetrics, 10(4):1037–1051, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716301560.

Choudhury:2021:MTE

[CFK21] Nazim Choudhury, Fahim Faisal, and Matloob Khushi. Mining temporal evolution of knowledge graphs and genealogical features for literature-based discovery prediction. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719304468.

Colavizza:2019:EIT

[CFvL19] Giovanni Colavizza, Thomas Franssen, and Thed van Leeuwen. An empirical investigation of the tribes and their territories: Are research specialisms rural and urban? Journal of Informetrics, 13(1):105–117, February 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302554.

Chen:2010:ISW

[CG10] Zifeng Chen and Jiancheng Guan. The impact of small world on innovation: an empirical study of 16 countries.

Journal of Informetrics, 4(1):97–106, January 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000741.

Chen:2011:BIR

[CG11] Kaihua Chen and Jiancheng Guan. A bibliometric investigation of research performance in emerging nanobiopharmaceuticals. Journal of Informetrics, 5(2):233–247, April 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157710000945.

Chen:2021:COG

[CGD^{+}21] Meiqian Chen, Zhaoxia Guo, Yucheng Dong, Francisco Chiclana, and Enrique Herrera-Viedma. Citations optimal growth path: a tool to analyze sensitivity to citations of h-like indexes. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721000869.

Clemente-Gallardo:2019:DRC

[CGFÍ^{+}19] J. Clemente-Gallardo, A. Ferrer, D. Íñiguez, A. Rivero, G. Ruiz, and A. Tarancón. Do researchers collaborate in a similar way to publish and to develop projects? Journal of Informetrics, 13(1):64–77, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717304431.

Cena:2015:PCI

[CGM15] Anna Cena, Marek Gagolewski, and Radko Mesiar. Problems and challenges of information resources producers’ clustering. Journal of Informetrics, 9(2):273–284, April 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715000231.

Cena:2022:VCM

[CGSZS22] Anna Cena, Marek Gagolewski, Grzegorz Siudem, and Barbara Zogała-Siudem. Validating citation models by proxy indices. Journal of Informetrics, 16(2):??, May 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000190.

Cao:2022:HDD

[CGXW22] Renmeng Cao, Yu Geng, Xiaoke Xu, and Xianwen Wang. How does duplicate tweeting boost social media exposure to scholarly articles? Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722000013.

Chuang:2014:BPT

[CH14] Kun-Yang Chuang and Yuh-Shan Ho. Bibliometric profile of top-cited single-author articles in the Science Citation Index Expanded. Journal of Informetrics, 8(4):951–962, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157714000893.

Comins:2015:CMS

[CH15] Jordan A. Comins and Thomas W. Hussey. Compressing multiple scales of impact detection by Reference Publication Year Spectroscopy. Journal of Informetrics, 9(3):449–454, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715200065.

Chang:2023:AFA

[CH23] Ying-Han Chang and Mu-Hsuan Huang. Analysis of factors affecting scientific migration move and distance by academic age, migrant type, and country: Migrant researchers in the field of business and management. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722001249.

Chang:2022:CNE

[Cha22] Yu-Wei Chang. Capability of non-English-speaking countries for securing a foothold in international journal publishing. Journal of Informetrics, 16(3):??, August 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000578.

Chen:2024:CSR

[CHD^{+}24] Guo Chen, Siqi Hong, Chenxin Du, Panting Wang, Zeyu Yang, and Lu Xiao. Comparing semantic representation methods for keyword analysis in bibliometric research. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000427.

Chen:2017:DPC

[Che17] Lixin Chen. Do patent citations indicate knowledge linkage? The evidence from text similarities between patents and their citations. Journal of Informetrics, 11(1):63–79, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715301711.

Chen:2011:IMR

[CHHL11] Dar-Zen Chen, Mu-Hsuan Huang, Hui-Chen Hsieh, and Chang-Pin Lin. Identifying missing relevant patent citation links by using bibliographic coupling in LED illuminating technology. Journal of Informetrics, 5(3):400–412, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711000307.

Chi:2016:DDC

[Chi16] Pei-Shan Chi. Differing disciplinary citation concentration patterns of book and journal literature? Journal of Informetrics, 10(3):814–829, August 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716300347.

Chowdhury:2021:FAG

[Cho21] K. P. Chowdhury. Functional analysis of generalized linear models under non-linear constraints with applications to identifying highly-cited papers. Journal of Informetrics, 15 (1):??, February 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157720306295.

Chowdhury:2023:NFA

[Cho23] K. P. Chowdhury. Nonparametric functional analysis under joint estimation with applications to identifying highly cited papers. Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723000718.

Cova:2015:UCR

[CJF^{+}15] Tânia F. G. G. Cova, Susana Jarmelo, Sebastião J. Formosinho, J. Sérgio Seixas de Melo, and Alberto A. C. C. Pais. Unsupervised characterization of research institutions with task-force estimation. Journal of Informetrics, 9(1):59– 68, January 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157714001072.

Chen:2014:EAJ

[CJW14] Kuan-Ming Chen, Tsung-Hau Jen, and Margaret Wu. Estimating the accuracies of journal impact factor through bootstrap. Journal of Informetrics, 8(1):181–196, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713001089.

Chekhovich:2022:ADP

[CK22] Yury V. Chekhovich and Andrey V. Khazov. Analysis of duplicated publications in Russian journals. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721001176.

Chung:2021:IPM

[CKKY21] Jaemin Chung, Namuk Ko, Hyeonsu Kim, and Janghyeok Yoon. Inventor profile mining approach for prospective human resource scouting. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115772030328X.

Chen:2023:ASP

[CKZ^{+}23] Ying Chen, Thorsten Koch, Nazgul Zakiyeva, Kailiang Liu, Zhitong Xu, Chun houh Chen, Junji Nakano, and Keisuke Honda. Article’s scientific prestige: Measuring the impact of individual articles in the Web of Science. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000044.

Claveau:2016:TSM

[Cla16] François Claveau. There should not be any mystery: a comment on sampling issues in bibliometrics. Journal of Informetrics, 10(4):1233–1240, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302589.

Chen:2011:NNB

[CLC11] Long-Sheng Chen, Cheng-Hsiang Liu, and Hui-Ju Chiu. A neural network based approach for sentiment classification in the blogosphere. Journal of Informetrics, 5(2):313–322, April 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711000046.

Cobo:2011:ADQ

[CLHHVH11] M. J. Cobo, A. G. López-Herrera, E. Herrera-Viedma, and F. Herrera. An approach for detecting, quantifying, and visualizing the evolution of a research field: a practical application to the Fuzzy Sets Theory field. Journal of Informetrics, 5(1):146–166, January 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000891.

Cassi:2017:IFM

[CLR^{+}17] Lorenzo Cassi, Agénor Lahatte, Ismael Rafols, Pierre Sautier, and Élisabeth de Turckheim. Improving fitness: Mapping research priorities against societal needs on obesity. Journal of Informetrics, 11(4):1095–1113, November 2017.

CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717301542.

Charemza:2024:JRR

[CLW24] Wojciech Charemza, Michał Lewandowski, and Łukasz Woźny. On journal rankings and researchers’ abilities. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000725.

Cerovsek:2014:CSC

[CM14] Tomo Cerovsek and Matjaz Mikos. A comparative study of cross-domain research output and citations: Research impact cubes and binary citation frequencies. Journal of Informetrics, 8(1):147–161, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713001053.

Cappelletti-Montano:2021:NPB

[CMCMM21] Beniamino Cappelletti-Montano, Silvia Columbu, Stefano Montaldo, and Monica Musio. New perspectives in bibliometric indicators: Moving from citations to citing authors. Journal of Informetrics, 15(3):??, August 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000353.

Cappelletti-Montano:2022:IOR

[CMCMM22] Beniamino Cappelletti-Montano, Silvia Columbu, Stefano Montaldo, and Monica Musio. Interpreting the outcomes of research assessments: a geometrical approach. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000062.

Chen:2024:CCA

[CML24] Xi Chen, Jin Mao, and Gang Li. A co-citation approach to the analysis on the interaction between scientific and technological knowledge. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157724000610.

Cormode:2013:SI

[CMMT13] Graham Cormode, Qiang Ma, S. Muthukrishnan, and Brian Thompson. Socializing the h-index. Journal of Informetrics, 7(3):718–721, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157713000448.

Calero-Medina:2008:CMC

[CMN08] Clara Calero-Medina and Ed C. M. Noyons. Combining mapping and citation network analysis for a better understanding of the scientific development: the case of the absorptive capacity field. Journal of Informetrics, 2(4):272–279, October 2008. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157708000497.

Coccia:2008:MSP

[Coc08] Mario Coccia. Measuring scientific performance of public research units for strategic change. Journal of Informetrics, 2(3):183–194, July 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157708000229.

Copiello:2019:PNE

[Cop19] Sergio Copiello. Peer and neighborhood effects: Citation analysis using a spatial autoregressive model and pseudospatial data. Journal of Informetrics, 13(1):238–254, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718300488.

Corsini:2023:DGF

[CP23] Alberto Corsini and Michele Pezzoni. Does grant funding foster research impact? Evidence from France. Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000731.

Craig:2007:DOA

[CPM^{+}07] Iain D. Craig, Andrew M. Plume, Marie E. McVeigh, James Pringle, and Mayur Amin. Do open access articles have greater citation impact?: a critical review of the literature. Journal of Informetrics, 1(3):239–248, July 2007. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157707000466.

Chakraborty:2024:MKM

[CPN24] Joyita Chakraborty, Dinesh K. Pradhan, and Subrata Nandi. A multiple k-means cluster ensemble framework for clustering citation trajectories. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157724000208.

Chen:2021:EIP

[CQAL21] Shiji Chen, Junping Qiu, Clément Arsenault, and Vincent Larivière. Exploring the interdisciplinarity patterns of highly cited papers. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157720306416.

Chen:2010:CSP

[CR10] P. Chen and S. Redner. Community structure of the physical review citation network. Journal of Informetrics, 4(3):278– 290, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000027.

Cimenler:2014:RAR

[CRS14] Oguz Cimenler, Kingsley A. Reeves, and John Skvoretz. A regression analysis of researchers’ social network metrics on their citation performance in a college of engineering. Journal of Informetrics, 8(3):667–682, July 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000571.

Cimenler:2015:ECR

[CRS15] Oguz Cimenler, Kingsley A. Reeves, and John Skvoretz. An evaluation of collaborative research in a college of engineering. Journal of Informetrics, 9(3):577–590, July 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000516.

Croft:2022:PCC

[CS22] William L. Croft and Jörg-Rüdiger Sack. Predicting the citation count and CiteScore of journals one year in advance. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722001018.

Csato:2021:JRS

[Csa21] László Csató. Journal ranking should depend on the level of aggregation. Journal of Informetrics, 13(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719301397.

Correa:2017:PAC

[CSdFCA17] Edilson A. Corrêa, Jr., Filipi N. Silva, Luciano da F. Costa, and Diego R. Amancio. Patterns of authors contribution in scientific manuscripts. Journal of Informetrics, 11(2):498– 510, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716302693.

Chen:2018:SAM

[CSH18] Chaomei Chen, Min Song, and Go Eun Heo. A scalable and adaptive method for finding semantically equivalent cue words of uncertainty. Journal of Informetrics, 12(1):158– 180, February 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157717301712.

Csomos:2018:SSA

[Cso18] György Csomós. A spatial scientometric analysis of the publication output of cities worldwide. Journal of Informetrics, 12(2):547–566, May 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S1751157718301949.

Csomos:2021:IRA

[Cso21] György Csomós. Introducing recalibrated academic performance indicators in the evaluation of individuals’ research performance: a case study from Eastern Europe. Journal of Informetrics, 14(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720301139.

Csomos:2016:EPC

[CT16] György Csomós and Géza Tóth. Exploring the position of cities in global corporate research and development: a bibliometric analysis by two different geographical approaches. Journal of Informetrics, 10(2):516–532, May 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715301723.

Chen:2017:UTE

[CTDM17] Baitong Chen, Satoshi Tsutsui, Ying Ding, and Feicheng Ma. Understanding the topic evolution in a scientific domain: an exploratory study for the field of information retrieval. Journal of Informetrics, 11(4):1175–1189, November 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717300536.

Chakraborty:2015:UMD

[CTGM15] Tanmoy Chakraborty, Vihar Tammana, Niloy Ganguly, and Animesh Mukherjee. Understanding and modeling diverse scientific careers of researchers. Journal of Informetrics, 9(1): 69–78, January 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157714001102.

Chi:2022:ESA

[CTL22] Yuxue Chi, Xianyi Tang, and Yijun Liu. Exploring the “awakening effect” in knowledge diffusion: a case study of publications in the library and information science domain. Journal of Informetrics, 16(4):??, November 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000943.

Chen:2021:NAA

[CtRlY21] Kun Chen, Xian tong Ren, and Guo liang Yang. A novel approach for assessing academic journals: Application of integer DEA model for management science and operations research field. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115772100047X.

Citron:2018:NAS

[CW18] Daniel T. Citron and Samuel F. Way. Network assembly of scientific communities of varying size and specificity. Journal of Informetrics, 12(1):181–190, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717301797.

Cheng:2024:OPR

[CWT^{+}24] Xi Cheng, Haoran Wang, Li Tang, Weiyan Jiang, Maotian Zhou, and Guoyan Wang. Open peer review correlates with altmetrics but not with citations: Evidence from Nature Communications and PLoS One. Journal of Informetrics, 18 (3):??, August 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000531.

Chen:2016:SPK

[CX16] Guo Chen and Lu Xiao. Selecting publication keywords for domain analysis in bibliometrics: a comparison of three methods. Journal of Informetrics, 10(1):212–223, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115771600002X.

Chen:2007:FSG

[CXMR07] P. Chen, H. Xie, S. Maslov, and S. Redner. Finding scientific gems with Google’s PageRank algorithm. Journal of Informetrics, 1(1):8–15, January 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157706000034.

Chen:2022:SMP

[CXZ^{+}22a] Liang Chen, Shuo Xu, Lijun Zhu, Jing Zhang, Haiyun Xu, and Guancan Yang. A semantic main path analysis method to identify multiple developmental trajectories. Journal of Informetrics, 16(2):??, May 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000335.

Chen:2022:DLB

[CXZ^{+}22b] Liang Chen, Shuo Xu, Lijun Zhu, Jing Zhang, Guancan Yang, and Haiyun Xu. A deep learning based method benefiting from characteristics of patents for semantic relation classification. Journal of Informetrics, 16(3):??, August 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722000645.

Choi:2022:MKE

[CY22] Jaewoong Choi and Janghyeok Yoon. Measuring knowledge exploration distance at the patent level: Application of network embedding and citation analysis. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157722000384.

Chen:2023:NCC

[CY23] Wei Chen and Yan Yan. New components and combinations: the perspective of the internal collaboration networks of scientific teams. Journal of Informetrics, 17(2):??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723000329.

Chen:2018:QQA

[CySlY18] Kun Chen, Yao yao Song, and Guo liang Yang. Quality and quantity are not always positively correlated: a case study of Chinese economics journals. Journal of Informetrics, 12(4):1178–1181, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718303985.

Cui:2021:QIT

[CZFD21] Haochuan Cui, An Zeng, Ying Fan, and Zengru Di. Quantifying the impact of a teamwork publication. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000882.

Cimini:2016:IIB

[CZG16] Giulio Cimini, Andrea Zaccaria, and Andrea Gabrielli. Investigating the interplay between fundamentals of national research systems: Performance, investments and international collaborations. Journal of Informetrics, 10(1):200–211, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715301048.

Chang:2024:SWC

[CZZ24] Le Chang, Huiying Zhang, and Chao Zhang. Should we circumvent knowledge path dependency? The impact of conventional learning and collaboration diversity on knowledge creation. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724001093.

Batista-Jr:2021:PJR

[dABJGMC21] Antônio de Abreu Batista-Jr, Fábio Castro Gouveia, and Jesús P. Mena-Chalco. Predicting the Q of junior researchers using data from the first years of publication. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000018.

Danell:2016:ERO

[Dan16] Rickard Danell. Evaluating research organizations’ contribution to science is not the same task as evaluating the performance of their scientists. Journal of Informetrics, 10(2): 655–657, May 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157716300979.

Daraio:2015:EES

[DBS15] Cinzia Daraio, Andrea Bonaccorsi, and Léopold Simar. Efficiency and economies of scale and specialization in European universities: a directional distance approach. Journal of Informetrics, 9(3):430–448, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000437.

Dondio:2019:IHP

[DCG^{+}19] Pierpaolo Dondio, Niccolò Casnici, Francisco Grimaldo, Nigel Gilbert, and Flaminio Squazzoni. The “invisible hand” of peer review: the implications of author–referee networks on peer review in a scholarly journal. Journal of Informetrics, 13(2):708–716, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718304206.

DellAnno:2021:THP

[DCM21] Roberto Dell’Anno, Rocco Caferra, and Andrea Morone. A “Trojan Horse” in the peer-review process of fee-charging economic journals. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157720300080.

DAngelo:2019:TUM

[DD19] Ciriaco Andrea D’Angelo and Samuele Di Russo. Testing for universality of Mendeley readership distributions. Journal of Informetrics, 13(2):726–737, May 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302876.

deDiego:2021:SER

[dDGFFI^{+}21] Isaac Martín de Diego, César González-Fernández, Alberto Fernández-Isabel, Rubén R. Fernández, and Javier Cabezas. System for evaluating the reliability and novelty of medical scientific papers. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721000596.

Demir:2018:PJW

[Dem18] Selcuk Besir Demir. Predatory journals: Who publishes in them and why? Journal of Informetrics, 12(4):1296–1311, November 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157718301962.

Diaz-Faes:2015:UPI

[DFCGB15] Adrián A. Díaz-Faes, Rodrigo Costas, M. Purificación Galindo, and María Bordons. Unravelling the performance of individual scholars: Use of canonical biplot analysis to explore the performance of scientists by academic rank and scientific field. Journal of Informetrics, 9(4):722–733, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115771520020X.

Diniz-Filho:2016:DAP

[DFFBR16] José Alexandre F. Diniz-Filho, Maria Clorinda S. Fioravanti, Luis Mauricio Bini, and Thiago Fernando Rangel. Drivers of academic performance in a Brazilian university under a government-restructuring program. Journal of Informetrics, 10(1):151–161, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715300213.

Delbianco:2021:MSA

[DFHCT21] Fernando Delbianco, Andrés Fioriti, Allan Hernandez-Chanto, and Fernando Tohmé. A Markov-switching approach to the study of citations in academic journals. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720301851.

DeFilippo:2021:ESC

[DG21] Daniela De Filippo and Juan Gorraiz. Is the Emerging Source Citation Index an aid to assess the citation impact in social science and humanities? Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157719304067.

Dorta-Gonzalez:2013:IMT

[DGDG13] P. Dorta-González and M. I. Dorta-González. Impact maturity times and citation time windows: the 2-year maximum journal impact factor. Journal of Informetrics, 7(3): 593–602, July 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000254.

Dorta-Gonzalez:2014:JTC

[DGDGSPSV14] Pablo Dorta-González, María Isabel Dorta-González, Dolores Rosa Santos-Peñate, and Rafael Suárez-Vega. Journal topic citation potential and between-field comparisons: the topic normalized impact factor. Journal of Informetrics, 8(2):406–418, April 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714000261.

Dorta-Gonzalez:2024:CSS

[DGRCDG24a] Pablo Dorta-González, Alejandro Rodríguez-Caro, and María Isabel Dorta-González. Corrigendum to “Societal and scientific impact of policy research: a large-scale empirical study of some explanatory factors using Altmetric and Overton” [Journal of Informetrics 18/3 (2024) 101530]. Journal of Informetrics, 18(4):??, November 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772400110X. See [DGRCDG24b].

Dorta-Gonzalez:2024:SSI

[DGRCDG24b] Pablo Dorta-González, Alejandro Rodríguez-Caro, and María Isabel Dorta-González. Societal and scientific impact of policy research: a large-scale empirical study of some explanatory factors using altmetric and Overton. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S1751157724000439. See corrigendum [DGRCDG24a].

Dunaiski:2018:ARE

[DGV18a] Marcel Dunaiski, Jaco Geldenhuys, and Willem Visser. Author ranking evaluation at scale. Journal of Informetrics, 12(3):679–702, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157717303528.

Dunaiski:2018:HER

[DGV18b] Marcel Dunaiski, Jaco Geldenhuys, and Willem Visser. How to evaluate rankings of academic entities using test data. Journal of Informetrics, 12(3):631–655, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718301469.

Dunaiski:2019:GVA

[DGV19a] Marcel Dunaiski, Jaco Geldenhuys, and Willem Visser. Globalised vs averaged: Bias and ranking performance on the author level. Journal of Informetrics, 13(1):299–313, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718304498.

Dunaiski:2019:IBN

[DGV19b] Marcel Dunaiski, Jaco Geldenhuys, and Willem Visser. On the interplay between normalisation, bias, and performance of paper impact metrics. Journal of Informetrics, 13(1):270– 290, February 2019. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157718304164.

Ding:2017:EPC

[DHLW17] Cherng G. Ding, Wen-Chi Hung, Meng-Che Lee, and Hung-Jui Wang. Exploring paper characteristics that facilitate the knowledge flow from science to technology. Journal of Informetrics, 11(1):244–256, February 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716301705.

Dienes:2015:C

[Die15] Keith R. Dienes. Completing h. Journal of Informetrics, 9(2):385–397, April 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1751157715000048.

Ding:2011:CDT

[Din11a] Ying Ding. Community detection: Topological vs. topical. Journal of Informetrics, 5(4):498–514, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000319.

Ding:2011:SCE

[Din11b] Ying Ding. Scientific collaboration and endorsement: Network analysis of coauthorship and citation networks. Journal of Informetrics, 5(1):187–203, January 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000957.

Drivas:2021:RIC

[DK21] Kyriakos Drivas and Ioannis Kaplanis. The role of international collaborations in securing the patent grant. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720303011.

Dziezyc:2022:ERG

[DK22] Maciej Dziezyc and Przemysław Kazienko. Effectiveness of research grants funded by European Research Council and Polish National Science Centre. Journal of Informetrics, 16 (1):??, February 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721001140.

Drozdz:2017:HOH

[DKK^{+}17] Stanisław Drozdz, Andrzej Kulig, Jarosław Kwapień, Artur Niewiarowski, and Marek Stanuszek. Hierarchical organization of H. Eugene Stanley scientific collaboration community in weighted network representation. Journal of Informetrics, 11(4):1114–1127, November 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717301803.

Ding:2013:DRA

[DLGC13] Ying Ding, Xiaozhong Liu, Chun Guo, and Blaise Cronin. The distribution of references across texts: Some implications for citation analysis. Journal of Informetrics, 7(3):583–592, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157713000230.

Du:2019:MKT

[DLGT19] Jian Du, Peixin Li, Qianying Guo, and Xiaoli Tang. Measuring the knowledge translation and convergence in pharmaceutical innovation by funding science–technology–innovation linkages analysis. Journal of Informetrics, 13(1):132–148, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718301779.

Du:2021:PPC

[DLH^{+}21] Jian Du, Peixin Li, Robin Haunschild, Yinan Sun, and Xiaoli Tang. Paper-patent citation linkages as early signs for predicting delayed recognized knowledge: Macro and micro evidence. Journal of Informetrics, 14(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719301555.

DeSordi:2017:PIS

[DLM^{+}17] José Osvaldo De Sordi, Wanderlei Lima de Paulo, Manuel Antonio Meireles, Marcia Carvalho de Azevedo, and Luis Hernan Contreras Pinochet. Proposal of indicators for the structural analysis of scientific articles. Journal of Informetrics, 11(2):483–497, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157716303042.

Demetrescu:2019:SVA

[DLM^{+}19] Camil Demetrescu, Francesco Lupia, Angelo Mendicelli, Andrea Ribichini, Francesco Scarcello, and Marco Schaerf. On the Shapley value and its application to the Italian VQR research assessment exercise. Journal of Informetrics, 13(1):87–104, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157718302141.

delaPena:2011:IFC

[dlP11] José A. de la Peña. Impact functions on the citation network of scientific articles. Journal of Informetrics, 5(4):565–573, October 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711000459.

delaVega:2018:RSM

[dlVV18] Casilda Lasso de la Vega and Oscar Volij. Ranking scholars: a measure representation. Journal of Informetrics, 12(2): 510–517, May 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157718300415.

deMoya-Anegon:2018:SRB

[dMAGBLIM18] Felix de Moya-Anegon, Vicente P. Guerrero-Bote, Carmen Lopez-Illescas, and Henk F. Moed. Statistical relationships between corresponding authorship, international coauthorship and citation impact of national research systems. Journal of Informetrics, 12(4):1251–1262, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718300725.

Duarte-Martinez:2022:UPS

[DMCLH22] V. Duarte-Martínez, M. J. Cobo, and A. G. López-Herrera. Uncovering patterns in the supervision of Spanish theses: a comprehensive analysis. Journal of Informetrics, 16(3):??, August 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722000712.

Didegah:2018:IQI

[DMS18] Fereshteh Didegah, Niels Mejlgaard, and Mads P. Sørensen. Investigating the quality of interactions and public engagement around scientific papers on Twitter. Journal of Informetrics, 12(3):960–971, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717302572.

Dehdarirad:2017:RIC

[DN17] Tahereh Dehdarirad and Stefano Nasini. Research impact in co-authorship networks: a two-mode analysis. Journal of Informetrics, 11(2):371–388, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302450.

deNooy:2015:DTA

[dNL15] Wouter de Nooy and Loet Leydesdorff. The dynamics of triads in aggregated journal-journal citation relations: Specialty developments at the above-journal level. Journal of Informetrics, 9(3):542–554, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771520034X.

Donoso:2017:SII

[Don17] Jose Fernandez Donoso. A simple index of innovation with complexity. Journal of Informetrics, 11(1):1–17, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716300050.

Donner:2018:EPM

[Don18] Paul Donner. Effect of publication month on citation impact. Journal of Informetrics, 12(1):330–343, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717303929.

Donner:2024:RMF

[Don24] Paul Donner. Remarks on modified fractional counting. Journal of Informetrics, 18(4):??, November 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772400097X.

Dosso:2021:DCD

[DS21] Dennis Dosso and Gianmaria Silvello. Data credit distribution: a new method to estimate databases impact. Journal of Informetrics, 14(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720301954.

daSilva:2007:MQS

[dSSOH07] Roberto da Silva, Raquel Stasiu, Viviane Moreira Orengo, and Carlos A. Heuser. Measuring quality of similarity functions in approximate data matching. Journal of Informetrics, 1(1):35–46, January 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S175115770600006X.

Didegah:2013:WFH

[DT13] Fereshteh Didegah and Mike Thelwall. Which factors help authors produce the highest impact research? Collaboration, journal and document properties. Journal of Informetrics, 7(4):861–873, October 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000709.

Didegah:2012:ICJ

[DTG12] Fereshteh Didegah, Mike Thelwall, and Ali Gazni. An international comparison of journal publishing and citing behaviours. Journal of Informetrics, 6(4):516–531, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157712000314.

Dogan:2021:NTW

[DU21] Turgut Dogan and Alper Kursat Uysal. A novel term weighting scheme for text classification: TF-MONO. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720300705.

Dunaiski:2016:EPA

[DVG16] Marcel Dunaiski, Willem Visser, and Jaco Geldenhuys. Evaluating paper and author ranking algorithms using impact and contribution awards. Journal of Informetrics, 10(2): 392–407, May 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715301425.

Dong:2021:ICC

[DWW21] Ke Dong, Jiang Wu, and Kaili Wang. On the inequality of citation counts of all publications of individual authors. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157721000742.

Ebadi:2022:DET

[EAG22] Ashkan Ebadi, Alain Auger, and Yvan Gauthier. Detecting emerging technologies and their evolution using deep learning and weak signal analysis. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722000967.

Egghe:2011:PFC

[EBG11] L. Egghe, L. Bornmann, and R. Guns. A proposal for a First-Citation-Speed-Index. Journal of Informetrics, 5(1):181–186, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000933.

Ennas:2015:FTR

[ED15] Gianfranco Ennas and Maria Chiara Di Guardo. Features of top-rated gold open access journals: an analysis of the Scopus database. Journal of Informetrics, 9(1):79–89, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157714001096.

Egghe:2007:GET

[Egg07a] L. Egghe. General evolutionary theory of information production processes and applications to the evolution of networks. Journal of Informetrics, 1(2):115–122, April 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157706000241.

Egghe:2007:WBI

[Egg07b] Leo Egghe. Welcome to the Journal of Informetrics. Journal of Informetrics, 1(1):1, January 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157706000083.

Egghe:2008:EST

[Egg08a] L. Egghe. Examples of simple transformations of the hindex: Qualitative and quantitative conclusions and consequences for other indices. Journal of Informetrics, 2(2):136– 148, April 2008. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157707000776.

Egghe:2008:IMT

[Egg08b] L. Egghe. The influence of merging on h-type indices. Journal of Informetrics, 2(3):252–262, July 2008. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157708000308.

Egghe:2009:MDI

[Egg09] L. Egghe. Mathematical derivation of the impact factor distribution. Journal of Informetrics, 3(4):290–295, October 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115770900011X.

Egghe:2010:CSS

[Egg10a] L. Egghe. Characteristic scores and scales based on h-type indices. Journal of Informetrics, 4(1):14–22, January 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157709000546.

Egghe:2010:CPI

[Egg10b] L. Egghe. Conjugate partitions in informetrics: Lorenz curves, h-type indices, Ferrers graphs and Durfee squares in a discrete and continuous setting. Journal of Informetrics, 4(3):320–330, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1751157710000088.

Egghe:2011:CGW

[Egg11] L. Egghe. Characterizations of the generalized Wuand Kosmulski-indices in Lotkaian systems. Journal of Informetrics, 5(3):439–445, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S175115771100037X.

Egghe:2012:ARC

[Egg12a] L. Egghe. Averages of ratios compared to ratios of averages: Mathematical results. Journal of Informetrics, 6(2):307–317, April 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711001192.

Egghe:2012:FYB

[Egg12b] L. Egghe. Five years “Journal of Informetrics”. Journal of Informetrics, 6(3):422–426, July 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000120.

Egghe:2013:FRB

[Egg13a] L. Egghe. The functional relation between the impact factor and the uncitedness factor revisited. Journal of Informetrics, 7(1):183–189, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157712000909.

Egghe:2013:MCH

[Egg13b] L. Egghe. A mathematical characterization of the Hirschindex by means of minimal increments. Journal of Informetrics, 7(2):388–393, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157713000072.

Egghe:2014:ICS

[Egg14] L. Egghe. Impact coverage of the success-index. Journal of Informetrics, 8(2):384–389, April 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000066. See comments [Rou14a].

Egghe:2015:MRE

[Egg15] Leo Egghe. Message from the retiring Editor-in-Chief. Journal of Informetrics, 9(1):A1–A2, January 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000085.

Egghe:2021:TPD

[Egg21] Leo Egghe. A theory of pointwise defined impact measures. Journal of Informetrics, 15(3):??, August 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000407.

Egghe:2024:MIH

[Egg24a] Leo Egghe. Mathematical informetrics: Hirsch-type equations and bundles. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723001049.

Egghe:2024:NTD

[Egg24b] Leo Egghe. Networks and their degree distribution, leading to a new concept of small worlds. Journal of Informetrics, 18 (3):??, August 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000671.

ElGibari:2018:EUP

[EGR18] Samira El Gibari, Trinidad Gómez, and Francisco Ruiz. Evaluating university performance using reference point based composite indicators. Journal of Informetrics, 12(4):1235– 1250, November 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157717303450.

Elleby:2010:PPI

[EI10] Anita Elleby and Peter Ingwersen. Publication point indicators: a comparative case study of two publication point systems and citation impact in an interdisciplinary context. Journal of Informetrics, 4(4):512–523, October 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000556.

Eom:2008:AAC

[Eom08] Sean Eom. All author cocitation analysis and first author cocitation analysis: a comparative empirical investigation. Journal of Informetrics, 2(1):53–64, January 2008. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157707000624.

Egghe:2007:ENL

[EP07] L. Egghe and G. Proot. The estimation of the number of lost multi-copy documents: a new type of informetrics theory. Journal of Informetrics, 1(4):257–268, ???? 2007. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157707000351.

Egghe:2019:IST

[ER19a] Leo Egghe and Ronald Rousseau. Infinite sequences and their h-type indices. Journal of Informetrics, 13(1):291–298, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718304486.

Egghe:2019:SSF

[ER19b] Leo Egghe and Ronald Rousseau. Solution by step functions of a minimum problem in L^{2}[0, T], using generalized hand g-indices. Journal of Informetrics, 13(3):785–792, August 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719301087.

Egghe:2021:PCG

[ER21] Leo Egghe and Ronald Rousseau. Polar coordinates and generalized h-type indices. Journal of Informetrics, 14(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719303943. See comments

[Smo21b].

Egghe:2022:RFD

[ER22] Leo Egghe and Ronald Rousseau. Rank-frequency data and impact in a continuous model: Introducing impact bundles. Journal of Informetrics, 16(3):??, August 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000499.

Egghe:2023:GII

[ER23] Leo Egghe and Ronald Rousseau. Global informetric impact: a description and definition using bundles. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722001195.

Ebadi:2015:HBI

[ES15] Ashkan Ebadi and Andrea Schiffauerova. How to become an important player in scientific collaboration networks? Journal of Informetrics, 9(4):809–825, October 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000565.

Ebadi:2021:AML

[ETGS21] Ashkan Ebadi, Stéphane Tremblay, Cyril Goutte, and Andrea Schiffauerova. Application of machine learning techniques to assess the trends and alignment of the funded research output. Journal of Informetrics, 14(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718301901.

Fassin:2024:IDJ

[Fas24] Yves Fassin. The internal dynamics of journals’ h-cores over time. Journal of Informetrics, 18(2):??, May 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000312.

Franceschet:2010:ESC

[FC10] Massimo Franceschet and Antonio Costantini. The effect of scholar collaboration on impact and quality of academic papers. Journal of Informetrics, 4(4):540–553, October 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115771000057X.

Franceschet:2011:FIR

[FC11] Massimo Franceschet and Antonio Costantini. The first Italian research assessment exercise: a bibliometric perspective. Journal of Informetrics, 5(2):275–291, April 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710001008.

Franceschet:2017:TDA

[FC17] Massimo Franceschet and Giovanni Colavizza. TimeRank: a dynamic approach to rate scholars using citations. Journal of Informetrics, 11(4):1128–1141, November 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717301025.

Ficcadenti:2021:WRH

[FCAD21] Valerio Ficcadenti, Roy Cerqueti, Marcel Ausloos, and Gurjeet Dhesi. Words ranking and Hirsch index for identifying the core of the hapaxes in political texts. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719303037.

Fu:2024:IKE

[FCZ24] Zhongmeng Fu, Yuan Cao, and Yong Zhao. Identifying knowledge evolution in computer science from the perspective of academic genealogy. Journal of Informetrics, 18(2): ??, May 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157724000361.

Franceschini:2012:FCA

[FGMM12] Fiorenzo Franceschini, Maurizio Galetto, Domenico Maisano, and Luca Mastrogiacomo. Further clarifications about the success-index. Journal of Informetrics, 6(4):669–673, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000569.

Franceschini:2013:IMS

[FGMM13] Fiorenzo Franceschini, Maurizio Galetto, Domenico Maisano, and Luca Mastrogiacomo. An informetric model for the success-index. Journal of Informetrics, 7(1):109–116, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157712000880.

Fan:2022:DAC

[FGW^{+}22] Lingxu Fan, Lei Guo, Xinhua Wang, Liancheng Xu, and Fangai Liu. Does the author’s collaboration mode lead to papers’ different citation impacts? An empirical analysis based on propensity score matching. Journal of Informetrics, 16(4): ??, November 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115772200102X.

Fu:2013:IRC

[FH13] Hui-Zhen Fu and Yuh-Shan Ho. Independent research of China in Science Citation Index Expanded during 1980–2011. Journal of Informetrics, 7(1):210–222, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000958.

Frenken:2009:SST

[FHH09] Koen Frenken, Sjoerd Hardeman, and Jarno Hoekman. Spatial scientometrics: Towards a cumulative research program. Journal of Informetrics, 3(3):222–232, July 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000303.

Frenken:2017:WDU

[FHH17] Koen Frenken, Gaston J. Heimeriks, and Jarno Hoekman. What drives university research performance? An analysis using the CWTS Leiden Ranking data. Journal of Informetrics, 11(3):859–872, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717301785.

Fiala:2012:TAP

[Fia12] Dalibor Fiala. Time-aware PageRank for bibliographic networks. Journal of Informetrics, 6(3):370–388, July 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000119.

Finardi:2013:CBJ

[Fin13] Ugo Finardi. Correlation between Journal Impact Factor and Citation Performance: an experimental study. Journal of Informetrics, 7(2):357–370, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712001150.

Finardi:2014:TER

[Fin14] Ugo Finardi. On the time evolution of received citations, in different scientific fields: an empirical study. Journal of Informetrics, 8(1):13–24, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000801.

Font-Julian:2022:WTO

[FJOROMT22] Cristina I Font-Julián, José-Antonio Ontalba-Ruipérez, Enrique Orduña-Malea, and Mike Thelwall. Which types of online resource support US patent claims? Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721001188.

Frandsen:2015:GDS

[FJW^{+}15] Tove Faber Frandsen, Rasmus Højbjerg Jacobsen, Johan A. Wallin, Kim Brixen, and Jakob Ousager. Gender differences in scientific performance: a bibliometric matching analysis of Danish health sciences graduates. Journal of Informetrics, 9(4):1007–1017, October 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715300316.

Fan:2022:EIS

[FLB22] Yangliu Fan, Sune Lehmann, and Anders Blok. Extracting the interdisciplinary specialty structures in social media data-based research: a clustering-based network approach. Journal of Informetrics, 16(3):??, August 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000621.

Feng:2023:HDS

[FLQ23] Sida Feng, Huajiao Li, and Yabin Qi. How to detect the sleeping beauty papers and princes in technology considering indirect citations? Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000561.

Franceschini:2010:CTO

[FM10a] Fiorenzo Franceschini and Domenico Maisano. The citation triad: an overview of a scientist’s publication output based on Ferrers diagrams. Journal of Informetrics, 4(4):503–511, October 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000428.

Franceschini:2010:HSN

[FM10b] Fiorenzo Franceschini and Domenico Maisano. The Hirsch spectrum: a novel tool for analyzing scientific journals. Journal of Informetrics, 4(1):64–73, January 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000704.

Franceschini:2011:RRO

[FM11a] Fiorenzo Franceschini and Domenico Maisano. Regularity in the research output of individual scientists: an empirical analysis by recent bibliometric tools. Journal of Informetrics, 5(3):458–468, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000411.

Franceschini:2011:SES

[FM11b] Fiorenzo Franceschini and Domenico Maisano. Structured evaluation of the scientific output of academic research groups by recent h-based indicators. Journal of Informetrics, 5(1):64–74, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157710000830.

Franceschini:2014:SFN

[FM14] Fiorenzo Franceschini and Domenico Maisano. Sub-field normalization of the IEEE scientific journals based on their connection with technical societies. Journal of Informetrics, 8(3):508–533, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714000443.

Franceschini:2017:CRI

[FM17a] Fiorenzo Franceschini and Domenico Maisano. Critical remarks on the Italian research assessment exercise VQR 2011– 2014. Journal of Informetrics, 11(2):337–357, May 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S1751157716303005. See comments [BCGM17b] and rejoinder [BCGM17b].

Franceschini:2017:RCB

[FM17b] Fiorenzo Franceschini and Domenico Maisano. A rejoinder to the comments of Benedetto et al. on the paper “Critical remarks on the Italian research assessment exercise VQR 2011–2014” (Journal of Informetrics, 11(2): 337–357). Journal of Informetrics, 11(3):645–646, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717301839. See [BCGM17b].

Franceschini:2014:SJP

[FMM14] Fiorenzo Franceschini, Domenico Maisano, and Luca Mastrogiacomo. Scientific journal publishers and omitted citations in bibliometric databases: Any relationship? Journal of Informetrics, 8(3):751–765, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000637.

Franceschini:2016:EAC

[FMM16a] Fiorenzo Franceschini, Domenico Maisano, and Luca Mastrogiacomo. Empirical analysis and classification of database errors in Scopus and Web of Science. Journal of Informetrics, 10(4):933–953, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S175115771630061X.

Franceschini:2016:MEH

[FMM16b] Fiorenzo Franceschini, Domenico Maisano, and Luca Mastrogiacomo. The museum of errors/horrors in Scopus. Journal of Informetrics, 10(1):174–182, February 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715301462.

Frandsen:2010:WNC

[FN10] Tove Faber Frandsen and Jeppe Nicolaisen. What is in a name? Credit assignment practices in different disciplines. Journal of Informetrics, 4(4):608–617, October 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000647.

Frandsen:2007:JSC

[Fra07] Tove Faber Frandsen. Journal self-citations — analysing the JIF mechanism. Journal of Informetrics, 1(1):47–58, January 2007. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157706000071.

Frandsen:2009:EOA

[Fra09] Tove Faber Frandsen. The effects of open access on unpublished documents: a case study of economics working papers. Journal of Informetrics, 3(2):124–133, April 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157708000746.

Franceschet:2010:DBP

[Fra10a] Massimo Franceschet. The difference between popularity and prestige in the sciences and in the social sciences: a bibliometric analysis. Journal of Informetrics, 4(1):55–63, January 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157709000698.

Franceschet:2010:JIF

[Fra10b] Massimo Franceschet. Journal influence factors. Journal of Informetrics, 4(3):239–248, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000984.

Fiala:2015:DPB

[FSZB15] Dalibor Fiala, Lovro Subelj, Slavko Zitnik, and Marko Bajec. Do PageRank-based author rankings outperform simple citation counts? Journal of Informetrics, 9(2):334–348, April 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715000267.

Fairclough:2015:MPM

[FT15a] Ruth Fairclough and Mike Thelwall. More precise methods for national research citation impact comparisons. Journal of Informetrics, 9(4):895–906, October 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715300894.

Fairclough:2015:NRI

[FT15b] Ruth Fairclough and Mike Thelwall. National research impact indicators from Mendeley readers. Journal of Informetrics, 9(4):845–859, October 2015. CODEN ????

ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157715300596.

Fiala:2017:PBP

[FT17] Dalibor Fiala and Gabriel Tutoky. PageRank-based prediction of award-winning researchers and the impact of citations. Journal of Informetrics, 11(4):1044–1068, November 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S175115771730038X.

Feng:2015:VQS

[FZDW15] Feng Feng, Leiyong Zhang, Yuneng Du, and Weiguang Wang. Visualization and quantitative study in bibliographic databases: a case in the field of university-industry cooperation. Journal of Informetrics, 9(1):118–134, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714001114.

Feng:2023:TPP

[FZZ23] Xiaodong Feng, Qihang Zhao, and RuiJie Zhu. Towards popularity prediction of information cascades via degree distribution and deep neural networks. Journal of Informetrics, 17 (3):??, August 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115772300038X.

Gonzalez-Albo:2011:AVP

[GAB11] Borja González-Albo and María Bordons. Articles vs. proceedings papers: Do they differ in research relevance and impact? A case study in the Library and Information Science field. Journal of Informetrics, 5(3):369–381, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157711000125.

Gonzalez-Alvarez:2017:RPH

[GÁCC17] Julio González-Álvarez and Teresa Cervera-Crespo. Research production in high-impact journals of contemporary neuroscience: a gender analysis. Journal of Informetrics, 11(1):232–243, February 2017. CODEN ????

ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157716302462.

Giuffrida:2019:ACW

[GAD19] Cristiano Giuffrida, Giovanni Abramo, and Ciriaco Andrea D’Angelo. Are all citations worth the same? Valuing citations by the value of the citing items. Journal of Informetrics, 13(2):500–514, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302426.

Gagolewski:2009:RLP

[Gag09] Marek Gagolewski. A remark on limit properties of generalized h- and g-indices. Journal of Informetrics, 3(4):367– 368, October 2009. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157709000121.

Gagolewski:2011:BIA

[Gag11] Marek Gagolewski. Bibliometric impact assessment with R and the CITAN package. Journal of Informetrics, 5(4):678– 692, October 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711000708.

Gagolewski:2013:SIA

[Gag13] Marek Gagolewski. Scientific impact assessment cannot be fair. Journal of Informetrics, 7(4):792–802, October 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000540.

Garfield:2009:SSS

[Gar09] Eugene Garfield. From the science of science to scientometrics visualizing the history of science with HistCite software. Journal of Informetrics, 3(3):173–179, July 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000297.

Gauffriau:2017:CAC

[Gau17] Marianne Gauffriau. A categorization of arguments for counting methods for publication and citation indicators. Journal of Informetrics, 11(3):672–684, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302760.

Ghimire:2021:DND

[GAW21] Sarad Ghimire, Saman Hassanzadeh Amin, and Leslie J. Wardley. Developing new data envelopment analysis models to evaluate the efficiency in Ontario universities. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000432.

Gonzalez-Betancor:2023:DSS

[GBDG23] Sara M. González-Betancor and Pablo Dorta-González. Does society show differential attention to researchers based on gender and field? Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000779.

Guerrero-Bote:2012:FSF

[GBMA12] Vicente P. Guerrero-Bote and Félix Moya-Anegón. A further step forward in measuring journals’ scientific prestige: the SJR2 indicator. Journal of Informetrics, 6(4):674–688, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000521.

Gomez-Deniz:2024:MCC

[GDDG24] Emilio Gómez-Déniz and Pablo Dorta-González. Modeling citation concentration through a mixture of Leimkuhler curves. Journal of Informetrics, 18(2):??, May 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000324.

Gomez-Ferri:2021:MDC

[GFGALG21] Javier Gómez-Ferri, Gregorio González-Alcaide, and Ramón LLopis-Goig. Measuring dissatisfaction with coauthorship: an empirical approach based on the researchers’ perception. Journal of Informetrics, 13(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719301336.

Gao:2009:SIA

[GG09] Xia Gao and Jiancheng Guan. A scale-independent analysis of the performance of the Chinese innovation system. Journal of Informetrics, 3(4):321–331, October 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000364.

Galiani:2019:EAB

[GG19] Sebastian Galiani and Ramiro H. Gálvez. An empirical approach based on quantile regression for estimating citation ageing. Journal of Informetrics, 13(2):738–750, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718303419.

Gao:2010:CIS

[GGSG10] Xia Gao, Xiaochuan Guo, Katz J. Sylvan, and Jiancheng Guan. The Chinese innovation system during economic transition: a scale-independent view. Journal of Informetrics, 4(4):618–628, October 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157710000659.

Gomez:2021:MMC

[GHP21] Charles J. Gomez, Andrew C. Herman, and Paolo Parigi. Moving more, but closer: Mapping the growing regionalization of global scientific mobility using ORCID. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719304389.

Gyorffy:2021:RFP

[GHS21] Balázs Győrffy, Péter Herman, and István Szabó. Research funding: past performance is a stronger predictor of future scientific output than reviewer scores. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719303694.

Gurney:2014:AKC

[GHvdB^{+}14] Thomas Gurney, Edwin Horlings, Peter van den Besselaar, Koichi Sumikura, Antoine Schoen, Patricia Laurens, and Daniel Pardo. Analysing knowledge capture mechanisms: Methods and a stylised bioventure case. Journal of Informetrics, 8(1):259–272, January 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713001168.

Gizinski:2024:BTI

[GKR^{+}24] Stanisław Giziński, Paulina Kaczyńska, Hubert Ruczyński, Emilia Wiśnios, Bartosz Pieliński, Przemysław Biecek, and Julian Sienkiewicz. Big tech influence over AI research revisited: Memetic analysis of attribution of ideas to affiliation. Journal of Informetrics, 18(4):??, November 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000841.

Gingras:2011:TNK

[GL11] Yves Gingras and Vincent Larivière. There are neither “king” nor “crown” in scientometrics: Comments on a supposed “alternative” method of normalization. Journal of Informetrics, 5(1):226–227, January 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000921.

Glanzel:2007:CSS

[Glä07] Wolfgang Glänzel. Characteristic scores and scales: a bibliometric analysis of subject characteristics based on long-term citation observation. Journal of Informetrics, 1(1):92–102, January 2007. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157706000125.

Glanzel:2010:RRS

[Glä10] Wolfgang Glänzel. On reliability and robustness of scientometrics indicators based on stochastic models. An evidencebased opinion paper. Journal of Informetrics, 4(3):313–319, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000064.

Glaser:2017:FEQ

[Glä17] Jochen Gläser. A fight on epistemological quicksand: Comment on the dispute between van den Besselaar et al. and Butler. Journal of Informetrics, 11(3):927–932, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157717301979.

Gao:2021:PIR

[GLW^{+}21] Qiang Gao, Zhentao Liang, Ping Wang, Jingrui Hou, Xiuxiu Chen, and Manman Liu. Potential index: Revealing the future impact of research topics based on current knowledge networks. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157721000365.

Guo:2021:PDA

[GLY21] Xiaolong Guo, Xiaoxiao Li, and Yugang Yu. Publication delay adjusted impact factor: the effect of publication delay of articles on journal impact factor. Journal of Informetrics, 15 (1):??, February 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157720301474.

Gagolewski:2012:ADP

[GM12] Marek Gagolewski and Radko Mesiar. Aggregating different paper quality measures with a generalized h-index. Journal of Informetrics, 6(4):566–579, October 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000417.

Gorraiz:2016:ADO

[GMFGVZ16] Juan Gorraiz, David Melero-Fuentes, Christian Gumpenberger, and Juan-Carlos Valderrama-Zurián. Availability of digital object identifiers (DOIs) in Web of Science and Scopus. Journal of Informetrics, 10(1):98–109, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715301176.

Garcia:2023:HMR

[GMPRSFV23] J. A. García, J. J. Montero-Parodi, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. How to motivate a reviewer with a present bias to work harder. Journal of Informetrics, 17(4): ??, November 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000871.

Gomez-Nunez:2014:OSJ

[GNBVQ^{+}14] Antonio J. Gómez-Núñez, Vladimir Batagelj, Benjamín Vargas-Quesada, Félix Moya-Anegón, and Zaida Chinchilla-Rodríguez. Optimizing SCImago Journal & Country Rank classification by community detection. Journal of Informetrics, 8(2):369–383, April 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714000182.

Giatsidis:2019:RCG

[GNZ^{+}19] Christos Giatsidis, Giannis Nikolentzos, Chenhui Zhang, Jie Tang, and Michalis Vazirgiannis. Rooted citation graphs density metrics for research papers influence evaluation. Journal of Informetrics, 13(2):757–768, May 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718303626.

Goldman:2014:CID

[Gol14] Alyssa W. Goldman. Conceptualizing the interdisciplinary diffusion and evolution of emerging fields: the case of systems biology. Journal of Informetrics, 8(1):43–58, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000862.

Garcia-Perez:2011:SAW

[GP11] Miguel A. García-Pérez. Strange attractors in the Web of Science database. Journal of Informetrics, 5(1):214–218, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000702.

Garcia-Perez:2012:EIC

[GP12] Miguel A. García-Pérez. An extension of the h index that covers the tail and the top of the citation curve and allows ranking researchers with similar h. Journal of Informetrics, 6(4):689–699, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000557.

Gonzalez-Pereira:2010:NAM

[GPGBMA10] Borja González-Pereira, Vicente P. Guerrero-Bote, and Félix Moya-Anegón. A new approach to the metric of journals’ scientific prestige: the SJR indicator. Journal of Informetrics, 4(3):379–391, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000246.

Garcia-Perez:2013:CBV

[GPNA13] Miguel A. García-Pérez and Vicente Núñez-Antón. Correlation between variables subject to an order restriction, with application to scientometric indices. Journal of Informetrics, 7(2):542–554, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000126.

Gu:2024:IMM

[GPZC24] Jiangwei Gu, Xuelian Pan, Shuxin Zhang, and Jiaoyu Chen. International mobility matters: Research collaboration and scientific productivity. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115772400035X.

Guns:2009:RRV

[GR09] Raf Guns and Ronald Rousseau. Real and rational variants of the h-index and the g-index. Journal of Informetrics, 3(1): 64–71, January 2009. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157708000692.

Glanzel:2010:HTC

[GS10] Wolfgang Glänzel and András Schubert. Hirsch-type characteristics of the tail of distributions. The generalised h-index. Journal of Informetrics, 4(1):118–123, January 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000765.

GomesdeCarvalho:2021:BSR

[GSV^{+}21] Gustavo Dambiski Gomes de Carvalho, Carla Cristiane Sokulski, Wesley Vieira da Silva, Hélio Gomes de Carvalho, Rafael Vignoli de Moura, Antonio Carlos de Francisco, and Claudimar Pereira da Veiga. Bibliometrics and systematic reviews: a comparison between the Proknow-C and the Methodi Ordinatio. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719303682.

Glanzel:2016:PPE

[GTD16] Wolfgang Glänzel, Bart Thijs, and Koenraad Debackere. Productivity, performance, efficiency, impact — what do we measure anyway?: Some comments on the paper “A farewell to the MNCS and like size-independent indicators” by Abramo and D’Angelo. Journal of Informetrics, 10(2):658– 660, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716300980. See [AD16b,

Zit16, AD16c].

Guns:2017:DEN

[GW17] Raf Guns and Lili Wang. Detecting the emergence of new scientific collaboration links in Africa: a comparison of expected and realized collaboration intensities. Journal of Informetrics, 11(3):892–903, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303790.

Guo:2024:EEF

[GWL24] Liying Guo, Yang Wang, and Meiling Li. Exploration, exploitation and funding success: Evidence from junior scientists supported by the Chinese Young Scientists Fund. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000051.

Guler:2016:ABA

[GWMP16] Arzu Tugce Guler, Cathelijn J. F. Waaijer, Yassene Mohammed, and Magnus Palmblad. Automating bibliometric analyses using Taverna scientific workflows: a tutorial on integrating Web services. Journal of Informetrics, 10(3):830– 841, August 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157716300141.

Guan:2017:ICK

[GYZ17] Jiancheng Guan, Yan Yan, and Jing Jing Zhang. The impact of collaboration and knowledge networks on citations. Journal of Informetrics, 11(2):407–422, May 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716301444.

Garcia-Zorita:2018:RDV

[GZRMLSC18] Carlos Garcia-Zorita, Ronald Rousseau, Sergio Marugan-Lazaro, and Elias Sanz-Casado. Ranking dynamics and volatility. Journal of Informetrics, 12(3):567–578, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157717303553.

Hagen:2013:HCC

[Hag13] Nils T. Hagen. Harmonic coauthor credit: a parsimonious quantification of the byline hierarchy. Journal of Informetrics, 7(4):784–791, October 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000527.

Hagen:2014:CCP

[Hag14a] Nils T. Hagen. Counting and comparing publication output with and without equalizing and inflationary bias. Journal of Informetrics, 8(2):310–317, April 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000042.

Hagen:2014:RBH

[Hag14b] Nils T. Hagen. Reversing the byline hierarchy: the effect of equalizing bias on the accreditation of primary, secondary and senior authors. Journal of Informetrics, 8(3):618–627, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157714000510.

Hagen:2015:CIA

[Hag15] Nils T. Hagen. Contributory inequality alters assessment of academic output gap between comparable countries. Journal of Informetrics, 9(3):629–641, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715200302.

Han:2007:MIK

[Han07] Yoo-Jin Han. Measuring industrial knowledge stocks with patents and papers. Journal of Informetrics, 1(4):269–276, ???? 2007. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157707000478.

Haunschild:2016:NMR

[HB16] Robin Haunschild and Lutz Bornmann. Normalization of Mendeley reader counts for impact assessment. Journal of Informetrics, 10(1):62–73, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771520051X.

Haunschild:2023:IPY

[HB23a] Robin Haunschild and Lutz Bornmann. Identification of potential young talented individuals in the natural and life sciences: a bibliometric approach. Journal of Informetrics, 17 (3):??, August 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000196.

Haunschild:2023:WPC

[HB23b] Robin Haunschild and Lutz Bornmann. Which papers cited which tweets? An exploratory analysis based on Scopus data. Journal of Informetrics, 17(2):??, May 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000081.

Huang:2023:WMS

[HbHB^{+}23] Junjie Huang, Win bin Huang, Yi Bu, Qi Cao, Huawei Shen, and Xueqi Cheng. What makes a successful rebuttal in computer science conferences?: a perspective on social interaction. Journal of Informetrics, 17(3):??, August 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000524.

Huang:2012:TCS

[HCC12] Mu-Hsuan Huang, Han-Wen Chang, and Dar-Zen Chen. The trend of concentration in scientific research and technological innovation: a reduction of the predominant role of the U.S. in world research & technology. Journal of Informetrics, 6(4):457–468, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000259.

Hu:2013:WCL

[HCL13] Zhigang Hu, Chaomei Chen, and Zeyuan Liu. Where are citations located in the body of scientific articles? A study of the distributions of citation locations. Journal of Informetrics, 7(4):887–896, October 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000692.

Hwang:2019:CRA

[HCS19] Yan-An Hwang, Chih-Hao Chiu, and Jian-Ming Shih. A correction: Ranking authors using fractional counting of citations: an axiomatic approach. Journal of Informetrics, 13(3):769–770, August 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157719300653. See [BM16].

Huang:2021:PCI

[HCZ21] Ying Huang, Lixin Chen, and Lin Zhang. Patent citation inflation: the phenomenon, its measurement, and relative indicators to temper its effects. Journal of Informetrics, 14 (2):??, May 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771830484X.

Haunschild:2022:SSF

[HDB22] Robin Haunschild, Angela D. Daniels, and Lutz Bornmann. Scores of a specific field-normalized indicator calculated with different approaches of field-categorization: Are the scores different or similar? Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721001127.

Huang:2012:GCC

[HDC12] Mu-Hsuan Huang, Huei-Ru Dong, and Dar-Zen Chen. Globalization of collaborative creativity through cross-border patent activities. Journal of Informetrics, 6(2):226–236, April 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115771100099X.

Hu:2021:RBD

[HDD^{+}21] Beibei Hu, Yang Ding, Xianlei Dong, Yi Bu, and Ying Ding. On the relationship between download and citation counts: an introduction of Granger-causality inference. Journal of Informetrics, 15(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720306428.

He:2013:MDS

[HDT^{+}13] Bing He, Ying Ding, Jie Tang, Vignesh Reguramalingam, and Johan Bollen. Mining diversity subgraph in multidisciplinary scientific collaboration networks: a meso perspective. Journal of Informetrics, 7(1):117–128, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157712000806.

He:2012:MPA

[HDY12] Bing He, Ying Ding, and Erjia Yan. Mining patterns of author orders in scientific publications. Journal of Informetrics, 6(3):359–367, July 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000028.

Harzing:2014:CAN

[HG14] Anne-Wil Harzing and Axèle Giroud. The competitive advantage of nations: an application to academia. Journal of Informetrics, 8(1):29–42, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000849.

Higham:2017:UDG

[HGJZ17] K. W. Higham, M. Governale, A. B. Jaffe, and U. Zülicke. Unraveling the dynamics of growth, aging and inflation for citations to scientific articles from specific research fields. Journal of Informetrics, 11(4):1190–1200, November 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717302699.

Haddawy:2016:CER

[HHAA16] Peter Haddawy, Saeed-Ul Hassan, Awais Asghar, and Sarah Amin. A comprehensive examination of the relation of three citation-based journal metrics to expert judgment of journal quality. Journal of Informetrics, 10(1):162–173, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715301164.

Haddawy:2017:UFG

[HHAL17] Peter Haddawy, Saeed-Ul Hassan, Craig W. Abbey, and Inn Beng Lee. Uncovering fine-grained research excellence: the global research benchmarking system. Journal of Informetrics, 11(2):389–406, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715300146.

Huang:2023:DIM

[HHB^{+}23] Shengzhi Huang, Yong Huang, Yi Bu, Zhuoran Luo, and Wei Lu. Disclosing the interactive mechanism behind scientists’ topic selection behavior from the perspective of the productivity and the impact. Journal of Informetrics, 17 (2):??, May 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000342.

Huang:2014:TIF

[HHC14] Mu-Hsuan Huang, Wei-Tzu Huang, and Dar-Zen Chen. Technological impact factor: an indicator to measure the impact of academic publications on practical innovation. Journal of Informetrics, 8(1):241–251, January 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713001132.

Hicks:2017:WYD

[Hic17] Diana Hicks. What year? Difficulties in identifying the effect of policy on university output. Journal of Informetrics, 11(3):933–936, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157717301980.

Hanssen:2015:VER

[HJ15] Thor-Erik Sandberg Hanssen and Finn Jørgensen. The value of experience in research. Journal of Informetrics, 9(1):16– 24, January 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157714000972.

Henneken:2009:UAL

[HKA^{+}09] Edwin A. Henneken, Michael J. Kurtz, Alberto Accomazzi, Carolyn S. Grant, Donna Thompson, Elizabeth Bohlen, and Stephen S. Murray. Use of astronomical literature — a report on usage patterns. Journal of Informetrics, 3(1):1–8, January 2009. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157708000515.

Hric:2018:SBM

[HKK18] Darko Hric, Kimmo Kaski, and Mikko Kivelä. Stochastic block model reveals maps of citation patterns and their evolution in time. Journal of Informetrics, 12(3):757–783, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157717301499.

Heo:2023:HAI

[HKXS23] Go Eun Heo, Young Soo Ko, Qing Xie, and Min Song. High acknowledgement index: Characterizing research supporters with factors of acknowledgement affecting paper citation counts. Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115772300072X.

Huang:2019:SNE

[HL19] Meiting Huang and Weishu Liu. Substantial numbers of easily identifiable illegal DOIs still exist in Scopus. Journal of Informetrics, 13(3):901–903, August 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719300938.

Haunschild:2019:DPD

[HLB^{+}19] Robin Haunschild, Loet Leydesdorff, Lutz Bornmann, Iina Hellsten, and Werner Marx. Does the public discuss other topics on climate change than researchers? A comparison of explorative networks based on author keywords and hashtags. Journal of Informetrics, 13(2):695–707, May 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S175115771830419X. See corrigendum [HLB^{+}21].

Haunschild:2021:CDP

[HLB^{+}21] Robin Haunschild, Loet Leydesdorff, Lutz Bornmann, Iina Hellsten, and Werner Marx. Corrigendum to “Does the public discuss other topics on climate change than researchers? A comparison of explorative networks based on author keywords and hashtags” [J. Informetrics 13 (2019) 695–707]. Journal of Informetrics, 14(1):??, February 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720300390. See [HLB^{+}19].

He:2023:RFM

[HLD^{+}23] Chaocheng He, Fuzhen Liu, Ke Dong, Jiang Wu, and Qingpeng Zhang. Research on the formation mechanism of research leadership relations: an exponential random graph model analysis approach. Journal of Informetrics, 17(2):??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000263.

Huang:2022:TMC

[HLHC22] Chen-Hao Huang, John S. Liu, Mei Hsiu-Ching Ho, and Tzu-Chuan Chou. Towards more convergent main paths: a relevance-based approach. Journal of Informetrics, 16(3): ??, August 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722000694.

Hu:2017:HUN

[HLR17] Xiaojun Hu, Loet Leydesdorff, and Ronald Rousseau. Heterogeneity in an undirected network: Definition and measurement. Journal of Informetrics, 11(2):669–682, May 2017.

CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157717301074.

He:2024:PED

[HLRD24] Guoxiu He, Chenxi Lin, Jiayu Ren, and Peichen Duan. Predicting the emergence of disruptive technologies by comparing with references via soft prompt-aware shared BERT. Journal of Informetrics, 18(4):??, November 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724001081.

Hu:2017:UMM

[HLSH17] Zhigang Hu, Gege Lin, Taian Sun, and Haiyan Hou. Understanding multiply mentioned references. Journal of Informetrics, 11(4):948–958, November 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717300482.

Hou:2024:IIS

[HLZ24] Jianhua Hou, Hao Li, and Yang Zhang. Influence of interdisciplinarity of scientific papers on the durability of citation diffusion: a perspective from citation discontinuance. Journal of Informetrics, 18(3):??, August 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000385.

Haley:2021:JVJ

[HM21a] M. Ryan Haley and M. Kevin McGee. Jointly valuing journal visibility and author citation count: an axiomatic approach. Journal of Informetrics, 14(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115771930152X.

Hladchenko:2021:EPT

[HM21b] Myroslava Hladchenko and Henk F. Moed. The effect of publication traditions and requirements in research assessment and funding policies upon the use of national journals in 28 post-socialist countries. Journal of Informetrics, 15(4):

??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000614.

Halevi:2017:SGS

[HMBI17] Gali Halevi, Henk Moed, and Judit Bar-Ilan. Suitability of Google Scholar as a source of scientific information and as a source of data for scientific evaluation — review of the literature. Journal of Informetrics, 11(3):823–834, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157717300676.

Hossain:2013:CND

[HMU13] Liaquat Hossain, Shahriar Tanvir Murshed, and Shahadat Uddin. Communication network dynamics during organizational crisis. Journal of Informetrics, 7(1):16–35, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157712000570.

Hu:2022:DIN

[HNT22] Guangyuan Hu, Rong Ni, and Li Tang. Do international nonstop flights foster influential research? Evidence from Sino– US scientific collaboration. Journal of Informetrics, 16(4): ??, November 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722001006.

Hren:2022:WMB

[HPNM22] Darko Hren, David G. Pina, Christopher R. Norman, and Ana Marusić. What makes or breaks competitive research proposals? a mixed-methods analysis of research grant evaluation reports. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722000414.

Hou:2021:ISE

[HPZ21] Lei Hou, Yueling Pan, and Jonathan J. H. Zhu. Impact of scientific, economic, geopolitical, and cultural factors on international research collaboration. Journal of Informetrics, 15 (3):??, August 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000651.

Hu:2016:SIA

[HR16] Xiaojun Hu and Ronald Rousseau. Scientific influence is not always visible: the phenomenon of under-cited influential publications. Journal of Informetrics, 10(4):1079–1091, November 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157716301900.

Hu:2018:NAE

[HR18] Xiaojun Hu and Ronald Rousseau. A new approach to explore the knowledge transition path in the evolution of science and technology: From the biology of restriction enzymes to their application in biotechnology. Journal of Informetrics, 12(3):842–857, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718301718.

Hu:2011:DFB

[HRC11a] Xiaojun Hu, Ronald Rousseau, and Jin Chen. On the definition of forward and backward citation generations. Journal of Informetrics, 5(1):27–36, January 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000684.

Hu:2011:TSO

[HRC11b] Xiaojun Hu, Ronald Rousseau, and Jin Chen. Time series of outgrow indices. Journal of Informetrics, 5(3):413–421, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711000290.

Herranz:2012:SFN

[HRC12] Neus Herranz and Javier Ruiz-Castillo. Sub-field normalization in the multiplicative case: Average-based citation indicators. Journal of Informetrics, 6(4):543–556, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157712000223.

Hennemann:2012:MGS

[HRL12] Stefan Hennemann, Diego Rybski, and Ingo Liefner. The myth of global science collaboration-collaboration patterns in epistemic communities. Journal of Informetrics, 6(2):217– 225, April 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711001088.

Haustein:2011:ASB

[HS11] Stefanie Haustein and Tobias Siebenlist. Applying social bookmarking data to evaluate journal usage. Journal of Informetrics, 5(3):446–457, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000393.

Haunschild:2016:PMC

[HSB16] Robin Haunschild, Hermann Schier, and Lutz Bornmann. Proposal of a minimum constraint for indicators based on means or averages. Journal of Informetrics, 10(2):485–486, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716300694.

Hajibabaei:2022:GSP

[HSE22] Anahita Hajibabaei, Andrea Schiffauerova, and Ashkan Ebadi. Gender-specific patterns in the artificial intelligence scientific ecosystem. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115772200027X.

Haunschild:2018:AGS

[HSMB18] Robin Haunschild, Hermann Schier, Werner Marx, and Lutz Bornmann. Algorithmically generated subject categories based on citation relations: an empirical micro study using papers on overall water splitting. Journal of Informetrics, 12(2):436–447, May 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S175115771730278X.

Horbach:2022:UWW

[HSSM22] Serge P. J. M. Horbach, Jesper W. Schneider, and Maxime Sainte-Marie. Ungendered writing: Writing styles are unlikely to account for gender differences in funding rates in the natural and technical sciences. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000840.

Hu:2021:IHC

[HTLC21] Ya-Han Hu, Chun-Tien Tai, Kang Ernest Liu, and Cheng-Fang Cai. Identification of highly-cited papers using topicmodel-based and bibliometric features: the consideration of keyword popularity. Journal of Informetrics, 14(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719301099.

Hu:2018:FPN

[HTX^{+}18] Zhigang Hu, Wencan Tian, Shenmeng Xu, Chunbo Zhang, and Xianwen Wang. Four pitfalls in normalizing citation indicators: an investigation of ESI’s selection of highly cited papers. Journal of Informetrics, 12(4):1133–1145, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718301147.

Hou:2023:DPK

[HTZS23] Jianhua Hou, Shiqi Tang, Yang Zhang, and Haoyang Song. Does prior knowledge affect patent technology diffusion? A semantic-based patent citation contribution analysis. Journal of Informetrics, 17(2):??, May 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000184.

Hurlimann:2015:URU

[Hür15] Werner Hürlimann. On the uniform random upper bound family of first significant digit distributions. Journal of Informetrics, 9(2):349–358, April 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157715000255.

Hur:2024:EHT

[Hur24] Wonchang Hur. Entropy, heterogeneity, and their impact on technology progress. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157724000191.

Hu:2014:RTD

[HW14] Zewen Hu and Yishan Wu. Regularity in the time-dependent distribution of the percentage of never-cited papers: an empirical pilot study based on the six journals. Journal of Informetrics, 8(1):136–146, January 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000904.

Hou:2022:NMM

[HWL22] Jianhua Hou, Dongyi Wang, and Jing Li. A new method for measuring the originality of academic articles based on knowledge units in semantic networks. Journal of Informetrics, 16(3):??, August 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772200058X.

Hou:2024:DOI

[HWX24] Li Hou, Qiang Wu, and Yundong Xie. Does open identity of peer reviewers positively relate to citations? Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000026.

Hou:2022:HDS

[HWZW22] Jianhua Hou, Yuanyuan Wang, Yang Zhang, and Dongyi Wang. How do scholars and non-scholars participate in dataset dissemination on Twitter. Journal of Informetrics, 16 (1):??, February 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000948.

Habibzadeh:2008:JWI

[HY08] Farrokh Habibzadeh and Mahboobeh Yadollahie. Journal weighted impact factor: a proposal. Journal of Informetrics, 2(2):164–172, April 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S175115770800014X. See comments [WvE08].

Hou:2021:SMB

[HY21] Jianhua Hou and Xiucai Yang. Social media-based sleeping beauties: Defining, identifying and features. Journal of Informetrics, 14(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157719303049.

Huang:2015:IST

[HYC15] Mu-Hsuan Huang, Hsiao-Wen Yang, and Dar-Zen Chen. Increasing science and technology linkage in fuel cells: a cross citation analysis of papers and patents. Journal of Informetrics, 9(2):237–249, April 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771500019X.

Huang:2021:ICP

[HYS21] Cui Huang, Chao Yang, and Jun Su. Identifying core policy instruments based on structural holes: a case study of China’s nuclear energy policy. Journal of Informetrics, 15 (2):??, May 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115772100016X.

Hu:2023:EIP

[HZL23] Zewen Hu, Xiji Zhou, and Angela Lin. Evaluation and identification of potential high-value patents in the field of integrated circuits using a multidimensional patent indicators pre-screening strategy and machine learning approaches. Journal of Informetrics, 17(2):??, May 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000317.

He:2019:MEG

[HZW19] Zekai He, Ni Zhen, and Chaojiang Wu. Measuring and exploring the geographic mobility of American professors from graduating institutions: Differences across disciplines, academic ranks, and genders. Journal of Informetrics, 13(3):771–784, August 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157718302785.

Hou:2023:HBS

[HZW^{+}23] Jianhua Hou, Bili Zheng, Dongyi Wang, Yang Zhang, and Chaomei Chen. How boundary-spanning paper sparkles citation: From citation count to citation network. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000597.

Han:2024:IMC

[HZZY24] Fang Han, Ruhao Zhang, Shengtai Zhang, and Junpeng Yuan. International mobility characteristics, effects of, and effects on elite scientists. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723001104.

Inoue:2021:JEN

[IPS21] Masaaki Inoue, Thong Pham, and Hidetoshi Shimodaira. Joint estimation of non-parametric transitivity and preferential attachment functions in scientific co-authorship networks. Journal of Informetrics, 14(3):??, August 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771930269X.

Inoue:2010:ACR

[IST10] Hiroyasu Inoue, Wataru Souma, and Schumpeter Tamada. Analysis of cooperative research and development networks on Japanese patents. Journal of Informetrics, 4(1):89–96, January 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115770900073X.

Jalili:2011:EAT

[Jal11] Mahdi Jalili. Error and attack tolerance of small-worldness in complex networks. Journal of Informetrics, 5(3):422–430, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711000332.

Jarneving:2007:BCA

[Jar07a] Bo Jarneving. Bibliographic coupling and its application to research-front and other core documents. Journal of Informetrics, 1(4):287–307, ???? 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157707000594.

Jarneving:2007:CGB

[Jar07b] Bo Jarneving. Complete graphs and bibliographic coupling: a test of the applicability of bibliographic coupling for the identification of cognitive cores on the field level. Journal of Informetrics, 1(4):338–356, ???? 2007. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157707000612.

Jin:2024:HDN

[JCWX24] Qianqian Jin, Hongshu Chen, Xuefeng Wang, and Fei Xiong. How do network embeddedness and knowledge stock influence collaboration dynamics? Evidence from patents. Journal of Informetrics, 18(4):??, November 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772400066X.

Jiang:2023:DLP

[JFZZ23] Hongxun Jiang, Shaokun Fan, Nan Zhang, and Bin Zhu. Deep learning for predicting patent application outcome: the fusion of text and network embeddings. Journal of Informetrics, 17(2):??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157723000275.

Jeong:2016:TAD

[JHK^{+}16] Yoo Kyung Jeong, Go Eun Heo, Keun Young Kang, Dong Sup Yoon, and Min Song. Trajectory analysis of drugresearch trends in pancreatic cancer on PubMed and ClinicalTrials.gov. Journal of Informetrics, 10(1):273–285, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715301280.

Jeon:2024:ARL

[JJ24] Jong Hee Jeon and Jason J. Jung. Article ranking with location-based weight in contextual citation network. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724001032.

Jee:2021:EFC

[JKHS21] Su Jung Jee, Minji Kwon, Jung Moon Ha, and So Young Sohn. Exploring the forward citation patterns of patents based on the evolution of technology fields. Journal of Informetrics, 13(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157719301191.

Jeon:2023:MNS

[JLAL23] Daeseong Jeon, Junyoup Lee, Joon Mo Ahn, and Changyong Lee. Measuring the novelty of scientific publications: a fastText and local outlier factor approach. Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000755.

Jiang:2023:DRL

[JLH23] Zhuoren Jiang, Tianqianjin Lin, and Cui Huang. Deep representation learning of scientific paper reveals its potential scholarly impact. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000019.

Jensen:2016:GTE

[JLYM16] Scott Jensen, Xiaozhong Liu, Yingying Yu, and Stasa Milojevic. Generation of topic evolution trees from heterogeneous bibliographic networks. Journal of Informetrics, 10(2):606– 621, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715302145.

Jee:2022:PPI

[JPL22] Jeonghun Jee, Sanghyun Park, and Sungjoo Lee. Potential of patent image data as technology intelligence source. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000153.

Jeong:2014:TGA

[JS14] Do-Heon Jeong and Min Song. Time gap analysis by the topic model-based temporal technique. Journal of Informetrics, 8(3):776–790, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714000650.

Jung:2022:DDA

[JS22] Sukhwan Jung and Aviv Segev. DAC: Descendant-aware clustering algorithm for network-based topic emergence prediction. Journal of Informetrics, 16(3):??, August 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000724.

Jeong:2014:CBA

[JSD14] Yoo Kyung Jeong, Min Song, and Ying Ding. Contentbased author co-citation analysis. Journal of Informetrics, 8(1):197–211, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157713001107.

Jablonska-Sabuka:2014:FRT

[JSSK14] Matylda Jablońska-Sabuka, Robert Sitarz, and Andrzej Kraslawski. Forecasting research trends using population dynamics model with Burgers’ type interaction. Journal of Informetrics, 8(1):111–122, January 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713001041.

Jang:2017:HPB

[JWL17] Hyun Jin Jang, Han-Gyun Woo, and Changyong Lee. Hawkes process-based technology impact analysis. Journal of Informetrics, 11(2):511–529, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303169.

Jeong:2021:EDS

[JXYS21] Yoo Kyung Jeong, Qing Xie, Erjia Yan, and Min Song. Examining drug and side effect relation using author-entity pair bipartite networks. Journal of Informetrics, 14(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719302354.

Jung:2021:ATM

[JY21] Sukhwan Jung and Wan Chul Yoon. An alternative topic model based on Common Interest Authors for topic evolution analysis. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719303517.

Jiang:2024:MII

[JYG24] Man Jiang, Siluo Yang, and Qiang Gao. Multidimensional indicators to identify emerging technologies: Perspective of technological knowledge flow. Journal of Informetrics, 18(1): ??, February 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723001086.

Jiang:2021:FSP

[JZ21] Xiaorui Jiang and Hai Zhuge. Forward search path count as an alternative indirect citation impact indicator. Journal of Informetrics, 13(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719300094.

Jiang:2024:ORD

[JZDZ24] Huihuang Jiang, Jianlin Zhou, Yiming Ding, and An Zeng. Overcoming recognition delays in disruptive research: the impact of team size, familiarity, and reputation. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000622.

Jiang:2018:RBD

[JZY^{+}18] Ling (Alice) Jiang, Nibing Zhu, Zhilin Yang, Shen Xu, and Minjoon Jun. The relationships between distance factors and international collaborative research outcomes: a bibliometric examination. Journal of Informetrics, 12(3):618–630, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157717300469.

Kakushadze:2016:ISD

[Kak16] Zura Kakushadze. An index for SSRN downloads. Journal of Informetrics, 10(1):9–28, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715301851.

Karpov:2014:EWC

[Kar14] Alexander Karpov. Equal weights coauthorship sharing and the Shapley value are equivalent. Journal of Informetrics, 8(1):71–76, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157713000850.

Klavans:2014:MA

[KB14] Richard Klavans and Kevin W. Boyack. Mapping altruism. Journal of Informetrics, 8(2):431–447, April 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000285.

Klavans:2017:RPA

[KB17] Richard Klavans and Kevin W. Boyack. Research portfolio analysis and topic prominence. Journal of Informetrics, 11(4):1158–1174, November 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S1751157717302110.

Khosrowjerdi:2021:CRS

[KB21] Mahmood Khosrowjerdi and Lutz Bornmann. Is culture related to strong science? An empirical investigation. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000316.

Kiss:2010:CEM

[KBCR10] Istvan Z. Kiss, Mark Broom, Paul G. Craze, and Ismael Rafols. Can epidemic models describe the diffusion of topics across disciplines? Journal of Informetrics, 4(1):74–82, January 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157709000716.

Kumar:2023:RIP

[KBP23] Dhananjay Kumar, Plaban Kumar Bhowmick, and Jiaul H. Paik. Researcher influence prediction (ResIP) using academic genealogy network. Journal of Informetrics, 17(2): ??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000172.

Kuan:2021:BCP

[KCH21a] Chung-Huei Kuan, Dar-Zen Chen, and Mu-Hsuan Huang. Bibliographically coupled patents: Their temporal pattern and combined relevance. Journal of Informetrics, 13(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157719302561.

Kuan:2021:OCI

[KCH21b] Chung-Huei Kuan, Dar-Zen Chen, and Mu-Hsuan Huang. The overlooked citations: Investigating the impact of ignoring citations to published patent applications. Journal of Informetrics, 14(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719303098.

Kuan:2024:DCN

[KCH24] Chung-Huei Kuan, Dar-Zen Chen, and Mu-Hsuan Huang. Dubious cross-national affiliations obscure the assessment of international research collaboration. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000099.

Kosztyan:2021:SSI

[KCK21] Zsolt T. Kosztyán, Tibor Csizmadia, and Attila I. Katona. SIMILAR — systematic iterative multilayer literature review method. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157720306283.

Kim:2015:EDP

[KD15] Jinseok Kim and Jana Diesner. The effect of data preprocessing on understanding the evolution of collaboration networks. Journal of Informetrics, 9(1):226–236, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715000036.

Kretschmer:2015:WCWa

[KdBOK15a] Hildrun Kretschmer, Donald deB. Beaver, Bulent Ozel, and Theo Kretschmer. Who is collaborating with whom? Part I. Mathematical model and methods for empirical testing. Journal of Informetrics, 9(2):359–372, April 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771500005X.

Kretschmer:2015:WCWb

[KdBOK15b] Hildrun Kretschmer, Donald deB. Beaver, Bulent Ozel, and Theo Kretschmer. Who is collaborating with whom? Part II. Application of the methods to male and to female networks. Journal of Informetrics, 9(2):373–384, April 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715000188.

Khoveyni:2017:CSU

[KE17] Mohammad Khoveyni and Robabeh Eslami. Comment on “Using multi-level frontiers in DEA models to grade countries/territories” by G.-l. Yang et al. [Journal of Informetrics 10(1) (2016), 238–253]. Journal of Informetrics, 11(3):649–654, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S1751157717301815. See [YAY^{+}16] and reply [YAY^{+}17].

Ke:2018:CST

[Ke18] Qing Ke. Comparing scientific and technological impact of biomedical research. Journal of Informetrics, 12(3):706–717, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718300981.

Ke:2021:AES

[Ke21a] Qing Ke. An analysis of the evolution of science–technology linkage in biomedicine. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771930166X.

Ke:2021:CDC

[Ke21b] Qing Ke. The citation disadvantage of clinical research. Journal of Informetrics, 14(1):??, February 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719303025.

Kaur:2015:QVQ

[KFM^{+}15] Jasleen Kaur, Emilio Ferrara, Filippo Menczer, Alessandro Flammini, and Filippo Radicchi. Quality versus quantity in scientific impact. Journal of Informetrics, 9(4):800–808, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715200120.

Kostoff:2007:UCM

[KG07] Ronald N. Kostoff and Elie Geisler. The unintended consequences of metrics in technology evaluation. Journal of Informetrics, 1(2):103–114, April 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115770700034X.

Kuan:2011:PRI

[KHC11a] Chung-Huei Kuan, Mu-Hsuan Huang, and Dar-Zen Chen. Positioning research and innovation performance using shape centroids of h-core and h-tail. Journal of Informetrics, 5(4):515–528, October 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S175115771100040X.

Kuan:2011:RPA

[KHC11b] Chung-Huei Kuan, Mu-Hsuan Huang, and Dar-Zen Chen. Ranking patent assignee performance by h-index and shape descriptors. Journal of Informetrics, 5(2):303–312, April 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157711000034.

Kuan:2013:CFE

[KHC13] Chung-Huei Kuan, Mu-Hsuan Huang, and Dar-Zen Chen. Cross-field evaluation of publications of research institutes using their contributions to the fields’ MVPs determined by h-index. Journal of Informetrics, 7(2):455–468, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000102.

Kuan:2018:MLT

[KHC18] Chung-Huei Kuan, Mu-Hsuan Huang, and Dar-Zen Chen. Missing links: Timing characteristics and their implications for capturing contemporaneous technological developments. Journal of Informetrics, 12(1):259–270, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717302730.

Khreisat:2009:MLA

[Khr09] Laila Khreisat. A machine learning approach for Arabic text classification using N-gram frequency statistics. Journal of Informetrics, 3(1):72–77, January 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157708000709.

Kim:2022:EST

[KJKS22] Erin H. J. Kim, Yoo Kyung Jeong, YongHwan Kim, and Min Song. Exploring scientific trajectories of a largescale dataset using topic-integrated path extraction. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721001139.

Kim:2016:CPB

[KJS16] Ha Jin Kim, Yoo Kyung Jeong, and Min Song. Content- and proximity-based author co-citation analysis using citation sentences. Journal of Informetrics, 10(4):954–966, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716300529.

Kretschmer:2007:LDD

[KK07] Hildrun Kretschmer and Theo Kretschmer. Lotka’s distribution and distribution of co-author pairs’ frequencies. Journal of Informetrics, 1(4):308–337, ???? 2007. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157707000600.

Kortelainen:2012:EPE

[KK12] Terttu Kortelainen and Mari Katvala. “Everything is plentiful — except attention”. Attention data of scientific journals on social web tools. Journal of Informetrics, 6(4):661–668, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000478.

Kumar:2013:HPF

[KK13] B. T. Sampath Kumar and D. Vinay Kumar. HTTP 404-page (not) found: Recovery of decayed URL citations. Journal of Informetrics, 7(1):145–157, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771200082X.

Kim:2015:RCC

[KK15] Jinseok Kim and Jinmo Kim. Rethinking the comparison of coauthorship credit allocation schemes. Journal of Informetrics, 9(3):667–673, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715200454.

Korytkowski:2019:PCM

[KK19] Przemyslaw Korytkowski and Emanuel Kulczycki. Publication counting methods for a national research evaluation exercise. Journal of Informetrics, 13(3):804–816, August 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718305029.

Korytkowski:2021:GBP

[KK21] Przemyslaw Korytkowski and Emanuel Kulczycki. The gap between plan S requirements and grantees’ publication practices. Journal of Informetrics, 15(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000274.

Kulczycki:2017:TEB

[KKK17] Emanuel Kulczycki, Marcin Korzeń, and Przemyslaw Korytkowski. Toward an excellence-based research funding system: Evidence from Poland. Journal of Informetrics, 11(1):282–298, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157716302231.

Kleminski:2021:WSP

[KKK21] Rajmund Klemiński, Przemyslaw Kazienko, and Tomasz Kajdanowicz. Where should I publish? Heterogeneous, networks-based prediction of paper’s citation success. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000717.

Kuan:2021:CPA

[KLC21] Chung-Huei Kuan, Jia-Tian Lin, and Dar-Zen Chen. Characterizing patent assignees by their structural positions relative to a Field’s evolutionary trajectory. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157721000584.

Karimi:2021:CGL

[KLI21] Fatemeh Karimi, Shahriar Lotfi, and Habib Izadkhah. Community-guided link prediction in multiplex networks. Journal of Informetrics, 15(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000493.

Katchanov:2022:DSN

[KM22] Yurij L. Katchanov and Yulia V. Markova. Dynamics of senses of new physics discourse: Co-keywords analysis. Journal of Informetrics, 16(1):??, February 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721001164.

Kong:2019:HDC

[KMJ^{+}19] Xiangjie Kong, Mengyi Mao, Huizhen Jiang, Shuo Yu, and Liangtian Wan. How does collaboration affect researchers’ positions in co-authorship networks? Journal of Informetrics, 13(3):887–900, August 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718304218.

Katchanov:2021:DMP

[KMS21] Yurij L. Katchanov, Yulia V. Markova, and Natalia A. Shmatko. The distinction machine: Physics journals from the perspective of the Kolmogorov–Smirnov statistic. Journal of Informetrics, 13(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719303657.

Katchanov:2023:EDM

[KMS23a] Yurij L. Katchanov, Yulia V. Markova, and Natalia A. Shmatko. Empirical demonstration of the Matthew effect in scientific research careers. Journal of Informetrics, 17(4): ??, November 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000901.

Katchanov:2023:UPS

[KMS23b] Yurij L. Katchanov, Yulia V. Markova, and Natalia A. Shmatko. Uncited papers in the structure of scientific communication. Journal of Informetrics, 17(2):??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723000160.

Kongo:2014:AAH

[Kon14] Takumi Kongo. An alternative axiomatization of the Hirsch index. Journal of Informetrics, 8(1):252–258, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713001144.

Kostoff:2008:CCU

[Kos08] Ronald N. Kostoff. Comparison of China/USA science and technology performance. Journal of Informetrics, 2(4):354– 363, October 2008. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157708000321.

Kosmulski:2009:NSI

[Kos09] Marek Kosmulski. New seniority-independent Hirsch-type index. Journal of Informetrics, 3(4):341–347, October 2009.

CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115770900039X.

Kosmulski:2010:HTA

[Kos10a] Marek Kosmulski. Hirsch-type approach to the 2nd generation citations. Journal of Informetrics, 4(3):257–264, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157710000040.

Kosmulski:2010:HTI

[Kos10b] Marek Kosmulski. Hirsch-type index of international recognition. Journal of Informetrics, 4(3):351–357, July 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000210.

Kosmulski:2011:SPN

[Kos11] Marek Kosmulski. Successful papers: a new idea in evaluation of scientific output. Journal of Informetrics, 5(3): 481–485, July 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711000320.

Kosmulski:2012:CAR

[Kos12a] Marek Kosmulski. Calibration against a reference set: a quantitative approach to assessment of the methods of assessment of scientific output. Journal of Informetrics, 6(3): 451–456, July 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712000260.

Kosmulski:2012:MI

[Kos12b] Marek Kosmulski. Modesty-index. Journal of Informetrics, 6(3):368–369, July 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1751157712000132.

Kosmulski:2012:NIC

[Kos12c] Marek Kosmulski. Nemo iudex in causa sua? Journal of Informetrics, 6(4):611–614, October 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000466.

Kosmulski:2012:OLA

[Kos12d] Marek Kosmulski. The order in the lists of authors in multiauthor papers revisited. Journal of Informetrics, 6(4):639– 644, October 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712000491.

Kosmulski:2013:Y

[Kos13a]  Marek Kosmulski. Are you in h ?  Journal of

Informetrics, 7(3):693–698, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157713000436.

Kosmulski:2013:FTB

[Kos13b] Marek Kosmulski. Family-tree of bibliometric indices. Journal of Informetrics, 7(2):313–317, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712001058.

Kosmulski:2015:GDP

[Kos15] Marek Kosmulski. Gender disparity in Polish science by year (1975–2014) and by discipline. Journal of Informetrics, 9(3): 658–666, July 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715200521.

Ko:2013:IEJ

[KP13] Young Man Ko and Ji Young Park. An index for evaluating journals in a small domestic citation index database whose citation rate is generally very low: a test based on the Korea Citation Index (KCI) database. Journal of Informetrics, 7(2):404–411, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000047.

Kim:2021:GFR

[KP21] Sangil Kim and Keon Chul Park. Government funded R&D collaboration and it’s impact on SME’s business performance. Journal of Informetrics, 15(3):??, August 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000687.

Kosyakov:2024:SIL

[KP24] Denis Kosyakov and Vladimir Pislyakov. “I’d like to publish in Q1, but there’s no Q1 to be found”: Study of journal quartile distributions across subject categories and topics. Journal of Informetrics, 18(1):??, February 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000075.

Keyhanipour:2009:GAW

[KPB09] Amir Hosein Keyhanipour, Maryam Piroozmand, and Kambiz Badie. A GP-adaptive web ranking discovery framework based on combinative content and context features. Journal of Informetrics, 3(1):78–89, January 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157708000710.

Kim:2022:DTD

[KPS22] Hyeyoung Kim, Hyelin Park, and Min Song. Developing a topic-driven method for interdisciplinarity analysis. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000074.

Kwiek:2021:GBH

[KR21] Marek Kwiek and Wojciech Roszka. Gender-based homophily in research: a large-scale study of man-woman collaboration. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157721000420.

Kwiek:2024:TRP

[KR24] Marek Kwiek and Wojciech Roszka. Top research performance in Poland over three decades: a multidimensional micro-data approach. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115772400107X.

Kaur:2013:USI

[KRM13] Jasleen Kaur, Filippo Radicchi, and Filippo Menczer. Universality of scholarly impact metrics. Journal of Informetrics, 7(4):924–932, October 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157713000746.

Kaur:2014:USS

[KRM14] Jasleen Kaur, Filippo Radicchi, and Filippo Menczer. On the use of sampling statistics to advance bibliometrics. Journal of Informetrics, 8(2):419–420, April 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000170.

Kang:2016:EPA

[KS16] Kiyeon Kang and So Young Sohn. Evaluating the patenting activities of pharmaceutical research organizations based on new technology indices. Journal of Informetrics, 10(1):74– 81, February 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715300559.

Kraker:2015:VCR

[KSJL15] Peter Kraker, Christian Schlögl, Kris Jack, and Stefanie Lindstaedt. Visualization of co-readership patterns from an online reference management system. Journal of Informetrics, 9(1):169–182, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714001151.

Koski:2016:TFA

[KSS16] Timo Koski, Erik Sandström, and Ulf Sandström. Towards field-adjusted production: Estimating research productivity from a zero-truncated distribution. Journal of Informetrics, 10(4):1143–1152, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716300888.

Kousha:2018:CMAb

[KT18] Kayvan Kousha and Mike Thelwall. Can Microsoft Academic help to assess the citation impact of academic books? Journal of Informetrics, 12(3):972–984, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302487.

Kousha:2019:CGS

[KT19] Kayvan Kousha and Mike Thelwall. Can Google Scholar and Mendeley help to assess the scholarly impacts of dissertations? Journal of Informetrics, 13(2):467–484, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718304991.

Kousha:2018:CMAa

[KTA18] Kayvan Kousha, Mike Thelwall, and Mahshid Abdoli. Can Microsoft Academic assess the early citation impact of inpress articles? A multi-discipline exploratory analysis. Journal of Informetrics, 12(1):287–298, February 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717304303.

Kousha:2010:UWR

[KTR10] Kayvan Kousha, Mike Thelwall, and Somayeh Rezaie. Using the Web for research evaluation: the Integrated Online Impact indicator. Journal of Informetrics, 4(1):124–135, January 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157709000777.

Kuan:2024:IPF

[Kua24] Chung-Huei Kuan. Integrating prior field knowledge as key documents with main path analysis utilizing key-node path search. Journal of Informetrics, 18(3):??, August 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000804.

Kong:2021:CCA

[KW21] Ling Kong and Dongbo Wang. Comparison of citations and attention of cover and non-cover papers. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719305048.

Kim:2012:NAT

[KYC12] Hyoungshick Kim, Ji Won Yoon, and Jon Crowcroft. Network analysis of temporal trends in scholarly research productivity. Journal of Informetrics, 6(1):97–110, January 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157711000629.

Kwon:2023:ECP

[KYhK23] Eunrang Kwon, Jinhyuk Yun, and Jeong han Kang. The effect of the COVID-19 pandemic on gendered research productivity and its correlates. Journal of Informetrics, 17(1): ??, February 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000056.

Kong:2023:KCS

[KZDK23] Jia Kong, Jiarui Zhang, Sanhong Deng, and Lele Kang. Knowledge convergence of science and technology in patent inventions. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723000603.

Kong:2024:TFG

[KZH^{+}24] Ling Kong, Wei Zhang, Haotian Hu, Zhu Liang, Yonggang Han, Dongbo Wang, and Min Song. Transdisciplinary fine-grained citation content analysis: a multi-task learning perspective for citation aspect and sentiment classification. Journal of Informetrics, 18(3):??, August 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000555.

Lafouge:2015:SEC

[LA15] Thierry Lafouge and Abdelatif Agouzal. The source-effort coverage of an exponential informetric process. Journal of Informetrics, 9(1):156–168, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714001163.

Lafouge:2007:SIC

[Laf07] Thierry Lafouge. The source-item coverage of the exponential function. Journal of Informetrics, 1(1):59–67, January 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157706000095.

Lucio-Arias:2009:DER

[LAL09] Diana Lucio-Arias and Loet Leydesdorff. The dynamics of exchanges and references among scientific texts, and the autopoiesis of discursive knowledge. Journal of Informetrics, 3(3):261–271, July 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S175115770900025X.

Lambiotte:2007:WSB

[LAT07] R. Lambiotte, M. Ausloos, and M. Thelwall. Word statistics in blogs and RSS feeds: Towards empirical universal evidence. Journal of Informetrics, 1(4):277–286, ???? 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157707000582.

Lathabai:2021:INO

[Lat21] Hiran H. Lathabai. ψ-index: a new overall productivity index for actors of science and technology. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719305012.

Laakso:2016:HOA

[LB16] Mikael Laakso and Bo-Christer Björk. Hybrid open access — a longitudinal study. Journal of Informetrics, 10(4):919– 932, November 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157716301523.

Leydesdorff:2021:DIT

[LB21] Loet Leydesdorff and Lutz Bornmann. Disruption indices and their calculation using Web-of-Science data: Indicators of historical developments or evolutionary dynamics? Journal of Informetrics, 15(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000900.

Lando:2017:MCI

[LBB17] Tommaso Lando and Lucio Bertoli-Barsotti. Measuring the citation impact of journals with generalized Lorenz curves. Journal of Informetrics, 11(3):689–703, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302401.

Lu:2019:ALC

[LBD^{+}19] Chao Lu, Yi Bu, Xianlei Dong, Jie Wang, Ying Ding, Vincent Larivière, Cassidy R. Sugimoto, Logan Paul, and Chengzhi Zhang. Analyzing linguistic complexity and scientific impact. Journal of Informetrics, 13(3):817–829, August 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115771830227X.

Lai:2021:IIP

[LBK^{+}21] Kuei-Kuei Lai, Priyanka C. Bhatt, Vimal Kumar, Hsueh-Chen Chen, Yu-Hsin Chang, and Fang-Pei Su. Identifying the impact of patent family on the patent trajectory: a case of thin film solar cells technological trajectories. Journal of Informetrics, 15(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000146.

Liang:2023:RCI

[LBML23] Zhentao Liang, Zhichao Ba, Jin Mao, and Gang Li. Research complexity increases with scientists’ academic age: Evidence from library and information science. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722001286.

Leydesdorff:2014:RPY

[LBMM14] Loet Leydesdorff, Lutz Bornmann, Werner Marx, and Stasa Milojević. Referenced publication years spectroscopy applied to iMetrics: Scientometrics, Journal of Informetrics, and a relevant subset of JASIST. Journal of Informetrics, 8(1):162– 174, January 2014. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713001077.

Liu:2024:HDP

[LBP24] Leilei Liu, Zhichao Ba, and Lei Pei. How does policy information shape its adoption? A citation analysis of large-scale energy policies in China. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724001019.

Lepori:2013:CPS

[LBSA13] Benedetto Lepori, Vitaliano Barberio, Marco Seeber, and Isidro Aguillo. Core-periphery structures in national higher education systems. A cross-country analysis using interlinking data. Journal of Informetrics, 7(3):622–634, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000242.

Leydesdorff:2016:CPB

[LBZ16] Loet Leydesdorff, Lutz Bornmann, and Ping Zhou. Construction of a pragmatic base line for journal classifications and maps based on aggregated journal–journal citation relations. Journal of Informetrics, 10(4):902–918, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115771630102X.

Lyu:2022:CBM

[LBZ^{+}22] Haihua Lyu, Yi Bu, Zhenyue Zhao, Jiarong Zhang, and Jiang Li. Citation bias in measuring knowledge flow: Evidence from the Web of Science at the discipline level. Journal of Informetrics, 16(4):??, November 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000906.

Liu:2022:NLB

[LCL^{+}22] Jialin Liu, Hongkan Chen, Zhibo Liu, Yi Bu, and Weiye Gu. Non-linearity between referencing behavior and citation impact: a large-scale, discipline-level analysis. Journal of Informetrics, 16(3):??, August 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000700.

Li:2019:CMS

[LCY19] Kai Li, Pei-Ying Chen, and Erjia Yan. Challenges of measuring software impact through citations: an examination of the lme4 R package. Journal of Informetrics, 13(1):449– 461, February 2019. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157718304796.

Li:2012:ACD

[LDS^{+}12] Daifeng Li, Ying Ding, Xin Shuai, Johan Bollen, Jie Tang, Shanshan Chen, Jiayi Zhu, and Guilherme Rocha. Adding community and dynamic to topic models. Journal of Informetrics, 6(2):237–253, April 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711001039.

Lv:2018:TDT

[LDSD18] Yanhua Lv, Ying Ding, Min Song, and Zhiguang Duan. Topology-driven trend analysis for drug discovery. Journal of Informetrics, 12(3):893–905, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718300920.

Letina:2016:NAA

[Let16] Srebrenka Letina. Network and actor attribute effects on the performance of researchers in two fields of social science in a small peripheral community. Journal of Informetrics, 10(2): 571–595, May 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715302182.

Lin:2022:NDS

[LEW22] Yiling Lin, James A. Evans, and Lingfei Wu. New directions in science emerge from disconnection and discord. Journal of Informetrics, 16(1):??, February 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772100105X.

Leydesdorff:2012:MAT

[Ley12] Loet Leydesdorff. McCall’s area transformation versus the integrated impact indicator (I3). Journal of Informetrics, 6(4):513–514, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157712000247.

Leydesdorff:2013:DSU

[Ley13a] Loet Leydesdorff. Does the specification of uncertainty hurt the progress of scientometrics? Journal of Informetrics, 7(2): 292–293, April 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712000971.

Leydesdorff:2013:RSI

[Ley13b] Loet Leydesdorff. The revised SNIP indicator of Elsevier’s Scopus. Journal of Informetrics, 7(4):859–860, October 2013.

CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000679.

Leydesdorff:2017:PB

[Ley17a] Loet Leydesdorff. In praise of bibliometrics. Journal of Informetrics, 11(2):595–597, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717300755.

Leydesdorff:2017:PSD

[Ley17b] Loet Leydesdorff. The positive side of discursive disagreements in the social sciences. Journal of Informetrics, 11(4): 1043, November 2017. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157717302468.

Liu:2012:MIA

[LF12] Xuan Zhen Liu and Hui Fang. Modifying h-index by allocating credit of multi-authored papers whose author names rank based on contribution. Journal of Informetrics, 6(4):557– 565, October 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712000429.

Liu:2021:CAC

[LF21] Xuan Zhen Liu and Hui Fang. A comparison among citationbased journal indicators and their relative changes with time. Journal of Informetrics, 14(1):??, February 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719302901.

Liu:2023:GCM

[LF23] Xuan Zhen Liu and Hui Fang. A geometric counting method adaptive to the author number. Journal of Informetrics, 17 (2):??, May 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000299.

Lariviere:2011:ARV

[LG11] Vincent Larivière and Yves Gingras. Averages of ratios vs. ratios of averages: an empirical analysis of four levels of aggregation. Journal of Informetrics, 5(3):392–399, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157711000149.

Liu:2024:DSC

[LGS^{+}24] Qiuling Liu, Lei Guo, Yiping Sun, Linlin Ren, Xinhua Wang, and Xiaohui Han. Do scholars’ collaborative tendencies impact the quality of their publications? A generalized propensity score matching analysis. Journal of Informetrics, 18(1): ??, February 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723001128.

Liu:2023:NIS

[LGX^{+}23] Junwan Liu, Xiaofei Guo, Shuo Xu, Yinglu Song, and Kaiyue Ding. A new interpretation of scientific collaboration patterns from the perspective of symbiosis: an investigation for long-term collaboration in publications. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722001250.

Liu:2021:WCM

[LH21] Meijun Liu and Xiao Hu. Will collaborators make scientists move? A Generalized Propensity Score analysis. Journal of Informetrics, 15(1):??, February 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720306301.

Liu:2022:MAE

[LH22] Meijun Liu and Xiao Hu. Movers’ advantages: the effect of mobility on scientists’ productivity and collaboration. Journal of Informetrics, 16(3):??, August 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000633.

Li:2024:RAA

[LH24] Hao Li and Jianhua Hou. Revalidation of the applicability of altmetrics indicators in article-level evaluation: an empirical analysis of papers of different types of citation trajectories. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724000853.

Lin:2013:ICM

[LHC13] Chi-Shiou Lin, Mu-Hsuan Huang, and Dar-Zen Chen. The influences of counting methods on university rankings based on paper count and citation count. Journal of Informetrics, 7(3):611–621, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000278.

Lungeanu:2014:UAI

[LHC14] Alina Lungeanu, Yun Huang, and Noshir S. Contractor. Understanding the assembly of interdisciplinary teams and its impact on performance. Journal of Informetrics, 8(1):59– 70, January 2014. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000837.

Liang:2021:KRB

[LHDH21] Guoqiang Liang, Haiyan Hou, Ying Ding, and Zhigang Hu. Knowledge recency to the birth of Nobel Prizewinning articles: Gender, career stage, and country. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719301804.

Liu:2018:MAA

[LHT18] Weishu Liu, Guangyuan Hu, and Li Tang. Missing author address information in Web of Science — an explorative study. Journal of Informetrics, 12(3):985–997, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771730353X.

Liu:2015:CGG

[LHTW15] Weishu Liu, Guangyuan Hu, Li Tang, and Yuandi Wang. China’s global growth in social science research: Uncovering evidence from bibliometric analyses of SSCI publications (1978–2013). Journal of Informetrics, 9(3):555–569, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715000528.

Liu:2018:SPI

[LHWS18] Meijun Liu, Xiao Hu, Yuandi Wang, and Dongbo Shi. Survive or perish: Investigating the life cycle of academic journals from 1950 to 2013 using survival analysis methods. Journal of Informetrics, 12(1):344–364, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S175115771730281X.

Liao:2021:MEH

[Lia21] Chien Hsiang Liao. The Matthew effect and the halo effect in research funding. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157720306258.

Lopez-Illescas:2008:CCI

[LIdMAM08] Carmen López-Illescas, Félix de Moya-Anegón, and Henk F. Moed. Coverage and citation impact of oncological journals in the Web of Science and Scopus. Journal of Informetrics, 2(4):304–316, October 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S175115770800045X.

Lindahl:2018:EPR

[Lin18a] Jonas Lindahl. Erratum to “Predicting research excellence at the individual level: the importance of publication rate, top journal publications, and top 10 publications in the case of early career mathematicians” [J. Informetr. 12 (2) (2018) 518–533]. Journal of Informetrics, 12(4):1327–1329, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718302347. See [Lin18b].

Lindahl:2018:PREa

[Lin18b] Jonas Lindahl. Predicting research excellence at the individual level: the importance of publication rate, top journal publications, and top 10 publications in the case of early career mathematicians. Journal of Informetrics, 12(2):518–533, May 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S1751157717301633. See comment

[vdBS18], rejoinder [Lin18c], and erratum [Lin18a].

Lindahl:2018:PREb

[Lin18c] Jonas Lindahl. Predicting research excellence at the individual level with bibliometric indicators: a rejoinder on a comment by van den Besselaar and Sandström (2018). Journal of Informetrics, 12(4):1194–1198, November 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718303912. See [Lin18b, vdBS18].

Lindahl:2023:CPD

[Lin23] Jonas Lindahl. Conscientiousness predicts doctoral students’ research productivity. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722001067. See corrigen-

dum [Lin24].

Lindahl:2024:CCP

[Lin24] Jonas Lindahl. Corrigendum to “Conscientiousness predicts doctoral students’ research productivity” [Journal of Informetrics 17/1 (2023) 101353]. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157724001111. See [Lin23].

Liu:2022:TFT

[LJB^{+}22] Meijun Liu, Ajay Jaiswal, Yi Bu, Chao Min, Sijie Yang, Zhibo Liu, Daniel Acuña, and Ying Ding. Team formation and team impact: the balance between team freshness and repeat collaboration. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722000888.

Liu:2023:SNI

[LJHH23] Yan Liu, Mudi Jiang, Lianyu Hu, and Zengyou He. The statistical nature of h-index of a network node and its extensions. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723000494.

Lee:2021:LMR

[LJJ21] O-Joun Lee, Hyeon-Ju Jeon, and Jason J. Jung. Learning multi-resolution representations of research patterns in bibliographic networks. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115772030643X.

Liu:2013:CDK

[LJM13] Xiang Liu, Tingting Jiang, and Feicheng Ma. Collective dynamics in knowledge networks: Emerging trends analysis. Journal of Informetrics, 7(2):425–438, April 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000059.

Loizides:2017:EQC

[LK17] Orestis-Stavros Loizides and Polychronis Koutsakis. On evaluating the quality of a computer science/computer engineering conference. Journal of Informetrics, 11(2):541–552, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716301808.

Lachance:2014:CLP

[LL14] Christian Lachance and Vincent Larivière. On the citation lifecycle of papers with delayed recognition. Journal of Informetrics, 8(4):863–872, October 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000789.

Lee:2015:MCR

[LL15] Seonghee Lee and Hakyeon Lee. Measuring and comparing the R&D performance of government research institutes: a bottom-up data envelopment analysis approach. Journal of Informetrics, 9(4):942–953, October 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715301127.

Lin:2021:UTE

[LLGM21] Deming Lin, Wenbin Liu, Yinxin Guo, and Martin Meyer. Using technological entropy to identify technology life cycle. Journal of Informetrics, 15(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000080.

Liu:2012:TIM

[LLH12] John S. Liu, Louis Y. Y. Lu, and Mei Hsiu-Ching Ho. Total influence and mainstream measures for scientific researchers. Journal of Informetrics, 6(4):496–504, October 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000338.

Lu:2021:HDA

[LLH^{+}21] Wei Lu, Zhifeng Liu, Yong Huang, Yi Bu, Xin Li, and Qikai Cheng. How do authors select keywords? A preliminary study of author keyword selection behavior. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720300134.

Luo:2022:CRQ

[LLHW22] Zhuoran Luo, Wei Lu, Jiangen He, and Yuqi Wang. Combination of research questions and methods: a new measurement of scientific novelty. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722000347.

Li:2021:OAE

[LLW21] Huixu Li, Lanjian Liu, and Xianwen Wang. The open access effect in social media exposure of scholarly articles: a matched-pair analysis. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721000250.

Li:2024:EIB

[LMF24] Xin Li, Xiaodi Ma, and Ye Feng. Early identification of breakthrough research from sleeping beauties using machine learning. Journal of Informetrics, 18(2):??, May 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000300.

Leydesdorff:2010:NFL

[LO10] Loet Leydesdorff and Tobias Opthof. Normalization at the field level: Fractional counting of citations. Journal of Informetrics, 4(4):644–646, October 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000416.

Leydesdorff:2011:RPN

[LO11] Loet Leydesdorff and Tobias Opthof. Remaining problems with the “New Crown Indicator” (MNCS) of the CWTS. Journal of Informetrics, 5(1):224–225, January 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000908.

Lambiotte:2009:CKC

[LP09] R. Lambiotte and P. Panzarasa. Communities, knowledge creation, and information diffusion. Journal of Informetrics, 3(3):180–190, July 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157709000248.

Leydesdorff:2017:FFC

[LP17] Loet Leydesdorff and Han Woo Park. Full and fractional counting in bibliometric networks. Journal of Informetrics, 11(1):117–120, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157716303133.

Letchford:2016:ASP

[LPM16] Adrian Letchford, Tobias Preis, and Helen Susannah Moat. The advantage of simple paper abstracts. Journal of Informetrics, 10(1):1–8, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715300766.

Li:2014:CLA

[LQLJ14] Jiang Li, Lili Qiao, Wenyuze Li, and Yidan Jin. Chineselanguage articles are not biased in citations: Evidences from Chinese–English bilingual journals in Scopus and Web of Science. Journal of Informetrics, 8(4):912–916, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714000844.

Li:2015:RCL

[LQLJ15] Jiang Li, Lili Qiao, Wenyuze Li, and Yidan Jin. Rejoinder to “Chinese-language articles are biased in citations”. Journal of Informetrics, 9(3):591, July 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715300365. See [SL15].

Liang:2007:TBP

[LR07] Liming Liang and Ronald Rousseau. Transformations of basic publication-citation matrices. Journal of Informetrics, 1(3):249–255, July 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1751157707000326.

Liu:2008:DTS

[LR08] Yuxian Liu and Ronald Rousseau. Definitions of time series in citation analysis with special attention to the hindex. Journal of Informetrics, 2(3):202–210, July 2008. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157708000242.

Liang:2010:MJI

[LR10] Liming Liang and Ronald Rousseau. Measuring a journal’s input rhythm based on its publication-reference matrix. Journal of Informetrics, 4(2):201–209, April 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000807.

Leydesdorff:2011:IIJ

[LR11] Loet Leydesdorff and Ismael Rafols. Indicators of the interdisciplinarity of journals: Diversity, centrality, and citations. Journal of Informetrics, 5(1):87–100, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157710000854.

Leydesdorff:2012:ION

[LR12] Loet Leydesdorff and Ismael Rafols. Interactive overlays: a new method for generating global journal maps from Webof-Science data. Journal of Informetrics, 6(2):318–332, April 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157711001027.

Li:2013:CNP

[LRC13] Yunrong Li and Javier Ruiz-Castillo. The comparison of normalization procedures based on different classification systems. Journal of Informetrics, 7(4):945–958, October 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000771.

Li:2013:QEA

[LRCRC13] Yunrong Li, Filippo Radicchi, Claudio Castellano, and Javier Ruiz-Castillo. Quantitative evaluation of alternative field normalization procedures. Journal of Informetrics, 7(3):746– 755, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157713000485.

Liu:2017:POZ

[LRE17] Yuxian Liu, Ronald Rousseau, and Leo Egghe. Partial orders for zero-sum arrays with applications to network theory. Journal of Informetrics, 11(1):257–274, February 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302383.

Liu:2013:LFS

[LRG13] Yuxian Liu, Ronald Rousseau, and Raf Guns. A layered framework to study collaboration as a form of knowledge sharing and diffusion. Journal of Informetrics, 7(3):651–664, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157713000382.

Lu:2021:SCC

[LRH^{+}21] Wei Lu, Yan Ren, Yong Huang, Yi Bu, and Yuehan Zhang. Scientific collaboration and career stages: an ego-centric perspective. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115772100078X.

Liu:2013:RCT

[LRWY13] Joan Q. Liu, Ronald Rousseau, Mona S. Wang, and Fred Y. Ye. Ratios of h-cores, h-tails and uncited sources in sets of scientific papers and technical patents. Journal of Informetrics, 7(1):190–197, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000922.

Lee:2014:MJP

[LS14] Hakyeon Lee and Juneseuk Shin. Measuring journal performance for multidisciplinary research: an efficiency perspective. Journal of Informetrics, 8(1):77–88, January 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000813.

Li:2024:EPD

[LSSP24] Yingqun Li, Ningyuan Song, Yu Shen, and Lei Pei. Exploring the potential of disruptive innovation in the social sciences: a quantitative study of its impact on societal visibility. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724000968.

Li:2010:RLI

[LSW^{+}10] Jiang Li, Mark Sanderson, Peter Willett, Michael Norris, and Charles Oppenheim. Ranking of library and information science researchers: Comparison of data sources for correlating citation data, and expert judgments. Journal of Informetrics, 4(4):554–563, October 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000593.

Li:2014:SHS

[LSZY14] Jiang Li, Dongbo Shi, Star X. Zhao, and Fred Y. Ye. A study of the “heartbeat spectra” for “sleeping beauties”. Journal of Informetrics, 8(3):493–502, July 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000418.

Levitt:2013:ASF

[LT13] Jonathan M. Levitt and Mike Thelwall. Alphabetization and the skewing of first authorship towards last names early in the alphabet. Journal of Informetrics, 7(3):575–582, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000229.

Li:2021:CID

[LT21] Xin Li and Xuli Tang. Characterizing interdisciplinarity in drug research: a translational science perspective. Journal of Informetrics, 15(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000870.

Li:2022:PCC

[LTC22] Xin Li, Xuli Tang, and Qikai Cheng. Predicting the clinical citation count of biomedical papers using multilayer perceptron neural network. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722000852.

Lepori:2018:WUE

[LTH18] Benedetto Lepori, Michael Thelwall, and Bareerah Hafeez Hoorani. Which US and European higher education institutions are visible in ResearchGate and what affects their RG score? Journal of Informetrics, 12(3):806–818, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718301068.

Li:2024:ICL

[LTL24] Xin Li, Xuli Tang, and Wei Lu. Investigating clinical links in edge-labeled citation networks of biomedical research: a translational science perspective. Journal of Informetrics, 18 (3):??, August 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000713.

Lundberg:2007:LCC

[Lun07] Jonas Lundberg. Lifting the crown-citation z-score. Journal of Informetrics, 1(2):145–154, April 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157706000265.

Luoto:2021:SDP

[Luo21] Severi Luoto. Sex differences in people and things orientation are reflected in sex differences in academic publishing. Journal of Informetrics, 14(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720300043.

Leydesdorff:2008:ICS

[LW08] Loet Leydesdorff and Caroline S. Wagner. International collaboration in science and the formation of a core group.

Journal of Informetrics, 2(4):317–325, October 2008. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157708000448.

Leydesdorff:2009:MLI

[LW09] Loet Leydesdorff and Caroline Wagner. Macro-level indicators of the relations between research funding and research output. Journal of Informetrics, 3(4):353–362, October 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157709000418.

Lu:2010:GCG

[LW10] Kun Lu and Dietmar Wolfram. Geographic characteristics of the growth of informetrics literature 1987–2008. Journal of Informetrics, 4(4):591–601, October 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000623.

Leydesdorff:2011:SMW

[LW11] Loet Leydesdorff and Kasper Welbers. The semantic mapping of words and co-words in contexts. Journal of Informetrics, 5(3):469–475, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157711000095.

Lehmann:2017:WJG

[LW17] Robert Lehmann and Klaus Wohlrabe. Who is the ‘Journal Grand Master’? A new ranking based on the Elo rating system. Journal of Informetrics, 11(3):800–809, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716303625.

Leydesdorff:2014:EUC

[LWB14] Loet Leydesdorff, Caroline S. Wagner, and Lutz Bornmann. The European Union, China, and the United States in the top-1 and top-10% layers of most-frequently cited publications: Competition and collaborations. Journal of Informetrics, 8(3):606–617, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000509.

Leydesdorff:2019:DMS

[LWB19a] Loet Leydesdorff, Caroline S. Wagner, and Lutz Bornmann. Diversity measurement: Steps towards the measurement of interdisciplinarity? Journal of Informetrics, 13(3):904–905, August 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719300999.

Leydesdorff:2019:IDC

[LWB19b] Loet Leydesdorff, Caroline S. Wagner, and Lutz Bornmann. Interdisciplinarity as diversity in citation patterns among journals: Rao–Stirling diversity, relative variety, and the Gini coefficient. Journal of Informetrics, 13(1):255–269, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718303535.

Liu:2022:EPR

[LWCH22] Xiaojuan Liu, Chenlin Wang, Dar-Zen Chen, and Mu-Hsuan Huang. Exploring perception of retraction based on mentioned status in post-retraction citations. Journal of Informetrics, 16(3):??, August 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000566.

Li:2014:NBM

[LWWL14] Yongli Li, Chong Wu, Xiaoyu Wang, and Peng Luo. A network-based and multi-parameter model for finding influential authors. Journal of Informetrics, 8(3):791–799, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714000674.

Li:2022:BCN

[LWWZ22] Heyang Li, Meijun Wu, Yougui Wang, and An Zeng. Bibliographic coupling networks reveal the advantage of diversification in scientific projects. Journal of Informetrics, 16 (3):??, August 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722000736.

Liao:2012:QDR

[LY12] Chien Hsiang Liao and Hsiuju Rebecca Yen. Quantifying the degree of research collaboration: a comparative study of collaborative measures. Journal of Informetrics, 6(1):27– 33, January 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711000800.

Liu:2014:ESS

[LY14] Yu Liu and Yongliang Yang. Empirical study of L-sequence: the basic h-index sequence for cumulative publications with consideration of the yearly citation performance. Journal of Informetrics, 8(3):478–485, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000303.

Li:2018:CMN

[LY18] Kai Li and Erjia Yan. Co-mention network of R packages: Scientific impact and clustering structure. Journal of Informetrics, 12(1):87–100, February 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717304108.

Li:2019:NFP

[LY19] Kai Li and Erjia Yan. Are NIH-funded publications fulfilling the proposed research? An examination of conceptmatchedness between NIH research grants and their supported publications. Journal of Informetrics, 13(1):226–237, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718303717.

Li:2024:SDC

[LY24] Jing Li and Qian Yu. Scientists’ disciplinary characteristics and collaboration behaviour under the convergence paradigm: a multilevel network perspective. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772400004X.

Liu:2021:NCC

[LYC21] Yunmei Liu, Liu Yang, and Min Chen. A new citation concept: Triangular citation in the literature. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157721000122.

Li:2017:HCR

[LYF17] Kai Li, Erjia Yan, and Yuanyuan Feng. How is R cited in research outputs? Structure, impacts, and citation standard. Journal of Informetrics, 11(4):989–1002, November 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717300329.

Yang:2018:MIC

[lYFyS18] Guo liang Yang, Hirofumi Fukuyama, and Yao yao Song. Measuring the inefficiency of Chinese research universities based on a two-stage network DEA model. Journal of Informetrics, 12(1):10–30, February 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717302766.

Yang:2014:SDR

[lYRyYbL14] Guo liang Yang, Ronald Rousseau, Li ying Yang, and Wen bin Liu. A study on directional returns to scale. Journal of Informetrics, 8(3):628–641, July 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000522.

Liping:2009:REA

[LYYY09] Yu Liping, Chen Yuqing, Pan Yuntao, and Wu Yishan. Research on the evaluation of academic journals based on structural equation modeling. Journal of Informetrics, 3(4):304– 311, October 2009. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157709000340.

Li:2017:DME

[LYZ^{+}17] Menghui Li, Liying Yang, Huina Zhang, Zhesi Shen, Chensheng Wu, and Jinshan Wu. Do mathematicians, economists and biomedical scientists trace large topics more strongly than physicists? Journal of Informetrics, 11(2):598–607, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716303753.

Liu:2022:FDG

[LZH^{+}22] Meijun Liu, Ning Zhang, Xiao Hu, Ajay Jaiswal, Jian Xu, Hong Chen, Ying Ding, and Yi Bu. Further divided gender gaps in research productivity and collaboration during the COVID-19 pandemic: Evidence from coronavirus-related literature. Journal of Informetrics, 16(2):??, May 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000475.

Liu:2023:CTW

[LZL23] Xin Liu, Chengjing Zhang, and Jiang Li. Conceptual and technical work: Who will disrupt science? Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000573.

Liang:2015:UPU

[LZR15] Liming Liang, Zhen Zhong, and Ronald Rousseau. Uncited papers, uncited authors and uncited topics: a case study in library and information science. Journal of Informetrics, 9(1): 50–58, January 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157714000959.

Ma:2012:ABC

[Ma12] Ruimin Ma. Author bibliographic coupling analysis: a test based on a Chinese academic database. Journal of Informetrics, 6(4):532–542, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000405.

Masood:2021:UGE

[MA21] Muhammad Ali Masood and Rabeeh Ayaz Abbasi. Using graph embedding and machine learning to identify rebels on Twitter. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157720306386.

Magnone:2012:AES

[Mag12] Edoardo Magnone. An analysis for estimating the shortterm effects of Japan’s triple disaster on progress in materials science. Journal of Informetrics, 6(2):289–297, April 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157712000041.

Magnone:2013:SLC

[Mag13] Edoardo Magnone. A scientometric look at calendar events. Journal of Informetrics, 7(1):101–108, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000818.

Mallig:2010:RDB

[Mal10] Nicolai Mallig. A relational database for bibliometric analysis. Journal of Informetrics, 4(4):564–580, October 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157710000611.

Malesios:2016:MRJ

[Mal16] C. Malesios. Measuring the robustness of the journal h-index with respect to publication and citation values: a Bayesian sensitivity analysis. Journal of Informetrics, 10(3):719–731, August 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715302303.

Marzolla:2015:QAI

[Mar15] Moreno Marzolla. Quantitative analysis of the Italian National Scientific Qualification. Journal of Informetrics, 9(2): 285–316, April 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715000243.

Marzolla:2016:AEP

[Mar16] Moreno Marzolla. Assessing evaluation procedures for individual researchers: the case of the Italian National Scientific Qualification. Journal of Informetrics, 10(2):408–438, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715300961.

Martin:2017:WSS

[Mar17] Ben R. Martin. When social scientists disagree: Comments on the Butler–van den Besselaar debate. Journal of Informetrics, 11(3):937–940, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717301992.

Manganote:2014:VRD

[MAS14] Edmilson J. T. Manganote, Mariana S. Araujo, and Peter A. Schulz. Visualization of ranking data: Geographical signatures in international collaboration, leadership and research impact. Journal of Informetrics, 8(3):642–649, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714000534.

Miura:2023:RUI

[MAS23] Takahiro Miura, Kimitaka Asatani, and Ichiro Sakata. Revisiting the uniformity and inconsistency of slow-cited papers in science. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000032.

Mutz:2017:TFR

[MBD17] Rüdiger Mutz, Lutz Bornmann, and Hans-Dieter Daniel. Are there any frontiers of research performance? Efficiency measurement of funded research projects with the Bayesian stochastic frontier analysis for count data. Journal of Informetrics, 11(3):613–628, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303741.

Moed:2016:NMC

[MBIH16] Henk F. Moed, Judit Bar-Ilan, and Gali Halevi. A new methodology for comparing Google Scholar and Scopus. Journal of Informetrics, 10(2):533–551, May 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715302285.

Ma:2023:HCI

[MBZS23] Yaxue Ma, Zhichao Ba, Haiping Zhao, and Jianjun Sun. How to configure intellectual capital of research teams for triggering scientific breakthroughs: Exploratory study in the field of gene editing. Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000846.

McCain:2010:CJL

[McC10] Katherine W. McCain. Core journal literatures and persistent research themes in an emerging interdisciplinary field: Exploring the literature of evolutionary developmental biology. Journal of Informetrics, 4(2):157–165, April 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000820.

Meester:2016:RSM

[MCD16] Wim J. N. Meester, Lisa Colledge, and Elizabeth E. Dyas. A response to “The museum of errors/horrors in Scopus” by Franceschini et al. Journal of Informetrics, 10(2):569–570, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716301080.

Mogoutov:2008:BIL

[MCKM08] Andrei Mogoutov, Alberto Cambrosio, Peter Keating, and Philippe Mustar. Biomedical innovation at the laboratory, clinical and commercial interface: a new method for mapping research projects, publications and patents in the field of microarrays. Journal of Informetrics, 2(4):341–353, October 2008. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157708000333.

Meier:2023:CPE

[MCLV23] Max Meier, Julie Callaert, Paolo Landoni, and Bart Van Looy. Co-patents in Europe: Methodological concerns, unfolding trends. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115772300055X.

Mutz:2012:SCD

[MD12] Rüdiger Mutz and Hans-Dieter Daniel. Skewed citation distributions and bias factors: Solutions to two core problems with the journal impact factor. Journal of Informetrics, 6(2): 169–176, April 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711001180.

Mutz:2018:BQB

[MD18] Rüdiger Mutz and Hans-Dieter Daniel. The bibliometric quotient (BQ), or how to measure a researcher’s performance capacity: a Bayesian Poisson Rasch model. Journal of Informetrics, 12(4):1282–1295, November 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718300798.

Massucci:2019:MAR

[MD19a] Francesco Alessandro Massucci and Domingo Docampo. Measuring the academic reputation through citation networks via PageRank. Journal of Informetrics, 13(1):185– 201, February 2019. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771830110X.

Mutz:2019:HCF

[MD19b] Rüdiger Mutz and Hans-Dieter Daniel. How to consider fractional counting and field normalization in the statistical modeling of bibliometric data: a multilevel Poisson regression approach. Journal of Informetrics, 13(2):643–657, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718302852.

Munoz:2022:MET

[MdGNS22] María M. Muñoz, María-Mercedes Rojas de Gracia, and Carlos Navas-Sarasola. Measuring engagement on Twitter using a composite index: an application to social media influencers. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115772200075X.

Moed:2021:NOJ

[MdMAGBLI21a] Henk F. Moed, Felix de Moya-Anegon, Vicente Guerrero-Bote, and Carmen Lopez-Illescas. Are nationally oriented journals indexed in Scopus becoming more international? The effect of publication language and access modality. Journal of Informetrics, 14(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S1751157719304419. See correction [MdMAGBLI21b].

Moed:2021:CSN

[MdMAGBLI21b] Henk F. Moed, Felix de Moya-Anegon, Vicente Guerrero-Bote, and Carmen Lopez-Illescas. Corrigendum to “Are nationally oriented journals indexed in Scopus becoming more international? The effect of publication language and access modality” [J. Informetrics 14 (2020) 101011]. Journal of Informetrics, 15(1):??, February 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720304855. See [MdMAGBLI21a].

Moed:2011:CUR

[MdMALIV11] Henk F. Moed, Félix de Moya-Anegón, Carmen López-Illescas, and Martijn Visser. Is concentration of university research associated with better research performance? Journal of Informetrics, 5(4):649–658, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000678.

Megnigbeto:2014:EUC

[Mêg14] Eustache Mêgnigbêto. Efficiency, unused capacity and transmission power as indicators of the Triple Helix of university–industry–government relationships. Journal of Informetrics, 8(1):284–294, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157713001181.

Maisonobe:2016:WNS

[MEG^{+}16] Marion Maisonobe, Denis Eckert, Michel Grossetti, Laurent Jégou, and Béatrice Milard. The world network of scientific collaborations between cities: domestic or international dynamics? Journal of Informetrics, 10(4):1025–1036, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716300657.

Megnigbeto:2018:MTH

[Mêg18] Eustache Mêgnigbêto. Modelling the Triple Helix of university–industry–government relationships with game theory: Core, Shapley value and nucleolus as indicators of synergy within an innovation system. Journal of Informetrics, 12(4):1118–1132, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717304315.

Meho:2019:USC

[Meh19] Lokman I. Meho. Using Scopus’s CiteScore for assessing the quality of computer science conferences. Journal of Informetrics, 13(1):419–433, February 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718304176.

Munoz-Ecija:2021:CMD

[MÉVQR21] Teresa Muñoz-Écija, Benjamín Vargas-Quesada, and Zaida Chinchilla Rodríguez. Coping with methods for delineating emerging fields: Nanoscience and nanotechnology as a case study. Journal of Informetrics, 13(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718305030.

Moehrle:2021:BSF

[MF21] Martin G. Moehrle and Jonas Frischkorn. Bridge strongly or focus — an analysis of bridging patents in four application fields of carbon fiber reinforcements. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000092.

Miranda:2018:OOR

[MGC18] Ruben Miranda and Esther Garcia-Carpintero. Overcitation and overrepresentation of review papers in the most cited papers. Journal of Informetrics, 12(4):1015–1030, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718300555.

Moreira:2009:AED

[MGLF09] Bárbara L. Moreira, Marcos André Gonçalves, Alberto H. F. Laender, and Edward A. Fox. Automatic evaluation of digital libraries with 5SQual. Journal of Informetrics, 3(2):102–123, April 2009. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157708000734.

Mrowinski:2022:AJC

[MGS22] Maciej J. Mrowinski, Marek Gagolewski, and Grzegorz Siudem. Accidentality in journal citation patterns. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000931.

Milojevic:2013:ASI

[Mil13] Stasa Milojević. Accuracy of simple, initials-based methods for author name disambiguation. Journal of Informetrics, 7(4):767–773, October 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157713000539.

Milojevic:2015:QCE

[Mil15] Stasa Milojević. Quantifying the cognitive extent of science. Journal of Informetrics, 9(4):962–973, October 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771530081X.

Mingers:2014:PSI

[Min14] John Mingers. Problems with the SNIP indicator. Journal of Informetrics, 8(4):890–894, October 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000856.

Miroiu:2013:AHI

[Mir13] Adrian Miroiu. Axiomatizing the Hirsch index: Quantity and quality disjoined. Journal of Informetrics, 7(1):10–15, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000636.

McCarty:2013:AAP

[MJ13] Christopher McCarty and James W. Jawitz. Attitudes about publishing and normal science advancement. Journal of Informetrics, 7(4):850–858, October 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000655.

Marsh:2011:GDP

[MJB11] Herbert W. Marsh, Upali W. Jayasinghe, and Nigel W. Bond. Gender differences in peer reviews of grant applications: a substantive-methodological synergy in support of the null hypothesis model. Journal of Informetrics, 5(1):167–180, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115771000091X.

Ma:2024:DOD

[MJZ^{+}24] Xiaowei Ma, Hong Jiao, Yang Zhao, Shan Huang, and Bo Yang. Does open data have the potential to improve the response of science to public health emergencies? Journal of Informetrics, 18(2):??, May 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772400018X.

Mansilla:2007:BJI

[MKCM07] R. Mansilla, E. Köppen, G. Cocho, and P. Miramontes. On the behavior of journal impact factor rank-order distribution. Journal of Informetrics, 1(2):155–160, April 2007. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157707000296.

Momeni:2022:MFA

[MKM^{+}22] Fakhri Momeni, Fariba Karimi, Philipp Mayr, Isabella Peters, and Stefan Dietze. The many facets of academic mobility and its impact on scholars’ career. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157722000323.

Mao:2021:QCD

[MLCL21] Jin Mao, Zhentao Liang, Yujie Cao, and Gang Li. Quantifying cross-disciplinary knowledge flow from the perspective of content: Introducing an approach based on knowledge memes. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157720301516.

Martinez:2021:UCT

[MLDP21] Roberto Fernandez Martinez, Ruben Lostado Lorza, Ana Alexandra Santos Delgado, and Nelson Piedra. Use of classification trees and rule-based models to optimize the funding assignment to research projects: a case study of UTPL. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157720306246.

Mason:2021:IPH

[MMCR21] Shannon Mason, Margaret K. Merga, Manuel S. González Canché, and Saiyidi Mat Roni. The internationality of published higher education scholarship: How do the ‘top’ journals compare? Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157721000262.

Martin-Martin:2018:EOA

[MMCvLLC18] Alberto Martín-Martín, Rodrigo Costas, Thed van Leeuwen, and Emilio Delgado López-Cózar. Evidence of open access of scientific publications in Google Scholar: a large-scale analysis. Journal of Informetrics, 12(3):819–841, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718301457.

Martin-Martin:2017:CWU

[MMOMHLC17] Alberto Martin-Martin, Enrique Orduna-Malea, Anne-Wil Harzing, and Emilio Delgado López-Cózar. Can we use Google Scholar to identify highly-cited documents? Journal of Informetrics, 11(1):152–163, February 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771630298X.

Martin-Martin:2018:ALM

[MMOMLC18] Alberto Martín-Martín, Enrique Orduna-Malea, and Emilio Delgado López-Cózar. Author-level metrics in the new academic profile platforms: the online behaviour of the Bibliometrics community. Journal of Informetrics, 12(2):494–509, May 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717302316.

Martin-Martin:2018:GSW

[MMOMTLC18] Alberto Martín-Martín, Enrique Orduna-Malea, Mike Thelwall, and Emilio Delgado López-Cózar. Google Scholar, Web of Science, and Scopus: a systematic comparison of citations in 252 subject categories. Journal of Informetrics, 12(4):1160–1177, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718303249.

McLevey:2017:IMS

[MMY17] John McLevey and Reid McIlroy-Young. Introducing metaknowledge: Software for computational research in information science, network analysis, and science of science. Journal of Informetrics, 11(1):176–197, February 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302000.

Mariani:2016:IMP

[MMZ16] Manuel Sebastian Mariani, Matús Medo, and Yi-Cheng Zhang. Identification of milestone papers through timebalanced network centrality. Journal of Informetrics, 10(4): 1207–1223, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1751157716301729.

Mund:2015:TES

[MN15] Carolin Mund and Peter Neuhäusler. Towards an earlystage identification of emerging topics in science — the usability of bibliometric characteristics. Journal of Informetrics, 9(4):1018–1033, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157715200405.

Mathis:2024:TET

[MO24] Bryan Mathis and Ryosuke L. Ohniwa. Trends in emerging topics generation across countries in life science and medicine. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724000658.

Moed:2010:CCI

[Moe10a] Henk F. Moed. CWTS crown indicator measures citation impact of a research group’s publication oeuvre. Journal of Informetrics, 4(3):436–438, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000313.

Moed:2010:MCC

[Moe10b] Henk F. Moed. Measuring contextual citation impact of scientific journals. Journal of Informetrics, 4(3):265–277, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157710000039.

Moed:2018:TMP

[Moe18] Henk F. Moed. Towards a multi-paradigmatic, value free informetrics: a reply to Paul Wouters’ book review “The failure of a paradigm”. Journal of Informetrics, 12(2):541– 546, May 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S1751157718300944. See [Wou18].

Mannella:2013:TDI

[MR13] Riccardo Mannella and Paolo Rossi. On the time dependence of the h-index. Journal of Informetrics, 7(1):176–182, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000855.

Milojevic:2017:CSI

[MRBI17] Stasa Milojević, Filippo Radicchi, and Judit Bar-Ilan. Citation success index — an intuitive pair-wise journal comparison metric. Journal of Informetrics, 11(1):223–231, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716301936.

Moghadasi:2013:LCE

[MRR13] Shiva Imani Moghadasi, Sri Devi Ravana, and Sudharshan N. Raman. Low-cost evaluation techniques for information retrieval systems: a review. Journal of Informetrics, 7(2):301– 312, April 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712001046.

Meyer:2009:UWR

[MS09] Eric T. Meyer and Ralph Schroeder. Untangling the web of e-research: Towards a sociology of online knowledge. Journal of Informetrics, 3(3):246–260, July 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000273.

Milojevic:2014:RHK

[MSL^{+}14] Stasa Milojević, Cassidy R. Sugimoto, Vincent Larivière, Mike Thelwall, and Ying Ding. The role of handbooks in knowledge creation and diffusion: a case of science and technology studies. Journal of Informetrics, 8(3):693–709, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115771400056X.

Min:2016:MDR

[MSPD16] Chao Min, Jianjun Sun, Lei Pei, and Ying Ding. Measuring delayed recognition for papers: Uneven weighted summation and total citations. Journal of Informetrics, 10(4):1153–1165, November 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157716300530.

Matveeva:2021:ERU

[MSY21] Nataliya Matveeva, Ivan Sterligov, and Maria Yudkevich. The effect of Russian University Excellence Initiative on publications and collaboration patterns. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720306271.

Moussa:2010:RMJ

[MT10] Salim Moussa and Mourad Touzani. Ranking marketing journals using the Google Scholar-based hg-index. Journal of Informetrics, 4(1):107–117, January 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000753.

Mutz:2016:SFA

[Mut16] Rüdiger Mutz. Some further aspects of sampling: Comment on Williams and Bornmann. Journal of Informetrics, 10(4):1241–1242, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S1751157716302565. See [WB16a, WB16b].

Ma:2022:RKR

[MYZ^{+}22] Guoshuai Ma, Qian Yuhua, Yayu Zhang, Hongren Yan, Honghong Cheng, and Zhiguo Hu. The recognition of kernel research team. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722000918.

Nane:2016:IIC

[Nan16] Gabriela F. Nane. To infer or not to infer? A comment on Williams and Bornmann. Journal of Informetrics, 10(4):1131–1134, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S1751157716302590. See [WB16a, WB16b].

Nykl:2015:ARB

[NCJ15] Michal Nykl, Michal Campr, and Karel Jezek. Author ranking based on personalized PageRank. Journal of Informetrics, 9(4):777–799, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715200181.

Nicolaisen:2008:RRR

[NF08] Jeppe Nicolaisen and Tove Faber Frandsen. The Reference Return Ratio. Journal of Informetrics, 2(2):128–135, April 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157707000764.

Nicolaisen:2012:CFS

[NF12] Jeppe Nicolaisen and Tove Faber Frandsen. Consensus formation in science modeled by aggregated bibliographic coupling. Journal of Informetrics, 6(2):276–284, April 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000733.

Nomaler:2013:DMD

[NFH13] Önder Nomaler, Koen Frenken, and Gaston Heimeriks. Do more distant collaborations have more citation impact? Journal of Informetrics, 7(4):966–971, October 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000783.

Ng:2017:CCG

[Ng17] Yew-Kwang Ng. Counting citations: Generalizing the Perry– Reny index. Journal of Informetrics, 11(3):685–688, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115771730175X.

Nielsen:2017:GCI

[Nie17] Mathias Wullum Nielsen. Gender and citation impact in management research. Journal of Informetrics, 11(4):1213– 1228, November 2017. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S175115771730216X.

Nykl:2014:PVE

[NJFD14] Michal Nykl, Karel Jezek, Dalibor Fiala, and Martin Dostal. PageRank variants in the evaluation of citation networks. Journal of Informetrics, 8(3):683–692, July 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000583.

Nane:2017:PAR

[NLC17] Gabriela F. Nane, Vincent Larivière, and Rodrigo Costas. Predicting the age of researchers using bibliometric data. Journal of Informetrics, 11(3):713–729, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303443.

Nguyen:2021:GEG

[NLK^{+}21] Ai Linh Nguyen, Wenyuan Liu, Khiam Aik Khor, Andrea Nanetti, and Siew Ann Cheong. The golden eras of graphene science and technology: Bibliographic evidences from journal and patent publications. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157719303542.

Nakajima:2023:QGI

[NLSM23] Kazuki Nakajima, Ruodan Liu, Kazuyuki Shudo, and Naoki Masuda. Quantifying gender imbalance in East Asian academia: Research career and citation practice. Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000858.

Nassiri:2013:NSI

[NMNJM13] Isar Nassiri, Ali Masoudi-Nejad, Mahdi Jalili, and Ali Moeini. Normalized similarity index: an adjusted index to prioritize article citations. Journal of Informetrics, 7(1):91– 98, January 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712000661.

Norris:2007:CAW

[NO07] Michael Norris and Charles Oppenheim. Comparing alternatives to the Web of Science for coverage of the social sciences’ literature. Journal of Informetrics, 1(2):161–169, April 2007. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157706000228.

Norris:2010:PRI

[NO10] Michael Norris and Charles Oppenheim. Peer review and the h-index: Two studies. Journal of Informetrics, 4(3): 221–232, July 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157709000790.

Nieminen:2013:RLC

[NPS13] Paavo Nieminen, Ilkka Pölönen, and Tuomo Sipola. Research literature clustering using diffusion maps. Journal of Informetrics, 7(4):874–886, October 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000680.

Nair:2012:SI

[NT12] Gopalan M. Nair and Berwin A. Turlach. The stochastic hindex. Journal of Informetrics, 6(1):80–87, January 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157711000836.

Nair:2022:MID

[NV22] N. Unnikrishnan Nair and B. Vineshkumar. Modelling informetric data using quantile functions. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722000189.

Niu:2016:WPY

[NZZ^{+}16] Qikai Niu, Jianlin Zhou, An Zeng, Ying Fan, and Zengru Di. Which publication is your representative work? Journal of Informetrics, 10(3):842–853, August 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716300359.

Ortega:2010:DBW

[OA10] José Luis Ortega and Isidro Aguillo. Differences between web sessions according to the origin of their visits. Journal of Informetrics, 4(3):331–337, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000076.

Ortega:2013:ICC

[OA13] José Luis Ortega and Isidro F. Aguillo. Institutional and country collaboration in an online service of scientific profiles:

Google Scholar Citations. Journal of Informetrics, 7(2):394– 403, April 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000023.

Ortega:2013:MPE

[OG13] Francisco J. Ortega and Jose M. Gavilan. The measurement of production efficiency in scientific journals through stochastic frontier analysis models: Application to quantitative economics journals. Journal of Informetrics, 7(4):959– 965, October 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771300076X.

Opthof:2010:CJF

[OL10] Tobias Opthof and Loet Leydesdorff. Caveats for the journal and field normalizations in the CWTS (“Leiden”) evaluations of research performance. Journal of Informetrics, 4(3): 423–430, July 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000106. See reply

[vRvLV^{+}10].

Ortega:2011:MAC

[OLRF11] José Luis Ortega, Elena López-Romero, and Inés Fernández. Multivariate approach to classify research institutes according to their outputs: the case of the CSIC’s institutes. Journal of Informetrics, 5(3):323–332, July 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000058.

Orzechowski:2023:ASI

[OMFF23] Kamil P. Orzechowski, Maciej J. Mrowinski, Agata Fronczak, and Piotr Fronczak. Asymmetry of social interactions and its role in link predictability: the case of coauthorship networks. Journal of Informetrics, 17(2):??, May 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000305.

Onodera:2016:PIC

[Ono16] Natsuo Onodera. Properties of an index of citation durability of an article. Journal of Informetrics, 10(4):981–1004, November 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157716300086.

Ocholla:2010:CIE

[OOB10] Dennis N. Ocholla, Omwoyo Bosire Onyancha, and Johannes Britz. Can information ethics be conceptualized by using the core/periphery model? Journal of Informetrics, 4(4):492– 502, October 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000398.

Ortega:2014:ICA

[Ort14] José Luis Ortega. Influence of co-authorship networks in the research impact: Ego network analyses from Microsoft Academic Search. Journal of Informetrics, 8(3):728–737, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157714000613.

Ortega:2015:RBA

[Ort15] José Luis Ortega. Relationship between altmetric and bibliometric indicators across academic social sites: the case of CSIC’s members. Journal of Informetrics, 9(1):39–49, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157714000984.

Ortega:2018:LCA

[Ort18] José Luis Ortega. The life cycle of altmetric impact: a longitudinal study of six metrics from PlumX. Journal of Informetrics, 12(3):579–589, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717302870.

Ortega:2021:HDM

[Ort21a] José Luis Ortega. How do media mention research papers? Structural analysis of blogs and news networks using citation coupling. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157721000468.

Ortega:2021:PCA

[Ort21b] José Luis Ortega. Proposal of composed altmetric indicators based on prevalence and impact dimensions. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719304845.

Onder:2021:SFD

[ÖSY21] Ali Sina Önder, Sascha Schweitzer, and Hakan Yilmazkuday. Specialization, field distance, and quality in economists’ collaborations. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721000936.

Ou:2024:ERF

[OZZW24] Guiyan Ou, Kang Zhao, Renxian Zuo, and Jiang Wu. Effects of research funding on the academic impact and societal visibility of scientific research. Journal of Informetrics, 18(4): ??, November 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724001044.

Pajic:2015:SCB

[Paj15] Dejan Pajić. On the stability of citation-based journal rankings. Journal of Informetrics, 9(4):990–1006, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115771530033X.

Pooladian:2016:LSB

[PB16] Aida Pooladian and Ángel Borrego. A longitudinal study of the bookmarking of library and information science literature in Mendeley. Journal of Informetrics, 10(4):1135–1142, November 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157716301857.

Perlin:2024:DIR

[PBIR24] Marcelo Perlin, Denis Borenstein, Takeyoshi Imasato, and Marcos Reichert. The determinants and impact of research grants: the case of Brazilian productivity scholarships. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724000762.

Piwowar:2010:PSR

[PC10] Heather A. Piwowar and Wendy W. Chapman. Public sharing of research datasets: a pilot study of associations. Journal of Informetrics, 4(2):148–156, April 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000881.

Pradhan:2021:ACI

[PCCN21] Dinesh K. Pradhan, Joyita Chakraborty, Prasenjit Choudhary, and Subrata Nandi. An automated conflict of interest based greedy approach for conference paper assignment system. Journal of Informetrics, 14(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719301373.

Patelli:2017:SIN

[PCPG17] Aurelio Patelli, Giulio Cimini, Emanuele Pugliese, and Andrea Gabrielli. The scientific influence of nations on global scientific and technological development. Journal of Informetrics, 11(4):1229–1237, November 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717301335.

Pech:2021:SMH

[PD21] Gerson Pech and Catarina Delgado. Screening the most highly cited papers in longitudinal bibliometric studies and systematic literature reviews of a research field or journal: Widespread used metrics vs a percentile citation-based approach. Journal of Informetrics, 15(3):??, August 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000328.

Porwal:2024:SIA

[PD24] Priya Porwal and Manoj H. Devare. Scientific impact analysis: Unraveling the link between linguistic properties and citations. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724000397.

Peng:2015:AMP

[Pen15] Tai-Quan Peng. Assortative mixing, preferential attachment, and triadic closure: a longitudinal study of tiegenerative mechanisms in journal citation networks. Journal of Informetrics, 9(2):250–262, April 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000206.

Perc:2010:GSS

[Per10a] Matjaz Perc. Growth and structure of Slovenia’s scientific collaboration network. Journal of Informetrics, 4(4):475– 482, October 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000362.

Perc:2010:ZLL

[Per10b] Matjaz Perc. Zipf’s law and log-normal distributions in measures of scientific output across fields and institutions: 40 years of Slovenia’s research as an example. Journal of Informetrics, 4(3):358–364, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000234.

Persson:2010:IRT

[Per10c] Olle Persson. Identifying research themes with weighted direct citation links. Journal of Informetrics, 4(3):415–422, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000283.

Persson:2017:BAE

[Per17a] Rasmus A. X. Persson. Bibliometric author evaluation through linear regression on the coauthor network. Journal of Informetrics, 11(1):299–306, February 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303194.

Persson:2017:NTI

[Per17b] Rasmus A. X. Persson. Note on the theory of the w-index. Journal of Informetrics, 11(3):745–747, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771630390X.

Peters:2017:WUJ

[Pet17] Gjalt-Jorn Ygram Peters. Why not to use the journal impact factor as a criterion for the selection of junior researchers: a comment on Bornmann and Williams (2017). Journal of Informetrics, 11(3):888–891, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157717302365. See [BW17b] and rejoin-

der [BW17a].

Peter:2018:BRB

[Pet18] Ingwersen Peter. Book review: Becoming Metric-Wise: a Bibliometric Guide for Researchers, Ronald Rousseau, Leo Egghe and Raf Guns, Elsevier/Chandos Publishing (2018). 385 pages. Journal of Informetrics, 12(3):703–705, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718302335.

Petersen:2021:MMM

[Pet21] Alexander M. Petersen. Megajournal mismanagement: Manuscript decision bias and anomalous editor activity at PLOS ONE. Journal of Informetrics, 13(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718302724.

Perez-Hornero:2013:AJI

[PHANPP13] Patricia Pérez-Hornero, José Pablo Arias-Nicolás, Antonio A. Pulgarín, and Antonio Pulgarín. An annual JCR impact factor calculation based on Bayesian credibility formulas. Journal of Informetrics, 7(1):1–9, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000648.

Paul-Hus:2017:SIA

[PHMSML17] Adèle Paul-Hus, Philippe Mongeon, Maxime Sainte-Marie, and Vincent Larivière. The sum of it all: Revealing collaboration patterns by combining authorship and acknowledgements. Journal of Informetrics, 11(1):80–87, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716302723.

Pislyakov:2022:SPM

[Pis22] Vladimir Pislyakov. On some properties of medians, percentiles, baselines, and thresholds in empirical bibliometric analysis. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722000748.

Park:2013:DSS

[PL13] Han Woo Park and Loet Leydesdorff. Decomposing social and semantic networks in emerging “big data” research. Journal of Informetrics, 7(3):756–765, July 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000473.

Petridis:2013:EAF

[PMAT13] Konstantinos Petridis, Chrisovalantis Malesios, Garyfallos Arabatzis, and Emmanuel Thanassoulis. Efficiency analysis of forestry journals: Suggestions for improving journals’ quality. Journal of Informetrics, 7(2):505–521, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000175.

Prathap:2021:WJS

[PML21] Gangan Prathap, Somenath Mukherjee, and Loet Leydesdorff. Within-journal self-citations and the Pinski–Narin influence weights. Journal of Informetrics, 14(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719302950.

Patelli:2023:GSC

[PNCG23] Aurelio Patelli, Lorenzo Napolitano, Giulio Cimini, and Andrea Gabrielli. Geography of science: Competitiveness and inequality. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722001109.

Purkayastha:2019:CTA

[PPFKB19] Amrita Purkayastha, Eleonora Palmaro, Holly J. Falk-Krzesinski, and Jeroen Baas. Comparison of two articlelevel, field-independent citation metrics: Field-Weighted Citation Impact (FWCI) and Relative Citation Ratio (RCR). Journal of Informetrics, 13(2):635–642, May 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718303559.

Parolo:2015:ADS

[PPG^{+}15] Pietro Della Briotta Parolo, Raj Kumar Pan, Rumi Ghosh, Bernardo A. Huberman, Kimmo Kaski, and Santo Fortunato. Attention decay in science. Journal of Informetrics, 9(4):734–745, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157715200442.

Pan:2018:MSI

[PPPF18] Raj K. Pan, Alexander M. Petersen, Fabio Pammolli, and Santo Fortunato. The memory of science: Inflation, myopia, and the knowledge network. Journal of Informetrics, 12(3):656–678, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717303139.

Prathap:2013:BAI

[Pra13a] Gangan Prathap. Bornmann’s alternative to the h index. Journal of Informetrics, 7(1):209, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000910. See [Bor13a].

Prathap:2013:EB

[Pra13b] Gangan Prathap. Energetics of the h-bubble. Journal of Informetrics, 7(2):439–441, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000084.

Prathap:2013:IC

[Pra13c] Gangan Prathap. The S-index and consistency. Journal of Informetrics, 7(2):540–541, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000199.

Perianes-Rodriguez:2009:SHI

[PRCRVQ^{+}09] Antonio Perianes-Rodríguez, Zaida Chinchilla-Rodríguez, Benjamín Vargas-Quesada, Carlos Olmeda Gómez, and Félix Moya-Anegón. Synthetic hybrid indicators based on scientific collaboration to quantify and evaluate individual research results. Journal of Informetrics, 3(2):91–101, April 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157708000722.

Perianes-Rodriguez:2015:MVF

[PRRC15] Antonio Perianes-Rodriguez and Javier Ruiz-Castillo. Multiplicative versus fractional counting methods for co-authored publications. The case of the 500 universities in the Leiden Ranking. Journal of Informetrics, 9(4):974–989, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115771530050X.

Perianes-Rodriguez:2017:CWS

[PRRC17] Antonio Perianes-Rodriguez and Javier Ruiz-Castillo. A comparison of the Web of Science and publication-level classification systems of science. Journal of Informetrics, 11(1): 32–45, February 2017. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157716302371.

Perianes-Rodriguez:2016:CBN

[PRWvE16] Antonio Perianes-Rodriguez, Ludo Waltman, and Nees Jan van Eck. Constructing bibliometric networks: a comparison between full and fractional counting. Journal of Informetrics, 10(4):1178–1195, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302036.

Petersen:2013:IGR

[PS13] Alexander M. Petersen and Sauro Succi. The Z-index: a geometric representation of productivity and impact which accounts for information in the entire rank-citation profile. Journal of Informetrics, 7(4):823–832, October 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000564.

Potter:2021:ICC

[PSA21] Ross W. K. Potter, Martin Szomszor, and Jonathan Adams. Interpreting CNCIs on a country-scale: the effect of domestic and international collaboration type. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720301188.

Pandey:2021:ARC

[PSG^{+}21] Pradumn Kumar Pandey, Mayank Singh, Pawan Goyal, Animesh Mukherjee, and Soumen Chakrabarti. Analysis of reference and citation copying in evolving bibliographic networks. Journal of Informetrics, 14(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719303384.

Perlin:2017:BSO

[PSI^{+}17] Marcelo S. Perlin, André A. P. Santos, Takeyoshi Imasato, Denis Borenstein, and Sergio Da Silva. The Brazilian scientific output published in journals: a study based on a large CV database. Journal of Informetrics, 11(1):18–31, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716301559.

Perme:2012:CCD

[PSZZ12] Maja Pohar Perme, Janez Stare, Rok Zaucer, and Matjaz Zaucer. Comparison of the citation distribution and h-index between groups of different sizes. Journal of Informetrics, 6(4):712–720, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157712000582.

Prabowo:2008:FTS

[PT08] Rudy Prabowo and Mike Thelwall. Finding and tracking subjects within an ongoing debate. Journal of Informetrics, 2(2): 107–127, April 2008. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157707000752.

Prabowo:2009:SAC

[PT09] Rudy Prabowo and Mike Thelwall. Sentiment analysis: a combined approach. Journal of Informetrics, 3(2):143–157, April 2009. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157709000108.

Prabowo:2007:GOT

[PTA07] Rudy Prabowo, M. Thelwall, and Mikhail Alexandrov. Generating overview timelines for major events in an RSS corpus. Journal of Informetrics, 1(2):131–144, April 2007. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115770600023X.

Panagopoulos:2017:DRS

[PTV17] George Panagopoulos, George Tsatsaronis, and Iraklis Varlamis. Detecting rising stars in dynamic collaborative networks. Journal of Informetrics, 11(1):198–222, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716300645.

Prathap:2021:SRR

[PUKR21] Gangan Prathap, Ephrance Abu Ujum, Sameer Kumar, and Kuru Ratnavelu. Scoring the resourcefulness of researchers using bibliographic coupling patterns. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000390.

Pinto:2021:SCN

[PVH21] Pablo E. Pinto, Andres Vallone, and Guillermo Honores. The structure of collaboration networks: Findings from three decades of co-invention patents in Chile. Journal of Informetrics, 13(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719302202.

Park:2019:RSC

[PW19] Hyoungjoo Park and Dietmar Wolfram. Research software citation in the Data Citation Index: Current practices and implications for research software sharing and reuse. Journal of Informetrics, 13(2):574–582, May 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302372.

Pak:2021:ATC

[PWY21] Chol Myong Pak, Weibin Wang, and Guang Yu. An analysis of in-text citations based on fractional counting. Journal of Informetrics, 14(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719301567.

Peng:2024:BGP

[PXS24] Xianzhe Peng, Huixin Xu, and Jin Shi. Are the bibliometric growth patterns of excellent scholars similar? From the analysis of ACM Fellows. Journal of Informetrics, 18(3): ??, August 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000567.

Park:2018:TOD

[PY18] Inchae Park and Byungun Yoon. Technological opportunity discovery for technological convergence based on the prediction of technology knowledge flow in a citation network. Journal of Informetrics, 12(4):1199–1222, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718300907.

Pan:2018:EUC

[PYCH18] Xuelian Pan, Erjia Yan, Ming Cui, and Weina Hua. Examining the usage, citation, and diffusion patterns of bibliometric mapping software: a comparative study of three tools. Journal of Informetrics, 12(2):481–493, May 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717303292.

Pan:2019:HIS

[PYCH19] Xuelian Pan, Erjia Yan, Ming Cui, and Weina Hua. How important is software to library and information science research? A content analysis of full-text publications. Journal of Informetrics, 13(1):397–406, February 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718304504.

Peng:2024:RAI

[PYSM24] Wen Peng, Mingliang Yue, Mingyue Sun, and Tingcan Ma. Revision and academic impact: a case study of bioRxiv preprint papers. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723001098.

Pan:2015:AIS

[PYWH15] Xuelian Pan, Erjia Yan, Qianqian Wang, and Weina Hua. Assessing the impact of software on science: a bootstrapped learning of software entities in full-text papers. Journal of Informetrics, 9(4):860–871, October 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715300602.

Quirin:2010:GBD

[QCVQdMA10] Arnaud Quirin, Oscar Cordón, Benjamín Vargas-Quesada, and Félix de Moya-Anegón. Graph-based data mining: a new tool for the analysis and comparison of scientific domains represented as scientograms. Journal of Informetrics, 4(3): 291–312, July 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000052.

Qian:2021:UHS

[QLS21] Yue Qian, Yu Liu, and Quan Z. Sheng. Understanding hierarchical structural evolution in a scientific discipline: a case study of artificial intelligence. Journal of Informetrics, 14 (3):??, August 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157719302925.

Qun:2023:EES

[QM23] Zhao Qun and Yang Menghui. An efficient entropy of sum approach for measuring diversity and interdisciplinarity. Journal of Informetrics, 17(3):??, August 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000500.

Quesada:2009:MHI

[Que09] Antonio Quesada. Monotonicity and the Hirsch index. Journal of Informetrics, 3(2):158–160, April 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000091.

Quesada:2011:AHI

[Que11] Antonio Quesada. Axiomatics for the Hirsch index and the Egghe index. Journal of Informetrics, 5(3):476–480, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157711000101.

Qian:2024:NAE

[QZY^{+}24] Minghui Qian, Mengchun Zhao, Jianliang Yang, Guancan Yang, Jiayuan Xu, and Xusen Cheng. A novel approach to enterprise technical collaboration: Recommending R&D partners through technological similarity and complementarity. Journal of Informetrics, 18(4):??, November 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772400083X.

Rorstad:2015:PRE

[RA15] Kristoffer Rørstad and Dag W. Aksnes. Publication rate expressed by age, gender and academic position — a large-scale analysis of Norwegian academic staff. Journal of Informetrics, 9(2):317–333, April 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000218.

Rodriguez:2007:MBB

[RBV07] Marko A. Rodriguez, Johan Bollen, and Herbert Van de Sompel. Mapping the bid behavior of conference referees. Journal of Informetrics, 1(1):68–82, January 2007. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157706000101.

Radicchi:2012:TFC

[RC12a] Filippo Radicchi and Claudio Castellano. Testing the fairness of citation indicators for comparison across scientific domains: the case of fractional citation counts. Journal of Informetrics, 6(1):121–130, January 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000794. See comments [Sir12].

Radicchi:2012:WSC

[RC12b] Filippo Radicchi and Claudio Castellano. Why Sirtes’s claims (Sirtes, 2012) do not square with reality. Journal of Informetrics, 6(4):615–618, October 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000442. See [Sir12].

Ruiz-Castillo:2014:CCS

[RC14] Javier Ruiz-Castillo. The comparison of classificationsystem-based normalization procedures with source normalization alternatives in Waltman and Van Eck (2013). Journal of Informetrics, 8(1):25–28, January 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000795. See [WvE13].

Ruiz-Castillo:2016:ROI

[RC16] Javier Ruiz-Castillo. Research output indicators are not productivity indicators. Journal of Informetrics, 10(2):661–663, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115771630092X.

Ruiz-Castillo:2014:SSP

[RCC14] Javier Ruiz-Castillo and Rodrigo Costas. The skewness of scientific productivity. Journal of Informetrics, 8(4):917–934, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115771400087X.

Ruiz-Castillo:2018:IFC

[RCC18] Javier Ruiz-Castillo and Rodrigo Costas. Individual and field citation distributions in 29 broad scientific fields. Journal of Informetrics, 12(3):868–892, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718300853.

Ruiz-Castillo:2015:FNC

[RCW15] Javier Ruiz-Castillo and Ludo Waltman. Field-normalized citation impact indicators using algorithmically constructed classification systems of science. Journal of Informetrics, 9(1):102–117, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1751157714001126.

Rossi:2018:TMA

[RDF^{+}18] Luciano Rossi, Rafael J. P. Damaceno, Igor L. Freire, Etelvino J. H. Bechara, and Jesús P. Mena-Chalco. Topological metrics in academic genealogy graphs. Journal of Informetrics, 12(4):1042–1058, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718300750.

Rehs:2021:SML

[Reh21] Andreas Rehs. A supervised machine learning approach to author disambiguation in the Web of Science. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000377.

Rossi:2017:GIM

[RFMC17] Luciano Rossi, Igor L. Freire, and Jesús P. Mena-Chalco. Genealogical index: a metric to analyze advisor-advisee relationships. Journal of Informetrics, 11(2):564–582, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716302632.

Robinson-Garcia:2017:DNB

[RGMJC17] Nicolas Robinson-Garcia, Philippe Mongeon, Wei Jeng, and Rodrigo Costas. DataCite as a novel bibliometric source: Coverage, strengths and limitations. Journal of Informetrics, 11(3):841–854, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717300834.

Rahman:2015:EEP

[RGRE15] A. I. M. Jakaria Rahman, Raf Guns, Ronald Rousseau, and Tim C. E. Engels. Is the expertise of evaluation panels congruent with the research interests of the research groups: a quantitative approach based on barycenters. Journal of Informetrics, 9(4):704–721, October 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S1751157715000553. See corrigendum [RGRE16].

Rahman:2016:CSE

[RGRE16] A. I. M. Jakaria Rahman, Raf Guns, Ronald Rousseau, and Tim C. E. Engels. Corrigendum to “Is the expertise of evaluation panels congruent with the research interests of the research groups: a quantitative approach based on barycenters” [Journal of Informetrics 9 (4) (2015) 704–721]. Journal of Informetrics, 10(4):1052–1054, November 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302346. See [RGRE15].

Rousseau:2017:MCD

[RGRE17] Ronald Rousseau, Raf Guns, A. I. M. Jakaria Rahman, and Tim C. E. Engels. Measuring cognitive distance between publication portfolios. Journal of Informetrics, 11(2):583– 594, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157717300743.

Robinson-Garcia:2019:MFM

[RGSM^{+}19] Nicolás Robinson-Garcia, Cassidy R. Sugimoto, Dakota Murray, Alfredo Yegros-Yegros, Vincent Larivière, and Rodrigo Costas. The many faces of mobility: Using bibliometric data to measure the movement of scientists. Journal of Informetrics, 13(1):50–63, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718300865.

Rousseau:2013:B

[RGZSC13] Ronald Rousseau, Carlos García-Zorita, and Elias Sanz-Casado. The h-bubble. Journal of Informetrics, 7(2):294– 300, April 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712001022.

Rousseau:2023:PDC

[RGZSC23] Ronald Rousseau, Carlos Garcia-Zorita, and Elías Sanz-Casado. Publications during COVID-19 times: an unexpected overall increase. Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115772300086X.

Rousseau:2013:TTS

[RH13] Ronald Rousseau and Xiaojun Hu. Two time series, their meaning and some applications. Journal of Informetrics, 7(3):603–610, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1751157713000266.

Rodriguez:2008:MTA

[RJDD08] Victor Rodriguez, Frizo Janssens, Koenraad Debackere, and Bart De Moor. On material transfer agreements and visibility of researchers in biotechnology. Journal of Informetrics, 2(1): 89–100, January 2008. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157707000739.

Reingewertz:2018:AGB

[RL18] Yaniv Reingewertz and Carmela Lutmar. Academic in-group bias: an empirical examination of the link between author and journal affiliation. Journal of Informetrics, 12(1):74– 86, February 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157717302456.

Rojko:2022:SPA

[RL22] Katarina Rojko and Borut Luzar. Scientific performance across research disciplines: Trends and differences in the case of Slovenia. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115772200013X.

Rousseau:2013:MPM

[RLG13] Ronald Rousseau, Yuxian Liu, and Raf Guns. Mathematical properties of Q-measures. Journal of Informetrics, 7(3):737– 745, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157713000497. See addendum

and correction [RLG14].

Rousseau:2014:ACM

[RLG14] Ronald Rousseau, Yuxian Liu, and Raf Guns. An addendum and correction to “Mathematical properties of Q-measures” (vol. 7, issue 3, pp. 737–745). Journal of Informetrics, 8(3):486–490, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714000054. See [RLG13].

Rousseau:2012:PID

[RLY12] Ronald Rousseau, Yuxian Liu, and Fred Y. Ye. A preliminary investigation on diffusion through a layered system. Journal of Informetrics, 6(2):177–191, April 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000065.

Rodriguez-Navarro:2018:DRA

[RNB18a] Alonso Rodríguez-Navarro and Ricardo Brito. Double rank analysis for research assessment. Journal of Informetrics, 12(1):31–41, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S1751157717302419.

Rodriguez-Navarro:2018:TRE

[RNB18b] Alonso Rodríguez-Navarro and Ricardo Brito. Technological research in the EU is less efficient than the world average. EU research policy risks Europeans’ future. Journal of Informetrics, 12(3):718–731, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718300701.

Rodriguez-Navarro:2024:RAM

[RNB24] Alonso Rodríguez-Navarro and Ricardo Brito. Rank analysis of most cited publications, a new approach for research assessments. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724000166.

Rons:2012:PBF

[Ron12] Nadine Rons. Partition-based Field Normalization: an approach to highly specialized publication records. Journal of Informetrics, 6(1):1–10, January 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S175115771100085X. See corrigendum [Ron13].

Rons:2013:CPB

[Ron13] Nadine Rons. Corrigendum to “Partition-based Field Normalization: an approach to highly specialized publication records” J. Informetrics 6 (2012) 1–10. Journal of Informetrics, 7(1):99, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000764. See [Ron12].

Rons:2018:BAS

[Ron18] Nadine Rons. Bibliometric approximation of a scientific specialty by combining key sources, title words, authors and references. Journal of Informetrics, 12(1):113–132, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157716303571.

Rousseau:2007:IMP

[Rou07] Ronald Rousseau. The influence of missing publications on the Hirsch index. Journal of Informetrics, 1(1):2–7, January 2007. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157706000022.

Rousseau:2008:WAI

[Rou08] Ronald Rousseau. Woeginger’s axiomatisation of the h-index and its relation to the g-index, the h^{(}2)-index and the R^{2}- index. Journal of Informetrics, 2(4):335–340, October 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157708000345.

Rousseau:2014:CIC

[Rou14a] Ronald Rousseau. Comments on “Impact coverage of the success-index” by Leo Egghe. Journal of Informetrics, 8(3): 491–492, July 2014. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771400042X. See [Egg14].

Rousseau:2014:REI

[Rou14b] Ronald Rousseau. A refinement of Egghe’s increment studies. Journal of Informetrics, 8(1):212–216, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713001120.

Rousseau:2016:PCB

[Rou16] Ronald Rousseau. Positive correlation between journal production and journal impact factors. Journal of Informetrics, 10(2):567–568, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157716301031.

Rousseau:2019:LWB

[Rou19] Ronald Rousseau. On the Leydesdorff–Wagner–Bornmann proposal for diversity measurement. Journal of Informetrics, 13(3):906–907, August 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157719300987.

Rodriguez:2008:RBS

[RP08] Marko A. Rodriguez and Alberto Pepe. On the relationship between the structural and socioacademic communities of a coauthorship network. Journal of Informetrics, 2(3):195–201, July 2008. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157708000230.

Rahman:2017:NQA

[RRKH17] Mohammad Tariqur Rahman, Joe Mac Regenstein, Noor Lide Abu Kassim, and Nazmul Haque. The need to quantify authors’ relative intellectual contributions in a multiauthor paper. Journal of Informetrics, 11(1):275–281, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157717300111.

Rao:2010:GJA

[RS10a] I. K. Ravichandra Rao and Divya Srivastava. Growth of journals, articles and authors in malaria research. Journal of Informetrics, 4(3):249–256, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000996.

Rodriguez:2010:EMR

[RS10b] Marko A. Rodriguez and Joshua Shinavier. Exposing multirelational networks to single-relational network analysis algorithms. Journal of Informetrics, 4(1):29–41, January 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157709000558.

Reznik:2016:HRH

[RS16] Ilia Reznik and Vladimir Shatalov. Hidden revolution of human priorities: an analysis of biographical data from Wikipedia. Journal of Informetrics, 10(1):124–131, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715301334.

Righetto:2019:LTC

[RSRP19] Lorenzo Righetto, Alessandro Spelta, Emanuele Rabosio, and Fabio Pammolli. Long-term correlations in short, nonstationary time series: an application to international R&D collaborations. Journal of Informetrics, 13(2):583–592, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718302499.

Raghuvanshi:2024:PTE

[RV24] Adarsh Raghuvanshi and Vinayak. Profiling team exploration strategies of collaborating authors from artificial intelligence in computer science. Journal of Informetrics, 18(4): ??, November 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000981.

Rousseau:2010:AIM

[RW10] Ronald Rousseau and Dietmar Wolfram. The ASIS&T– ISSI “metrics” pre-conference seminar and the Global Alliance. Journal of Informetrics, 4(2):139–140, April 2010.

CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115770900087X.

Radicchi:2017:QPI

[RWB17] Filippo Radicchi, Alexander Weissman, and Johan Bollen. Quantifying perceived impact of scientific publications. Journal of Informetrics, 11(3):704–712, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717300846.

Roth:2012:AIQ

[RWL12] Camille Roth, Jiang Wu, and Sergi Lozano. Assessing impact and quality from local dynamics of citation networks. Journal of Informetrics, 6(1):111–120, January 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000770.

Rousseau:2015:ISI

[RXL15] Ronald Rousseau, Fang Xu, and Wenbin Liu. Interpolated sub-impact factor (SIF) sequences for journal rankings. Journal of Informetrics, 9(4):907–914, October 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771530078X.

Rousseau:2012:RAI

[RY12a]  Ronald Rousseau and Liying Yang.  Reflections on

the activity index and related indicators. Journal of Informetrics, 6(3):413–421, July 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157712000053.

Rousseau:2012:FRB

[RY12b] Ronald Rousseau and Fred Y. Ye. A formal relation between the h-index of a set of articles and their I3 score. Journal of Informetrics, 6(1):34–35, January 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000812.

Roh:2023:DTS

[RY23] Taeyeoun Roh and Byungun Yoon. Discovering technology and science innovation opportunity based on sentence generation algorithm. Journal of Informetrics, 17(2):??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723000287.

Rousseau:2010:DPI

[RYY10] Ronald Rousseau, Liying Yang, and Ting Yue. A discussion of Prathap’s h _{2}-index for institutional evaluation with an application in the field of HIV infection and therapy. Journal of Informetrics, 4(2):175–184, April 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000856.

Rousseau:2015:GCF

[RZ15] Ronald Rousseau and Star X. Zhao. A general conceptual framework for characterizing the ego in a network. Journal of Informetrics, 9(1):145–149, January 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771400114X.

Ruan:2021:PCC

[RZLC21] Xuanmin Ruan, Yuanyang Zhu, Jiang Li, and Ying Cheng. Predicting the citation counts of individual papers via a BP neural network. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719303979.

Sjogaarde:2018:GAC

[SA18] Peter Sjögårde and Per Ahlgren. Granularity of algorithmically constructed publication-level classifications of research publications: Identification of topics. Journal of Informetrics, 12(1):133–152, February 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717303371.

Silva:2016:UNS

[SAB^{+}16] Filipi N. Silva, Diego R. Amancio, Maria Bardosova, Luciano da F. Costa, and Osvaldo N. Oliveira, Jr. Using network science and text analytics to produce surveys in a scientific topic. Journal of Informetrics, 10(2):487–502, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715301966.

Shahzad:2022:QOL

[SAFR22] Murtuza Shahzad, Hamed Alhoori, Reva Freedman, and Shaikh Abdul Rahman. Quantifying the online long-term interest in research. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722000402.

Sangwal:2011:GCP

[San11a] K. Sangwal. On the growth of citations of publication output of individual authors. Journal of Informetrics, 5(4):554–564, October 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711000447.

Sangwal:2011:PNM

[San11b] Keshra Sangwal. Progressive nucleation mechanism and its application to the growth of journals, articles and authors in scientific fields. Journal of Informetrics, 5(4):529–536, October 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711000423.

Sangwal:2013:CIF

[San13a] Keshra Sangwal. Citation and impact factor distributions of scientific journals published in individual countries. Journal of Informetrics, 7(2):487–504, April 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000138.

Sangwal:2013:CDM

[San13b] Keshra Sangwal. Comparison of different mathematical functions for the analysis of citation distribution of papers of individual authors. Journal of Informetrics, 7(1):36–49, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000776.

Sangwal:2014:DCP

[San14] Keshra Sangwal. Distributions of citations of papers of individual authors publishing in different scientific disciplines: Application of Langmuir-type function. Journal of Informetrics, 8(4):972–984, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771400090X.

Sangwal:2015:GDC

[San15] Keshra Sangwal. On the growth dynamics of citations of articles by some Nobel Prize winners. Journal of Informetrics, 9(3):466–476, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000449.

Subochev:2018:RJU

[SAP18] Andrey Subochev, Fuad Aleskerov, and Vladimir Pislyakov. Ranking journals using social choice theory methods: a novel approach in bibliometrics. Journal of Informetrics, 12(2): 416–429, May 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157718300919.

Sarabia:2008:GDL

[Sar08] José María Sarabia. A general definition of the Leimkuhler curve. Journal of Informetrics, 2(2):156–163, April 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157708000138.

Serenko:2011:WFE

[SB11] Alexander Serenko and Nick Bontis. What’s familiar is excellent: the impact of exposure effect on perceived journal quality. Journal of Informetrics, 5(1):219–223, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157710000696.

Serenko:2013:FBD

[SB13] Alexander Serenko and Nick Bontis. First in, best dressed: the presence of order-effect bias in journal ranking surveys. Journal of Informetrics, 7(1):138–144, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000879.

Simoes:2021:SCS

[SC21] Nadia Simoes and Nuno Crespo. Self-citations and scientific evaluation: Leadership, influence, and performance. Journal of Informetrics, 14(1):??, February 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719302640.

Serenko:2011:SPK

[SCBB11] Alexander Serenko, Raymond A. K. Cox, Nick Bontis, and Lorne D. Booker. The superstar phenomenon in the knowledge management and intellectual capital academic discipline. Journal of Informetrics, 5(3):333–345, July 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771100006X.

Schreiber:2008:MII

[Sch08] Michael Schreiber. A modification of the h-index: the h m-index accounts for multi-authored manuscripts. Journal of Informetrics, 2(3):211–216, July 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157708000254.

Schreiber:2010:CSM

[Sch10a] Michael Schreiber. A case study of the modified g index: Counting multi-author publications fractionally. Journal of Informetrics, 4(4):636–643, October 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000672.

Schreiber:2010:HMI

[Sch10b] Michael Schreiber. How to modify the g-index for multiauthored manuscripts. Journal of Informetrics, 4(1):42–54, January 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157709000571.

Schreiber:2010:NFO

[Sch10c] Michael Schreiber. A new family of old Hirsch index variants. Journal of Informetrics, 4(4):647–651, October 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157710000404.

Schubert:2012:JDN

[Sch12] András Schubert. Jazz discometrics — a network approach. Journal of Informetrics, 6(4):480–484, October 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000326.

Schall:2013:MCP

[Sch13a] Daniel Schall. Measuring contextual partner importance in scientific collaboration networks. Journal of Informetrics, 7(3):730–736, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157713000461.

Schneider:2013:CUS

[Sch13b] Jesper W. Schneider. Caveats for using statistical significance tests in research assessments. Journal of Informetrics, 7(1): 50–62, January 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771200065X.

Schreiber:2013:CSA

[Sch13c] Michael Schreiber. A case study of the arbitrariness of the h-index and the highly-cited-publications indicator. Journal of Informetrics, 7(2):379–387, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712001174.

Schreiber:2013:HRP

[Sch13d] Michael Schreiber. How relevant is the predictive power of the h-index? A case study of the time-dependent Hirsch index. Journal of Informetrics, 7(2):325–329, April 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000035.

Schreiber:2013:HDA

[Sch13e] Michael Schreiber. How to derive an advantage from the arbitrariness of the g-index. Journal of Informetrics, 7(2): 555–561, April 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000187.

Schreiber:2014:ECB

[Sch14a] Michael Schreiber. Examples for counterintuitive behavior of the new citation-rank indicator P100 for bibliometric evaluations. Journal of Informetrics, 8(3):738–748, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714000601.

Schreiber:2014:HIO

[Sch14b] Michael Schreiber. How to improve the outcome of performance evaluations in terms of percentiles for citation frequencies of my papers. Journal of Informetrics, 8(4):873– 879, October 2014. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157714000832.

Schreiber:2014:NCR

[Sch14c] Michael Schreiber. Is the new citation-rank approach P100’ in bibliometrics really new? Journal of Informetrics, 8(4):997–1004, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714000935.

Schreiber:2015:RIP

[Sch15] Michael Schreiber. Restricting the h-index to a publication and citation time window: a case study of a timed Hirsch index. Journal of Informetrics, 9(1):150–155, January 2015.

CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714001175.

Schneider:2016:ISI

[Sch16] Jesper W. Schneider. The imaginarium of statistical inference when data are the population: Comments to Williams and Bornmann. Journal of Informetrics, 10(4):1243–1248, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716302607. See [WB16a,

WB16b].

Shen:2022:REI

[SCJX22] Hongquan Shen, Ying Cheng, Xiufang Ju, and Juan Xie. Rethinking the effect of inter-gender collaboration on research performance for scholars. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722001055.

Singh:2016:CTT

[SCMG16] Mayank Singh, Tanmoy Chakraborty, Animesh Mukherjee, and Pawan Goyal. Is this conference a top-tier? ConfAssist: an assistive conflict resolution framework for conference categorization. Journal of Informetrics, 10(4):1005–1022, November 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715301255.

Sun:2023:ASR

[SCML23] Zhuanlan Sun, C. Clark Cao, Chao Ma, and Yiwei Li. The academic status of reviewers predicts their language use. Journal of Informetrics, 17(4):??, November 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000743.

Song:2023:UWS

[SCZ23] Ningyuan Song, Kejun Chen, and Yuehua Zhao. Understanding writing styles of scientific papers in the IS-LS domain: Evidence from abstracts over the past three decades.

Journal of Informetrics, 17(1):??, February 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000020.

Serenko:2011:CES

[SD11] Alexander Serenko and Michael Dohan. Comparing the expert survey and citation impact journal ranking methods: Example from the field of artificial intelligence. Journal of Informetrics, 5(4):629–648, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000666.

Shi:2022:DTI

[SD22] Xuanyu Shi and Jian Du. Distinguishing transformative from incremental clinical evidence: a classifier of clinical research using textual features from abstracts and citing sentences. Journal of Informetrics, 16(2):??, May 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000141.

Shu:2017:MSU

[SDAJ17] Fei Shu, Jesse David Dinneen, Banafsheh Asadi, and Charles-Antoine Julien. Mapping science using Library of Congress Subject Headings. Journal of Informetrics, 11(4):1080–1094, November 2017. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157717300950.

Sun:2016:MES

[SDL16] Xiaoling Sun, Kun Ding, and Yuan Lin. Mapping the evolution of scientific fields based on cross-field authors. Journal of Informetrics, 10(3):750–761, August 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715302352.

Serenko:2010:DAJ

[Ser10] Alexander Serenko. The development of an AI journal ranking based on the revealed preference approach. Journal of Informetrics, 4(4):447–459, October 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000349.

Subelj:2019:CSC

[SFCK19] Lovro Subelj, Dalibor Fiala, Tadej Ciglaric, and Luka Kronegger. Convexity in scientific collaboration networks. Journal of Informetrics, 13(1):10–31, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718301950.

Song:2023:IIS

[SFF23] Yijia Song, Zhichao Fang, and Hui-Zhen Fu. Influence of interdisciplinarity on scientific impact: the case of climate change field. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723000354.

Schubert:2007:SAH

[SG07] András Schubert and Wolfgang Glänzel. A systematic analysis of Hirsch-type indices for journals. Journal of Informetrics, 1(3):179–184, July 2007. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157707000284.

Squazzoni:2012:SMS

[SG12] Flaminio Squazzoni and Claudio Gandelli. Saint Matthew strikes again: an agent-based model of peer review and the scientific community structure. Journal of Informetrics, 6(2): 265–275, April 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711001179.

Sarabia:2010:GMG

[SGDSP10] José María Sarabia, Emilio Gómez-Déniz, María Sarabia, and Faustino Prieto. A general method for generating parametric Lorenz and Leimkuhler curves. Journal of Informetrics, 4(4):524–539, October 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000568.

Sanchez-Gil:2018:RDT

[SGGMF18] Susana Sánchez-Gil, Juan Gorraiz, and David Melero-Fuentes. Reference density trends in the major disciplines. Journal of Informetrics, 12(1):42–58, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717301487.

Sidiropoulos:2016:GSS

[SGKM16] A. Sidiropoulos, A. Gogoglou, D. Katsaros, and Y. Manolopoulos. Gazing at the skyline for star scientists. Journal of Informetrics, 10(3):789–813, August 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715200399.

Song:2015:SSP

[SHD15] Min Song, Go Eun Heo, and Ying Ding. SemPathFinder: Semantic path analysis for discovering publicly unknown knowledge. Journal of Informetrics, 9(4):686–703, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715200533.

Sun:2024:HRM

[SHL24] Zhuanlan Sun, Dongjin He, and Yiwei Li. How the readability of manuscript before journal submission advantages peer review process: Evidence from biomedical scientific publications. Journal of Informetrics, 18(3):??, August 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000609.

Song:2023:MDP

[SHZ23] Haoyang Song, Jianhua Hou, and Yang Zhang. The measurements and determinants of patent technological value: Lifetime, strength, breadth, and dispersion from the technology diffusion perspective. Journal of Informetrics, 17(1): ??, February 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722001237.

Silagadze:2023:AMS

[Sil23] Z. K. Silagadze. On arXiv moderation system. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000585.

Sirtes:2012:FEE

[Sir12] Daniel Sirtes. Finding the Easter eggs hidden by oneself: Why Radicchi and Castellano’s (2012) fairness test for citation indicators is not fair. Journal of Informetrics, 6(3):448– 450, July 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000090. See [RC12a]

and reply [RC12b].

Sivertsen:2016:WMP

[Siv16] Gunnar Sivertsen. A welcome to methodological pragmatism. Journal of Informetrics, 10(2):664–666, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716300943.

Sanchez-Jimenez:2017:RGS

[SJGBMA17] Rodrigo Sánchez-Jiménez, Vicente P. Guerrero-Bote, and Félix Moya-Anegón. The role of guarantor in scientific collaboration: the neighbourhood matters. Journal of Informetrics, 11(1):103–116, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302504.

Sanchez-Jimenez:2024:ADA

[SJGCGB^{+}24] Rodrigo Sánchez-Jiménez, Pablo Guerrero-Castillo, Vicente P. Guerrero-Bote, Gali Halevi, and Félix De-Moya-Anegón. Analysis of the distribution of authorship by gender in scientific output: a global perspective. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000695.

Shu:2019:CJP

[SJZ^{+}19] Fei Shu, Charles-Antoine Julien, Lin Zhang, Junping Qiu, Jing Zhang, and Vincent Larivière. Comparing journal and paper level classifications of science. Journal of Informetrics, 13(1):202–225, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718303298.

Stare:2014:SAI

[SK14] Janez Stare and Natasa Kejzar. On standardization of the Activity Index. Journal of Informetrics, 8(3):503–507, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714000431.

Saarela:2021:CWA

[SK21] Mirka Saarela and Tommi Kärkkäinen. Can we automate expert-based journal rankings? Analysis of the Finnish publication indicator. Journal of Informetrics, 14(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719302305.

Song:2015:EAN

[SKK15] Min Song, Erin Hea-Jin Kim, and Ha Jin Kim. Exploring author name disambiguation on PubMed-scale. Journal of Informetrics, 9(4):924–941, October 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715200259.

Saarela:2016:EBV

[SKLR16] Mirka Saarela, Tommi Kärkkäinen, Tommi Lahtonen, and Tuomo Rossi. Expert-based versus citation-based ranking of scholarly and scientific publication channels. Journal of Informetrics, 10(3):693–718, August 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715302194.

Smolinsky:2022:IIA

[SKM22] Lawrence Smolinsky, Bernhard Klingenberg, and Brian D. Marx. Interpretation and inference for altmetric indicators arising from sparse data statistics. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157722000025.

Skupin:2009:DCC

[Sku09] André Skupin. Discrete and continuous conceptualizations of science: Implications for knowledge domain visualization. Journal of Informetrics, 3(3):233–245, July 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000261.

Shi:2019:DSC

[SKY^{+}19] Gui-Yuan Shi, Yi-Xiu Kong, Guang-Hui Yuan, Rui-Jie Wu, An Zeng, and Matús Medo. Discoverers in scientific citation data. Journal of Informetrics, 13(2):717–725, May 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718303146.

Su:2012:FSG

[SL12] Hsin-Ning Su and Pei-Chun Lee. Framing the structure of global open innovation research. Journal of Informetrics, 6(2):202–216, April 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157711001076.

Shu:2015:CLA

[SL15] Fei Shu and Vincent Larivière. Chinese-language articles are biased in citations. Journal of Informetrics, 9(3):526–528, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715300341. See rejoinder

[LQLJ15].

Solomon:2013:LCC

[SLB13] David J. Solomon, Mikael Laakso, and Bo-Christer Björk. A longitudinal comparison of citation rates and growth among open access journals. Journal of Informetrics, 7(3):642–650, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115771300028X.

Seol:2023:TFS

[SLK^{+}23] Youngjin Seol, Seunghyun Lee, Cheolhan Kim, Janghyeok Yoon, and Jaewoong Choi. Towards firm-specific technology opportunities: a rule-based machine learning approach to technology portfolio analysis. Journal of Informetrics, 17(4): ??, November 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000895.

Seeber:2012:FAW

[SLL^{+}12] Marco Seeber, Benedetto Lepori, Alessandro Lomi, Isidro Aguillo, and Vitaliano Barberio. Factors affecting web links between European higher education institutions. Journal of Informetrics, 6(3):435–447, July 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000235.

Sun:2023:EED

[SLLM23] Zhuanlan Sun, Sheng Liu, Yiwei Li, and Chao Ma. Expedited editorial decision in COVID-19 pandemic. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772300007X.

Sugimoto:2013:JAR

[SLNC13] Cassidy R. Sugimoto, Vincent Larivière, Chaoqun Ni, and Blaise Cronin. Journal acceptance rates: a cross-disciplinary analysis of variability and relationships with journal measures. Journal of Informetrics, 7(4):897–906, October 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000710.

Stegehuis:2015:PLT

[SLW15] Clara Stegehuis, Nelly Litvak, and Ludo Waltman. Predicting the long-term citation impact of recent publications. Journal of Informetrics, 9(3):642–657, July 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715300080.

Smolinsky:2018:ORR

[SM18] Lawrence Smolinsky and Brian D. Marx. Odds ratios, risk ratios, and Bornmann and Haunschild’s new indicators. Journal of Informetrics, 12(3):732–735, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302414. See response [BHM18].

Small:2009:CTC

[Sma09] Henry Small. Critical thresholds for co-citation clusters and emergence of the giant component. Journal of Informetrics, 3(4):332–340, October 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000376.

Small:2017:TEG

[Sma17] Henry Small. A tribute to Eugene Garfield: Information innovator and idealist. Journal of Informetrics, 11(3):599– 612, August 2017. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157717301463.

Small:2018:CHC

[Sma18] Henry Small. Characterizing highly cited method and nonmethod papers using citation contexts: the role of uncertainty. Journal of Informetrics, 12(2):461–480, May 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157718300440.

Smolinsky:2016:ENC

[Smo16] Lawrence Smolinsky. Expected number of citations and the crown indicator. Journal of Informetrics, 10(1):43–47, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715301073.

Smolinsky:2019:ORM

[Smo19] Lawrence Smolinsky. Odds ratios and Mantel–Haenszel quotients. Journal of Informetrics, 13(1):462–463, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719300306.

Smolinsky:2021:AOP

[Smo21a] Lawrence Smolinsky. Arbitrage opportunities in publication and ghost authors. Journal of Informetrics, 14(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719304328.

Smolinsky:2021:TTP

[Smo21b] Lawrence Smolinsky. Two theorems on “Polar coordinates and generalized h-type indices”. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157720301826. See [ER21].

Schreiber:2012:EFA

[SMP12] M. Schreiber, C. C. Malesios, and S. Psarakis. Exploratory factor analysis for the Hirsch index, 17 h-type variants, and some traditional bibliometric indicators. Journal of Informetrics, 6(3):347–358, July 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000107.

Soler:2007:RIS

[Sol07] José M. Soler. A rational indicator of scientific creativity. Journal of Informetrics, 1(2):123–130, April 2007. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157706000253.

Spaan:2010:DPI

[Spa10] Jos A. E. Spaan. The danger of pseudoscience in informetrics. Journal of Informetrics, 4(3):439–440, July 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000325.

Sun:2024:FSB

[SPL24] Zhuanlan Sun, Ka Lok Pang, and Yiwei Li. The fading of status bias during the open peer review process. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000415.

Sarabia:2012:MPD

[SPT12] José María Sarabia, Faustino Prieto, and Carmen Trueba. Modeling the probabilistic distribution of the impact factor. Journal of Informetrics, 6(1):66–79, January 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000824.

Su:2011:PNE

[SPZ^{+}11] Cheng Su, YunTao Pan, YanNing Zhen, Zheng Ma, Jun-Peng Yuan, Hong Guo, ZhengLu Yu, CaiFeng Ma, and YiShan Wu. PrestigeRank: a new evaluation method for papers and journals. Journal of Informetrics, 5(1):1–13, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000337.

Silva:2013:QIS

[SROdFC13] F. N. Silva, F. A. Rodrigues, O. N. Oliveira, Jr., and L. da F. Costa. Quantifying the interdisciplinarity of scientific journals and fields. Journal of Informetrics, 7(2):469– 477, April 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000096.

Sasson:2015:ISM

[SRP15] Elan Sasson, Gilad Ravid, and Nava Pliskin. Improving similarity measures of relatedness proximity: Toward augmented concept maps. Journal of Informetrics, 9(3):618–628, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715000541.

Sivertsen:2019:MSC

[SRZ19] Gunnar Sivertsen, Ronald Rousseau, and Lin Zhang. Measuring scientific contributions with modified fractional counting. Journal of Informetrics, 13(2):679–694, May 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771830350X.

Sicilia:2011:CIF

[SSAGB11] Miguel-Angel Sicilia, Salvador Sánchez-Alonso, and Elena García-Barriocanal. Comparing impact factors from two different citation databases: the case of computer science. Journal of Informetrics, 5(4):698–704, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000083.

Strumia:2019:BFP

[ST19] Alessandro Strumia and Riccardo Torre. Biblioranking fundamental physics. Journal of Informetrics, 13(2):515–539, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157718301512.

Stacey:2021:RPA

[Sta21] Anthony G. Stacey. Robust parameterisation of ages of references in published research. Journal of Informetrics, 14 (3):??, August 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157720300328.

Small:2017:DDI

[STP17] Henry Small, Hung Tseng, and Mike Patek. Discovering discoveries: Identifying biomedical discoveries using citation contexts. Journal of Informetrics, 11(1):46–62, February 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716302255.

Sun:2024:TFP

[Sun24] Zhuanlan Sun. Textual features of peer review predict topcited papers: an interpretable machine learning perspective. Journal of Informetrics, 18(2):??, May 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000142.

Sandstrom:2018:FEP

[SV18] Ulf Sandström and Peter Van den Besselaar. Funding, evaluation, and the performance of national research systems. Journal of Informetrics, 12(1):365–384, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717302882. See corrigendum [SV24].

Sandstrom:2024:CSF

[SV24] Ulf Sandström and Peter Van den Besselaar. Corrigendum to “Funding, evaluation, and the performance of national research systems [J. Informetrics, 12/1 (2018) 365– 384]. Journal of Informetrics, 18(3):??, August 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772400052X. See [SV18].

Smith:2021:GMT

[SVKM21] Thomas Bryan Smith, Raffaele Vacca, Till Krenz, and Christopher McCarty. Great minds think alike, or do they often differ? Research topic overlap and the formation of scientific teams. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157720306210.

Silva:2011:IRW

[SVTdFC11] F. N. Silva, M. P. Viana, B. A. N. Travençolo, and L. da F. Costa. Investigating relationships within and between category networks in Wikipedia. Journal of Informetrics, 5(3):431–438, July 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157711000344.

Seiler:2012:REB

[SW12] Christian Seiler and Klaus Wohlrabe. Ranking economists on the basis of many indicators: an alternative approach using RePEc data. Journal of Informetrics, 6(3):389–402, July 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000089.

Seiler:2013:AS

[SW13] Christian Seiler and Klaus Wohlrabe. Archetypal scientists. Journal of Informetrics, 7(2):345–356, April 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712001034.

Seiler:2014:HRJ

[SW14] Christian Seiler and Klaus Wohlrabe. How robust are journal rankings based on the impact factor? Evidence from the economic sciences. Journal of Informetrics, 8(4):904–911, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157714000820.

Sung:2015:MSB

[SWHC15] Hui-Yun Sung, Chun-Chieh Wang, Mu-Hsuan Huang, and Dar-Zen Chen. Measuring science-based science linkage and non-science-based linkage of patents through non-patent references. Journal of Informetrics, 9(3):488–498, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715000486.

Song:2021:CPD

[SX21] Min Song and Qing Xie. Characterizing the psychiatric drug responses of Reddit users from a socialomics perspective. Journal of Informetrics, 14(3):??, August 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720300766.

Shen:2022:CCI

[SXAC22] Hongquan Shen, Juan Xie, Weiyi Ao, and Ying Cheng. The continuity and citation impact of scientific collaboration with different gender composition. Journal of Informetrics, 16(1): ??, February 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115772100119X.

Smolinsky:2024:VBI

[SY24] Lawrence Smolinsky and Seungwon Yang. Validity and bias of indicators of international collaboration: a theoretical analysis with an empirical study of Ukraine–Russia–United States and China–United States. Journal of Informetrics, 18 (1):??, February 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000014.

Sato:2022:PRP

[SYK22] Ryoma Sato, Makoto Yamada, and Hisashi Kashima. Poincaré: Recommending publication venues via treatment effect estimation. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722000359.

Shen:2016:IAS

[SYP^{+}16] Zhesi Shen, Liying Yang, Jiansuo Pei, Menghui Li, Chensheng Wu, Jianzhang Bao, Tian Wei, Zengru Di, Ronald Rousseau, and Jinshan Wu. Interrelations among scientific fields and their relative influences revealed by an inputoutput analysis. Journal of Informetrics, 10(1):82–97, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715300717.

Shen:2018:LDC

[SYW18] Zhesi Shen, Liying Yang, and Jinshan Wu. Lognormal distribution of citation counts is the reason for the relation between Impact Factors and Citation Success Index. Journal of Informetrics, 12(1):153–157, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717303334.

Strotmann:2010:CCC

[SZ10] Andreas Strotmann and Dangzhi Zhao. Combining commercial citation indexes and open-access bibliographic databases to delimit highly interdisciplinary research fields for citation analysis. Journal of Informetrics, 4(2):194–200, April 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157709000972.

Sziklai:2021:RIW

[Szi21] Balázs R. Sziklai. Ranking institutions within a discipline: the steep mountain of academic excellence. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157721000043.

Shen:2019:RMC

[SZR^{+}19] Si Shen, Danhao Zhu, Ronald Rousseau, Xinning Su, and Dongbo Wang. A refined method for computing bibliographic coupling strengths. Journal of Informetrics, 13(2):605–615, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716300244.

Shang:2022:IME

[SZZ22] Jing Shang, Mingbin Zeng, and Gupeng Zhang. Investigating the mentorship effect on the academic success of young scientists: an empirical study of the 985 project universities of China. Journal of Informetrics, 16(2):??, May 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000372.

Shen:2023:EJC

[SZZ23] Chaohai Shen, Star X. Zhao, and Xiaolan Zhou. The effect of journal competition on research quality with endogenous choices of open access or restricted access. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000548.

Tohalino:2022:PRG

[TA22] Jorge A. V. Tohalino and Diego R. Amancio. On predicting research grants productivity via machine learning. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000128.

Tang:2013:DBF

[Tan13] Li Tang. Does “birds of a feather flock together” matter — evidence from a longitudinal study on US–China scientific collaboration. Journal of Informetrics, 7(2):330–344, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712001009.

Tang:2024:HOD

[Tan24] Li Tang. Halt the ongoing decoupling and reboot US–China scientific collaboration. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157724000348.

Tahamtan:2018:CEP

[TB18a] Iman Tahamtan and Lutz Bornmann. Core elements in the process of citing publications: Conceptual overview of the literature. Journal of Informetrics, 12(1):203–216, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717302687.

Tahamtan:2018:CSL

[TB18b] Iman Tahamtan and Lutz Bornmann. Creativity in science and the link to cited references: Is the creative potential of papers reflected in their cited references? Journal of Informetrics, 12(3):906–930, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302219.

Thelwall:2019:GRP

[TBM^{+}19] Mike Thelwall, Carol Bailey, Meiko Makita, Pardeep Sud, and Devika P. Madalli. Gender and research publishing in India: Uniformly high inequality? Journal of Informetrics, 13(1):118–131, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302980.

Thelwall:2019:GDR

[TBTB19] Mike Thelwall, Carol Bailey, Catherine Tobin, and Noel-Ann Bradshaw. Gender differences in research areas, methods and topics: Can people and thing orientations explain the results? Journal of Informetrics, 13(1):149–169, February 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718303596.

Tonta:2010:DLS

[TD10] Yasar Tonta and Hamid R. Darvish. Diffusion of latent semantic analysis as a research tool: a social network analysis approach. Journal of Informetrics, 4(2):166–174, April 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157709000819.

Toth:2024:WCB

[TDCM24] Tamás Tóth, Márton Demeter, Sándor Csuhai, and Zsolt Balázs Major. When career-boosting is on the line: Equity and inequality in grant evaluation, productivity, and the educational backgrounds of Marie Skłodowska-Curie actions individual fellows in social sciences and humanities. Journal of Informetrics, 18(2):??, May 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000294.

Taskin:2021:SCP

[TDKZ21] Zehra Taskin, Güleda Dogan, Emanuel Kulczycki, and Alesia Ann Zuccala. Self-citation patterns of journals indexed in the Journal Citation Reports. Journal of Informetrics, 15(4): ??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000924.

Tosi:2021:SKG

[TdR21] Mauro Dalle Lucca Tosi and Julio Cesar dos Reis. SciK-Graph: a knowledge graph approach to structure a scientific field. Journal of Informetrics, 15(1):??, February 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772030626X.

Thelwall:2015:GJI

[TF15a] Mike Thelwall and Ruth Fairclough. Geometric journal impact factors correcting for individual highly cited articles. Journal of Informetrics, 9(2):263–272, April 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771500022X.

Thelwall:2015:ITD

[TF15b] Mike Thelwall and Ruth Fairclough. The influence of time and discipline on the magnitude of correlations between citation counts and quality scores. Journal of Informetrics, 9(3): 529–541, July 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771520017X.

Thelwall:2017:ACI

[TF17a] Mike Thelwall and Ruth Fairclough. The accuracy of confidence intervals for field normalised indicators. Journal of Informetrics, 11(2):530–540, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717300408.

Thelwall:2017:RPN

[TF17b] Mike Thelwall and Ruth Fairclough. The research production of nations and departments: a statistical model for the share of publications. Journal of Informetrics, 11(4):1142–1157, November 2017. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157717301669.

Thelwall:2016:DLHb

[The16a] Mike Thelwall. Are the discretised lognormal and hooked power law distributions plausible for citation data? Journal of Informetrics, 10(2):454–470, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716300074.

Thelwall:2016:TTM

[The16b] Mike Thelwall. Are there too many uncited articles? Zero inflated variants of the discretised lognormal and hooked power law distributions. Journal of Informetrics, 10(2):622–633, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716300153.

Thelwall:2016:CCD

[The16c] Mike Thelwall. Citation count distributions for large monodisciplinary journals. Journal of Informetrics, 10 (3):863–874, August 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1751157716300517.

Thelwall:2016:DLHa

[The16d] Mike Thelwall. The discretised lognormal and hooked power law distributions for complete citation data: Best options for modelling and regression. Journal of Informetrics, 10(2): 336–346, May 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715301449.

Thelwall:2016:DJR

[The16e] Mike Thelwall. Not dead, just resting: the practical value of per publication citation indicators. Journal of Informetrics, 10(2):667–670, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716300931.

Thelwall:2016:PAM

[The16f] Mike Thelwall. The precision of the arithmetic mean, geometric mean and percentiles for citation data: an experimental simulation modelling approach. Journal of Informetrics, 10(1):110–123, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157715301437.

Thelwall:2017:CIN

[The17a] Mike Thelwall. Confidence intervals for normalised citation counts: Can they delimit underlying research capability? Journal of Informetrics, 11(4):1069–1079, November 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717301505.

Thelwall:2017:MAM

[The17b] Mike Thelwall. Microsoft Academic: a multidisciplinary comparison of citation counts with Scopus and Mendeley for 29 journals. Journal of Informetrics, 11(4):1201–1212, November 2017. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157717302900.

Thelwall:2017:TPF

[The17c] Mike Thelwall. Three practical field normalised alternative indicator formulae for research evaluation. Journal of Informetrics, 11(1):128–151, February 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771630205X.

Thelwall:2018:DCS

[The18a] Mike Thelwall. Dimensions: a competitor to Scopus and the Web of Science? Journal of Informetrics, 12(2):430–435, May 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S175115771830066X.

Thelwall:2018:DFC

[The18b] Mike Thelwall. Do females create higher impact research? Scopus citations and Mendeley readers for articles from five countries. Journal of Informetrics, 12(4):1031–1041, November 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115771830258X.

Thelwall:2018:MAA

[The18c] Mike Thelwall. Microsoft Academic automatic document searches: Accuracy for journal articles and suitability for citation analysis. Journal of Informetrics, 12(1):1–9, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717303346.

Thelwall:2019:RSS

[The19a] Mike Thelwall. The rhetorical structure of science? A multidisciplinary analysis of article headings. Journal of Informetrics, 13(2):555–563, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157718304942.

Thelwall:2019:SCC

[The19b] Mike Thelwall. Should citations be counted separately from each originating section? Journal of Informetrics, 13(2): 658–678, May 2019. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157718305121.

Tian:2024:ESI

[THQL24] Dan Tian, Xiao Hu, Yuchen Qian, and Jiang Li. Exploring the scientific impact of negative results. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723001062.

Tang:2021:PAI

[TLD^{+}21] Xuli Tang, Xin Li, Ying Ding, Min Song, and Yi Bu. The pace of artificial intelligence innovations: Speed, talent, and trial-and-error. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157720301991.

Tian:2023:PES

[TLM23] Yunpei Tian, Gang Li, and Jin Mao. Predicting the evolution of scientific communities by interpretable machine learning approaches. Journal of Informetrics, 17(2):??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115772300024X.

Toth:2021:MAB

[TLV^{+}21] István Tóth, Zsolt I. Lázár, Levente Varga, Ferenc Járai-Szabó, István Papp, Răzvan V. Florian, and Mária Ercsey-Ravasz. Mitigating ageing bias in article level metrics using citation network analysis. Journal of Informetrics, 15(1): ??, February 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157720306222.

Tang:2024:DCC

[TLZM24] Kun Tang, Baiyang Li, Qiyu Zhu, and Lecun Ma. Disruptive content, cross agglomeration interaction, and agglomeration replacement: Does cohesion foster strength? Journal of Informetrics, 18(4):??, November 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000828.

Takano:2016:UCI

[TMK16] Yasutomo Takano, Cristian Mejia, and Yuya Kajikawa. Unconnected component inclusion technique for patent network analysis: Case study of Internet of Things-related technologies. Journal of Informetrics, 10(4):967–980, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716300207.

Thor:2016:ICC

[TMLB16] Andreas Thor, Werner Marx, Loet Leydesdorff, and Lutz Bornmann. Introducing CitedReferencesExplorer (CRExplorer): a program for reference publication year spectroscopy with cited references standardization. Journal of Informetrics, 10(2):503–515, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716300049.

Thelwall:2018:CSU

[TN18] Mike Thelwall and Tamara Nevill. Could scientists use Altmetric.com scores to predict longer term citation counts? Journal of Informetrics, 12(1):237–248, February 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717303905.

Tokmachev:2023:HSS

[Tok23] Andrey M. Tokmachev. Hidden scales in statistics of citation indicators. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722001092.

Tol:2008:RSI

[Tol08] Richard S. J. Tol. A rational, successive g-index applied to economics departments in Ireland. Journal of Informetrics, 2(2):149–155, April 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157708000023.

Tol:2013:IER

[Tol13a] Richard S. J. Tol. Identifying excellent researchers: a new approach. Journal of Informetrics, 7(4):803–810, October 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000503.

Tol:2013:MEC

[Tol13b] Richard S. J. Tol. The Matthew effect for cohorts of economists. Journal of Informetrics, 7(2):522–527, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000163.

Tol:2023:NBN

[Tol23] Richard S. J. Tol. Nobel begets Nobel in economics. Journal of Informetrics, 17(4):??, November 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000822.

Thelwall:2012:WRB

[TS12] Mike Thelwall and Pardeep Sud. Webometric research with the Bing Search API 2.0. Journal of Informetrics, 6(1):44– 52, January 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711000988.

Thelwall:2014:NCA

[TS14] Mike Thelwall and Pardeep Sud. No citation advantage for monograph-based collaborations? Journal of Informetrics, 8(1):276–283, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S175115771300117X.

Thelwall:2016:NDT

[TS16] Mike Thelwall and Pardeep Sud. National, disciplinary and temporal variations in the extent to which articles with more authors have more impact: Evidence from a geometric field normalised citation indicator. Journal of Informetrics, 10(1): 48–61, February 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715301619.

Tsai:2014:CIA

[Tsa14] Chih-Fong Tsai. Citation impact analysis of top ranked computer science journals and their rankings. Journal of Informetrics, 8(2):318–328, April 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000030.

Tsouchnika:2022:PCS

[TSAH22] Maria Tsouchnika, Alex Smolyak, Panos Argyrakis, and Shlomo Havlin. Patent collaborations: From segregation to globalization. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721001097.

Torres-Salinas:2009:LCA

[TSM09] Daniel Torres-Salinas and Henk F. Moed. Library Catalog Analysis as a tool in studies of social sciences and humanities: an exploratory study of published book titles in economics. Journal of Informetrics, 3(1):9–26, January 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157708000527.

Torres-Salinas:2024:FNB

[TSOMDV^{+}24] Daniel Torres-Salinas, Enrique Orduña-Malea, Ángel Delgado-Vázquez, Juan Gorraiz, and Wenceslao Arroyo-Machado. Foundations of narrative bibliometrics. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1751157724000592.

Torres-Salinas:2019:MBH

[TSRFAM19] Daniel Torres-Salinas, Esteban Romero-Frías, and Wenceslao Arroyo-Machado. Mapping the backbone of the Humanities through the eyes of Wikipedia. Journal of Informetrics, 13(3):793–803, August 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157718302955.

Torres-Salinas:2013:MCP

[TSRSRG^{+}13] Daniel Torres-Salinas, Rosa Rodríguez-Sánchez, Nicolás Robinson-García, J. Fdez-Valdivia, and J. A. García. Mapping citation patterns of book chapters in the Book Citation Index. Journal of Informetrics, 7(2):412–424, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157713000060.

Torres-Salinas:2022:TNN

[TSVBAM22] Daniel Torres-Salinas, Pilar Valderrama-Baca, and Wenceslao Arroyo-Machado. Is there a need for a new journal metric? Correlations between JCR Impact Factor metrics and the Journal Citation Indicator-JCI. Journal of Informetrics, 16(3):??, August 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000670.

Tahmooresnejad:2022:FIT

[TT22] Leila Tahmooresnejad and Ekaterina Turkina. Female inventors over time: Factors affecting female inventors’ innovation performance. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722000086.

Tu:2021:RDS

[Tu21] Jing Tu. The role of dyadic social capital in enhancing collaborative knowledge creation. Journal of Informetrics, 14 (2):??, May 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157719304122.

Tu:2024:OIC

[Tu24] Jing Tu. Openness to international collaboration and tie strength in enhancing knowledge creation. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723001074.

Tosic:2021:UBL

[TV21] Aleksandar Tošić and Jernej Vičič. Use of Benford’s law on academic publishing networks. Journal of Informetrics, 15 (3):??, August 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000341.

Thelwall:2014:DCA

[TW14a] Mike Thelwall and Paul Wilson. Distributions for cited articles from individual subjects and years. Journal of Informetrics, 8(4):824–839, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714000698.

Thelwall:2014:RCD

[TW14b] Mike Thelwall and Paul Wilson. Regression for citation data: an evaluation of different methods. Journal of Informetrics, 8(4):963–971, October 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000923.

Trappey:2023:PLK

[TWC^{+}23] Amy J. C. Trappey, Ann Y. E. Wei, Neil K. T. Chen, Kuo-An Li, L. P. Hung, and Charles V. Trappey. Patent landscape and key technology interaction roadmap using graph convolutional network — case of mobile communication technologies beyond 5G. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722001079.

Tian:2008:GSP

[TWH08] Yangge Tian, Cheng Wen, and Song Hong. Global scientific production on GIS research by bibliometric analysis from 1997 to 2006. Journal of Informetrics, 2(1):65–74, January 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157707000715.

Tong:2024:CND

[TWY24] Tong Tong, Wanru Wang, and Fred Y. Ye. A complement to the novel disruption indicator based on knowledge entities. Journal of Informetrics, 18(2):??, May 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000373.

Teng:2024:EST

[TWZ^{+}24] Hao Teng, Nan Wang, Hongyu Zhao, Yingtong Hu, and Haitao Jin. Enhancing semantic text similarity with functional semantic knowledge (FOP) in patents. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000925.

Tang:2024:MIC

[TYWG24] Li Tang, Defang Yang, Mingxing Wang, and Ying Guo. The mediating impact of citation scope: Evidence from China’s ESI publications. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724000543.

Urdiales:2024:AAB

[UG24] Cristina Urdiales and Eduardo Guzmán. An automatic and association-based procedure for hierarchical publication subject categorization. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000913.

Uddin:2016:IAS

[UK16] Shahadat Uddin and Arif Khan. The impact of authorselected keywords on citation counts. Journal of Informetrics, 10(4):1166–1177, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716301146.

Uddin:2023:ICJ

[UKL23] Shahadat Uddin, Arif Khan, and Haohui Lu. Impact of COVID-19 on Journal Impact Factor. Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157723000834.

Viana:2013:TVC

[VAdFC13] Matheus P. Viana, Diego R. Amancio, and Luciano da F. Costa. On time-varying collaboration networks. Journal of Informetrics, 7(2):371–378, April 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712001162.

Vanclay:2008:RFJ

[Van08] Jerome K. Vanclay. Ranking forestry journals using the hindex. Journal of Informetrics, 2(4):326–334, October 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157708000357.

Vanclay:2011:EAR

[Van11] Jerome K. Vanclay. An evaluation of the Australian Research Council’s journal ranking. Journal of Informetrics, 5(2):265– 274, April 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000994.

Vanclay:2012:PPA

[Van12a] Jerome K. Vanclay. Publication patterns of award-winning forest scientists and implications for the Australian ERA journal ranking. Journal of Informetrics, 6(1):19–26, January 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157711000757.

Vanclay:2012:WWW

[Van12b] Jerome K. Vanclay. What was wrong with Australia’s journal ranking? Journal of Informetrics, 6(1):53–54, January 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157711000848.

Vanclay:2013:FAC

[Van13] Jerome K. Vanclay. Factors affecting citation rates in environmental science. Journal of Informetrics, 7(2):265–271, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712000995.

Vasileiadou:2009:SOU

[Vas09] Eleftheria Vasileiadou. Stabilisation operationalised: Using time series analysis to understand the dynamics of research collaboration. Journal of Informetrics, 3(1):36–48, January 2009. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157708000679.

Vieira:2014:HGM

[VCG14] Elizabeth S. Vieira, José A. S. Cabral, and José A. N. F. Gomes. How good is a model based on bibliometric indicators in predicting the final decisions made by peers?

Journal of Informetrics, 8(2):390–405, April 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771400025X.

Vieira:2022:WDD

[VCT22] Elizabeth S. Vieira, Jorge Cerdeira, and Aurora A. C. Teixeira. Which distance dimensions matter in international research collaboration? A cross-country analysis by scientific domain. Journal of Informetrics, 16(2):??, May 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000116.

vandenBesselaar:2012:SCM

[vdB12] Peter van den Besselaar. Selection committee membership: Service or self-service. Journal of Informetrics, 6(4):580– 585, October 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712000430.

vandenBesselaar:2014:C

[vdBBL14] Peter van den Besselaar, Lutz Bornmann, and Loet Leydesdorff. Correction. Journal of Informetrics, 8(4):801, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157714000686.

vandenBesselaar:2017:DOR

[vdBHS17a] Peter van den Besselaar, Ulf Heyman, and Ulf Sandström. Do observations have any role in science policy studies? A reply. Journal of Informetrics, 11(3):941–944, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157717302006.

vandenBesselaar:2017:PEO

[vdBHS17b] Peter van den Besselaar, Ulf Heyman, and Ulf Sandström. Perverse effects of output-based research funding? Butler’s Australian case revisited. Journal of Informetrics, 11(3):905– 918, August 2017. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157717301943.

vandenBesselaar:2022:EWS

[vdBM22] Peter van den Besselaar and Charlie Mom. The effect of writing style on success in grant applications. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000098.

vandenBesselaar:2015:ECG

[vdBS15] Peter van den Besselaar and Ulf Sandström. Early career grants, performance, and careers: a study on predictive validity of grant decisions. Journal of Informetrics, 9(4):826– 838, October 2015. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715300067.

vandenBesselaar:2018:QMH

[vdBS18] Peter van den Besselaar and Ulf Sandström. Quantity matters, but how does it work?: a comment on Lindahl. Journal of Informetrics, 12(4):1059–1062, November 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302967. See [Lin18b, Lin18c].

Verleysen:2014:BRB

[VE14] Frederik T. Verleysen and Tim C. E. Engels. Barycenter representation of book publishing internationalization in the Social Sciences and Humanities. Journal of Informetrics, 8(1):234–240, January 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157713001090.

vanEck:2008:GI

[vEW08] Nees Jan van Eck and Ludo Waltman. Generalizing the hand g-indices. Journal of Informetrics, 2(4):263–271, October 2008. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157708000503.

vanEck:2009:PUL

[vEW09] Nees Jan van Eck and Ludo Waltman. On the proper understanding of the limiting behavior of generalizations of the h- and g-indices. Journal of Informetrics, 3(4):369–370, October 2009. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157709000133.

vanEck:2014:CNS

[vEW14] Nees Jan van Eck and Ludo Waltman. CitNetExplorer: a new software tool for analyzing and visualizing citation networks. Journal of Informetrics, 8(4):802–823, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714000662.

Vieira:2010:CSA

[VG10a] E. S. Vieira and J. A. N. F. Gomes. Citations to scientific articles: Its distribution and dependence on the article features. Journal of Informetrics, 4(1):1–13, January 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157709000534.

Vieira:2010:RII

[VG10b] E. S. Vieira and J. A. N. F. Gomes. A research impact indicator for institutions. Journal of Informetrics, 4(4):581– 590, October 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771000060X.

Volkovich:2016:MVM

[VGRB16] Zeev Volkovich, Oleg Granichin, Oleg Redkin, and Olga Bernikova. Modeling and visualization of media in Arabic. Journal of Informetrics, 10(2):439–453, May 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715302297.

Viiu:2016:TEH

[Vîi16] Gabriel-Alexandru Vîiu. A theoretical evaluation of Hirschtype bibliometric indicators confronted with extreme selfcitation. Journal of Informetrics, 10(2):552–566, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716300542.

Viiu:2017:DRE

[Vîi17] Gabriel-Alexandru Vîiu. Disaggregated research evaluation through median-based characteristic scores and scales: a comparison with the mean-based approach. Journal of Informetrics, 11(3):748–765, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717300901.

Viiu:2018:LDE

[Vîi18] Gabriel-Alexandru Vîiu. The lognormal distribution explains the remarkable pattern documented by characteristic scores and scales in scientometrics. Journal of Informetrics, 12(2): 401–415, May 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1751157717303887.

Vinkler:2012:CSA

[Vin12] Péter Vinkler. The case of scientometricians with the “absolute relative” impact indicator. Journal of Informetrics, 6(2): 254–264, April 2012. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711001167.

Vinkler:2013:CRA

[Vin13a] Péter Vinkler. Comparative rank assessment of journal articles. Journal of Informetrics, 7(3):712–717, July 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000424.

Vinkler:2013:WIP

[Vin13b] Péter Vinkler. Would it be possible to increase the Hirsch-index, π-index or CDS-index by increasing the number of publications or citations only by unity? Journal of Informetrics, 7(1):72–83, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000612.

Vinkler:2014:UPR

[Vin14] Péter Vinkler. The use of the percentage rank position index for comparative evaluation of journals. Journal of Informetrics, 8(2):340–348, April 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000029.

Vieira:2016:GPH

[VL16] Elizabeth S. Vieira and Benedetto Lepori. The growth process of higher education institutions and public policies. Journal of Informetrics, 10(1):286–298, February 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715301802.

vanLeeuwen:2018:BRM

[vL18] Thed van Leeuwen. Book review: Measuring Research: What Everyone Needs to Know, Cassidy R. Sugimoto, Vincent Larivière, Oxford University Press (2018). Journal of Informetrics, 12(4):1232–1234, November 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718303420.

Vaccario:2017:QSR

[VMWM17] Giacomo Vaccario, Matús Medo, Nicolas Wider, and Manuel Sebastian Mariani. Quantifying and suppressing ranking bias in a large citation network. Journal of Informetrics, 11(3):766–782, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717300974.

Vancraeynest:2024:NAC

[VPAE24] Bram Vancraeynest, Hoang-Son Pham, and Amr Ali-Eldin. A new approach to computing the distances between research disciplines based on researcher collaborations and similarity measurement techniques. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157724000403.

Vinayak:2023:SCD

[VRK23] Vinayak, Adarsh Raghuvanshi, and Avinash Kshitij. Signatures of capacity development through research collaborations in artificial intelligence and machine learning. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722001110.

vanRaan:2010:RCR

[vRvLV^{+}10] Anthony F. J. van Raan, Thed N. van Leeuwen, Martijn S. Visser, Nees Jan van Eck, and Ludo Waltman. Rivals for the crown: Reply to Opthof and Leydesdorff. Journal of Informetrics, 4(3):431–435, July 2010. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000301. See [OL10].

Vanderelst:2013:SRF

[VS13] Dieter Vanderelst and Niko Speybroeck. Scientometrics reveals funding priorities in medical research policy. Journal of Informetrics, 7(1):240–247, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000867.

Vardakas:2015:AFC

[VTPF15] Konstantinos Z. Vardakas, Grigorios Tsopanakis, Alexandra Poulopoulou, and Matthew E. Falagas. An analysis of factors contributing to PubMed’s growth. Journal of Informetrics, 9(3):592–617, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771500053X.

vanVlokhoven:2019:EOA

[vV19] Has van Vlokhoven. The effect of open access on research quality. Journal of Informetrics, 13(2):751–756, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718304814.

Verleysen:2016:CPP

[VW16] Frederik T. Verleysen and Arie Weeren. Clustering by publication patterns of senior authors in the social sciences and humanities. Journal of Informetrics, 10(1):254–272, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715301747.

Vaughan:2010:WCO

[VY10] Liwen Vaughan and Justin You. Word co-occurrences on Webpages as a measure of the relatedness of organizations: a new Webometrics concept. Journal of Informetrics, 4(4):483–491, October 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157710000386.

Vaughan:2013:WTO

[VY13] Liwen Vaughan and Rongbin Yang. Web traffic and organization performance measures: Relationships and data sources examined. Journal of Informetrics, 7(3):699–711, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157713000412.

vanZyl:2013:SSI

[vZ13] J. M. van Zyl. A simulation study to investigate the accuracy of approximating averages of ratios using ratios of averages. Journal of Informetrics, 7(4):907–913, October 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000722.

Valderrama-Zurian:2015:SAD

[VZAMMFAB15] Juan-Carlos Valderrama-Zurián, Remedios Aguilar-Moya, David Melero-Fuentes, and Rafael Aleixandre-Benavent. A systematic analysis of duplicate records in Scopus. Journal of Informetrics, 9(3):570–576, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000504.

Valderrama-Zurian:2016:TTT

[VZMFAB16] Juan-Carlos Valderrama-Zurián, David Melero-Fuentes, and Rafael Aleixandre-Benavent. Towards transparency trends in academic databases? Journal of Informetrics, 10(4):1023– 1024, November 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157716300827.

Valderrama-Zurian:2019:OCP

[VZMFAB19] J. C. Valderrama-Zurian, D. Melero-Fuentes, and R. Aleixandre-Benavent. Origin, characteristics, predominance and conceptual networks of eponyms in the bibliometric literature. Journal of Informetrics, 13(1):434–448, February 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302906.

Waltman:2012:EAU

[Wal12] Ludo Waltman. An empirical analysis of the use of alphabetical authorship in scientific publishing. Journal of Informetrics, 6(4):700–711, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000594.

Walters:2014:DAI

[Wal14] William H. Walters. Do Article Influence scores overestimate the citation impact of social science journals in subfields that are related to higher-impact natural science disciplines? Journal of Informetrics, 8(2):421–430, April 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714000273.

Waltman:2015:MNE

[Wal15] Ludo Waltman. Message from the new Editor-in-Chief. Journal of Informetrics, 9(1):A3–A6, January 2015. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715000097.

Waltman:2016:CDU

[Wal16a] Ludo Waltman. Conceptual difficulties in the use of statistical inference in citation analysis. Journal of Informetrics, 10(4):1249–1252, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302619.

Waltman:2016:RLC

[Wal16b] Ludo Waltman. A review of the literature on citation impact indicators. Journal of Informetrics, 10(2):365–391, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715300900.

Waltman:2016:SSSa

[Wal16c] Ludo Waltman. Special section on size-independent indicators in citation analysis. Journal of Informetrics, 10(2):645, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716300864.

Waltman:2016:SSSb

[Wal16d] Ludo Waltman. Special section on statistical inference in citation analysis. Journal of Informetrics, 10(4):1224, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716302577.

Walters:2017:DSJ

[Wal17a] William H. Walters. Do subjective journal ratings represent whole journals or typical articles? Unweighted or weighted citation impact? Journal of Informetrics, 11(3):730–744, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157717300925.

Waltman:2017:SSP

[Wal17b] Ludo Waltman. Special section on performance-based research funding systems. Journal of Informetrics, 11(3):904, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157717301931.

Wang:2014:UME

[Wan14] Jian Wang. Unpacking the Matthew effect in citations. Journal of Informetrics, 8(2):329–339, April 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000078.

Wang:2024:MNF

[Wan24] Xing Wang. The misuse of the nonlinear field normalization method: Nonlinear field normalization citation counts at the paper level should not be added or averaged. Journal of Informetrics, 18(3):??, August 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000440.

Williams:2016:SIBa

[WB16a] Richard Williams and Lutz Bornmann. Sampling issues in bibliometric analysis. Journal of Informetrics, 10(4):1225– 1232, November 2016. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157715000656. See com-

ments [Nan16, Mut16, Sch16] and response [WB16b].

Williams:2016:SIBb

[WB16b] Richard Williams and Lutz Bornmann. Sampling issues in bibliometric analysis: Response to discussants. Journal of Informetrics, 10(4):1253–1257, November 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716302620.

Wei:2012:EDM

[WBL12] Jiuchang Wei, Bing Bu, and Liang Liang. Estimating the diffusion models of crisis information in micro blog. Journal of Informetrics, 6(4):600–610, October 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771200048X.

Wang:2021:PAS

[WCG21] Zhiqi Wang, Yue Chen, and Wolfgang Glänzel. Preprints as accelerator of scholarly communication: an empirical analysis in mathematics. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157720300729.

Wang:2023:CPB

[WDLX23] Feifei Wang, Jiaxin Dong, Wanzhao Lu, and Shuo Xu. Collaboration prediction based on multilayer all-author tripartite citation networks: a case study of gene editing. Journal of Informetrics, 17(1):??, February 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722001274.

Weber:2017:CSD

[Web17] Matthias Weber. Comment on “Does your surname affect the citability of your publications?”. Journal of Informetrics, 11(3):835–837, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S1751157717302092. See [AD17a] and response [AD17c].

Wu:2014:GPA

[WFC14] Yan Wu, Tom Z. J. Fu, and Dah Ming Chiu. Generalized preferential attachment considering aging. Journal of Informetrics, 8(3):650–658, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000558.

Wang:2024:ICL

[WFNH24] Jue Wang, Rainer Frietsch, Peter Neuhäusler, and Rosalie Hooi. International collaboration leading to high citations: Global impact or home country effect? Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000786.

Wang:2010:RPA

[WG10a] Gangbo Wang and Jiancheng Guan. The role of patenting activity for scientific research: a study of academic inventors from China’s nanotechnology. Journal of Informetrics, 4(3): 338–350, July 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771000009X.

Wong:2010:GBP

[WG10b] Chan-Yuan Wong and Kim-Leng Goh. Growth behavior of publications and patents: a comparative study on selected Asian economies. Journal of Informetrics, 4(4):460–474, October 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000350.

Wong:2012:SFD

[WG12] Chan-Yuan Wong and Kim-Leng Goh. The sustainability of functionality development of science and technology: Papers and patents of emerging economies. Journal of Informetrics, 6(1):55–65, January 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000721.

Wang:2015:STS

[WH15] Jian Wang and Diana Hicks. Scientific teams: Selfassembly, fluidness, and interdependence. Journal of Informetrics, 9(1):197–207, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714001187.

Huang:2016:PCB

[wH16] Ding wei Huang. Positive correlation between quality and quantity in academic journals. Journal of Informetrics, 10(2):329–335, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157715302054.

Wang:2021:TTS

[WHH21] Yajie Wang, Haiyan Hou, and Zhigang Hu. ‘To tweet or not to tweet?’ A study of the use of Twitter by scholarly book publishers in Social Sciences and Humanities. Journal of Informetrics, 15(3):??, August 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S1751157721000419. See corrigendum [WZHH23].

Wang:2022:MNM

[WHHW22] Xiaoguang Wang, Jing He, Han Huang, and Hongyu Wang. MatrixSim: a new method for detecting the evolution paths of research topics. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722000955.

Wang:2017:GDA

[WHL17] Yuandi Wang, Ruifeng Hu, and Meijun Liu. The geotemporal demographics of academic journals from 1950 to 2013 according to Ulrich’s database. Journal of Informetrics, 11(3):655– 671, August 2017. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157716302358.

Wang:2018:LLS

[WHLT18] Guoyan Wang, Guangyuan Hu, Chuanfeng Li, and Li Tang. Long live the scientists: Tracking the scientific fame of great minds in physics. Journal of Informetrics, 12(4):1089–1098, November 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157718300464.

Wang:2014:ICA

[WHW^{+}14] Xuefeng Wang, Meng Huang, Hongyuan Wang, Ming Lei, Donghua Zhu, Jie Ren, and Munazza Jabeen. International Collaboration Activity Index: Case study of dye-sensitized solar cells. Journal of Informetrics, 8(4):854–862, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714000807.

Wildgaard:2016:CCA

[Wil16] Lorna Wildgaard. A critical cluster analysis of 44 indicators of author-level performance. Journal of Informetrics, 10(4):1055–1078, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157716301171.

Wu:2023:IFS

[WJ23] Yingwen Wu and Yangjian Ji. Identifying firm-specific technology opportunities from the perspective of competitors by using association rule mining. Journal of Informetrics, 17 (2):??, May 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000238.

Wang:2019:EAA

[WJW^{+}19] Feifei Wang, Chenran Jia, Xiaohan Wang, Junwan Liu, Shuo Xu, Yang Liu, and Chenyuyan Yang. Exploring all-author tripartite citation networks: a case study of gene editing. Journal of Informetrics, 13(3):856–873, August 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771830244X.

Wu:2023:AAI

[WKCC23] Renli Wu, Donghyun Kang, Yi Chen, and Chuanfu Chen. Assessing academic impacts of machine learning applications on a social science: Bibliometric evidence from economics. Journal of Informetrics, 17(3):??, August 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000615.

Wu:2022:MMM

[WKY^{+}22] Lingfei Wu, Aniket Kittur, Hyejin Youn, Staša Milojević, Erin Leahey, Stephen M. Fiore, and Yong-Yeol Ahn. Metrics and mechanisms: Measuring the unmeasurable in the science of science. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722000426.

Wang:2023:NLN

[WLCL23] Chun-Chieh Wang, Jia-Tian Lin, Dar-Zen Chen, and Szu-Chia Lo. A new look at national diversity of inventor teams within organizations. Journal of Informetrics, 17(1): ??, February 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722001225.

Wallace:2009:MCC

[WLG09] Matthew L. Wallace, Vincent Larivière, and Yves Gingras. Modeling a century of citation distributions. Journal of Informetrics, 3(4):296–303, October 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000327.

Wu:2011:ESG

[WLH11] Jiang Wu, Sergi Lozano, and Dirk Helbing. Empirical study of the growth dynamics in real career h-index sequences. Journal of Informetrics, 5(4):489–497, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000289.

Wang:2019:EGR

[WLH19] Kai Wang, Xiaojuan Liu, and Yutong Han. Exploring Goodreads reviews for book impact assessment. Journal of Informetrics, 13(3):874–886, August 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718305078.

Wu:2018:JEI

[WLLL18] Dengsheng Wu, Jing Li, Xiaoli Lu, and Jianping Li. Journal editorship index for assessing the scholarly impact of academic institutions: an empirical analysis in the field of economics. Journal of Informetrics, 12(2):448–460, May 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717301530.

Wang:2024:DDP

[WLY^{+}24] Xiaoli Wang, Wenting Liang, Xuanting Ye, Lingdi Chen, and Yun Liu. Disruptive development path measurement for emerging technologies based on the patent citation network. Journal of Informetrics, 18(1):??, February 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000063.

Wang:2024:SCM

[WLZJ24] Zuzheng Wang, Yongxu Lu, Yuanyuan Zhou, and Jiaojiao Ji. SMIAltmetric: a comprehensive metric for evaluating social media impact of scientific papers on Twitter (X). Journal of Informetrics, 18(3):??, August 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000750.

Wang:2021:UIK

[WML^{+}21] Shiyun Wang, Jin Mao, Kun Lu, Yujie Cao, and Gang Li. Understanding interdisciplinary knowledge integration through citance analysis: a case study on eHealth. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000857.

Weiss:2023:TRP

[WNH^{+}23] Matthias Weiss, Lakshmi B. Nair, Bareerah H. Hoorani, Michael Gibbert, and Martin Hoegl. Transparency of reporting practices in quantitative field studies: the transparency sweet spot for article citations. Journal of Informetrics, 17 (2):??, May 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000214.

Woeginger:2008:AAE

[Woe08a] Gerhard J. Woeginger. An axiomatic analysis of Egghe’s gindex. Journal of Informetrics, 2(4):364–368, October 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157708000266.

Woeginger:2008:SAS

[Woe08b] Gerhard J. Woeginger. A symmetry axiom for scientific impact indices. Journal of Informetrics, 2(4):298–303, October 2008. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157708000461.

Woeginger:2014:ISB

[Woe14] Gerhard J. Woeginger. Investigations on the step-based research indices of Chambers and Miller. Journal of Informetrics, 8(3):659–666, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714000546.

Wu:2022:HDA

[WOLD22] Jiang Wu, Guiyan Ou, Xiaohui Liu, and Ke Dong. How does academic education background affect top researchers’ performance? Evidence from the field of artificial intelligence. Journal of Informetrics, 16(2):??, May 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772200044X.

Wouters:2018:FP

[Wou18] Paul Wouters. The failure of a paradigm. Journal of Informetrics, 12(2):534–540, May 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S1751157718300932. See reply [Moe18].

Wang:2013:ESW

[WPZ^{+}13] Xianwen Wang, Lian Peng, Chunbo Zhang, Shenmeng Xu, Zhi Wang, Chuanli Wang, and Xianbing Wang. Exploring scientists’ working timetable: a global survey. Journal of Informetrics, 7(3):665–675, July 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000394.

Wang:2021:ESR

[WRA^{+}21] Wei Wang, Jing Ren, Mubarak Alrashoud, Feng Xia, Mengyi Mao, and Amr Tolba. Early-stage reciprocity in sustainable scientific collaboration. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719302779.

Wagner:2011:AUM

[WRB^{+}11] Caroline S. Wagner, J. David Roessner, Kamau Bobb, Julie Thompson Klein, Kevin W. Boyack, Joann Keyton, Ismael Rafols, and Katy Börner. Approaches to understanding and measuring interdisciplinary scientific research (IDR): a review of the literature. Journal of Informetrics, 5(1):14– 26, January 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157710000581.

Wang:2023:ZLB

[WRGL23] Zhenhua Wang, Ming Ren, Dong Gao, and Zhuang Li. A Zipf’s law-based text generation approach for addressing imbalance in entity extraction. Journal of Informetrics, 17(4): ??, November 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000780.

Wang:2021:PPE

[WS21] Peiling Wang and Jing Su. Post-publication expert recommendations in faculty opinions (F1000Prime): Recommended articles and citations. Journal of Informetrics, 15 (3):??, August 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000456.

Wei:2022:RAI

[WTR^{+}22] Shelia X. Wei, Tong Tong, Ronald Rousseau, Wanru Wang, and Fred Y. Ye. Relations among the h-, g-, ψ-, and pindex and offset-ability. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115772200092X.

Waltman:2011:GSK

[WTvE11] Ludo Waltman, Robert J. W. Tijssen, and Nees Jan van Eck. Globalisation of science in kilometres. Journal of Informetrics, 5(4):574–582, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000599.

Wu:2013:IUD

[Wu13] Jiang Wu. Investigating the universal distributions of normalized indicators and developing field-independent index. Journal of Informetrics, 7(1):63–71, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000752.

Wu:2015:DSF

[Wu15] Jiang Wu. Distributions of scientific funding across universities and research disciplines. Journal of Informetrics, 9(1):183–196, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157714001199.

Wang:2024:CPT

[WV24] Jian Wang and Suzan Verberne. Comparing patent intext and front-page references to science. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157724000774.

Waltman:2008:SCJ

[WvE08] Ludo Waltman and Nees Jan van Eck. Some comments on the journal weighted impact factor proposed by Habibzadeh and Yadollahie. Journal of Informetrics, 2(4):369–372, October 2008. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115770800031X. See [HY08].

Waltman:2009:SCE

[WvE09] Ludo Waltman and Nees Jan van Eck. Some comments on Egghe’s derivation of the impact factor distribution. Journal of Informetrics, 3(4):363–366, October 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000388.

Waltman:2013:SEC

[WvE13] Ludo Waltman and Nees Jan van Eck. A systematic empirical comparison of different approaches for normalizing citation impact indicators. Journal of Informetrics, 7(4):833– 849, October 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000667. See com-

ments [RC14].

Waltman:2015:FNC

[WvE15] Ludo Waltman and Nees Jan van Eck. Field-normalized citation impact indicators and the choice of an appropriate counting method. Journal of Informetrics, 9(4):872–894, October 2015. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715300456.

Waltman:2010:UAM

[WvEN10] Ludo Waltman, Nees Jan van Eck, and Ed C. M. Noyons. A unified approach to mapping and clustering of bibliometric networks. Journal of Informetrics, 4(4):629–635, October 2010. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157710000660.

Waltman:2011:TNC

[WvEvL^{+}11] Ludo Waltman, Nees Jan van Eck, Thed N. van Leeuwen, Martijn S. Visser, and Anthony F. J. van Raan. Towards a new crown indicator: Some theoretical considerations. Journal of Informetrics, 5(1):37–47, January 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157710000817.

Waltman:2013:SMS

[WvEvLV13] Ludo Waltman, Nees Jan van Eck, Thed N. van Leeuwen, and Martijn S. Visser. Some modifications to the SNIP journal impact indicator. Journal of Informetrics, 7(2):272–285, April 2013. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157712001010.

Waltman:2016:ERP

[WvEVW16] Ludo Waltman, Nees Jan van Eck, Martijn Visser, and Paul Wouters. The elephant in the room: the problem of quantifying productivity in evaluative scientometrics.

Journal of Informetrics, 10(2):671–674, May 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716300918.

Waltman:2013:CPC

[WvEW13] Ludo Waltman, Nees Jan van Eck, and Paul Wouters. Counting publications and citations: Is more always better? Journal of Informetrics, 7(3):635–641, July 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000357.

Wong:2015:TST

[WW15] Chan-Yuan Wong and Lili Wang. Trajectories of science and technology and their co-evolution in BRICS: Insights from publication and patent analysis. Journal of Informetrics, 9(1):90–101, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714001084.

Wang:2016:LSA

[WW16] Qi Wang and Ludo Waltman. Large-scale analysis of the accuracy of the journal classification systems of Web of Science and Scopus. Journal of Informetrics, 10(2):347–364, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157715301930.

Wang:2014:HFD

[WWML14] Xianwen Wang, Zhi Wang, Wenli Mao, and Chen Liu. How far does scientific community look back? Journal of Informetrics, 8(3):562–568, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000480.

Wang:2010:ECB

[WWW10] Shuhua Wang, Hengjun Wang, and Paul R. Weldon. Effect of cooperation between Chinese scientific journals and international publishers on journals’ impact factor. Journal of Informetrics, 4(3):233–238, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000868.

Wu:2024:DCJ

[WWXL24] Dengsheng Wu, Shuwen Wang, Weixuan Xu, and Jianping Li. Do conference-journal articles receive more citations? A case study in physics. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724001020.

Wang:2021:ECS

[WWY21] Ruby W. Wang, Shelia X. Wei, and Fred Y. Ye. Extracting a core structure from heterogeneous information network using h-subnet and meta-path strength. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000444.

Wang:2021:LSC

[WXML21] Jingjing Wang, Shuqi Xu, Manuel S. Mariani, and Linyuan Lü. The local structure of citation networks uncovers expertselected milestone papers. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000912.

Wang:2012:ESW

[WXP^{+}12] Xianwen Wang, Shenmeng Xu, Lian Peng, Zhi Wang, Chuanli Wang, Chunbo Zhang, and Xianbing Wang. Exploring scientists’ working timetable: Do scientists often work overtime? Journal of Informetrics, 6(4):655–660, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157712000545.

Wang:2024:EMA

[WXZ24] Yuzhuo Wang, Yi Xiang, and Chengzhi Zhang. Exploring motivations for algorithm mention in the domain of natural language processing: a deep learning approach.

Journal of Informetrics, 18(4):??, November 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000634.

Wang:2021:PIB

[WY21] Jean J. Wang and Fred Y. Ye. Probing into the interactions between papers and patents of new CRISPR/CAS9 technology: a citation comparison. Journal of Informetrics, 15(4): ??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000602.

Wang:2023:UET

[WYC23] Cheng-Jun Wang, Lihan Yan, and Haochuan Cui. Unpacking the essential tension of knowledge recombination: Analyzing the impact of knowledge spanning on citation impact and disruptive innovation. Journal of Informetrics, 17(4): ??, November 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000767.

Wu:2017:DKD

[WYH17] Chaojiang Wu, Erjia Yan, and Chelsey Hill. Disciplinary knowledge diffusion in business research. Journal of Informetrics, 11(2):655–668, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303455.

Wu:2018:DIR

[WYLL18] Dengsheng Wu, Lili Yuan, Ruoyun Li, and Jianping Li. Decomposing inequality in research funding by university– institute sub-group: a three-stage nested Theil index. Journal of Informetrics, 12(4):1312–1326, November 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718301640.

Wang:2021:GBP

[WYLW21] Yukai Wang, Zhongkai Yang, Lanjian Liu, and Xianwen Wang. Gender bias in patenting process. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719303724.

Wang:2012:DCB

[WYX^{+}12] Mingyang Wang, Guang Yu, Jianzhong Xu, Huixin He, Daren Yu, and Shuang An. Development a case-based classifier for predicting highly cited papers. Journal of Informetrics, 6(4):586–599, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157712000454.

Wolfram:2014:CJS

[WZ14] Dietmar Wolfram and Yuehua Zhao. A comparison of journal similarity across six disciplines using citing discipline analysis. Journal of Informetrics, 8(4):840–853, October 2014. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157714000790.

Wang:2018:PAK

[WZ18] Jue Wang and Liwei Zhang. Proximal advantage in knowledge diffusion: the time dimension. Journal of Informetrics, 12(3):858–867, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157718301263.

Wang:2021:IRS

[WZ21a] Xing Wang and Zhihui Zhang. Improving the reliability of short-term citation impact indicators by taking into account the correlation between short- and long-term citation impact. Journal of Informetrics, 14(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719302603.

Wang:2021:UFT

[WZ21b] Yuzhuo Wang and Chengzhi Zhang. Using the full-text content of academic articles to identify and evaluate algorithm entities in the domain of natural language processing. Journal of Informetrics, 14(4):??, November 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720300985.

Wang:2024:EFM

[WZCC24] Zhongyi Wang, Haoxuan Zhang, Jiangping Chen, and Haihua Chen. An effective framework for measuring the novelty of scientific articles through integrated topic modeling and cloud model. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157724000993.

Wang:2023:CTT

[WZHH23] Yajie Wang, Alesia Zuccala, Haiyan Hou, and Zhigang Hu. Corrigendum to [‘To tweet or not to tweet?’ A study of the use of Twitter by scholarly book publishers in Social Sciences and Humanities’, Journal of Informetrics, Volume 15, (2021) 1–14/Article 101170]. Journal of Informetrics, 17(2): ??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722001043. See [WHH21].

Xu:2021:ERT

[XHA^{+}21] Shuo Xu, Liyuan Hao, Xin An, Guancan Yang, and Feifei Wang. Emerging research topics detection with multiple machine learning models. Journal of Informetrics, 13(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157719300367.

Xie:2021:PPP

[Xie21a] Zheng Xie. Predicting publication productivity for researchers: a piecewise Poisson model. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1751157719302676.

Xie:2021:PNC

[Xie21b] Zheng Xie. Predicting the number of coauthors for researchers: a learning model. Journal of Informetrics, 14 (2):??, May 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157719303852.

Xu:2021:ECI

[XLG21] Ran Xu, Arash Baghaei Lakeh, and Navid Ghaffarzadegan. Examining the characteristics of impactful research topics: a case of three decades of HIV–AIDS research. Journal of Informetrics, 15(1):??, February 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720306398.

Xiao:2014:KDP

[XLLZ14] Yu Xiao, Louis Y. Y. Lu, John S. Liu, and Zhili Zhou. Knowledge diffusion path analysis of data quality literature: a main path analysis. Journal of Informetrics, 8(3):594–605, July 2014. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157714000492.

Xie:2024:DLT

[XLSW24] Zheng Xie, Yiqin Lv, Yiping Song, and Qi Wang. Data labeling through the centralities of co-reference networks improves the classification accuracy of scientific papers. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000117.

Xiong:2018:ARE

[XlYcG18] Xi Xiong, Guo liang Yang, and Zhong cheng Guan. Assessing R&D efficiency using a two-stage dynamic DEA model: a case study of research institutes in the Chinese Academy of Sciences. Journal of Informetrics, 12(3):784–805, August 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718300105.

Xiong:2021:PDB

[XlYcG21] Xi Xiong, Guo liang Yang, and Zhong cheng Guan. A parallel DEA-based method for evaluating parallel independent subunits with heterogeneous outputs. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719302792.

Xu:2021:UER

[XMLM21] Shuqi Xu, Manuel Sebastian Mariani, Linyuan Lü, and Matús Medo. Unbiased evaluation of ranking metrics reveals consistent performance in science and technology citation data. Journal of Informetrics, 14(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719301646.

Xu:2014:CTS

[XMM14] Han Xu, Eric Martin, and Ashesh Mahidadia. Contents and time sensitive document ranking of scientific literature. Journal of Informetrics, 8(3):546–561, July 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000455.

Xu:2024:RAS

[XMW^{+}24] Shuo Xu, Xinyi Ma, Hong Wang, Xin An, and Ling Li. A recommendation approach of scientific non-patent literature on the basis of heterogeneous information network. Journal of Informetrics, 18(4):??, November 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000701.

Xie:2016:GGM

[XOL16] Zheng Xie, Zhenzheng Ouyang, and Jianping Li. A geometric graph model for coauthorship networks. Journal of Informetrics, 10(1):299–311, February 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715300912.

Xu:2021:CIP

[XOMW21] Fang Xu, Guiyan Ou, Tingcan Ma, and Xianwen Wang. The consistency of impact of preprints and their journal publications. Journal of Informetrics, 15(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000249.

Xie:2021:SAF

[XWK^{+}21] Qing Xie, Jiamin Wang, Giyeong Kim, Soobin Lee, and Min Song. A sensitivity analysis of factors influential to the popularity of shared data in data repositories. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000134.

Xie:2024:DHT

[XWW^{+}24] Yundong Xie, Qiang Wu, Yezhu Wang, Li Hou, and Yuanyuan Liu. Does the handling time of scientific papers relate to their academic impact and social attention? evidence from nature, science, and PNAS. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000178.

Xu:2021:TLI

[XWY^{+}21] Haiyun Xu, Jos Winnink, Zenghui Yue, Ziqiang Liu, and Guoting Yuan. Topic-linked innovation paths in science and technology. Journal of Informetrics, 14(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771930210X.

Xie:2021:ISL

[XWZL21] Yundong Xie, Qiang Wu, Peng Zhang, and Xingchen Li. Information Science and Library Science (IS-LS) journal subject categorisation and comparison based on editorship information. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719305024.

Xing:2021:SCS

[XWZY21] Yanmeng Xing, Fenghua Wang, An Zeng, and Fan Ying. Solving the cold-start problem in scientific credit allocation. Journal of Informetrics, 15(3):??, August 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000286.

Xu:2022:IMD

[XYP^{+}22] Haiyun Xu, Zenghui Yue, Hongshen Pang, Ehsan Elahi, Jing Li, and Lu Wang. Integrative model for discovering linked topics in science and technology. Journal of Informetrics, 16 (2):??, May 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722000177.

Xie:2021:MMT

[XZDS21] Qing Xie, Xinyuan Zhang, Ying Ding, and Min Song. Monolingual and multilingual topic analysis using LDA and BERT embeddings. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719305127.

Xie:2022:EIC

[XZKS22] Qing Xie, Xinyuan Zhang, Giyeong Kim, and Min Song. Exploring the influence of coauthorship with top scientists on researchers’ affiliation, research topic, productivity, and impact. Journal of Informetrics, 16(3):??, August 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000669.

Xie:2021:NEB

[XZS21] Qing Xie, Xinyuan Zhang, and Min Song. A network embedding-based scholar assessment indicator considering four facets: Research topic, author credit allocation, fieldnormalized journal impact, and published time. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000729.

Yun:2021:RBC

[YAL21] Jinhyuk Yun, Sejung Ahn, and June Young Lee. Return to basics: Clustering of scientific literature using structural information. Journal of Informetrics, 14(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157720301978.

Yan:2014:RDM

[Yan14] Erjia Yan. Research dynamics: Measuring the continuity and popularity of research topics. Journal of Informetrics, 8(1): 98–110, January 2014. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157713000874.

Yang:2024:TDK

[Yan24a] Alex J. Yang. On the temporal diversity of knowledge in science. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724001068.

Yang:2024:UID

[Yan24b] Alex J. Yang. Unveiling the impact and dual innovation of funded research. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723001050.

Yu:2022:CTR

[YAT^{+}22] Shuo Yu, Fayez Alqahtani, Amr Tolba, Ivan Lee, Tao Jia, and Feng Xia. Collaborative team recognition: a core plus extension structure. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722000980.

Yang:2016:UML

[YAY^{+}16] Guoliang Yang, Per Ahlgren, Liying Yang, Ronald Rousseau, and Jielan Ding. Using multi-level frontiers in DEA models to grade countries/territories. Journal of Informetrics, 10(1):238–253, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S1751157715301413. See comment [KE17] and reply [YAY^{+}17].

Yang:2017:RCS

[YAY^{+}17] Guoliang Yang, Per Ahlgren, Liying Yang, Ronald Rousseau, and Jielan Ding. Reply to ‘Comment on “Using multi-level frontiers in DEA models to grade countries/territories” by G.-l. Yang et al. [Journal of Informetrics 10(1) (2016), 238– 253]’. Journal of Informetrics, 11(3):647–648, August 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157717301827.

Yang:2022:IKS

[YBL^{+}22] Jinqing Yang, Yi Bu, Wei Lu, Yong Huang, Jiming Hu, Shengzhi Huang, and Li Zhang. Identifying keyword sleeping beauties: a perspective on the knowledge diffusion process. Journal of Informetrics, 16(1):??, February 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721001103.

Yan:2013:BEV

[YDCL13] Erjia Yan, Ying Ding, Blaise Cronin, and Loet Leydesdorff. A bird’s-eye view of scientific trading: Dependency relations among fields of science. Journal of Informetrics, 7(2):249– 264, April 2013. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157712000983.

Yan:2012:TDR

[YDMS12] Erjia Yan, Ying Ding, Stasa Milojević, and Cassidy R. Sugimoto. Topics in dynamic research communities: an exploratory study for the field of information retrieval. Journal of Informetrics, 6(1):140–153, January 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000976.

Yu:2015:TDU

[YDS^{+}15] Qi Yu, Ying Ding, Min Song, Sungjeon Song, Jianhua Liu, and Bin Zhang. Tracing database usage: Detecting main paths in database link networks. Journal of Informetrics, 9(1):1–15, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000947.

Yang:2023:DCS

[YDW^{+}23] Alex J. Yang, Sanhong Deng, Hao Wang, Yiqin Zhang, and Wenxia Yang. Disruptive coefficient and 2-step disruptive coefficient: Novel measures for identifying vital nodes in complex networks. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000366.

Yu:2008:ORD

[YDWC08] Hairong Yu, Mari Davis, Concepción S. Wilson, and Fletcher T. H. Cole. Object-relational data modelling for informetric databases. Journal of Informetrics, 2(3):240–251, July 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157708000291.

Yu:2023:DFE

[YF23] Dejian Yu and Anran Fang. Developing a four-entities reinforced rank model to evaluate the topic influence in academic networks. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723000470.

Yan:2024:MER

[YF24] Zhaoping Yan and Kaiyu Fan. A multi-entity reinforced main path analysis: Heterogeneous network embedding considering knowledge proximity. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724001056.

Yan:2014:PRC

[YG14] Erjia Yan and Raf Guns. Predicting and recommending collaborations: an author-, institution-, and country-level analysis. Journal of Informetrics, 8(2):295–309, April 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157714000091.

Yang:2018:ISN

[YHS18] Chao Yang, Cui Huang, and Jun Su. An improved SAO network-based method for technology trend analysis: a case study of graphene. Journal of Informetrics, 12(1):271–286, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S1751157717301645.

Yang:2016:VIS

[YHWZ16] Siluo Yang, Ruizhen Han, Dietmar Wolfram, and Yuehua Zhao. Visualizing the intellectual structure of information science (2006–2015): Introducing author keyword coupling analysis. Journal of Informetrics, 10(1):132–150, February 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715301103.

Yoon:2016:OPD

[YKKS16] Hyui Geon Yoon, Hyungjun Kim, Chang Ouk Kim, and Min Song. Opinion polarity detection in Twitter data combining shrinkage regression and topic modeling. Journal of Informetrics, 10(2):634–644, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715301012.

Yang:2022:ECB

[YL22] Jinqing Yang and Zhifeng Liu. The effect of citation behaviour on knowledge diffusion and intellectual structure. Journal of Informetrics, 16(1):??, February 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000961.

Yu:2022:ECP

[YLCC22] Houqiang Yu, Longfei Li, Xueting Cao, and Tao Chen. Exploring country’s preference over news mentions to academic papers. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722000992.

Yin:2018:DCS

[YLZ18] Zhifeng Yin, Zheng Liang, and Qiang Zhi. Does the concentration of scientific research funding in institutions promote knowledge output? Journal of Informetrics, 12(4):1146– 1159, November 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157717300421.

Yu:2021:TMP

[YP21] Dejian Yu and Tianxing Pan. Tracing the main path of interdisciplinary research considering citation preference: a case from blockchain domain. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721000079.

Yadav:2022:TEN

[YP22] Pratyush Yadav and Nargis Pervin. Towards efficient navigation in digital libraries: Leveraging popularity, semantics and communities to recommend scholarly articles. Journal of Informetrics, 16(4):??, November 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772200089X.

You:2022:DQP

[YPL^{+}22] Taekho You, Jinseo Park, June Young Lee, Jinhyuk Yun, and Woo-Sung Jung. Disturbance of questionable publishing to academia. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722000463.

Yoon:2023:QKS

[YPYJ23] Jisung Yoon, Jinseo Park, Jinhyuk Yun, and Woo-Sung Jung. Quantifying knowledge synchronization with the networkdriven approach. Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000809.

Ye:2008:PLM

[YR08] Fred Y. Ye and Ronald Rousseau. The power law model and total career h-index sequences. Journal of Informetrics, 2(4):288–297, October 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157708000473.

Yu:2021:IDM

[YS21] Dejian Yu and Libo Sheng. Influence difference main path analysis: Evidence from DNA and blockchain domain citation networks. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157721000572.

Yu:2021:BBY

[YSJ21] Xiaoyao Yu, Boleslaw K. Szymanski, and Tao Jia. Become a better you: Correlation between the change of research direction and the change of scientific performance. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772100064X.

Yu:2009:SJE

[YSPW09] Liping Yu, Xiaoming Shen, Yuntao Pan, and Yishan Wu. Scholarly journal evaluation based on panel data analysis. Journal of Informetrics, 3(4):312–320, October 2009. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157709000352.

Yoon:2023:EKC

[YSS23] Jeeyoung Yoon, Arida Ferti Syafiandini, and Min Song. Exploring the knowledge certainty shift: Metaknowledge analysis on drugs via assertion uncertainty burstiness. Journal of Informetrics, 17(2):??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000378.

Yun:2022:GBC

[Yun22] Jinhyuk Yun. Generalization of bibliographic coupling and co-citation using the node split network. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157722000438.

Yuret:2016:IEJ

[Yur16] Tolga Yuret. Interfield equality: Journals versus researchers. Journal of Informetrics, 10(4):1196–1206, November 2016. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716300670.

Yuret:2017:AFE

[Yur17a] Tolga Yuret. An analysis of the foreign-educated elite academics in the United States. Journal of Informetrics, 11(2): 358–370, May 2017. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S175115771630253X.

Yuret:2017:DRP

[Yur17b] Tolga Yuret. Do researchers pay attention to publication subsidies? Journal of Informetrics, 11(2):423–434, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157716302553.

Yang:2013:BAA

[YVW^{+}13] Jiansheng Yang, Michael W. Vannier, Fang Wang, Yan Deng, Fengrong Ou, James Bennett, Yang Liu, and Ge Wang. A bibliometric analysis of academic publication and NIH funding. Journal of Informetrics, 7(2):318–324, April 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771200096X.

Yang:2015:VIS

[YW15] Siluo Yang and Feifei Wang. Visualizing information science: Author direct citation analysis in China and around the world. Journal of Informetrics, 9(1):208–225, January 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715000024.

Yin:2017:TDS

[YW17] Yian Yin and Dashun Wang. The time dimension of science: Connecting the past to the future. Journal of Informetrics, 11(2):608–621, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157717300020.

Yang:2024:ETC

[YW24] Wenlong Yang and Yang Wang. Exploring team creativity: the nexus between freshness and experience. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724001007.

Yin:2023:MRS

[YWS23] Deyun Yin, Zhao Wu, and Sotaro Shibayama. Measuring risk in science. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723000512.

Ye:2023:RFD

[YWW^{+}23] Guanghui Ye, Cancan Wang, Chuan Wu, Ze Peng, Jinyu Wei, Xiaoying Song, Qitao Tan, and Lanqi Wu. Research frontier detection and analysis based on research grants information: a case study on health informatics in the US. Journal of Informetrics, 17(3):??, August 2023. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000469.

Yang:2023:KSH

[YWZ^{+}23] Alex Jie Yang, Linwei Wu, Qi Zhang, Hao Wang, and Sanhong Deng. The k-step h-index in citation networks at the paper, author, and institution levels. Journal of Informetrics, 17(4):??, November 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000810.

Yu:2024:EDR

[YX24] Dejian Yu and Bo Xiang. An ESTs detection research based on paper entity mapping: Combining scientific text modeling and neural prophet. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000646.

Yu:2017:GSD

[YXX^{+}17] Houqiang Yu, Shenmeng Xu, Tingting Xiao, Brad M. Hemminger, and Siluo Yang. Global science discussed in local altmetrics: Weibo and its comparison with Twitter. Journal of Informetrics, 11(2):466–482, May 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716301961.

Yu:2018:TLF

[YXX18] Houqiang Yu, Shenmeng Xu, and Tingting Xiao. Is there Lingua Franca in informal scientific communication? Evidence from language distribution of scientific tweets. Journal of Informetrics, 12(3):605–617, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771730442X.

Yu:2019:WPS

[YXXW19] Houqiang Yu, Tingting Xiao, Shenmeng Xu, and Yuefen Wang. Who posts scientific tweets? An investigation into the productivity, locations, and identities of scientific tweeters. Journal of Informetrics, 13(3):841–855, August 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157718303079.

Yang:2024:FAN

[YXXZ24] Guancan Yang, Jiaxin Xing, Shuo Xu, and Yuntian Zhao. A framework armed with node dynamics for predicting technology convergence. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157724000956.

Yu:2023:MPA

[YY23] Dejian Yu and Zhaoping Yan. Main path analysis considering citation structure and content: Case studies in different domains. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000068.

Yu:2014:CMD

[YYW14] Tian Yu, Guang Yu, and Ming-Yang Wang. Classification method for detecting coercive self-citation in journals. Journal of Informetrics, 8(1):123–135, January 2014. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157713000898.

Yan:2015:IES

[YZ15] Erjia Yan and Yongjun Zhu. Identifying entities from scientific publications: a comparison of vocabulary- and modelbased methods. Journal of Informetrics, 9(3):455–465, July 2015. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157715000474.

Yan:2018:RUR

[YZ18] Weiwei Yan and Yin Zhang. Research universities on the ResearchGate social networking site: an examination of institutional differences, research activity level, and social networks formed. Journal of Informetrics, 12(1):385–400, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717301384.

Yin:2022:ISM

[YZ22] Xingbo Yin and Xiaohua Zong. International student mobility spurs scientific research on foreign countries: Evidence from international students studying in China. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000985.

Yan:2013:CIW

[YZF13] Xiangbin Yan, Li Zhai, and Weiguo Fan. C-index: a weighted network node centrality measure for collaboration competence. Journal of Informetrics, 7(1):223–239, January 2013. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157712000946.

Yang:2021:ACS

[YZYW21] Siluo Yang, Mengxue Zheng, Yonghao Yu, and Dietmar Wolfram. Are Altmetric.com scores effective for research impact evaluation in the social sciences and humanities? Journal of Informetrics, 15(1):??, February 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157720306374.

Zhao:2021:IRB

[ZBK^{+}21] Zhenyue Zhao, Yi Bu, Lele Kang, Chao Min, Yiyang Bian, Li Tang, and Jiang Li. An investigation of the relationship between scientists’ mobility to/from China and their research performance. Journal of Informetrics, 14(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719301889.

Chen:2013:PDM

[zChHY13] Dar zen Chen, Mu hsuan Huang, and Fred Y. Ye. A probe into dynamic measures for h-core and h-tail. Journal of Informetrics, 7(1):129–137, January 2013. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000843.

Zhang:2024:UMP

[ZCM^{+}24] Wei Zhang, Juyang Cao, Manuel Sebastian Mariani, Zhen-Zhen Wang, Mingyang Zhou, Wei Chen, and Hao Liao. Uncovering milestone papers: a network diffusion and game theory approach. Journal of Informetrics, 18(3):??, August 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157724000580.

Zheng:2023:EPR

[ZCTN23] Xiang Zheng, Jiajing Chen, Alison Tollas, and Chaoqun Ni. The effectiveness of peer review in identifying issues leading to retractions. Journal of Informetrics, 17(3):??, August 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157723000482.

Zoller:2016:PVE

[ZDJ^{+}16] Daniel Zoller, Stephan Doerfel, Robert Jäschke, Gerd Stumme, and Andreas Hotho. Posted, visited, exported: Altmetrics in the social tagging system BibSonomy. Journal of Informetrics, 10(3):732–749, August 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715300936.

Zhao:2022:UCN

[ZF22] Qihang Zhao and Xiaodong Feng. Utilizing citation network structure to predict paper citation counts: a deep learning approach. Journal of Informetrics, 16(1):??, February 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157721001061.

Zhang:2012:PPJ

[ZG12] Lin Zhang and Wolfgang Glänzel. Proceeding papers in journals versus the “regular” journal publications. Journal of Informetrics, 6(1):88–96, January 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771100071X.

Zahedi:2018:RBB

[ZH18] Zohreh Zahedi and Stefanie Haustein. On the relationships between bibliographic characteristics of scientific documents and citation and Mendeley readership counts: a large-scale analysis of Web of Science publications. Journal of Informetrics, 12(1):191–202, February 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S1751157717302213.

Zhang:2024:OAD

[Zha24] Xijie Zhang. Is open access disrupting the journal business? A perspective from comparing full adopters, partial adopters, and non-adopters. Journal of Informetrics, 18(4): ??, November 2024. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157724000865.

Zhang:2022:DGN

[ZHS^{+}22] Ning Zhang, Guangye He, Dongbo Shi, Zhenyue Zhao, and Jiang Li. Does a gender-neutral name associate with the research impact of a scientist? Journal of Informetrics, 16 (1):??, February 2022. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157722000037.

Zitt:2010:CSN

[Zit10] Michel Zitt. Citing-side normalization of journal impact: a robust variant of the Audience Factor. Journal of Informetrics, 4(3):392–406, July 2010. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115771000026X.

Zitt:2016:PWP

[Zit16] Michel Zitt. Paving the way or pushing at open doors? A comment on Abramo and D’Angelo “Farewell to sizeindependent indicators”. Journal of Informetrics, 10(2):675– 678, May 2016. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157716300906. See [AD16b,

GTD16, AD16c].

Zhu:2023:ERS

[ZJMX23] Wanying Zhu, Ching Jin, Yifang Ma, and Cong Xu. Earlier recognition of scientific excellence enhances future achievements and promotes persistence. Journal of Informetrics, 17 (2):??, May 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000330.

Zhu:2024:DRI

[ZKJ^{+}24] Yongjun Zhu, Donghun Kim, Ting Jiang, Yi Zhao, Jiangen He, Xinyi Chen, and Wen Lou. Dependency, reciprocity, and informal mentorship in predicting long-term research collaboration: a co-authorship matrix-based multivariate time series analysis. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157723001116.

Zhou:2011:FCC

[ZL11] Ping Zhou and Loet Leydesdorff. Fractional counting of citations in research evaluation: a cross- and interdisciplinary assessment of the Tsinghua University in Beijing. Journal of Informetrics, 5(3):360–368, July 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000113.

Zhang:2010:SCA

[ZLJ^{+}10] Lin Zhang, Xinhai Liu, Frizo Janssens, Liming Liang, and Wolfgang Glänzel. Subject clustering analysis based on ISI category classification. Journal of Informetrics, 4(2):185– 193, April 2010. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157709000832.

Zhang:2018:DDL

[ZLL^{+}18] Yi Zhang, Jie Lu, Feng Liu, Qian Liu, Alan Porter, Hongshu Chen, and Guangquan Zhang. Does deep learning help topic extraction? A kernel k-means clustering method with word embedding. Journal of Informetrics, 12(4):1099–1117, November 2018. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157718300257.

Zhang:2016:FRS

[ZLR16] Jin Zhang, Guannan Liu, and Ming Ren. Finding a representative subset from large-scale documents. Journal of Informetrics, 10(3):762–775, August 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716300566.

Zhang:2021:CRD

[ZLW21] Chengzhi Zhang, Lifan Liu, and Yuzhuo Wang. Characterizing references from different disciplines: a perspective of citation content analysis. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721000055.

Zhu:2021:TSR

[ZLY21] Nibing Zhu, Chang Liu, and Zhilin Yang. Team size, research variety, and research performance: Do coauthors’ coauthors matter? Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157721000766.

Zhou:2021:PPS

[ZLYC21] Yanbo Zhou, Qu Li, Xuhua Yang, and Hongbing Cheng. Predicting the popularity of scientific publications by an agebased diffusion model. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000481.

Zheng:2024:TKS

[ZMBP24] Zhejun Zheng, Yaxue Ma, Zhichao Ba, and Lei Pei. Tree knowledge structure for better insight: Capturing biomedical science-technology knowledge linkage with MeSH. Journal of Informetrics, 18(4):??, November 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000816.

Zhang:2024:ASC

[ZMZ24] Dayong Zhang, Hao Men, and Zhaoxin Zhang. Assessing the stability of collaboration networks: a structural cohesion analysis perspective. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157724000038.

Zhao:2011:DBM

[ZRY11] Star X. Zhao, Ronald Rousseau, and Fred Y. Ye. hdegree as a basic measure in weighted networks. Journal of Informetrics, 5(4):668–677, October 2011. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157711000691.

Zhao:2008:CAA

[ZS08] Dangzhi Zhao and Andreas Strotmann. Comparing allauthor and first-author co-citation analyses of information science. Journal of Informetrics, 2(3):229–239, July 2008. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S175115770800028X.

Zogala-Siudem:2023:IRC

[ZSCSG23] Barbara Zogała-Siudem, Anna Cena, Grzegorz Siudem, and Marek Gagolewski. Interpretable reparameterisations of citation models. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157722001080.

Zhang:2023:UDD

[ZSF^{+}23] Tongyang Zhang, Ran Sun, Julia Fensel, Andrew Yu, Yi Bu, and Jian Xu. Understanding the domain development through a word status observation model. Journal of Informetrics, 17(2):??, May 2023. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S1751157723000202. See corrigendum [ZSF^{+}24].

Zhang:2024:CSU

[ZSF^{+}24] Tongyang Zhang, Ran Sun, Julia Fensel, Andrew Yu, Yi Bu, and Jian Xu. Corrigendum to “Understanding the domain development through a word status observation model” [Journal of Informetrics 17 (2023) 1–19/101395]. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000105. See [ZSF^{+}23].

Zhang:2016:HSM

[ZSH^{+}16] Yi Zhang, Lining Shang, Lu Huang, Alan L. Porter, Guangquan Zhang, Jie Lu, and Donghua Zhu. A hybrid similarity measure method for patent portfolio analysis. Journal of Informetrics, 10(4):1108–1130, November 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157715302169.

Zhang:2022:HPI

[ZSP^{+}22] Lin Zhang, Mengting Sun, Yujie Peng, Wenjing Zhao, Lixin Chen, and Ying Huang. How public investment fuels innovation: Clues from government-subsidized USPTO patents. Journal of Informetrics, 16(3):??, August 2022. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000657.

Zhang:2021:CBA

[ZSS21] Tianjiao Zhang, Jin Shi, and Lingyun Situ. The correlation between author–editorial cooperation and the author’s publications in journals. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157720306404.

Zhai:2017:MUM

[ZSWD17] Yujia Zhai, Shaojing Sun, Fang Wang, and Ying Ding. Multiplicity and uncertainty: Media coverage of autism causation. Journal of Informetrics, 11(3):873–887, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157717301323.

Zhang:2024:AIC

[ZSZ24] Jinzhu Zhang, Jialu Shi, and Peiyu Zhang. An approach for identifying complementary patents based on deep learning. Journal of Informetrics, 18(3):??, August 2024. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000749.

Zhang:2011:DRL

[ZTG11] Lin Zhang, Bart Thijs, and Wolfgang Glänzel. The diffusion of H-related literature. Journal of Informetrics, 5(4):583– 593, October 2011. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157711000605.

Zhang:2021:PFI

[ZW21] Fang Zhang and Shengli Wu. Predicting future influence of papers, researchers, and venues in a dynamic academic network. Journal of Informetrics, 14(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719301786.

Zeng:2021:ACS

[ZWBA21] Tong Zeng, Longfeng Wu, Sarah Bratt, and Daniel E. Acuna. Assigning credit to scientific datasets using article citation networks. Journal of Informetrics, 14(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719301841.

Zhang:2024:DCI

[ZWDH24] Yang Zhang, Yang Wang, Haifeng Du, and Shlomo Havlin. Delayed citation impact of interdisciplinary research. Journal of Informetrics, 18(1):??, February 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157723000937.

Zhang:2021:RMAa

[ZWDS21a] Baolong Zhang, Hao Wang, Sanhong Deng, and Xinning Su. RETRACTED: Measurement and analysis of Chinese journal discriminative capacity. Journal of Informetrics, 14(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719302652. See retraction

notice [ZWDS21b].

Zhang:2021:RMAb

[ZWDS21b] Baolong Zhang, Hao Wang, Sanhong Deng, and Xinning Su. Retracted: Measurement and analysis of Chinese journal discriminative capacity. Journal of Informetrics, 15(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157721000031. See [ZWDS21a].

Zhang:2019:RSA

[ZWG^{+}19] Yu Zhang, Min Wang, Florian Gottwalt, Morteza Saberi, and Elizabeth Chang. Ranking scientific articles based on bibliometric networks with a weighting scheme. Journal of Informetrics, 13(2):616–634, May 2019. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718302396.

Zhang:2024:IGC

[ZWL^{+}24] Ming-Ze Zhang, Tang-Rong Wang, Peng-Hui Lyu, Qi-Mei Chen, Ze-Xia Li, and Eric W. T. Ngai. Impact of gender composition of academic teams on disruptive output. Journal of Informetrics, 18(2):??, May 2024. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157724000336.

Zhang:2017:SES

[ZWM17] Ruizhi Zhang, Jian Wang, and Yajun Mei. Search for evergreens in science: a functional data analysis. Journal of Informetrics, 11(3):629–644, August 2017. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716303583.

Zhang:2021:BLN

[ZWM^{+}21] Yi Zhang, Mengjia Wu, Wen Miao, Lu Huang, and Jie Lu. Bi-layer network analytics: a methodology for characterizing emerging general-purpose technologies. Journal of Informetrics, 15(4):??, November 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000730.

Zhou:2021:IPW

[ZWZZ21] Yuhao Zhou, Ruijie Wang, An Zeng, and Yi-Cheng Zhang. Identifying prize-winning scientists by a competition-aware ranking. Journal of Informetrics, 14(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1751157719304766.

Zhang:2023:ARC

[ZXH^{+}23] Chengzhi Zhang, Yi Xiang, Wenke Hao, Zhicheng Li, Yuchen Qian, and Yuzhuo Wang. Automatic recognition and classification of future work sentences from academic articles in a specific domain. Journal of Informetrics, 17(1):??, February 2023. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722001262.

Zhang:2021:MIN

[ZXS21] Xinyuan Zhang, Qing Xie, and Min Song. Measuring the impact of novelty, bibliometric, and academicnetwork factors on citation count using a neural network. Journal of Informetrics, 15(2):??, May 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157721000110.

Zhang:2022:UPR

[ZXS^{+}22] Guangyao Zhang, Shenmeng Xu, Yao Sun, Chunlin Jiang, and Xianwen Wang. Understanding the peer review endeavor in scientific publishing. Journal of Informetrics, 16(2):??, May 2022. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157722000165.

Zhao:2012:EDD

[ZY12] Star X. Zhao and Fred Y. Ye. Exploring the directed h-degree in directed weighted networks. Journal of Informetrics, 6(4):619–630, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S1751157712000508.

Zhu:2016:SBD

[ZY16] Yongjun Zhu and Erjia Yan. Searching bibliographic data using graphs: a visual graph query interface. Journal of Informetrics, 10(4):1092–1107, November 2016. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157716301274.

Zhu:2017:EAR

[ZY17] Yongjun Zhu and Erjia Yan. Examining academic ranking and inequality in library and information science through faculty hiring networks. Journal of Informetrics, 11(2):641– 654, May 2017. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115771730041X.

Zhang:2021:MSC

[ZY21] Ronda J. Zhang and Fred Y. Ye. Measuring similarity for clarifying layer difference in multiplex ad hoc duplex information networks. Journal of Informetrics, 14(1):??, February 2021. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1751157719301890.

Zhai:2022:DCN

[ZY22] Li Zhai and Xiangbin Yan. A directed collaboration network for exploring the order of scientific collaboration. Journal of Informetrics, 16(4):??, November 2022. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157722000979.

Zhang:2012:ENE

[ZYCK12] Sifei Zhang, Chien-Chung Yuan, Ke-Chiun Chang, and Yun Ken. Exploring the nonlinear effects of patent H index, patent citations, and essential technological strength on corporate performance by using artificial neural network. Journal of Informetrics, 6(4):485–495, October 2012. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157712000284.

Zhai:2018:BDH

[ZYZ18] Li Zhai, Xiangbin Yan, and Guojing Zhang. Bi-directional h-index: a new measure of node centrality in weighted and directed networks. Journal of Informetrics, 12(1):299–314, February 2018. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S1751157717301591.

Zhou:2012:CBI

[ZZ12] Ping Zhou and Yongfeng Zhong. The citation-based indicator and combined impact indicator — new options for measuring impact. Journal of Informetrics, 6(4):631–638, October 2012. CODEN ???? ISSN 1751-1577 (print), 1875-

5879 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S175115771200051X.

Zuo:2018:MMB

[ZZ18] Zhiya Zuo and Kang Zhao. The more multidisciplinary the better? — the prevalence and interdisciplinarity of research collaborations in multidisciplinary institutions. Journal of Informetrics, 12(3):736–756, August 2018. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718300452.

Zhang:2021:ABD

[ZZ21a] Chengzhi Zhang and Qingqing Zhou. Assessing books’ depth and breadth via multi-level mining on tables of contents. Journal of Informetrics, 14(2):??, May 2021. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157719302238.

Zhou:2021:BCB

[ZZ21b] Qingqing Zhou and Chengzhi Zhang. Breaking community boundary: Comparing academic and social communication preferences regarding global pandemics. Journal of Informetrics, 15(3):??, August 2021. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S175115772100033X.

Zhou:2021:ITC

[ZZ21c] Qingqing Zhou and Chengzhi Zhang. Impacts towards a comprehensive assessment of the book impact by integrating multiple evaluation sources. Journal of Informetrics, 15 (3):??, August 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157721000663.

Zhang:2021:SGI

[ZZL^{+}21] Mengya Zhang, Gupeng Zhang, Yun Liu, Xiaorong Zhai, and Xinying Han. Scientists’ genders and international academic collaboration: an empirical study of Chinese universities and research institutes. Journal of Informetrics, 14(4): ??, November 2021. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157720300262.

Zuo:2019:SSG

[ZZN19] Zhiya Zuo, Kang Zhao, and Chaoqun Ni. Standing on the shoulders of giants? — Faculty hiring in information schools. Journal of Informetrics, 13(1):341–353, February 2019. CO-DEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1751157718304887.

Zhang:2023:EST

[ZZZ^{+}23] Min Zhang, Dongxin Zhang, Yin Zhang, Kristin Yeager, and Taylor N. Fields. An exploratory study of Twitter metrics for measuring user influence. Journal of Informetrics, 17(4): ??, November 2023. CODEN ???? ISSN 1751-1577 (print),

1875-5879 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1751157723000792.